A Bibliography of Publications by, and about, Edward Teller

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

10 September 2024 Version 1.189

Title word cross-reference

+ [KT48, TT35a]. $100 [Smi85b]. $12.95 [Edg91]. $19.50 [Oli69]. $24.95 [Car91]. $3.50 [Dys58]. $30.00 [Kev03]. $32.50 [Edg91]. $35 [Cas01]. $35.00 [Dys02a, Wat03]. $39.50 [Edg91]. $50 [Ano62]. $8.95 [Edg91]. = [TT35a]. ^{−} [BJT69]. ^{7} [CT41]. _{2} [SST71, ST39b, TT35a]. _{3} [HT39a]. _{4} [TT35a]. _{6} [TT35a]. β [GT36, GT37, HS19]. λ2000 [MT42]. SU(3) [GT85].

-Disintegration [GT36]. -Mesons [BJT69]. -Transformation [GT37].

0 [Cas01, Dys02a]. 0-7382-0532-X [Cas01, Dys02a]. 0-8047-1713-3 [Edg91]. 0-8047-1714-1 [Edg91]. 0-8047-1721-4 [Edg91]. 0-8047-1722-2 [Edg91].

1 [Har05]. 1-86094-419-1 [Har05]. 1-903985-12-9 [Tho03]. 100th [KRW05, Tel93d]. 17.25 [Pei87]. 1930 [BW05]. 1930/41 [Fer68]. 1939 [Sei90]. 1939-1945 [Sei90]. 1940 [TT40]. 1941 [TGF41]. 1942 [KW93].

1945 [Sei90]. 1947 [Sei90]. 1947-1977 [Sei90]. 1948 [Tel49a]. 1950s [Sei90]. 1957 [Tel57b]. 1960s [Mla98]. 1963 [Szi87]. 1973 [Kur73]. 1977 [Sei90]. 1979 [WT79]. 1990s [AB88, CT90a, Tel96b]. 1991 [MB92]. 1992 [GER^{+}92]. 1995 [Tel95a].

20 [Goe88]. 2003 [Dys09, LBB^{+}03]. 2008 [LV10]. 20th [Mar10, New03d]. 28 [Tel57b].

3 [Dic79].

40th [MKR87]. 411-415 [Ber03b].

50th [Tel93a].

60th [MF69, Oli69, Wig69, Wig96a].

70th [DB71, Tel95b]. 71 [Ber03b, Her03]. 7th [WD86].

80th [MW88].

90th [CI04].

A-Blasts [Kih67]. AAAS [Ano71, Bof71]. abandoned [NA91]. ABL [Paw02]. Absorption [BJT69, BJ41, MT42, NST40, ST39b]. Abundance [MT50]. abuse [Gre77]. Accelerates [Gon96d]. Accident [Cou79]. Accomplishments [Tel54b]. Accord [Tot63a]. according [Thi46]. Account [Ano15, Sei90, Nic87]. accuse [AA54]. Acheson [Tel46d]. Across [TG58]. Action [FT38, Tel75d]. Active [Gre46, WHIT95]. Adaptation [GEM^{+}65, Tel89a]. address [Tel62a, Tel94c]. Addresses [Tel91c]. Adiabatic [BT78, NT60, Tel67a]. adolescence [TW79b]. Adsorption [BET38, BDDT40, MT51b]. Advance [GBS^{+}80]. Advanced [EHJ70, GCB^{+}77, Tel76a, Tel77a, Dux10]. advanced/research [Dux10]. Advances [NK92, Tel92a]. Adventures [Har05]. adventurous [ABES12]. Advertisement [Cou79]. Advisors [Bad77b, Dys76, Edg91, Hig76, San77, Wig76, Zac76, Zie91, Yor76, Yor89]. Advisors/Oppenheimer [Zac76]. Advisory [Sei90, Syl87]. Advocate [Hou03]. Advocates [W.D75, Tel61c]. AEC [Wig62]. affair [Uda94, Uda98]. Affairs [Ano55, GR63]. Afraid [Tel94e, Gro04]. Africa [Flo88]. After [Tel83d, Wyd84, ESKR84, Mit97, Tel84c]. after-effects [Tel84c]. Aftermath [Tel86g, Tel87d, Tel87h, Wad87]. Again [Lau65, Tel85i, Han79]. Against [Ken63, TT94]. Age [Baz90, Car91, GR63, Her90, McM90, Rho90, Tel65b, Zie91, BP90, GA^{+}86a, GA^{+}86b, Mar96, McK84, TL58a, TZ91, Wag98, Tel49a, Tel93c, Wag81, Sei90]. Agency [Wal08]. Ages [TWH97]. ago [Cow03, TW79b]. Agrees [Ano65].

Aiming [RC87]. Air [Bet46, Flo88, Col94, New03b]. Airborne [Paw02]. Airmail [TZ90]. Alamos [MKR87, Sei99, Bet91, Hec10, Tel54b, Tel83d]. alapján [Vin03]. Alarm [Blu72a]. Alaska [Ano89a, O’N89a]. Albert [Dys58, Sel58, TG58, Ein39, Tel78a]. Alexander [Ano96]. Alexei [Har05]. Algorithm [BS00, CG95, Hit03, Blo84]. Alive [Bri00]. Alkali [LST41, ST38, ST40]. alle [GER^{+}92]. Allen [Fri62, Jon62, Kle62, Mon64, Mor62, Pic62, Rop62, Zuc88]. Alliance [Tel86e]. Allies [Wil86]. allowed [HS19]. along [Hax10]. Alpha [CT41, HT38a]. Alpha-Particle [HT38a]. Alpha-Particles [CT41]. Alpher [AM23]. Also [Bof85]. Alternate [Tel77b]. Alternative [BAA^{+}75, Gir75]. Alternatives [Tel58b, Tel85b]. Alwyn [Sei90]. Am [Her03, Tel86j]. Am. [Ber03b]. Amateur [Har01]. Amendment [Tel46d]. America [Sei90, Van03b, AB88, Bal86, Bar90, MK02, Nic87, Smi65, Smi70, TT^{+}77, Tel78d, TW79a, Tel86j, WT79]. American [Sei90, ALW^{+}88, Aue03, CHW91, DH83, Fit98, Gon96b, Ne’98, Tar18, Tit86, Tit01]. Amerika [Bar90]. Amounts [KMTW43]. Amundson [Boy05]. Analysis [SNST39]. Anders [Sei90]. Andrei [A^{+}91, Rhe95, Rhé11, Rhé11]. Andrew [WS92, WS03]. Angular [CT41]. anharmonic [PT33]. anharmonischen [PT33]. annihilation [Lap62, Mon64]. Anniversary [MKR87, Tel93a, Tel93d, ALW^{+}88]. Annual [TGF41]. answer [Tel73a]. Answers [Tel90b]. Anti [Bud83, Tel86f]. Anti-Missile [Bud83]. Anti-Weapons [Tel86f]. Antiprotons [DT56]. Antwort [Tel73a]. Any [GB89]. applicability [Tel82d]. Application [BT78, MBM^{+}83, MBM^{+}84, ST39b, Tel47b]. Applications [Has70, Tel91a, EGH86, HT39b, Tel91h]. Applied [Tel66b, Tel67b]. Appraisal [TTA^{+}90]. Appreciation [Wig96a, Gol99, Har01, Wig69]. apprentices [Riv96]. apprentis [Riv96]. Approach [Tel94d]. Approaches [WHIT95]. April [HPA97]. APS [Tel68d]. Architect [BS03a, Sul03, Wag81]. Architects [MK23]. Arena [BM04]. Aristid [Erm89]. Armageddon [Pow04, Mar97]. Arms [Sei90, Tel60b, Tel82a, Yor95, Zuc88, Bur84, DFH84, DFH85, HS83, HGS87, McM05]. Army [Sei90, Jon85]. Arnold [Ano38]. Art [Rab56, Jac10]. Articles [GR63, Sei90, Bur84]. Ask [Tot63a]. assessment [DFH84, DFH85, TT^{+}77]. association [AG80, AG81]. Assumptions [MT51a]. Asteroid [TT94]. Asteroids [Par92, GMS94]. astrophysics [MB92]. Asymmetric [Tel41]. Atanasoff [Smi10]. Athens [Har05]. Atmosphere [KMT46, Lei88]. Atmospheric [Bre53, TLCW01]. Atom [Bor57, Bor63, Dru58, Sch80, Sei90, Sul75, Tel56c, CHW91, Uda94, Uda98, Wil87, Zie91]. Atombombe [Tel01b, Thi46]. Atomfizika [Tel98f]. Atomic [Ano57b, Ano58b, Ber03a, Com56, Erm89, GR63, Ken63, Kev03, Ott85, Pal00, Rho86, Sei90, Syl87, TW38, Tel47a, Tel47c, Tel49a, Tel59, Tel68c, WM15, ZA04, Aue03, Bal86, Bet91, BMW89, Bur84, DA87, Dys02b, GA^{+}86a, GA^{+}86b, Gor09, Jon85, KMN^{+}97, Mac86, Mac87, McK84, Telxx, Tel01b, Thi46, Tit86, Tit01, Van03a, Edg91, ST99, Tel48b, Tel93c, Wag81, Boy05, Moo00, Sei90]. atomkor [TZ91, Wag98]. Atomphysik [Thi46]. Atomphysiker [Lin98, Her76]. Atoms [AT43b, BJT72, MB92]. Atomville [MK23]. Attack [Bof85, Tel64a, Wal08]. Attention [Ken54]. attributes [Ano79a]. Aufbaumöglichkeit [dEAH29]. August [WT79, Szi87, Uda94, Uda98]. Ausblick [TL58a]. Automated [TIW^{+}96]. autour [SB55b]. Award [Ano62, Tel92d, Wig62]. Awarded [Ano03a, Sea03c]. awards [New03f]. Away [HM90].

B [Dux10, Har05, Har11a, Sch80, Ano93, CP93, Dav91, Tel81g]. B.E.T. [MT51a]. Back [Bet01, Tel50a, TNR97]. Backs [Bir67]. Backyard [Sei90, Tit86, Tit01]. Bait [Ano95]. Ball [Sei90]. Ban [Ano60b, Ano63b, Tot63a, Tot63b, Ken63]. Bands [MT42, PT32a, PT32b]. Bankhead [dV08a]. barátjának [Tel03a]. Barbara [GG76]. Bardeen [HD02]. base [Ano95]. Based [FMT47, FMT49b, MT05, TWH97, FMT49a]. Basic [Tel60a, KPB73, Tel73a]. Basis [RTW32]. battlefields [MK02, Van03b]. Be [Tel87c, Bet46, Gro62, Gro83, Jon62, Tel70a]. Beams [LBF^{+}75]. because [Tel93e]. become [ALW^{+}88]. Bedrohung [SB55a]. Before [Wyd84]. beget [BCF^{+}97]. Begin [Ste74]. Beginnings [Gon96c, Ber87]. Behavior [ML07, Tel89a]. Behind [BJCP88, Dav91, Dav94, Dav95, Han92a, Bro92, Bro96, Car03]. Being [Ano15]. Beiträge [KRW05]. Bemerkung [Tel30a]. Bemerkungen [FST32, PT33]. Benefits [Bir67, Cou79]. Benzene [LT37, NST40, SNST39]. Berechnung [BT32]. Berichte [KRW05]. Berkeley [Sei99]. Bernstein [Ber03b]. Bersuker [Dux10]. beset [TW79b]. Best [Tel87f]. beszélgetések [Tel08]. BET [Dav94, Dav95, Tel83c]. Beta [HS19]. Bethe [Edg91, Ber80, Bet93, Bet97, Bro82, Goo99, Tel87j]. Betrachtungen [DB71]. Better [Ano60a, TW79a, Tel87e, WT79, Zuc88, Ber90, Hir88, Pei87, Sei90, Tre87]. Between [AT43b, Ano58a, Lin98, Ren34, Tel91d]. Beware [Ano57b]. Beyond [Tel86g, Tel87d, Tel87h, KMN^{+}97, Wad87]. Bibber [Har11a]. bibliográfiával [Wag98]. bibliographic [Ber84a]. bibliography [Wag81, Wag98]. Biennial [AG81, AG80]. Big [Hil15, Tel71d]. Bilder [Tel93b]. Bill [Day97]. binären [RTW32]. Binary [RTW32]. Biographical [Sei90, BMW89, LS09]. Biography [AH04, Cre10, Sei90, BP90, Law88, Lou02, Smi10]. Biological [ESKR84]. Birth [Anoxxa, Hil15, McM05]. Birthday [KRW05, Tar98, Wig96a, CI04, DB71, MF69, MW88, Sei98, Tel95b, Wig69, Oli69]. bis [BW05]. bizonytalansága [TZ91]. biztonság [TZ91]. Black [TG58]. Blast [BO72, Hol04, Ste73, Sul75]. Blasts [Kih67]. Block [Ano57a]. Blooms [Ano71]. blow [LT05]. blue [Fit00]. Blumberg [Bad77a, Ber78, Car91, Her90, McM90, Zie91, Baz90, Lak78, Rho90, She76]. Bohr [vW02, FK85, Tel69d, Tel80b, Telxx]. Boiler [Gre45]. Bomb [All56, Ano60a, Ano65, Ano93, Ano15, Ber03c, Bet82, Bet93, Bir02, Blu88b, BS03a, Dys76, FS95, Fit98, Gon96c, Her03, HM90, Ken54, Kev03, Lan93a, Lau65, Moo95, Rhe95, Rho86, Sch01, Sch80, Sch03a, Sei90, Sul03, Tel62e, TB82, Tel88b, Thi50, Tur79, WM15, Wol67, Wol09, Zor24, Bal03, Ban02, Ber03b, BHM91, Bro82, Bru62, Bur84, Car03, DeW79, Doy02, Dre02, Erm89, For15, Gol84, Han95, Har11b, Her02, HM97, Hol02, Jac10, Jon85, Kno90, KMN^{+}97, Kur86, Law88, Mac86, Mac87, Mag88, MSDP80, Mos68b, Mos68a, Mos72, Nor02, SB54, SB55b, SB55a, SB71, Telxx, Tel01b, Thi46, Van03a, WG17, ZA04, Kev03, MK23, Kni03, Boy05, Han95, Sei90]. Bombardment [MT47, WHIT95]. Bombe [Sch01]. Bombs [KMT46, Lan84, Ste73, Tel50b, Tit86, Tit01, O’N07, Sei90]. Book [Ano05a, Bad77b, Bad77a, Ber90, Ber03c, Bet01, Bir02, Bor03, Boy95, Boy05, Car91, Coo80, Cre10, Day97, Dow74, Dys58, Dys02a, Edg91, Eps92, Fra81, Fri62, Gob49, Hac04, Han92a, Han95, Har05, Har11a, Hed49, Hen05, Her90, Her01, Hig76, Hir88, Jac65, Jon62, Kau12, Kle62, Lib11, Loe05, Mar92, Mar81, Mar97, Mon64, Moo95, Moo00, Moo02, Mul50, Oli69, Pal02, Pei80, Pei87, Pen79, Pic62, PB62, Rop62, Ros80, Sch03a, Sei90, Sel58, Set02, TG58, Tel49a, Tel81e, Tel86g, Tel87h, Tho80, Tho03, Tre87, Van03b, Wan03, Wat03, Wig76, Zie91, Ber03b, Her03, Ano15]. Books [Ber78, Bor12, Cas01, Dys58, McM90, Mor62, Ore58, Pre62, Sei90, Zac76, Bur84]. Boorse [Sei90]. Boost [Poo89]. Bootlegger [Ano57b]. Boys [O’N94, O’N07, Boy95]. Branch [Sew88]. Braun [Riv96]. Breakdown [ST38, ST39a, ST40]. Breakthrough [Tel8xa]. Breeder [MBM^{+}83, MBM^{+}84]. Brian [Kev03, Sch80]. Brickbats [Lan84]. Brief [Tar10, Ano90a, Ano90b, Ano95]. Briefe [KRW05]. Brilliant [Tel91f, Ano90b, Tel91f]. brings [Ber87]. Broad [Bir67, Han92a]. Brotherhood [Ber03b, Doy02, Her02, Hol02, Ber03c, Her03, Kev03, Sch03a]. Brown [Kev03, Mon64, Fri62, Jon62, Kle62, Mor62, Pic62, Rop62]. Brunauer [Dav94, Dav95]. Build [GB89]. Building [For15, dEAH29]. Bulletin [Ano90a, Ano90b, Ano95, Isa93, Rot94, GR63]. Burns [Sei90]. Bursts [JT84].

C [Boy05, Sei90, TG58, TT35a, Tel95b, TT35a]. C. [Erm89]. CA [LV10, MB92]. Café [ST99]. Calculation [SST71, BT32]. Calculations [Bar65, MRR^{+}53, MRR^{+}52]. California [Kni03, San03]. Cambridge [Cas01, Dys02a, Wat03, BRS90]. Came [Tel91l, Mos68a]. Campus [Hal70]. Can [Bet46, Jon62, Tel70a, Gro62, Gro83]. Canal [Ano73]. Canceled [Gor92]. Candid [HH04a]. Capture [BJT72, FT47]. career [ST99, Wag81, Wag98]. Carl [CI04, Ros22]. Carlo [Bar65, BST66, Has70, Pes73]. Carnegie [HPA97]. Case [Joh13, Ken54, Kev03, NAH^{+}89, Rhe95, Tel85a, Gre77, Han79, Pei87, San79]. Case-Study [NAH^{+}89]. Cast [Smi85a]. Casualty [Bra79]. Catalysis [ET40, Nat40, DH83, KT35]. Catastrophe [Dud75, Sag83]. cathedral [Dys12a]. Celebrates [Gui04, ALW^{+}88, Sei98]. celebrating [MW88]. Celebration [KRW05]. Centenary [Gui04, FK85, Har08b]. Centennial [Ano08b, LV10, dV08a, Har11a]. Central [TIW^{+}96]. Century [Bet01, Bir02, Dys02a, Hac04, Her01, Pal02, Set02, Tho03, Wan03, Wat03, Har06b, Mar10, New03d, Alb10, Bor12, CD96, Har10, Sla12, TS01, TB02, LY95, Lib11, Moo02, Kau12, Cre10]. Ceremony [Ano03a, Sea03c]. Chadwell [Sei90]. Chain [BT32, And73]. Chain-Molecule [BT32]. Chains [Has70, Pes73, Tie94]. Challenge [Tel58d]. Change [Tel48c, TWH97, Ano90b, Tel85f]. changed [Har06b, Mar06]. Changes [JT82]. Characteristic [BT32]. Charged [NT60]. Chariot [Boy95, O’N89a]. Charles [MRTM95]. Charlton [Tel86d]. cheap [Ano90a]. Chemical [DTT37, HT39b]. Chemist [Har11b]. Chemistry [EHJ70, Joh13, KBY09, MT87, Tel73c, Tel91b, Ber84b]. Chemists [Joh13, Har08b]. Chen [LY95]. Chicago [AAF^{+}52, Shi91]. Chronicle [Tel78d]. Chronicles [Her01]. Chu [Poo89]. Chuck [Mar97]. Circles [Tay96]. Circus [Ano71]. Cites [Ano60a]. Cities [MTK46]. Citizen [Sei90, Bur84, Zuc88]. Citizens [Tel99a]. city [LT05]. Civil [TW79a, Tel84b, TWB73a, TWB73b, WT79]. Civilian [Ano03c]. Clarifying [LW04a]. Classical [JT55]. clean [Mag88]. Clearing [Col94]. Clerical [Wei75]. Climate [TTA^{+}90]. Climatic [Sag83, Tel85f]. cloning [BCF^{+}97]. Closer [Alb10, Bor12, Cre10, Lib11, Har10, Kau12, Sla12]. cloth [Wat03]. Cloud [Sei90, Gor09, Mil86, TG58]. Clusters [Tel85d]. Cold [HdH46, Poo89, Van03b, ESKR84, MK02, Fit00, Uda94, Uda98]. Collaboration [Ber10, AM23, Dür05]. colleagues [KGA^{+}03]. collected [Wig96b]. College [Har05]. Colliding [BT78]. Collier [Pei87]. Colloquium [Tel90a]. Combustion [Tel72c]. comets [GMS94]. Coming [dV08a]. Commemoration [Sea03a]. Comment [Ber03b, Dre02, Her03, Tel91h, Tel91a]. Commentary [Gol84, Han79, MSDP80, Bet97]. Comments [Bet82, Tel48a, Tel54a]. Commission [Sei90, Syl87, Tel48b]. Committee [Sei90, Syl87, Nat40]. Community [Hab69, Zie91]. Compelling [Tel58c]. competition [Rad04, SB55b, SB55b]. Competitiveness [GT87]. Complacency [Dic79]. Complementarity [Tel69d]. Complete [Gol99]. Completely [TIW^{+}96]. Complex [Hil15]. Component [BST66]. Components [AT43a, ALT43]. Composition [DTT37]. Compounds [KMTW43]. comprehensive [Wag81]. Compression [Tel67a, Tel72b]. Computational [Lei88]. Computer [Ano15, Mac92, Smi10]. Computing [MRR^{+}53, MRR^{+}52]. Concept [TIW^{+}96]. Concept-Level [TIW^{+}96]. conception [SB55b]. concepts [Tel60a]. Concerning [KT48, Tel72b]. Concert [Tel93a]. Condensed [ML07]. Condensed-Matter [ML07]. Conditions [Oli69, TJ72, MF69]. Condon [ST39b]. Conference [Kur73, SBTT39, TT40, TGF41, Tel90a, ESKR84]. Conferences [Anoxxb]. Confers [Ano61a]. Confinement [LBF^{+}75, Tel81f, Tel81g]. confirmed [DeW88]. Conflict [Bad77a, Ber78, Lak78, She76, BO76, BOG89, SB55a]. confrontations [CD96]. Congress [WD86, Mit97, Tot63c]. Congressional [WT79]. connection [Tel92d]. Connections [Lan93b, Lan99]. Consciousness [Bri00]. Consequences [ESKR84]. Conservationist [Tel70a]. Consider [Moh83]. Considerations [KT48]. Consistency [Tot63b]. Constant [Tel54a, Tel72a]. Constant-pressure [Tel54a].

Constantine [BBLS00]. Constants [Tel48c, Tel72a]. Constitution [Bor48, Tel48a]. construction [SB55a]. Constructive [TTHJ68]. Consults [Ano73]. contd [Nor83]. contentious [ST99]. Contributed [AG80, WD86]. Contributions [KRW05, ML07, Wig62, New03e, dV08b]. Control [Tel60b, DFH84, DFH85, HGS87, Zuc88]. Controlled [Tel56d, Tel56a, Tel56b, Fis10, Sem92]. Controversial [Bal03, Bof69, Kir98]. controversy [BHM91]. convenient [SH01]. Convention [Bof71]. Convergent [Tel72b]. Conversation [Goo99, Hei90, Wol67]. Conversations [Sch01, TTT91, TTT02, HH04a, Tel08, Bor03, Eps92, Mar92]. Conversion [Tel69a]. Cooled [TIW^{+}96]. cooling [HT36]. Cooperation [Tel99d]. Coral [Kur73]. Core [GT38b]. Corrections [Ano93, RT39]. Corvin [Byr01]. Cosmic [ART49, RT49b, WHIT95, Tel49b, Tel54c]. cosmological [AM23]. Cosmology [Rei72a, Rei72b]. Costandina [Sei90]. Could [Bof85, Sul75, LT05]. Council [Nat40]. Countdown [Sei90, Kur86]. Counting [Ano95]. Courage [Tel83e]. Courses [Ano60a]. Cover [Lew85]. Cover-Up [Lew85]. Creating [Zor24]. Creativity [NAH^{+}89]. Credibility [Sau75]. credit [Car03]. Crisis [Ano60a, Tel77b]. Critchfield [MRTM95]. Criterion [Dys58]. Critical [Gre46, KMTW43]. Critics [Bar88, Tel85e]. Crossing [Tel37]. Crucial [Tel84b]. Crusade [Zuc88, HGS87]. Cruz [Sei90]. Crystal [ST42, ALT43, FMRT34, Tel34a]. Crystals [FT38, ALT43, BP83, BP89, BP12, Eng72, KZ01, Dow74]. csillagháborús [Bro96]. CTR [Tel72b]. Culture [Boy05, Tel92h, Jac10, ZA04]. Curbs [Bir67]. customer [Ano95]. Cycle [Tel78c]. Cyclotron [Hil15].

D [HH08, Sei90, TG58, KT48]. Dan [Sei90]. Danger [Ken63, Moh83]. Dangerous [Tel47c, Tel82a]. Dangers [Dys58, Mar87, Ore58, Sel58, TG58, TL58b]. Daniel [Boy95, Tel49a]. dans [Dem15]. dargestellt [Thi46]. Dark [Bof85, Bor03, Eps92, Mar92, ESKR84, Han92a, TTT91, TTT02, Han95]. Darkness [Koe40]. Daughter [Moo00, Pal00]. dawn [Gor09]. Day [Kur86, TG58, Wyd84, Bru62, HHW99, Sei90]. Days [Flo88, WM15]. Dead [BS03a, Sul03, BS03b]. Dean [Sei90, DA87, Sei90]. Debate [Ano58a, Anoxxb, PT79, ST60]. Debates [Ros22]. Decades [Sei90, Mil86]. Decay [FTW47, FTW65]. December [Tel57b]. deception [Bro92, Bro96, Mit97, Han92a]. Decide [EN86]. Decision [Dys76, GB89, Tel95c, Tel98b, Fel79a]. decisions [Str63, Jon62]. Dedicated [Sea03b]. Deep [GT56]. Defence [Bud83, CT90a, Pei87]. Defense [Ber90, CT90b, DFH84, DFH85, Hal70, Hir88, IPB^{+}86, Pei87, Smi85a, Tel57a, TW79a, Tel83a, Tel84b, TB84, Tel85h, Tel86e, Tel86h, Tel87f, Tel87i, Tel87j, Tel88a, Tel91g, Tel99d, Tel02b, Tre87, WHIT95, Zuc88, Mit97, TWB73a, TWB73b, Tel77f, Tel87e, WT79, GBS^{+}80]. Defensive [TS84, Zam95]. Defining [Mla98]. degeneracy [JT37]. degenerate [JT37]. degrees [New03f]. del [GER^{+}92]. Delano [Ein39]. Delete [Dev62]. deleted [TdHK51]. della [Tel92f, Tel92g]. dell’Etica [Tel92f, Tel92g]. Demons [Dys12b]. Dempster [TG58]. Densities [Tel71c]. Density [CHL^{+}96, CHL^{+}97, CK71, Ros10]. Department [Ano15, GBS^{+}80]. depression [IT39]. Dep’t [Tel46c]. Derek [TG58]. deserve [Mit97]. Design [TIW^{+}96, Tel47b]. Designer [Zor24]. Desirable [ST60]. destroy [LT05]. destruction [Tay96]. Details [Mar97]. Deterrence [Mac92, Tel88a]. Detonations [TU51]. Deutung [TT32]. developer [BS03b]. Developing [Tel77g]. Development [Ano15, Fit98, Pre62, Tel50b, TS84, WN88, Ban02]. Deviations [BT41, BT45, BT51a, BT51b, BT53]. Device [Ano15, Tel51a]. Devices [Tel87g, Tel92j]. DeWolf [Edg91]. Dialogue [Blu88b]. Diamagnetism [Tel31]. Diamagnetismus [Tel31]. Diary [Sei90, DA87]. Dickson [Moo02]. Dictated [Tel03a]. Dictionary [DG99]. Did [Ano90a]. didn’t [Flo88, MHM90]. Dielectric [ST39a]. Dies [Gor92, Hou03, Bal03, San03]. digital [Dys12a, Smi10]. diktált [Tel03a]. Dimensional [AT43a, ALT43, TW33]. Dimensions [Gre46]. Diplon [KT34]. Dipole [GT48, TGF92]. Dirac [Tel73b]. Directions [MKR87]. Director [Sei90, Law88, Hou03]. Disagree [Bar88]. Disarmament [Ano58a, ST60, TP58, Tel88a]. Disaster [Tel81e, Tel81b, BCF^{+}97, Med79a, Med79b, Med79c]. disclosed [DeW79]. disclosure [BJCP88]. Discovery [BHT86, CHW91]. Discuss [Tel68d]. Discussion [GEM^{+}65, GCB^{+}77, IPB^{+}86, LBF^{+}75, NAH^{+}89]. Disease [Tel75a]. Disintegration [GT36]. Dispersal [MTK46]. Dispersing [WHIT95]. dispersion [LT36]. Dispute [Ano15]. Disputes [Ken54, Ken63]. Dissent [Ano71]. Distant [TT64]. Distinguished [Hou03]. Distributed [CT91, TLCW01]. Distribution [Bar65, CT41, JT54]. Distributions [Tie94]. Disturbing [Dys79a]. Ditfurth [vW02]. diverse [New03e]. dividend [Ano95]. Dividing [Les93]. documentary [CHW91]. does [GBS^{+}80]. dollar [CP93]. Domestic [Tel47d]. Don’t [Ano95, GM87]. Doom [Bof69]. Doppler [JT82]. Double [HT39a]. Doubleday [Tel49a]. Doubt [Sul75]. Doubts [Smi85a]. Downwind [Sei90, Bal86]. Dr [Ano05a, Hen05, Loe05, Sch80, Ber04a, Goo04a]. Dr. [Ano57a, Ano58b, Ano60a, Ano62, Ano79a, Ber04b, Blu72b, Byr01, Cou54, Goo04b, GTS^{+}06, Har08a, Ken54, Rho05, Sea03a, Set07, Tel90b, Wei04]. Drachkovitch [CD96]. Draft [Com48, Szi87, Tel48a]. Drama [Cou63, Her76, Ken54]. Drawing [Ber03d]. Dream [Lan83, Kir05]. dreiatomigen [Ren34]. dreiatomiger [FST32]. Dubious [Moh83]. due [GMS94]. Dug [Ano73]. dumb [Ano95]. During [Ano03a, Sea03c, Sew88]. Dynamical [PW84]. dynamite [Fin90].

E-Flat [Dys79b]. Early [BD82, MBM^{+}83, MBM^{+}84, Tel78e, Tel8xb, Wei94a, Tel91i, Tel49a]. Earth [IT40, NT60, TW79b, Tel98e, Coo80, Pei80, Pen79, Ros80, TG58, Tho80]. earthmoving [Kir05]. ed [GER^{+}92]. Ede [Vin03, Ano98, Ano03d, Bar90, Har04, Mál04, Rad04, Tel03b]. edited [Boy05, Zuc88]. edition [Tel91i]. Editor [GT37, GT38a, GT39c, JT54, Rab56, ST37a, ST40, ST39b, Tel02a, Asi62, Fer57, Hol04, Tel61c]. Editorial [Fel79b, Fel79a, ST81]. Eds [Oli69, Har11a]. Education [Hou03, Tel90c, Hol10]. Edward [Bad77a, Ber78, Ber90, Ber03c, Ber03b, Bir02, Blu72b, Bor03, Bor12, Bus11, Byr01, Coo80, Cou54, Cre10, Dür05, Dys58, Dys02a, Eps92, Fra81, Fri62, Gob49, Hac04, Har11a, Hed49, Her01, Her03, Hir88, Hou03, Jac65, Jon62, Kau12, Kev03, Kle62, LV10, Lib11, Lin98, Mar92, Mar81, Mon64, Mor62, Mul50, Oli69, Ore58, Pal02, Pei80, Pei87, Pen79, Pic62, PB62, Rop62, Ros80, Sch80, Sch03a, Sea03a, Sei90, Sel58, Set02, Tar98, TG58, Tel92d, Tre87, Zim92, Zuc88, vH83b, vW02, Alb10, Ano57c, Ano58a, Ano89a, Ano90a, Ano91, Anoxxb, Ano02, Ano03a, Ano03b, Ano05c, Ano08a, Ano15, Anoxxa, BD82, Bal03, Ban02, Bar90, Baz90, Bei04, Ber04b, Ber04a, Bet01, BG05, BW05, BO76, BOG89, BP90, BH89, Bro82]. Edward [BS03a, BM04, Bul83, Bun04, BS03b, Cas01, Cou79, Cow03, Cre02, Doy02, Dre02, Dür03, Dür05, Dys06a, Dys07, Dys09, Fra08, Goo04a, Goo04b, Gro04, GT87, HH97, HH04b, Har04, Har08a, Har10, Har11a, Har11b, Hax10, Hec10, Hei90, Hen05, Her02, Hna03, Hol02, Hol10, HM05, Hor97, Joh03, Kir98, KGA^{+}03, Lak78, Lan84, LL84, LBB^{+}03, Len05, LT05, LW04b, LS09, LV10, Lin98, Mad03, MF69, MW88, Mar98, McM03, Mil03a, Ne’98, Mál04, Nor83, Olt74a, Par92, Rad04, Rao03, Rhe95, Rho90, Rho05, Rob04, Ros22, San03, Sea03c, Sea03b, Sea03d, Sei98, Sel03, Sem92, Set07, She76, Sho01, Sla12, Smi85b, New03b, New03c, ST99, Sti03, Sul03, Szi87, Tar10, Tel90a, Tel91c, Tel94c, Tel98a, Tel98c, Tel03b, Tel10a, Tho03, Tis98, Vin03, Wan03, Wat03]. Edward [Wei04, Wig69, Wig96a, Woo03, Zie91, Zim92, dV08a, dV08b, vH83a, Ano05a, Car91, Her90, Loe05, McM90]. Effect [Dow74, JT82, KBY09, Ber84b, Ber84a, Ber06, Eng72, Ham71, HT39b, KZ01, PW84, Tel82d, Tel01a, Dux10]. Effective [SST71]. Effectiveness [CT90b]. Effects [HH09, Tel64a, HT36, Tel84c]. Effort [Ano15, Gon96b, Tel61c]. egyik [Wag98]. egyszeru [Tel93e]. Eigenschwingungen [BT32]. eindimensionalen [TW33]. Einführung [Thi46]. Einstein [Ein39, RTS^{+}xx, Tel78a]. Eiseley [TG58]. Electric [ST40, JT37]. Electrical [ST38]. electrographite [BJ41]. Electromagnetic [LBF^{+}75, TT64, Tel82b, MB92]. Electromagnetism [Tel77e]. Electron [Bar65, HT33, LBF^{+}75, TS70b, Tel91d, WCT39, IT39, Ren34]. Electron-Beams [LBF^{+}75]. Electron-Positron [WCT39]. Electron-Transfer [HT33]. Electronic [SNST39, ST41b, Tel41, TS70a]. Electrons [KT40, Tel31]. Electrostatic [LBF^{+}75]. Elektrographit [BJ41]. Elektronen [Ren34, Tel31]. Elektronen- [Ren34]. Elektronenübergänge [HT33]. Element [Tel78b, SKST69]. elementaren [Thi46]. Elementary [RT38, RT39, Thi46]. Elemente [dEAH29]. Elements [FMT47, FMT49b, MT49, MT50, dEAH29, FMT49a]. élete [BOG89]. eloadások [Tel08]. Emeritus [Hou03]. emigrations [Bar90]. Emission [AT44]. Emmett [Dav94, Dav95]. empty [RC87]. Encounters [Ano02]. Encyclopedia [Mey87, Mit05]. End [Dic79, Dru58, Sei90, Sul75, Fit00, GA^{+}86a, GA^{+}86b, Gor09]. ends [Bal03].

Enemy [Day97]. energia [Mál04]. Energy [BO76, CHL^{+}96, CHL^{+}97, Edg91, FT38, GT39c, KT41, Lak78, MW88, Sei90, She76, Syl87, Tel48b, Tel65a, Tel69b, Tel71c, Tel73d, Tel75a, Tel75d, Tel75b, Tel75c, Tel77b, Tel77g, Tel78e, TW79b, Tel79b, Tel79c, Tel94d, Tel97a, Tel99d, Ber80, CK71, HM05, Kir10, Ros10, Tel47b, Tel77h, Tel77f, Ano15, BOG89, Bad77a, Ber78, Coo80, Pei80, Pen79, Ros80, Tho80]. Energy-Crisis [Tel77b]. Energy-Density [CHL^{+}96, CHL^{+}97]. Engine [Tel72c]. Engineering [Tel70b]. Englman [Dow74]. enlightenment [Sch03b]. Enough [Gul75, Car03]. Enrico [Ano62, Gol99]. Entretiens [Rhé11]. Entropie [Tel61a]. entropy [Tel61a]. Environment [GEM^{+}65]. environmental [O’N07]. Envy [Cre10]. Eötvös [Rad04]. Eötvös-versenyen [Rad04]. episode [Mag88]. Equation [MRR^{+}52, MRR^{+}53, GS89]. Equations [FMT47, FMT49a, FMT49b, EGH86]. Equilibrium [BT41, BT45, BT51a, BT53, DTT37, TT35a, BT51b]. Era [Tel71d, PT81]. Erinnerungen [Dür05, Tel93b]. Ermenc [Sei90]. Ernest [Ber03c, Ber03b, Her03, Kev03, Sch03a, Bof69, Doy02, Her02, Hil15, Hol02]. Error [Wei75]. errors [Bro82]. Erwin [Day97]. erzählt [Lin98]. escape [Mar06]. escaped [O’N89a]. Eskimos [O’N07]. espionage [SH01]. essay [Edg91]. Essays [TG58, Zie91, CD96]. Essential [Tel61c]. est [Dem15]. Establishment [Bof71]. Esther [Moo95]. estimate [TdHK51]. Ethane [TT35b]. Ethics [Mit05, Tel92f, Tel92g]. Ethylene [TT35b]. etica [GER^{+}92]. Eugene [Hor97, BH89, HH04c, Tel95a, Wag81, Wag98, WS92, Wig96b, WS03, Wol67]. Europe [CD96, Fer68]. event [Tel86g, Tel87h]. events [New03d]. Ever [Fra05]. éves [Ano98]. Evil [Pon03]. Evolution [CHL^{+}97, Tel67b]. évszázad [TZ91]. excavation [O’N89a]. Exceptional [Wil98]. Excerpts [Ano63a, Sei90, DA87]. Exchanging [ML07]. Excitation [BT78, FT38]. Excited [CJTW73, SST71, Tel41]. Exiled [RB02]. Exiles [Fra05]. exist [NA91]. Exotic [Tel69c]. Expand [CT91]. Expanding [GT39a, GT39b, GEM^{+}65]. Expelled [MP01]. experiment [CP93, Tel10a]. Experiments [Kir98, TT64]. Expert [Tel76b]. Experts [Ano61a, Smi85a, Sul75]. Exploded [Bet46]. exploration [Uda94, Uda98]. Exploring [Tie94]. Explosion [TT64]. Explosions [LMT59, Tel61c, Tel82b, Fin90, Pre62]. Explosive [Tel87g, Tel92j]. Explosives [Tel59, Tel68c, TTHJ68, Tel70b, WN88]. Expresses [Tur79]. Extragalactic [JT84]. Eye [Tel82c]. eyed [Ano90a]. Eyes [Tel91f].

F [Dys76, Edg91, Pre62, Zac76, Zie91]. Face [TB84]. Face-Off [TB84]. faces [Hna03, Sel03]. facets [A^{+}91]. Facie [Tel46a]. Facts [Dys58, Ore58, Sel58, TG58, TvH83, TL58b]. Fails [Smi85a]. Failures [Tel78a]. fall [Boy95]. Fallout [Ano58a, MSDP80, TP58, BHM91, Han97]. Famous [Cou63, HH04a]. Fast [KT40, MRR^{+}53, MRR^{+}52]. Fateful [Tel98b]. Father [Ano65, Har11b]. Fathers [Blu88b, Les93]. Fear [Sei90, WM15, Wea12]. Feasibility [Tel60b]. February [Tel57b, ST81]. fél [TZ91]. Felder [DB71]. fénypontjai [Wag98]. Fermi [Ano62, Moo95, Sei90, FMT47, FMT49a, FMT49b, Fir05, Gol99, Mac86, Mac87, Tel62h, Wig62, vH01]. Fernbach [Oli69]. ferromagnetism [Tel30a]. Ferromagnetismus [Tel30a]. Fertile [WH98]. Festschrift [KRW05, MW88]. Feud [Ros22]. few [Tel85g]. Fewer [Tel92c]. fiasco [Fin90]. Field [JT55, NT60, WCT39]. Fields [DT56, Tel87b, DB71, New03e]. Fierce [BS03a, Sul03]. Fifth [SBTT39, ALW^{+}88]. fifty [TZ91]. Figures [Ano15]. File [McM94]. Filling [Jac10]. Film [BD82, Sch80]. films [New03b]. Final [Gor92, LT05]. Finds [Ano57a]. fire [Vol02]. Firecracker [Boy95, O’N94, O’N07]. First [Ano15, Flo88, Ken54, Tel48b, WM15, Zor24, Ano15]. First-Hand [Ano15]. Fission [BHT86, BW39, NBDG40, Tel79a, Tur40, CHW91]. Five [TK00, Wol67, ALW^{+}88, Har06b, Har06a]. fizika [Tel93e]. flap [Blu88a]. Flat [Dys79b]. fled [Mar06]. Flimsy [McM94]. Fluctuations [DTT37]. Focuses [Ken54]. foe [Cre02]. foes [Ano79a]. Following [Tel86d, Tel90a]. Fools [Pon03]. force [Flo88]. Forces [Bri65, CT38, JT55, WCT39]. Ford [Ano15]. Foreign [Tel47d, Tel92i, Tel94b, Tel96b]. Foreword [Pre62, Tel80a, Tel86a, Tel91e, Tel91b, Tel94a, Edg91]. Forging [DA87, Sei90]. Former [Fir05]. Foundations [WD86]. founders [Wag98]. Four [AT43a, Lan93b, Lan99, ALT43, Lan11]. Frage [dEAH29]. Fragments [Moo00, Pal00]. Francis [Gob49, Hed49, Mul50]. Franck [ST39b, Tel91j]. Franklin [Ein39]. Fred [TG58]. Free [Pei87, RT38, RT39, Ste73, Ano95, Tel31]. freedom [Har08a, Ano03a, Sea03c]. Freeze [BT82, Tel82e]. freien [Tel31]. French [Dem15, Rhé11, Riv96, SB55b]. Frequencies [TT35b]. Fresh [Tel94d]. Friedrich [CI04, ABES12]. Friend [Tel03a, Cre02]. friends [KGA^{+}03, RTS^{+}xx]. Fuchs [Sei90, Ber10, MHM90, Wil87]. Fuel [Tel72b]. Fueled [MT05]. Full [Ano15, Wag98]. Functions [Bar65, Blo84, Pup05]. Fundamental [Kur73]. Fundamentals [KBY09]. Funds [Ott85, Smi85b]. Fusion [Ano58b, W.D75, MBM^{+}83, MBM^{+}84, Poo89, Tel58a, Tel72b, Tel79a, Tel81f, Tel81g, Tel87g, Tel92j, TGF92, Fis10, HM05, Rei72a, Rei72b, Sem92]. Future [Dys58, MT94, Ore58, Sel58, TG58, Tel72c, Tel79b, Tel91g, TNR97, BT61, Jac10, TL58b, Tel73e]. Futurology [Tel74].

G [Baz90, Car91, Her90, McM90, Rho90, Zie91, Zuc88]. G. [Lak78, She76]. Gables [Kur73]. Gadamer [vW02]. gadgets [Moo95]. Game [Mac92]. Gamma [JT84]. Gamma-Ray [JT84]. Gamow [HPA97, AM23, Blo84, Blo88, GG76, Har01, Rei72a, Rei72b, Tel97b]. GAO [Bar88, Flo88]. Gardner [TG58]. Garwin [Zor24]. Gas [Kih67, MT51b, Ste73, TIW^{+}96]. Gas-Cooled [TIW^{+}96]. Gases [BET38, BDDT40, ST41a, KT35]. Gate [Smi86, Sei90]. Gathering [McG86]. Gatland [TG58]. Gave [HM90]. Geburtstag [DB71, KRW05]. geeignete [RTW32]. gehemmten [TW33]. gemeinverständlich [Thi46]. General [Ber03a, Sei90, Tel56a, Tel56b, TS70a, TS70b, TJ72, Bus11, Law88, Law88, Nor02, Syl87]. Generalizations [GT37, Tel86b, Tel86i]. Generalized [FMT47, FMT49b, FMT49a]. generation [Tel10b]. Generator [HS19]. Geneva [Yor87]. Genius [Mac92, HD02]. Geniuses [Lan93a, Pon03]. geographies [Kir98]. Georg [ABES12, vW02]. George [Tel81e, AM23, GG76, Har01, Rei72a, Rei72b, Tel97b]. Georgia [Har05]. Gerhard [Sto02]. German [Ano93, dEAH29, BT32, Bet93, BW05, BJ41, DB71, Dür05, FMRT34, FST32, Her76, HT33, Joh13, KRW05, LT36, Lin98, PT32a, PT32b, PT33, Ren34, RTW32, Sch01, SB55a, Tel30a, Tel30b, Tel31, TT31, TT32, TW33, Tel34a, Tel34b, TL58a, Tel61a, Tel65c, Tel71a, Tel73a, Tel86f, Tel93b, Tel01b, Thi46, Zim92]. Germany [Joh13]. Gertrud [Zuc88]. Geschichte [Thi46]. Gesellschaft [Tel71a]. Gespräche [Sch01]. Ghost [Ano05b]. Giant [Baz90, Ber88, Car91, Her90, McM90, Rho90, Zie91, BP90]. Giants [Blu88b, Mat98, GT39c]. Gift [MP01]. Gifts [WH98]. given [Car03]. Glasnost [Tel91k]. glass [BTW73, TWB73a, TWB73b]. Glimpses [Wil98]. Glitter [Smi85a]. Global [Lei88, TWH97, TLCW01, Ano90b]. glory [Les93]. Go [Bri00, Fra05, Mat98]. Goal [Tel60c]. God [Mos68b, Mos68a, Mos72]. Goeppert [Goe88]. goes [Vol02]. Golden [Baz90, Car91, Her90, McM90, Rho90, Zie91, BP90, Mar96, Sei99]. Golden-Age [Car91, Rho90]. Goldin [TW92]. Goldman [Sei90]. gondolatok [Tel08]. Good [Ano90a]. Goodbye [Ano03b]. Goodchild [Hen05, Ano05a, Loe05]. Gordon [Sei90, DA87]. government [DeW79]. GPB [Har05]. Grant [Sei90]. Graphite [HHT37]. Gratitude [Fir05]. Gravitation [Tel77e]. gravity [GT85]. gray [Pow04]. Great [Ano05b, Har05, SKST69, LY95, Mar06, Tel93e, GT39a, GT39b, GT39d, TG58]. Greb [Zuc88]. Greek [Vol02]. Gregg [Ber03c, Ber03b, Kev03, Sch03a]. Grosse [Erm89]. Ground [SST71, Ken54]. grounds [Kir05]. groups [Tel68d]. Groves [Sei90, Ber03a, Erm89, Jon62, Law88, Nor02]. grow [New03a]. Growth [Tot63c, Olt74b]. Grund [Thi46]. grundsätzliche [Tel73a]. Guardian [Har08a]. guerre [SB55b]. Guest [Tay96, Tel91h, Tel91a]. Guide [Sei90, Tot63c, Van03b, Bur84, MK02]. Gwinn [Bad77a, Ber78].

H [Kni03, Sul03, TT35a, Ano60a, Ano65, Ano15, Bal03, Bet82, Blu88b, BHM91, Bro82, Car03, DeW79, Dre02, For15, Gon96c, HM90, HM97, Ken54, Kno90, MSDP80, Mos68a, O’N07, SST71, Ste73, Tel88b, Tur79, Wol09]. H-Bomb [Sul03, Kni03, Ano60a, Ano65, Ano15, Bet82, Blu88b, Gon96c, HM90, Ken54, Tel88b, Tur79, Wol09, Bal03, BHM91, Bro82, Car03, DeW79, Dre02, HM97, Kno90, MSDP80, Mos68a]. H-Bombs [Ste73, O’N07]. H. [Hig76]. Haber [Riv96]. Habermas [vW02]. Habit [Tel73f]. háborúja [Bro96]. Hafemeister [Sei90]. hagyatéa [Tel65d]. halide [Har09]. Halides [LST41, ST38, ST40]. Hand [Ano15, Cun88]. Hane [Sei90]. Hangja [Bar90]. Hans [Edg91, Oli69, vW02, Ber80, Bet97, Bro82, Goo99, Tel87j]. Hans-Georg [vW02]. Hansen [Mar97]. Happy [Tar98]. Harass [Bof71]. hard [Ano90b]. hardback [Tho03]. hardcover [Edg91]. Hargittai [Bor12, Kau12, Bus11, Lib11]. Három [Bar90]. Harteck [Erm89]. Hastings [CG95, Hit03]. hate [Mil03c, Mil03d]. Hawkins [Zuc88]. Hazajöttem [Tel91l]. Hazard [MT94]. Hazards [GMS94, Tel62b, Tel62c, Tel62d]. HDR [Tel8xa]. Healing [Tel86g, Tel87d, Tel87h, Wad87]. health [Ano79a]. Hearing [Ano13]. Heat [HT39a, HHT37, TT35a, TW33]. Heaven [Coo80, Pei80, Pen79, Ros80, Tho80, TW79b]. Heavy [BKT41, JT54, KT41, Cun88]. Heidegger [vW02]. Heisenberg [Dür05, KW93, KRW05, Sei90, Tel93b, DB71, Dür05, Erm89, KRW05, Mac86, Mac87, Sch01, Tel76d, Tel93b, Telxx, Tel01b]. Helen [Zuc88]. Helium [BT40, KT40]. Hells [Cou63]. help [TW79b]. Helps [Tel86d]. Henry [Edg91, Kev03, Sei90]. Herbert [Bad77b, Dys76, Edg91, San77, Wig76, Zac76, Zie91]. Here [Ano60a]. Herken [Ber03b, Kev03, Ber03c, Sch03a]. Hermann [HH09]. Herzberg [Sto02]. Heterocatalytic [TU51]. Heterogeneous [DH83]. Hexadeuterobenzene [LT37]. HgCl [ST39b]. hideaway [Ano95]. hidrogénbombáig [BOG89]. High [BTW90, CHL^{+}96, CHL^{+}97, HS19, Tel62g, Tel71c, Tel74, Tel87b, Tel91i, TIW^{+}96, CK71, Ros10, Tel47b, Tel91d]. high-energy [Tel47b]. High-Temperature [TIW^{+}96, Tel91d]. higher [Hol10]. Highest [Ano03c]. Highlights [Ano08b, Wag81, Wag98]. Highly [CJTW73]. him [Vin03]. Himself [RB02]. hindered [Bro82]. Hiroshima [Kur86, TB62, Tel65d, TB75, Tel85b, Tel95c, Wyd84, Yor87, Fri62, Jon62, Kle62, Mon64, Mor62, Pic62, PB62, Rop62, Sei90]. Historical [Ber88, Edg91]. histories [DH83]. History [Ber03d, Bet82, Bro82, Har05, Her85, Hit03, Sei90, Tar18, TB82, Tel83c, BMW89, CHW91, Car03, For15, Fra08, Syl87, Tel82d, Thi46, Sei90]. Hitler [Sch01, Ano95, Gol04, Mar06, MP01, Sch01]. Hoimar [vW02]. hole [Jac10]. Holt [Kev03]. Home [Tel91l, Ano95]. Honor [Ano03c, Oli69, Sea03a, ST99, CD96, GT85, MF69, Tel95b]. honorary [New03f]. Honored [Byr01]. Honors [Wig62]. Hope [Ano89a, Tel46c, Smi65, Smi70]. hopeful [TW79b]. Hopes [TvH83]. Hospital [Ano79b]. hot [Ano95]. Hours [Tel83f]. House [Tau03, BTW73, TWB73a, TWB73b, Ano03a, Sea03c, Tel10a]. Houtermans [ABES12]. Howard [Sei90]. Hoyle [TG58]. Human [Tel78b, Tel85c]. Human-Rights [Tel85c]. humanistic [SKST69]. Hungarian [Bar90, BOG89, Bro96, Byr01, Har04, Har06a, Lan93a, Mar96, Mar10, Ne’98, Mál04, Rad04, New03a, Tel65d, TZ90, TZ91, Tel91l, Tel93e, Tel98f, TB02, Tel03a, Tel03b, Tel08, Vin03, Wag98]. Hungarians [Lan11]. Hungary [Ano91, NK92, Tel92a]. Hurricane [Tel82c]. Huszadik [TB02]. Hydrodynamic [dHT50, TU51]. Hydrodynamics [Tel71b]. Hydrogen [All56, BT40, BS03a, Dys76, Fit98, Han95, Ken54, Tel62e, TS70b, TB82, Zor24, Ban02, Har11b, Mos68b, Mos72, ST37a, SB54, SB55b, SB55a, SB71, Tel30b, WG17]. Hydrogen-Bomb [Dys76]. hype [DeW88]. hyperbaric [Jai94]. Hypothesis [Tel73b]. i.e [HPA97]. IAEA [Flo88]. Ice [TWH97, Tel73e]. Icy [Ros22]. Idea [Ano15, Bri00, Tel69d]. ideas [Hor97]. Identity [Ano58b]. Ideological [Ros22, CD96]. If [Dru58, NA91, LT05]. Ignition [Bre53, KMT46]. Ignorance [Tel68d, Gro04]. Igor [Rhe95]. II [Wig96b, ALT43, PT79, Tel92g, TIW^{+}96]. III [TLCW01]. ill [Ano79a]. Illies [vW02]. illus [Oli69]. Illusion [Tel60e]. Illustrious [Fer68, Sto02]. Images [Sei90]. Immense [TG58]. immigrants [Fer68]. Impact [KPB73, MT94]. Impacts [TT94]. Imperial [Har05]. Implications [KBY09, Sag83, Tel86g, Tel87h, SM86]. Implosion [Ano15]. importance [Tel92f, Tel92g]. Include [Ano03c]. indispensable [Nor02]. Industrial [Hil15]. Industries [MTK46]. Inelastic [ST42]. inertial [HM05]. INF [Flo88]. Infamous [Wel12]. Infamy [ST99]. Influential [Alb10, Bor12, Cre10, Lib11, Har10, Kau12, Sla12]. information [CI04]. infrared [TT32]. Infrastructure [Fit98]. Inhabited [TG58]. initiative [Mit97, DFH84, DFH85, Wes08]. Initiatives [Tel78e]. Insider [Tar18]. Insights [ML07]. Instabilities [Tel55a]. Institute [Har05, Tel90b]. Institution [HPA97]. intellectual [Fer68]. Intensities [LT37]. Intensity [JT82, Tel74, Tel87b]. Interaction [AT43b, DT56, MH92, Ren34, SH74]. Interactions [Kur73, BP83, Ber84b, BP89, BP12, Har09, KZ01, MB92]. Interdependence [Tel79c]. interest [AH04, Bus11]. Interference [HT42]. interjúk [Vin03]. Internal [Tel69a, Tel72c]. International [Tel93d, Tel99d, WD86, Sem92]. Internationalism [Rhe95]. Interpretation [MT42, Tel80b, TT32]. Interscience [Oli69]. Interview [Blu72b, LL84, Olt74a, Tel98c, LT05]. Interviews [BH89, Erm89, Vin03, Rhé11]. Introduction [Tel62f, Tel76a, Tel77a, Tel81a, Tel92a, Tel92b, EGH86, Moo95, Thi46]. Introductory [TW73]. Inupiat [O’N07]. Invariants [RTW32]. invented [Smi10]. inventor [Bal03]. Iodide [MT42]. Ion [TS70b, Tel30b]. Ionic [IT39]. Ions [BKT41, ET40, KT41]. Irks [Ano58b]. ISBN [Cas01, Dys02a, Edg91, Har05, Tho03]. Island [Bra79, Cou79, Dic79, Tel79d]. Isn’t [Smi85b]. Isotope [BT77, BT78]. Isotopes [HT38b, NBDG40]. Issue [Ano60b, Tel60d]. Issues [MT94, Tel68d, TK00, Bur84, Sei90]. Istvan [Bus11, Lib11, Bor12, Kau12]. Italian [Tel92f, Tel92g]. IV [HH04a].

J [Ber03b, Bof69, Han92a, Her03, Sei90, TG58, Zie91]. J. [Anoxxb, Goo80, Goo81, McM05, PC06]. Jahn [Dow74, Dux10, Ber84b, Ber84a, Ber06, Blo71, Eng72, Ham71, HH09, KZ01, KBY09, PW84, Pup05, Pup07, Tel82d, Tel01a]. James [Van03b, Tel91j]. January [Dys09, Kur73]. Jastrow [Tel85e]. Jefferson [Sei90]. Jeno [Wag98]. Jeremiah [Ros71]. Jew [Ne’98]. Jewel [Smi85a]. Jews [Mar06]. Joachim [vW02]. John [Zuc88, Ber10, HD02, Hor97, Mac92, NAH^{+}89, Réd05, Smi10, Tel57b]. Jones [Sei90]. Joseph [Sei90]. Journey [Bet01, Bir02, Dys02a, Hac04, Her01, Pal02, Set02, TS01, Tho03, Van03b, Wan03, Wat03, MK02, Sho01, TG58].

Jr. [ALT43]. Jubilees [Sei99]. Judging [Alb10, Har10, Sla12, Bor12, Kau12, Bus11, Cre10, Lib11]. Judith [Bir02, Dys02a, Hac04, Pal02, Set02, Tho03, Wan03, Wat03]. Judy [Her01]. Julius [ALT43]. Jürgen [vW02]. Just [Ano89b]. Justice [Bal86, Sei90].

K-Mesic [BJT72]. Kapitza [BRS90]. Karl [Erm89, Har11a]. Kármán [Hor97]. Karpin [Van03b]. karrier [Wag98]. Kaufman [Sch80]. Keepers [Kev03, Van03a]. Keeping [Tel92c]. Kennedy [Ano61a]. Kenneth [TG58, Ano15]. Kernbewegung [Ren34]. Kernzeitalter [TL58a]. készült [Vin03]. Kettenmoleküle [BT32]. Kettenreaktion [Her76]. Key [Ano15, Ken54, BS03b]. Keynote [Tel94c]. Kicking [Tel73f]. Kill [Lap62, Rot94, Mon64]. King [Tel87c]. Kirstallgitterspektren [Tel34a]. kivándorlás [Bar90]. Klaus [vW02, Ber10, Wil87, Sei90]. Knew [Woo03, LT05]. Knoll [Day97]. know [MHM90]. Knowledge [Tel57a, TG58]. Kojevnikov [Har05]. Konflikt [SB55a]. Konstruktion [SB55a]. kora [BOG89]. Kourtchatov [Rhe95]. Kowarski [Erm89]. kristallahk [BP83]. Kristallgitterspektren [FMRT34]. Kuczyński [vW02]. Kurzman [Sei90].

L [Bir02, Car91, Dys58, Hac04, Jon62, Pal02, Sei90, Sel58, TG58, Van03b, Wan03]. L. [Baz90, Rho90]. Lab [Smi85b, Tar18]. Laboratories [Tel50a, Sei99]. Laboratory [Lan93b, Sew88, Tel87a, Tel93c, Kni03, Lan99, Tar18, Tel76c, Yor75]. labs [Day77]. LabTV [New03b]. Lanczos [Blo84]. Landau [HH08]. Landed [Lan93b, Lan99]. Lang [Tel49a, Tel49a]. Langsdorf [Ano96]. langwelligen [Tel34b]. LANL [Tel93d]. Lapp [Mon64]. Larson [Erm89]. Laser [Gor92, MH92, Paw02, Smi85a, Smi85b, Tel72b, Wil86, Ano87, Blu88a, DeW88, Sem92, SH74]. Laser-Fusion [Tel72b]. Lasers [IPB^{+}86, Tel73c, Tel74, Tel75c, Tel83a, HM05]. last [New03g, Vin03]. late [Tel96a]. later [TZ91]. Latter [Dys58, Sel58, TG58]. lattice [FMRT34, Tel34a]. Lattices [AT43a, HT42, ST42, ALT43]. laureate [Bro82]. Laureates [Tot63a]. Law [LMT59, Gre77]. Lawren [Sei90]. Lawrence [Ber03c, Ber03b, Her03, Kev03, Sch03a, Doy02, Her02, Hil15, Hol02, Sei99, Tar18, Tel76c, Yor75]. Layer [HT42]. Layers [AT44, BET38, Tel91d]. Leading [Ano15]. Learned [Boy05, Tel81d, ZA04]. Leasing [Ste74]. Leben [Zim92]. Lecture [BCC^{+}46, Tel92d, Tel94c]. lectures [HM05, Tel08]. Lee [SH01]. Left [Mil03c, Mil03d]. Legacy [Fri62, Jon62, Kle62, Mon64, Mor62, Pic62, PB62, Rop62, Har11a, LV10, TB62, TB75, Tel65d]. Légiposta [TZ90]. Leipzig [BW05, KW93, Lin98, Lin98, Tel93b]. Lenses [TU51]. Leo [Zuc88, Fra05, HGS87, Hor97]. Leslie [Jon62, Law88, Nor02, Sei90, Ber03a, Erm89]. let [Flo88]. Letter [Asi62, BJCP88, Bet93, Cun88, DeW88, Ein39, Fer57, GT37, GT38a, GT39c, Han92b, Han97, HM97, Hol04, JT54, Kno90, Rab56, RC87, ST37a, ST40, Sho92, ST39b, Tel54b, Tel61c, Tel87j, Tel99a, Tel02a, Tur79, Wel77, Dem15]. Letters [BHM91, CI04, GM87, Ger75, Gir75, Gul75, KRW05, MHM90, Sau75, Smi75, SLM87, Wei75, vW02, BRS90, Flo88, Réd05]. Lev [HH08]. Level [CT91, TIW^{+}96, Bus11, Dux10]. Levels [BT40]. Lew [Erm89]. Lewis [Jon62, TG58]. Li [CT41]. Libby [Har11a, Pre62]. Life [Bad77a, Ber78, Day97, Lak78, LW04b, Sea03b, She76, Sul75, ABES12, A^{+}91, Bal03, Ber87, BO76, BOG89, BRS90, HD02, Joh03, PC06, New03c, Sto02, Tar10, Zim92]. Light [Ano15, GB89, Tel67a]. limited [Mar87]. limits [IT39]. Lines [LT37]. Linus [Ano58a]. literature [Thi46]. Little [Kev03]. Livable [Zuc88, HGS87]. live [Mos68a]. lived [New03d]. Livermore [LV10, MB92, Tar18, Tel76c, Yor75, Blu88a, Lin98, Sew88, Tar10, Zam95]. Lives [Ber03c, Her03, Kev03, Sch03a, Sei90, Ber03b, Doy02, GA^{+}86a, GA^{+}86b, Her02, Hol02]. LLNL [Tel90c]. Lloyd [Sei90]. Lobbies [Smi85b]. lobby [Ano87]. localized [PW84]. Logic [WD86, Zim92]. Logik [Zim92]. London [Har05, Pei87]. Long [Bet01, ESKR84, TIW^{+}96, TLCW01, Cow03, Tel34b]. long-ago [Cow03]. Long-Term [ESKR84, TIW^{+}96]. long-wave [Tel34b]. Look [Alb10, Bet01, Bor12, Cre10, Lib11, Tel57e, Har10, Kau12, Sla12]. looked [Sch03b]. Loren [TG58]. Lorus [TG58]. Loss [BKT41, KT41, Kni03]. Lost [Sau75]. Louis [Her90, McM90, Zie91]. Love [Boy05, Wal98, ZA04]. loves [Mil03c, Mil03d]. Low [CT91]. Low-Level [CT91]. Loyalties [Ber03c, Her03, Kev03, Sch03a, Ber03b, Doy02, Her02, Hol02]. Lueders [Day97]. Lunar [Tel87a].

M [CD96, Sei90, Van03b]. MA [Dys02a]. Machine [Lan83]. Machines [MRR^{+}53, MRR^{+}52, Tel47b]. Macmillan [Pei87]. MacPherson [Sei90]. Made [Bud84, Tel85h, Moo95, Nic87, New03e, Sei90]. Magnetic [KT34, Tel81f, Tel81g, TGF92, HT36]. Magnetism [IT40]. Magneto [dHT50]. Magneto-Hydrodynamic [dHT50]. Magnificent [Cou54]. main [Car03]. Make [BT82, Lau65, Tel61c]. Making [Ano15, Han95, Rho86, Sei90, Tel57e, Wol09, Yor87, McK84]. Malcolm [Sei90]. Man [Bor57, Bor63, GEM^{+}65, Tel78a, Wil98, Car03, Gro04, Nor02, RTS^{+}xx, Smi10]. Manhattan [Jon62, Moo95, Sei90, FS95, Gro62, Gro83, HHW99, Jon85, Lan93b, Lan99, Lan11, Law88, Nor02, Sei90]. Many [Ano55, Tel71e, Hna03, Sel03, Tel55b, Tel88a]. March [TT40]. Margery [TG58]. Maria [Goe88]. Mark [Oli69]. Markov [Has70, Pes73, Tie94]. Maroncelli [Van03b]. marslakó [Har06a]. marslakótól [Tel08]. Martian [Her01]. Martians [Har06b, Har06a, Har08b, Lan93b, Lan99, Mar96, Mar10]. Martin [Sei90, TG58, vW02]. Mary [Moo00]. Mass [Cas01]. Massachusetts [Wat03]. Material [Gre46]. Mathematics [Ber03d]. Matter [CJTW73, FT47, FTW47, FTW65, ML07, Oli69, Tel49c, MF69, RT49a, Gob49, Hed49, Mul50]. Matterhorn [Ano15]. matters [Rei72a, Rei72b]. Maximalist [Bor48]. May [LV10, TGF41, Flo88]. McKay [Sei90]. meaning [Hor97]. Means [Tel60e]. measurement [Tel61a]. Measurements [TLCW01]. mechanical [Tel61a]. Mechanics [Tel80b, ALT43, PT33]. Mechanism [BW39, SB54, SB71, All56]. Medal [Ano03a, Sea03c, Tel92d, Byr01, Sem92]. Medical [SM86, Tel86g, Tel87h]. medicine [Jai94]. Medvedev [Tel81e]. Meet [Blu88b]. Meeting [AG81, AG80]. megalapítója [Wag98]. mehratomige [TT31]. mehratomigen [HT33]. mehratomiger [PT32a, PT32b, TT32]. meine [Dür05]. Memoir [Ano15, LS09]. Memoirs [Sak95, Sei90, TS01, Erm89, Bet01, Cas01, Wan03, Wat03, Dys06a, Tho03, Bir02, Dys02a, Hac04, Her01, Pal02, Set02]. memorial [Rei72a, Rei72b]. Memoriam [BG05, Bun04, HH08, Dür03]. Memories [Dür05, McM94, Tel97b, Tel85g, Tel93b]. Memory [Fra08]. Men [All56, Ano55, Kev03, Mos68b, Mos68a, Mos72, Str63, SKST69, SB54, SB71, Van03a, Jon62]. Menace [All56, SB54, SB71]. mentor [AM23]. mert [Tel93e]. Mesic [BJT72]. Mesons [BJT69]. Mesotrons [FT47, FTW47, FTW65, MT47]. Message [Tel03b, Tel10b, Tel08]. Meßprozess [Tel61a]. metal [Har09]. Metastability [BT40]. Meter [TW80, Ano90a]. Method [Dav94, Dav95, HT36]. Methodology [WD86]. Methods [Has70]. Methyl [MT42]. Metropolis [BBLS00, BS00, CG95, Hit03]. Michael [Boy05]. mid [Tel96b]. mid-1990s [Tel96b]. Migration [FT38, Fer68]. Mile [Bra79, Cou79, Dic79, Tel79d]. Milestones [Gon96a, Gon96b, Gon96c, Gon96d, LBB^{+}03]. Military [Edg91, Hil15, Tel91h, Tel91a]. millennio [GER^{+}92]. Millennium [Tel97a]. Miller [Sei90]. Million [Smi85b]. Milne [TG58]. Milorad [CD96]. Mind [WH98]. Miniaturized [TLCW01]. Minimum [HT39a]. Minneapolis [Col94]. Minor [Dys79b]. Miracle [Tel92e]. Mirror [GEM^{+}65, Wil86]. Mirror-Reflected [Wil86]. Mirrors [TU51]. misinformation [Mit97]. Missile [Bri00, Bud83, Moh83, Smi85a, Mit97]. missiles [SLM87]. Missing [Joh13]. Mistake [Tel83b]. Mobilization [Joh13]. Model [HT38a]. Modellmässige [BT32]. Models [BT32, Joh13]. Modern [Anoxxa, Mac92, Ber84b, Har11a, LV10, McM05, Tel65c]. modernen [Tel65c]. Modernizing [Joh13]. Modulation [TWH97]. Molecular [Her85, Joh13, ST41a, TS70b, Pup07, FMRT34]. Molecule [BT32, Tel30b]. Molecules [BT78, DTT37, Dow74, FST32, HT33, PT32a, PT32b, ST41b, Tel62h, Tel69a, BP83, BP89, BP12, Eng72, Har09, JT37, KZ01, Ren34, TT31, TT32, Tel34b]. Molekül [FMRT34]. Molekül- [FMRT34]. molekulakh [BP83]. Moleküle [FST32, PT32a, PT32b, TT31, TT32]. Molekülen [HT33, Ren34]. Molekülspektren [Tel34b]. Molten [MT05]. Moment [Ano00, Cow03]. Moments [KT34]. Momentum [BKT41]. Monastero [GER^{+}92]. mondial [SB55b]. monograph [Dux10]. monopoly [Gor09]. Monte [Bar65, BST66, Has70, Pes73]. moon [O’N89b, Ric00]. moral [Zac76]. Morality [Tel98d]. Morrison [Edg91]. Moscow [Har05, BRS90]. Most [Alb10, Bor12, Cre10, Lib11, Har10, Kau12, Sla12]. Motion [NT60, TR60, Ren34]. Motz [Sei90]. mourns [Kni03, Sea03d]. Movement [Sei90, Tel85c, Bur84, O’N07, Smi65, Smi70]. movie [KMN^{+}97]. MSS [Goe88]. Much [GBS^{+}80, Ger75, Mac92]. múltán [TZ91]. Multilayer [MT51b]. Multimolecular [BET38]. Multiple [Fra05]. Multipole [TW80]. Muon [TW80]. Mute [Dev62]. MX [GM87]. my [Dür05]. myth [Mar96, TS82]. Myths [Bul83, Tel82a, vH83a, vH83b, Uda94, Uda98].

N [Tel95b]. nagyszeru [Tel93e]. narrative [Com56]. Nation [Sea03d]. National [Ano55, GR63, Lan93b, Nat40, Sei99, Tar18, Tel87j, Tel99d, Lan99]. Nations [Tel79c]. NATO [Tel86e]. Nature [Fer57, Ott85, Tel57c, Tel57d, Tel57f]. Navy [Tel90d, GM87]. Nazi [Joh13, MP01]. Near [CT90b, MT42, SNST39]. Near-term [CT90b]. Nebulae [GT39d, GT39a, GT39b]. Necessity [Tel86h]. Need [Smi75, Tel58c, Tel77h, TW79b]. Needs [TW79a, Tel94d, WT79]. Ne’eman [GT85]. Negative [FT47, FTW47, FTW65]. Nemeses [Fit98]. Neumann [Ber10, Hor97, Lan93b, Lan99, Mac92, NAH^{+}89, Réd05, Tel57b]. neutrino [Hax10]. Neutron [AT44, GT38b, HdH46]. Neutronen [BJ41]. Neutrons [ST41a, ST37b, ST42, BJ41, ST37a]. never [Kno90]. Newsletter [Tel90b]. Newspaper [Cou79]. next [Tel10b]. NH [HT39a]. Nicholas [BBLS00]. Nichols [Sei90]. Niels [vW02, FK85, Tel69d]. Nine [Kev03, Mar06, Van03a]. ninetieth [Sei98]. ninety [Mar98, Tis98]. Ning [LY95]. No [Bri00, Gir75, Ano90a, ST99, BAA^{+}75]. Nobel [Bro82, HD02, Tot63a, Wol67]. nodal [Pup05]. Non [Ano58b, Ott85]. Non-Atomic [Ano58b]. Non-Nuclear [Ott85]. Nonsense [McG85]. noon [Koe40]. nostalgia [Boy05]. Note [NST40]. notebook [NA91, NA93]. Notes [Bre53]. Nowhere [Tel81c]. Nuckolls [Wei94b]. Nuclear [Ano61a, Ano73, Ano15, BHT86, Ber88, BAA^{+}75, BT82, BCC^{+}46, BJT69, BJT72, BO72, BW39, Bri65, CHL^{+}96, CHL^{+}97, CJTW73, Cou79, CT38, Dic79, DT56, Dys58, Fer57, Fin90, GT48, HLL^{+}84, JT55, Kea80, KMT46, Mac92, MBM^{+}83, MBM^{+}84, MT47, Med79a, Med79b, Med79c, Mla98, NBDG40, NA91, NA93, NK03, Ore58, Ott85, Pre62, Ros22, Sag83, Sei90, Sel58, SM86, TT64, TG58, Tel57c, Tel57d, Tel58c, Tel61b, Tel64a, Tel65b, TTHJ68, Tel70b, Tel75b, Tel77g, Tel77d, Tel79b, Tel79c, Tel81b, Tel82a, Tel82b, TB84, Tel86g, Tel87d, Tel87h, Tel91b, Tel91k, Tel93d, Tel94d, TIW^{+}96, Tel97a, Tel99d, TTA^{+}90, Tur40, Van03b, Wal08, WCT39, Wig62, Wig96b, WN88, Zuc88, And73, Ano79a, Bal03, BT61, Bet91, Bur84, CHW91, Day77, ESKR84, Fis10, GS89, HS83]. nuclear [HGS87, Her76, Jac10, Kir05, Lin98, MK02, Mil86, Mil03c, Mil03d, Nic87, Mál04, O’N89a, Ren34, Sho92, SH01, Syl87, TL58a, TL58b, Tel77h, Tel82e, Tel84c, Tel85f, TZ91, Wad87, Wag98, Wea12, Her76, Tel98f, Sei90, Tel81e]. Nuclear-Dug [Ano73]. Nuclear-Energy [Tel75b]. Nuclear-Fusion [MBM^{+}84]. Nuclear-War [Tel87d]. nuclear-weapons [Bal03]. Nuclei [JT54, Hax10, MB92]. Nucleus [HT38a, TW38, Tel65a]. Nuke [Par92]. nukleáris [Mál04].

Obituary [BBLS00, Mad03, McM03, Tel76d, MRTM95]. Object [WHIT95].

Object-Dispersing [WHIT95]. Objections [Ano15]. Obsession [Cou54]. occasion [CI04, Har08b]. odyssey [Yor87]. Off [W.D75, TB84]. Offense [Wes08]. Office [Sei90, DA87, Ott85]. Officials [Smi85b]. Oil [BO72, Ste74, Tel65a]. One [Alb10, Bor12, BST66, Cre10, Lib11, Wyd84, Ber87, Har10, IT39, Kau12, LT05, Sla12, TW33, Wag98]. One-Component [BST66]. one-dimensional [TW33]. one-electron [IT39]. O’Neill [Boy95]. only [HD02, Tel79d]. Open [Ano60a, Tel87j, Tel99a]. Opening [Tel55a]. Openness [Tel60b, Mit97]. Operation [TIW^{+}96, Tel62a]. Opinion [Ano57b, Tay96]. Oppenheimer [Bad77b, Ber03c, Ber03b, Dys76, Edg91, Her03, Hig76, Kev03, San77, Sch03a, Wig76, Zac76, Zie91, Anoxxb, Ano13, Bro82, Cou63, Dev62, Doy02, Goo80, Goo81, Gre77, Gui04, Her02, Hol02, Ken54, Kev03, McM05, MS15, PC06, Rhe95, Yor76, Yor89]. opportunities [TL58b, Dys58, Ore58, Sel58, TG58]. Opposes [Tot63a]. Opposing [Tot63b]. opposition [Bro82]. Optimum [Pes73]. Opts [Bud83]. Oracles [Sei90, Syl87]. Orbital [JT37]. order [ALT43]. Ordered [HT42]. Organic [BT32, RT38, RT39]. organischer [BT32]. Origin [ART49, GT39a, GT39b, GT39d, IT40, JT84, MT49, MT50, RT49b, Tel49b, Tel54c]. original [Thi46]. Originalliteratur [Thi46]. Originated [Ano15]. Origins [Yor75, Dys12a, TW79b]. Orion [Dys02b]. Ortho [ST37b, KT35, ST37a]. Ortho- [ST37b, ST37a]. ortho-para [KT35]. Oscillation [HT33]. oscillators [PT33]. Oszillators [PT33]. öt [Har06a]. Other [Ano15, Rei72a, Rei72b]. Our [TL58b, Tel61c, Uda94, Uda98, Dys58, Ore58, Sel58, TG58]. Outlawing [Sho92]. Outlook [TL58a]. outrages [Isa93]. outstanding [Sem92]. Overkill [Mon64, Lap62]. oversight [Ano89b]. Owen [Gob49, Hed49, Mul50]. Owens [Bad77a, Ber78, Lak78, She76]. Oxford [Tho03]. Oxide [AT44]. Oxides [Tel89a].

P [BH89, HH04c, Wag81, WS92, WS03, Wol67]. Pacific [Ken54]. Packet [Tel67a]. Pact [Ken63, Ano60b]. pages [Wat03]. Pál [Wag98]. Palevsky [Moo00]. Pandora [Kev03, Van03a]. Panel [GEM^{+}65, IPB^{+}86, LBF^{+}75]. Pangloss [Ros71]. Panos [Car91, Zie91, Baz90, Her90, McM90, Rho90]. Papa [NA91]. Paper [RT39, Tel83c, New03g]. paperback [Edg91]. Papers [Goe88, Sei90, Bur84, GG76, GT85, WD86, Wig96b, AG80]. para [KT35, ST37a]. para-hydrogen [ST37a]. Paradoxes [Joh13]. Parahydrogen [HdH46, ST37b]. paramagnetic [KT35]. Part [Tel81f, Tel81g, Ano90b, Han95, Wig96b]. Partially [HT42]. Participant [Ano15]. Particle [HT38a]. Particles [CT41, NT60]. Past [Ano90b, Hol04]. Paths [TG58]. Patriot [Bei04]. Patrol [Par92]. Paul [Hor97, Moo02, Tel95a, Wag98, Erm89]. Pauling [Ano58a]. Peace [Ano95, Sei90, Tel46b, Tel60c, Tel60d, Tel85c, Bur84, MW88, Tel96b, Yor87]. Peaceful [Pre62, Tel58a, Tel59, Fin90]. Pebbles [Ano90b, Tel91f]. People [Tel55b, AH04]. Peril [Dru58, Smi65, Smi70]. Perilous [Tel60e]. Perils [Hal70]. Perseus [Cas01, Dys02a, Tho03, Wat03]. Personal [Bun04, Her85, Sei90, Tel97b, Bet91, Com56, For15, Nic87, Uda94, Uda98]. Personalities [Dav94, Dav95]. Perspective [Mar87, Bet91]. Perspectives [Ano15, Ber90, Hir88, Pei87, Tre87, Zuc88, Tel87e, Ano05b]. Peter [Ano05a, Loe05, Sei90, Hen05]. phenomena [MH92, SH74]. Phenomenology [LBF^{+}75]. Philip [Edg91]. philosophical [DB71]. philosophische [DB71]. Philosophy [AG80, AG81, WD86]. phone [Ano95]. Photochemical [FT38]. Photographs [FS95, Moo95]. Phys [Her03]. Phys. [Ber03b]. Physical [EHJ70, Tel48c, Tel64a, Tel67b, DB71, Mey87]. Physicist [Ano60a, Blu72a, BS03b, Gol99, Hou03, San03, Tel61c, ABES12, Ber87, BRS90, Har11b, Her76, Lin98, LY95, Yor87, Yor95, Zim92]. physicists [HH04a, Har06b, Tel81h, Har05]. Physics [Anoxxa, Anoxxb, Baz90, BCC^{+}46, CK71, CHL^{+}96, CHL^{+}97, HS83, KBY09, Kur73, Lei88, MKR87, ML07, Mla98, Rho90, TT40, TGF41, TWH97, Tel98f, BP90, Fis10, Har11a, HD02, LV10, Lon58, MW88, Ros10, SKST69, Tel60a, Tel65c, TTT91, TTT02, Thi46, WD86, Bor03, SBTT39, Tel93e, Zie91, Wig96b, Car91, Eps92, Her90, Mar92, McM90, Sei90]. Physics-Based [TWH97]. Physik [Tel65c]. Physikalische [DB71]. Physikers [Zim92]. picture [TR60]. pictures [Tel93b]. Picturing [FS95, Moo95]. pie [Aue03]. pill [Ano90a]. pioneer [Kni03, Smi10]. Pioneered [Mac92]. Pioneering [Ano15]. Plain [Bud84]. Plan [Kih67, Moh83, TT94, Tel75d]. Planet [Tel98e]. Planetary [Ros22, WHIT95]. Planners [MK23]. Plant [MT05]. Plants [Wal08]. Plasma [Bar65, BST66, Fis10, Lon58, Tel54a, Tel55a, MH92, SH74]. Plasmas [LBF^{+}75]. play [Mos68b, Mos68a, Mos72]. Plowshare [Pre62, Tel63a, Tel63b, Kir05, Tel62a]. poem [Tel10a]. Poet [TG58]. Point [TIW^{+}96, Ano89a]. Point-Design [TIW^{+}96]. points [Pup05]. Polar [LST41]. Policies [Sei90, CHW91, Nic87]. Policy [Sag83, Tel89b, Bal03, Tel96b]. Political [CD96, Sei90, Bro82, DFH84, DFH85, Syl87]. Politics [Hab69, Pow04, Tel47d, TS01, SH01, TB02, Tit86, Tit01, Wat03, Bet01, Bir02, Dys02a, Hac04, Her01, Pal02, Sei90, Set02, Tho03, Wan03]. politikában [TB02]. Pollution [Tel77h]. Polonium [AT44]. Polyatomic [HT33, PT32a, PT32b, ST41b, Tel69a, JT37, TT31, TT32]. Polymath [Har01]. popular [Jac10]. Portrait [Sch80]. Positron [WCT39]. possibility [dEAH29]. Possible [ST60]. Posterior [Tie94]. Postulate [Tel80b]. potato [Ano95]. Potential [SST71, Tel37, Tel61b, Tel75b]. poverty [Tel77h]. Power [Ano61b, BAA^{+}75, Cou79, Cre10, MT05, Sei90, Tel57a, Tel61b, TMF76, TIW^{+}96, Ano79a, Bet91, Bur84]. Pp [Edg91, Tho03, Car91, Cas01, Dys58, Dys02a, Har05, Kev03, Oli69, Pei87]. pp. [Bor12]. Prädissoziationsspektren [FST32]. Pragmatism [Zac76]. Preceded [Ken54]. precision [HS19]. preconceptions [Ano90a]. Prediction [TLCW01]. Predictions [Tel78d]. Predissociation [FST32]. Prehistoric [Tel68a]. Preliminary [Com48]. Prelude [Dys79b]. present [CHW91, TW79b]. presented [Thi46]. President [Bud83, Ein39, Ken63].

Presidential [Ano03a, Sea03c]. presidents [Pow04]. Press [Edg91, Har05, Pei87, Tel90a, Tho03]. Pressure [HT38b, Tel83b, Tel54a, Tel68d]. Pressures [Tel62g]. Prevalence [Blo88]. Prevention [Tel84a]. Price [Dys02a]. Prima [Tel46a]. Principle [ST39b, Tel60b]. Private [Cou63]. prize [Sem92, Wol67]. Prize-winner [Wol67]. Prizes [HD02]. Problem [HT39a]. Problems [Tel56a, Tel56b, Tel68c, Tel99d, ALT43, TW79b]. proceedings [AG80, AG81, Kur73]. Process [Tel56a, Tel56b, Tel90c, Tel61a]. Produced [CT41, Sch80]. Production [GT39c, MT47]. Professor [Ano38, Tel95b]. Profile [ST99]. profiles [AH04]. Profound [Bet91]. Program [Ano93, Ano15, Bet93, Gon96c, Ott85, Paw02, Sei90, Smi85a, TW79a, Bal86, WT79]. Progress [Tel65b, Tel85h, Kir98]. Progressive [Tel70a, Fel79a, Fel79b, Han79, MSDP80, San79]. Project [Tel62e, Ano15, Boy95, Dys02b, FS95, Gro62, Gro83, HHW99, Kir05, Lan93b, Lan99, Lan11, Law88, Nor02, O’N89a, Jon62, Moo95, Sei90, Pre62]. Projection [Tel80b]. Projects [Tot63c]. Prologue [Tel86g, Tel87d, Tel87h, Wad87]. Prometheus [Bor12]. promote [Tel96b]. Proof [Tel46a, Pup07]. propagation [ALT43]. Properties [MF69, Oli69]. Prophet [Bof69, Ber80]. propos [Rhé11]. Proposal [TT64]. Proposed [IT40]. Proposes [Ano60a]. Propulsion [TGF92]. Prospects [Tel67b, Tel79b, TWH97]. Protecting [SLM87]. Protection [TT94]. Protective [Tel85h]. Protective-Defense [Tel85h]. protégé [AM23]. Protest [AAF^{+}52]. Proton [Bar65, CT41, JT54, KT34]. Proton-Electron [Bar65]. Provided [Tel90b]. Providing [Tel85c]. Proving [Ken54, Kir05]. PSA [AG80, AG81]. Psychology [Tel95c]. Pt. [LT37]. Public [BM04, Tel68d, Mit97]. Published [Ano15]. Publishing [Wat03]. Pulses [Tel82b]. Pupil [Ano60a]. Purposes [Edg91]. Purse [NA91]. Pursuit [Fra81, Mar81, Tel80c]. Pursuits [NK03]. Put [Kih67].

Q [Cou79]. Qs [Cou79]. Quanta [DB71]. Quanten [DB71]. Quantenmechanik [PT33]. quantenmechanische [Tel61a]. Quantum [Joh13, ML07, Tel80b, CI04, PT33, Tel61a]. quantum-mechanical [Tel61a]. Quasars [Tel66a]. Quayle [TW92]. quest [Com56]. Question [Dev62, Tel87i, dEAH29]. Questioned [Ott85]. Questions [Pal00, Tel90b, Zor24, And73, Moo00]. Qui [Dem15].

R [Ber03a, Dow74, Erm89, Jon62, Law88, Nor02, Sei90]. Rabi [Zuc88]. Race [Ano61b, Gon96d, Sei90, WM15, Bur84, HS83, Mac86, Mac87, McM05, Sei90]. Rachel [Moo95]. Racing [Nor02]. Radial [Bar65]. Radiation [Ano15, TT64, TU51, Tel62b, Tel62c, Tel62d, Ano95, Tel49b, TS82]. radiation-free [Ano95]. Radiator [HdH46]. Radicals [Bof71, RT38, RT39]. Raised [Ano60b]. Raises [Cou79]. Rally [Ano65]. Ralph [Mon64, Pre62, AM23]. Raman [HT39b, LT37, PT32a, PT32b]. Ramanbanden [PT32a, PT32b]. Range [TLCW01]. rapped [Wei94b].

Rate [GT38a]. Ratify [Tot63a]. Ratio [KT34]. Ray [HT42, JT84, Smi85a, Tel46c, Blu88a, DeW88, Gor92, Smi85b]. Rays [ART49, RT49b, Tel54c]. Reaction [BHT86, CT41, TT35a, And73, BS03b, Her76]. Reactions [GT38a, KT48, RT38, RT39, Tel56d, Tel62g]. Reactor [Tel54a, Tel77c]. Reactors [Tel77d, TIW^{+}96]. read [Flo88]. readers [AH04]. readership [Bus11]. Readiness [Ken63]. Ready [Fra05]. Reagan [Bri00, Bud83, DFH84, DFH85, EN86, Fit00, Gar83, Tel83e]. Real [Ano05a, Hen05, Loe05, Ber04b, Ber04a, Goo04a, Goo04b, GTS^{+}06, Moo95, Rho05, Set07, Wei04]. reality [Hor97]. Really [BT82, HM90]. Reappraisal [Dys76]. Reascendancy [Nor83]. reasons [ALW^{+}88]. rebel [Dys06b]. Receives [Ano62]. Recipients [Ano03c]. recollections [WS92, WS03]. Reconsidering [Ber03a]. Record [WT79, Mat98]. Red [Gor09, GT39c]. Reflected [Wil86]. Reflecting [Joh03]. reflections [DB71, Tar10]. Reflects [nLM99]. Regime [MP01]. regrets [ST99]. Related [Sei90, Bur84, MH92, SH74]. Relations [Tel93d]. Relativistic [LBF^{+}75, Tel71b]. relativity [Tel57f]. Relevance [Tel75c]. Reluctant [Jac65, Tel64b]. rely [GM87]. Remark [PT33, Tel30a]. Remarks [ST39a, Tel55a, Tel68b, TS70a, TW73, Tel78c, Tel85d, Tel87b, Tel01a, Tel77f, FST32]. remembered [KGA^{+}03]. Remembering [Shi91]. Remembers [Fir05]. Remembrance [Bun04]. Reminiscences [Tel83f]. Remote [CT91]. Renewable [Tel86c]. Renner [HH09]. Reopen [Tel71f]. Report [Bar88, Dic79, Tel46c, Tel92i, Tel94b, Wel12, Flo88, Nat40, Tel46d]. Reports [KRW05]. Representatives [Tau03]. Research [Ano61a, Hal70, Smi85b, Tel70b, Tel85h, Bal03, Dux10, KPB73, Lon58, Mit97]. Reserves [Ste74]. resist [Tel82e]. resonance [BT77]. Resonances [Ber88]. Resource [Tel86c]. respond [Jac10]. response [Ber03b]. Responsibilities [Tel47a, Tel63c]. Responsibility [BT55, Tel76b, Tel71a]. Restraint [Smi75]. restricted [TW33]. retells [Her90]. retirement [Wei94a]. returns [Hec10]. revealed [NA93]. Reversal [Tel68b]. Review [All56, Ano05a, Bad77b, Bad77a, BD82, Ber88, Ber90, Ber03c, Bet01, Bir02, Bor03, Boy95, Boy05, Car91, Coo80, Cre10, Day97, Dow74, Dys58, Dys76, Dys02a, Eps92, Fra81, Fri62, Gob49, Hac04, Han92a, Han95, Har05, Har11a, Hed49, Hen05, Her90, Her01, Hig76, Hir88, Jac65, Jon62, Kau12, Kle62, Lib11, Loe05, Mar92, Mar81, Mar97, McM90, Mon64, Moo95, Moo00, Moo02, Mul50, Oli69, Pal02, Pei80, Pen79, Pic62, PB62, Rop62, Ros80, Sch03a, Sel58, Set02, TG58, Tel49a, Tel81e, Tel86g, Tel87h, Tho80, Tho03, Tre87, Van03b, Wan03, Wat03, Wig76, Zie91, AM23, Ber03b, Ber84a, Her03, Pei87, San77]. Reviews [Edg91, Sch80, Sei90, dV08b]. Revived [Byr01]. Revolutionary [Tel64b, Jac65]. Rewriting [Bro82]. Rhodes [Han95, Moo95, Sei90]. Rice [Gob49, Hed49, Mul50]. Richard [Han95, Moo95, Sei90, Zor24]. Rigden [Zuc88]. Rights [Tel85c]. RIP [Mil03a]. rise [Boy95, Wea12]. risk [SB55b]. Riskiness [Gar83]. risque [SB55b]. Road [Sei90, Tel81c, Bet91, Nic87, Van03b]. Roads [Tel88a]. Robert [Anoxxb, Ber03c, Ber03b, Goo80, Goo81, Her03, Kev03, McM05, Sch03a, Zie91, Bro82, Doy02, Her02, Hol02, Kev03, Rhe95, Tel85e]. rod [Ren34]. rod-shaped [Ren34]. Roger [Sei90]. Role [ET40, Ken54, MT51b, RT38, RT39, Tel66b, Tel70b, Tel81h, Tel90c]. Ronald [Bri00]. Roosevelt [Dem15, Ein39]. Roots [Ros22, O’N07, New03a]. Rotation [TW38, PT32a, PT32b]. Rotationsstruktur [PT32a, PT32b]. rotator [TW33]. Rotators [TW33]. Rudolf [HH09]. ruin [McM05]. Rules [GT36]. Russia [Har05]. Russian [BP83, BRS90, NA93, Tel58d]. Russian/Soviet [NA93]. Russians [Ano61b]. RVB [And08].

S [Car91, Zuc88]. S. [Baz90, Lak78, Rho90, She76]. SA [Ano05b]. sabotaged [Tel01b]. sabotiert [Tel01b]. Safe [Gul75]. Safer [BT82]. Safety [Tel68a]. Sagan [Ros22]. Sagittarius [Tel51b]. Sakharov [A^{+}91, Blu88b, Lou02, Rhe95, Rhé11]. Salt [MT05, PT79]. Sampling [Has70, Pes73]. Samra [Moo95]. Sanders [Pre62]. Sara [Tel02a]. Satellites [CT91]. Saturation [CT38]. Saunders [Tel81e]. Sausage [Zor24]. Say [Smi85b]. saying [ET20]. Says [Ano61b, Hal70, Kih67, Tel61c, Bro82, Lin98, NA91]. Scaling [LMT59]. Scattering [KT40, ST41a, ST37b, ST42, BT77, ST37a]. Schalldispersion [LT36]. Schemes [Tel72b]. scholars [Shi91]. Scholder [vW02]. Schreckens [Zim92]. Schriever [Ken63]. Schroeer [Sei90]. Schwingungsstruktur [HT33]. Science [Ano57a, Ano60a, AG81, Bet01, Bir02, Blu72b, Dav91, Dys02a, GBS^{+}80, Hab69, Hac04, Har05, Hel98, Her01, Hil15, LW04a, Mit05, Pal02, RB02, Ros71, Ros22, Sea03b, Set02, Tel66b, Tel67b, Tel71d, TT^{+}77, Tel78d, Tel89b, Tel94e, Tel98d, TS01, TB02, Tho03, Vol02, Wal98, Wan03, Wat03, Wig62, Zie91, AG80, Bet91, Blu88a, HH04a, HD02, Mar96, Mey87, PT81, Sto02, Tel92f, Tel92g, dV08b, Har06b, WD86, TG58]. Scientific [Ano65, Cas01, Har01, LW04b, Mac92, NAH^{+}89, Tot63c, Har11a, LV10, SKST69, New03g, New03e, Tar10, Wig96b]. scientifiques [Rhé11]. Scientist [Ano57c, Ano08a, Asi62, Tel87c, Zuc88, ALW^{+}88, Dys06b, Mil03c, Mil03d, Mál04, Sch03b, Smi65, Smi70, Tel71a, Vin03, ALW^{+}88]. Scientists [Alb10, AAF^{+}52, BAA^{+}75, BT55, Bor12, Cou63, Cre10, DG99, Dav91, GB89, Gol04, GR63, Lib11, MP01, Moh83, Rhe95, Sei90, Tel46b, Tel47a, Tel63c, Tel85e, Tel87j, Wea76, BMW89, ET20, Erm89, Har10, Kau12, Mar10, Rhé11, Shi91, Sla12, GR63]. Scienza [Tel92f, Tel92g, GER^{+}92]. Scope [Bus11, Dux10]. Scott [Boy05]. Scribner [Car91]. SDI [Blu88b, EN86, GT87, Mar87, Ott85, SLM87]. search [Tel77f, Vol02]. Secrecies [Tel71e]. Secrecy [Blu72b, Tel60e, Tel73f, Tel81c, Mit97, Zac76]. Secret [Ano95, FS95, Han92a, HM90, Moo95, Tel99b, Bro92, Bro96, DeW79, Fel79b, Kno90, Wea76, WG17]. Secret-Stealing [Tel99b]. secrète [SB55b]. Secrets [Bor03, Eps92, Mar92, Tel92c, Tel99c, HM97, NA93, TTT91, TTT02]. Section [WD86]. Security [Ano13, Ken54, Ken63, Tel58b, Tel60e, TMF76, TZ91]. Seeks [Ott85]. Sees [Dru58]. seinem [KRW05]. Seismic [LMT59]. Selected [Bet97, Sei90, Bur84, DH83, Réd05, Tel08]. Selection [GT36, WD86]. Selective [GT38a]. Senate [WT79, McG85, Tot63a]. Sensing [CT91]. Sensors [TLCW01]. Sep. [LBB^{+}03]. Separated [NBDG40]. Separation [BT78, BT77]. September [Dys09]. Sequel [Par99]. Series [BCC^{+}46, IT39, Tel86d]. set [Car03]. settembre [GER^{+}92]. Seven [Tel83f]. Seventh [TGF41]. Seventy [ALW^{+}88]. seventy-fifth [ALW^{+}88]. Seventy-five [ALW^{+}88]. Shaking [Hel98]. Shale [BO72, Ste74]. Shape [Ano61b]. shaped [Mar10, Ren34, New03d]. shapes [HS19]. shaping [Har06a]. shatterer [Goo80, Goo81]. Sheds [Ano15]. sheep [Ano95]. Shield [Ber90, Bri00, Hir88, Pei87, Sei90, Tre87, Wil86, Zuc88, DA87, Tel87e]. Shock [BT41, BT45, BT51a, BT53, Moo02, TJ72, BT51b, Han92b]. Shocks [dHT50]. Shoolery [Dys02a, Tho03, Wan03, Wat03, Bir02, Hac04, Her01, Pal02, Set02]. Shootin [Ric00]. Shooting [O’N89b]. Shortage [Tel75c]. Shots [GT56]. Should [Tel68d, Tel87c, Tel82e]. Show [Dev62]. Shows [Bri00, Tot63b]. Side [Bof85, Han92a]. Sidney [Oli69]. Sign [Bri00]. Silent [TW92]. Silk [NA91]. silos [RC87]. simple [Tel93e]. Simplest [Pup07]. simplicity [Tel80c, Fra81, Mar81]. Situation [Tel65c]. Six [Sei90, GA^{+}86a, GA^{+}86b]. Sixth [TT40]. size [Tel57f]. Skies [Tel71f]. skills [Kea80]. Sky [Par92]. Slabs [Gre46]. Slow [ST41a]. Small [SMM^{+}60, TR60, Ano90a]. Smith [Sei90]. Smog [Bir67]. Smoke [TTA^{+}90]. Smyth [Edg91]. Social [Tel76b]. society [Tel71a]. soft [Rot94]. soglie [GER^{+}92]. solution [Ano90b, Rot94]. Solutions [Tel77b, Tel77f]. Solved [Tel68c]. Some [GT37, Moh83, Sag83, Tel68b, Tel71c, Tel78d, Tel86b, Tel86i, Tel97b]. Sommerfeld [Ano38]. sorcerer [Riv96]. sorciers [Riv96]. SOT [Tel98a]. sound [LT36]. Sounds [Blu72a]. Source [Bir67]. Sources [Kih67]. South [Flo88]. Soviet [Har05, Gol84, Gon96c, NA93, Tel61c, WG17]. Soviets [Wes08]. Space [Ano61b, Tel61b, Tel85a, TGF92]. spaceship [Dys02b]. Speak [BAA^{+}75]. Specific [HT39a, TW33]. Spectra [FST32, ST41b, FMRT34, IT39, Tel34a, Tel34b]. Spectroscopy [Her85, Tel92a, NK92]. Spectrum [HdH46, HS19, ST39b, TT32]. Speed [Ott85, Tel62g]. Spektrums [TT32]. Spencer [Sei90]. spezifische [TW33]. Spheres [Gre46]. sponsored [HPA97]. Sputnik [Moo02]. spy [SH01, Wil87, Sei90]. Squabble [Smi85b]. Square [Ano57a]. stabförmigen [Ren34]. Stability [JT37, NT60, Tel54a, Tel62h, Tel86d, Tel88a]. stage [Car03]. Stalin [Gor09, Har05]. Stallion [Smi86, Sei90]. Stanford [Edg91]. Stanley [Bad77a, Ber78, Her90, McM90, Zie91]. Star [Han92a, Par92, Bof85, Bro92, Bro96, Bud84, Fit00, Gor92, Lew85, McG85, McG86, Par99, Smi85a, Tel85e, Tel86d, Vol02, Wes08]. Star-Wars [Tel86d]. Starry [Ano90a]. Stars [GT38b, dEAH29]. START [Tel91k]. State [Day97, FMT47, FMT49b, MRR^{+}53, SST71, Tel90b, EGH86, FMT49a, GS89, MRR^{+}52, Tel46c]. Statement [Szi87]. States [SST71, JT37, AB88, Tel67b]. Station [TIW^{+}96]. stationery [Tel10a]. statistical [ALT43]. Statistics [AT43a, ALT43]. Stealing [Tel99b, Tel99c]. Stellar [Tel85d]. Step [Ken63].

Stephen [Har11a]. Sternen [dEAH29]. Sternglass [Bof69]. Stewart [Tel02a]. Still [Ano90b, Bri00, Hel98, Flo88]. Stood [vH01]. Stop [Boy05, ZA04]. Story [Han92a, Jon62, LW04a, MP01, Bro92, Bro96, Dys02b, Gro62, Gro83, Mos68b, Mos68a, Mos72, RTS^{+}xx, TW79b]. Strange [Kev03]. Strangelove [Ber04b, Ber04a, Goo04a, Goo04b, GTS^{+}06, Har08a, Rho05, Set07, Wei04, Ano05a, Hen05, Loe05]. Strategic [Bud84, CT90a, CT90b, DFH84, DFH85, IPB^{+}86, Moh83, Tel86h, Tel87i, Mit97, Wes08]. Strategy [Mon64, Bud83, Lap62]. Strauss [Jon62]. strength [Blo84]. Strong [Tel72b, New03a, MB92]. Structural [Tel91b]. Structure [Gob49, Hed49, HT33, LT37, Mul50, PT32a, PT32b, Tel49c, TS70a, TS70b, RT49a]. Student [Fir05]. Students [Tel91c]. Studies [Tel68a, Tel08]. studio [Bar90]. stúdiójában [Bar90]. Study [BST66, NAH^{+}89, Gre77]. Sublimation [HHT37]. succeeded [LT05]. Successes [Tel78a]. Successful [Ken54]. Sudoplatov [McM94]. Sufficiency [Tel78e]. Suggested [Tel46d, Wil86]. Suggests [Tot63c]. Suitable [RTW32]. Sull’importanza [Tel92f, Tel92g]. Sun [Han95, HHW99]. Sunscreen [Tel98e]. Super [Tel46a, Thi50]. Superbomb [Edg91, GB89, Bro82, Yor76, Yor89, Zie91, Zac76, Bad77b, Dys76, Hig76, San77, Wig76]. superbombe [SB55b]. Superconductivity [BTW90, MT87, Tel89a, Tel90b, Tel92e, Tel91d, Tel91i]. Superlasers [CHL^{+}96, CHL^{+}97]. Supernovae [Tel71b]. Superpower [Tel86d]. Surface [BJT72, Dav91, ET40, MT51b]. Surfaces [BJT69, Tel37]. Surprise [Tur79]. Surprises [Ros10]. Survivability [CT90b]. survival [Kea80]. Survive [MK23]. sustained [And73]. switch [Ano95]. Sword [Ber90, Hir88, Pei87, Tre87, Zuc88, Tel87e, Sei90]. Swords [Mar97]. Sylves [Sei90]. Symposium [HPA97, LV10, Tel55a, dV08b, Har11a, dV08a]. System [Bof85, DTT37, TIW^{+}96]. Systems [GCB^{+}77, NST40, Tel76a, Tel77a, PW84, Pup07]. Szanton [WS92, WS03]. századi [TB02]. szigorúan [Bro96]. Szilard [Zuc88, Fra05, HGS87, Hor97, Lan93b, Lan99, SMM^{+}60].

T [Boy95, Sei90, Dav91]. take [Ano90a]. Taking [Ano95]. Talent [Ano65]. Tales [Tel49a]. Talk [Bre53]. talking [Yor87]. Talks [Ano60a]. Talley [Mar92]. Tamped [Gre45, Gre46]. Tangled [Ber03c, Cou63, Her03, Kev03, Sch03a, Ber03b, Doy02, Her02, Hol02]. tanulmányok [Tel08]. Tap [Kih67]. Tauscher [Tau03]. teacher [Ano08a]. teachers [Shi91]. Teaching [Ano60a]. Team [Ano93, CP93]. Team-B [CP93]. Technical [Fit98, Tel78e, Bro82, DFH84, DFH85]. Technically [Moh83]. Techniques [Tel91b]. Technological [Fit98]. Technologies [Ano15, Tel77c]. Technology [Har05, Mit05, MT05, Sei90, Tel78d, Tel84a, Tel89b, Tel8xa, Tel91g, Tel91a, HS83, KPB73, Mey87, TT^{+}77, Tel87e, Tel91h, Zuc88, Ber90, Hir88, Pei87, Tre87]. teljes [Wag98]. Tell [W.D75]. Teller [Ano05a, Ano08b, Bad77b, BD82, Ber03b, Bet01, Bir02, BW05, Bor12, Car91, Cas01, Cre10, Dow74, Dür05, Dux10, Dys58, Dys76, Dys02a, Dys06a, Dys09, Edg91, Fri62, Hac04, Har11a, Hen05, Her90, Her01, Hig76, Jon62, Kau12, Kle62, LV10, Lib11, Lin98, Loe05, Mad03, Mar92, McM90, Mil03a, Mon64, Mor62, Oli69, Pal02, Pei87, Pic62, Rop62, San77, Sem92, Set02, TG58, Tho03, Vin03, Wan03, Wat03, Wig76, Zac76, Zie91, dV08a, vW02, vH83b, Alb10, Ano57a, Ano57c, Ano58a, Ano58b, Ano60a, Ano61b, Ano62, Ano63a, Ano63b, Ano73, Ano79a, Ano79b, Ano87, Ano89a, Ano90a, Ano91, Ano95, Ano98, Anoxxb, Ano02, Ano03a, Ano03b, Ano03c, Ano03d, Ano05b, Ano05c, Ano08a, Ano15, Anoxxa, ALT43, Bal03, Ban02, Bar88]. Teller [Bar90, Baz90, Bei04, Ber90, Ber04b, Ber04a, Ber84b, Ber84a, Ber06, Bet01, Bir67, BG05, Blo84, Blo88, Blo71, Blu72a, Blu72b, BO76, BOG89, BP90, Blu88b, BH89, Bro82, Bro92, Bro96, BS03a, BM04, Bul83, Bun04, BS03b, Byr01, Coo80, Cou54, Cou63, Cou79, Cow03, Cre02, Cun88, Dav94, Dav95, Dev62, Doy02, Dre02, Dru58, Dür03, Dür05, Dys07, Eng72, Eps92, Fir05, Fit98, Flo88, Fra08, Fra81, Gob49, Goo04a, Goo04b, Gro04, GT87, Hal70, Ham71, HH97, HH04b, Har04, Har08a, Har08b, HH09, Har10, Har11b, Har11c, Hax10, Hec10, Hed49, Hei90, Her90, Her02, Hir88, Hna03, Hol02, Hol10, HM05, Hor97, Hou03, Jac65, Joh03, KZ01, Ken63, Kih67, Kir98, Kir10, KBY09, KGA^{+}03, Lak78, Lan84, Lan93b, Lan99, LL84]. Teller [LBB^{+}03, Len05, LT05, LW04a, LW04b, LS09, Lin98, W.D75, Mag88, Mar81, MF69, MW88, Mar98, McM03, ML07, Mil02, Mil03b, Mil03c, Mil03d, Mul50, nLM99, Ne’98, Mál04, Nor83, NA91, Olt74a, Ore58, Pal02, Par92, Pei80, Pen79, PW84, PB62, Poo89, Pup05, Pup07, Rad04, Rao03, Rhe95, Rho90, Rho05, Riv96, Rob04, RB02, Ros80, Ros22, San03, Sch03b, Sea03a, Sea03c, Sea03b, Sea03d, Sei90, Sei98, Sel03, Set07, She76, Sho01, Sla12, Smi85b, New03b, New03c, New03f, New03d, New03g, New03e, ST99, Sti03, Sul03, Szi87, Tar10, Tau03, Tel61c, Tel82d, Tel90a, Tel90b, Tel91c, TW92, Tel98c, Tel01a, Tel03a, Tel03b, Tel10a, Tho80, Tis98, Tot63a, Tot63b, Tot63c, Tre87, Vin03, Wei04, Wel12, Wig62, Wig69, Wig96a, Woo03, Yor76, Yor89, Zim92, Zuc88]. Teller [dV08b, vH83a, Bad77a, Ber78, Ber03c, Bor03, Bus11, Har11a, Her03, Kev03, Sch80, Sch03a, Han92a]. Tellerand [Sel58]. telltale [Flo88]. Temperature [TIW^{+}96, Tel91d, Tel91i]. Temperatures [BTW90, TdHK51]. Tension [MT51b]. Term [ESKR84, TIW^{+}96, CT90b]. terror [Zim92]. terzo [GER^{+}92]. Test [Ano63b, Ano15, Gor92, GT56, Ken54, Tot63b, Ano60b, Ken63, Tot63a]. Test-Ban [Ano63b]. Tested [Ano79b]. Testimony [Ano63a]. Testing [CHL^{+}96, CHL^{+}97, Sei90, Tel61c, Bal86, Mil86, Tit86, Tit01, Sei90]. Tests [Dru58, Ott85, Tel58c, BT61]. tévútjának [Bro96]. Textbook [Jai94]. Thais [Ano73]. Theater [dV08a]. Their [Has70, HHW99, AM23, Hor97, Lan93a, Tel47b]. them [NA91]. Theodore [Hor97]. Theorem [Tel86b, Blo71, Pup05, Pup07, Tel86i]. Theoretical [Anoxxb, KT48, SBTT39, TT40, TGF41]. Theorie [LT36, Ren34, Tel30a, Tel34a, Tel34b]. Theory [BTW90, BDDT40, FMT47, FMT49b, GT37, Gre45, JT55, Mac92, MT51a, RTW32, Tel54c, Tel62h, Tel66a, TS70b, Tel89a, WCT39, EGH86, FMT49a, KT35, LT36, Ren34, Tel30a, Tel57f, Tel10a, Tel34a, Tel34b]. there [Fit00, Gol84, NA91]. Thermal [BT41, BT45, BT51a, BT53, BT51b, BJ41]. thermischer [BJ41]. Thermoelectric [IT40]. Thermonuclear [GT38a, Gon96a, Gon96b, Gon96c, Gon96d, Tel50b, Tel51a, Tel54a, Tel56d, Tel56a, Tel56b, Tel69b, Tel72b, Tel73d, BS03b, Lon58, Sem92, Tel73e]. things [Ano90b]. Think [Ano90a, Tel85i]. Thinking [AB88]. Third [Tel97a]. Thomas [FMT47, FMT49a, FMT49b, Tel62h]. Thorium [MT05, Tel78c]. Thorium-Fueled [MT05]. Though [Bri00]. thought [Tel96a]. Thoughts [Tel71c, Tel78a, Ano90a, Tel08]. Thousand [WM15]. Threat [Wal08, WHIT95, SB55a]. Three [And73, AT43b, Isa93, Tel99d, Bar90, Bra79, Cou79, Tel79d]. Three-Mile [Tel79d]. Three-Mile-Island [Cou79]. Thrower [Bir02]. ties [New03a]. Time [Ano00, BD82, Bri00, CI04, HT36, Mac86, Mac87, Tel68b, Tel8xb, Sei90]. Times [Bad77a, Ber78, Har05, Lak78, She76, BO76, BOG89, New03c]. timing [Ano90a]. Timothy [Van03b]. titkos [Bro96]. Titus [Sei90]. Today [Hel98, AH04]. Told [Jon62, Gro62, Gro83, TW79b, WS92, WS03]. Too [Ano90a, Ger75, Ne’98]. took [Ano89a]. Top [Han92a, Bro92, Bro96]. Top-Secret [Han92a, Bro92, Bro96]. története [Bro96]. tragédiája [Har04]. tragedy [Har04]. tragic [Uda94, Uda98]. trail [Hax10]. Transfer [HT33, Tel91d]. Transformation [GT37, KT35]. Transition [Blo88, SNST39, Tel41]. trap [Boy05]. Travel [TB02]. Traveler [Van03b, MK02]. Treasure [BO72]. Treatise [EHJ70]. Treaty [Tot63a]. Trefort [BOG89]. Triatomic [FST32, Ren34]. Tribute [Bei04, Tau03]. trillion [CP93]. Trinity [Sei90, KMN^{+}97, Nic87]. Trip [Tel92i, Tel94b, Van03b]. True [HD02, MP01, Dys02b]. Truman [Gor09]. Truth [Mil02, Mil03c, Mil03d]. Tudományban [TB02]. tudós [Vin03]. tudósa [Mál04]. Tuesday [San03]. turbulent [TW79b]. Turing [Dys12a]. Turn [Tel91a, Tel91h]. TV [Dev62]. Twelfth [Nat40]. Twentieth [Bet01, Bir02, Bor12, Cre10, Dys02a, Hac04, Her01, Lib11, Pal02, Set02, Tho03, Wan03, Wat03, Har06b, Kau12, LY95, Alb10, CD96, Har10, Sla12, TS01, TB02]. Twentieth-Century [Bet01, Bir02, Dys02a, Hac04, Her01, Pal02, Tho03, Wan03, Wat03, CD96, TS01]. Twenty [Wol67]. Twenty-Five [Wol67]. Twinkle [Moo02]. twins [Flo88]. Two [AT43a, ALT43, Cou63, Tel47a, Tel63c, HD02]. Two-Dimensional [AT43a, ALT43]. Type [AT43b]. typical [Van03b].

U.S. [Ano15, Ken63, Tau03, Tel94d, Tel96b]. Ulam [Car03, Dys12b, Han92b, Wel12]. Ultimate [Dud75, Tel86c, Tel02b, Hor97, Vol02]. ultraroten [TT32]. Ultraviolet [NST40, SNST39]. uncertainty [TZ91]. Unconventional [Tel78a]. Underground [GT56, LMT59, MT05]. understandably [Thi46]. Understanding [CG95]. United [AB88, Tel67b]. Universe [GT39a, GT39b, Dys12a, Tel57f, Dys79a, TG58]. University [Edg91, Har05, HPA97, Tel70b, Wel77, AAF^{+}52, Shi91]. unknown [TK79].

Unlikely [Ber10]. unrealized [Kir05]. Unrest [Hal70]. Unsinkable [Tel90d]. Unsolved [Tel68c]. Unusual [Oli69, MF69]. Unworkable [Moh83]. update [Ano95]. Uproar [Tur79]. Urals [Med79a, Med79b, Med79c, Tel81b, Tel81e]. Uranium [KMTW43, NBDG40]. Urged [Kih67]. US$19.95 [Pei87]. USA [Har05, LV10, Bro96]. Use [Ano15, Kih67]. Uses [Pre62, Tel58a, Tel59, TTHJ68]. Using [Has70, Pes73]. utazás [TB02]. utcától [BOG89]. utolsó [Vin03]. üzenete [Tel03b]. Üzenetek [Tel08].

v [BP83, WHIT95]. Valence [RTW32]. Valency [EHJ70]. Valenztheorie [RTW32]. válogatott [Tel08]. values [Zac76]. VanDeMark [Kev03]. Vapor [HT38b, NST40]. Varenna [GER^{+}92]. varied [dV08b]. Vast [Ste74]. Vector [RTW32]. Vehicles [Tel61b]. Vektorinvarianten [RTW32]. vele [Vin03]. Verantwortlichkeit [Tel71a]. versenyen [Rad04]. Vibration [TT35b]. Vibrations [BT32, GT48, LST41, Tel41]. Vibronic [BP83, BP89, BP12, Har09, KZ01, Ber84b]. Vibronnie [BP83]. victim [Tel79d]. View [Moh83, Sul75]. viewpoints [Bet91]. világa [Vin03]. világformáló [Har06a]. Villa [GER^{+}92]. Vincent [Sei90]. Vindicates [Bar88]. Virial [Tel86b, Tel86i]. Visit [Ano91]. Visiting [Fra08]. voice [Mar10, Bar90]. Void [Zie91]. Vol [Wig96b]. volume [FK85, Rei72a, Rei72b, MKR87]. Vows [Ken63]. vzaimodeistviya [BP83].

W [Ano15, Sei90, TG58]. Waals [AT43b, BDDT40]. Waffenabwehr [Tel86f]. Walker [Moo02]. Wants [Par92, Ano95]. War [Han92a, Sag83, SM86, Tel46b, Tel57e, Tel84a, Tel87d, Van03b, Bro92, Bro96, ESKR84, Kea80, MW88, MK02, Tel84c, Tel85f, Vol02, Wad87, ESKR84, Fit00, Tel86g, Tel87h, Uda94, Uda98, Tel86g, Tel87h]. War-Making [Tel57e]. Warfare [Fer57, Tel57c, Tel57d]. Wärme [TW33]. Warming [TWH97, Ano90b]. Warning [Ano61b, Ano63b]. Warranted [Smi85b]. Warrior [Cas01, Par92]. Wars [Han92a, Bof85, Bro92, Bro96, Bud84, Fit00, Gor92, Lew85, McG85, McG86, Par99, Smi85a, Tel85e, Tel86d, Vol02, Wes08]. Was [Sei90, GA^{+}86a, GA^{+}86b, Gro04, KAS96, Kno90, Tel79d]. Washington [HPA97, SBTT39, TT40, TGF41, Anoxxb, Flo88]. Wasserstoffbombe [SB55a]. Wasserstoffmolekülion [Tel30b]. watch [NA91]. Water [Bet46, Gre45, Gre46]. Water-Tamped [Gre45, Gre46]. wave [Pup05, Tel34b]. Waves [BT41, BT45, BT51a, BT53, BT51b, Han92b]. Way [Fit00, TG58, KAS96]. weak [MB92]. Weaknesses [Bud84]. Weapon [Gor92, Vol02]. weaponeers [Wei94a]. Weapons [Ano15, NA91, Tel47c, TS84, Tel86f, Tel92h, Tel93d, Van03b, Wel77, Bal03, Bet91, Day77, MK02, Sho92, Tel69c, Yor87, NA93]. Weart [Sei90]. Weather [TLCW01]. Weaver [Sei90]. Wechselwirkung [Ren34]. Week [Tel02a]. Weighs [Wal08]. Weird [Blu88a]. Weiss [Zuc88]. Weizsäcker [CI04]. Wells [Kih67]. Wen [SH01]. Wendy [Mar92]. Were [Sei90, Nic87]. Werk [Zim92]. Werner [DB71, Dür05, DB71, Dür05, Erm89, KW93, KRW05, Tel76d]. West [Mar10, Ste74]. Whatever [Tel85i]. Where [Lan93b, Lan99, Tel77d, vH01]. Whew [Ano90b]. which [TW79b]. White [Ano03a, Sea03c, Tel10a]. Who [And08, Ano15, HM90, Mac92, Tel94e, Dem15, Gro04, Har06b, Mar06, Mar10, Mos68b, Mos68a, Mos72, Smi10, Dem15]. Whose [Bri00]. Wichita [Gol84]. Widening [Ano95]. Widespread [Tel84c]. Wiesner [SMM^{+}60]. Wigner [Tel95a, BH89, HH04c, Hor97, Lan93b, Lan99, Wag81, Wag98, WS92, Wig96b, WS03, Wol67]. Wiley [Oli69]. Will [BCF^{+}97, BT82, BO72, Bri00, EN86, Ken63]. Willard [Pre62]. William [Han92a, Sei90]. Williams [Sei90]. Wilson [Mar92]. winner [HD02, Wol67]. Winter [HLL^{+}84, Ros22, TTA^{+}90]. Wissenschaftlers [Tel71a]. withdrawn [Kev03]. Within [Fit98]. Without [TWB73a, TWB73b]. witnesses [Sch01]. Wizards [Zuc88]. women [HHW99]. Wood [TW92]. Woodruff [BJCP88]. word [Ano95]. Work [Ano55, Tel55b, Tel56c, Wel77, Bro82, SKST69, Sem92, Zim92]. works [Bet97, Wig96b]. World [Ano61b, Ano71, FS95, GR63, Moo95, Sei90, SMM^{+}60, Tel48a, Tel77g, Tel86g, Tel87d, Tel87h, WM15, Zuc88, ESKR84, GA^{+}86a, GA^{+}86b, Har06a, HGS87, Mar06, Mos68a, SB55b, TR60, Vin03, Wad87, Com48]. world-shaping [Har06a]. worlds [Goo80, Goo81]. Worldwide [ESKR84, TT94]. Worrying [Boy05, ZA04]. worst [Tel86g, Tel87h]. Would [Kih67, Lau65]. Writer [Asi62, Tur79].

x [Oli69, Blu88a, Cas01, DeW88, Dys02a, Gor92, HT42, LT37, Smi85a, Smi85b]. X-Ray [HT42, Smi85a, Blu88a, DeW88, Gor92, Smi85b]. xii [Tho03, Wat03]. xiv [Car91, Edg91]. xvi [Edg91]. xxiii [Har05].

Yang [LY95, Tel95b]. Year [Tel48b, Tel99a]. Years [Mla98, Tel83d, Wol67, Mit97, TW79b, TZ91]. Yield [BO72]. York [Bor12, Car91, Dys58, Edg91, Hig76, Oli69, Pei87, Tel49a, Tel90b, Tel02a, Zie91, Bad77b, Dys76, San77, Wig76, Zac76]. York/Collier [Pei87]. Young [Sch03b, AH04]. Yuval [GT85].

Zeitzeugen [Sch01]. Zeman [Boy05]. Zhores [Tel81e]. zum [DB71]. zur [PT33, Tel30a, dEAH29, LT36, Ren34, TT32]. Zusammenarbeit [Dür05]. zwischen [Lin98, Ren34].

References

Anonymous:1991:ASF

[A^{+}91] Boris L. Altshuler et al., editors. Andrei Sakharov: facets of a life. Editions Frontières, Gif-sur-Yvette, France, 1991. ISBN 2-86332-096-3 (hardcover). 730 pp. LCCN QC16 .S25 A54 1991.

Alsop:1954:WA

[AA54] Joseph Alsop and Stewart Alsop. We accuse! Harper’s Magazine, ?? (??):25–45, October 1954. CODEN HAMAA3. ISSN 1045-7143. URL

http://harpers.org/archive/1954/10/we-accuse/.

Anderson:1952:PUC

[AAF^{+}52] Herbert L. Anderson, Samuel K. Allison, Enrico Fermi, Willard F. Libby, Joseph E. Mayer, Edward Teller, and Harold C. Urey. Protest from University of Chicago scientists. Bulletin of the Atomic Scientists, 8(7):255–256, October 1952. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Issue devoted to the restrictions of the McCarran Acts — the Internal Security Act of 1950, and the Immigration and Nationality Act of 1952 — on freedom of travel by scientists across US borders.

Anderson:1988:TAA

[AB88] Annelise Graebner Anderson and Dennis L. Bark, editors. Thinking about America: the United States in the 1990s, volume 375 of Hoover Press publication. Hoover Institution Press, Stanford, CA, USA, 1988. ISBN 0-8179-8751-7, 0-8179-8752-5 (paperback). xlvii + 590 pp. LCCN E876 .T47 1988.

Amaldi:2012:ALF

[ABES12] Edoardo Amaldi, S. (Saverio) Braccini, Antonio Ereditato, and Paola Scampoli, editors. The adventurous life of Friedrich Georg Houtermans, physicist (1903–1966). Springerbriefs in Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-32854-7 (paperback), 3-642-32855-5 (e-book). ISSN 2191-5423. xiv + 152 pp. LCCN Q124.6-

127.2; QC16.H688 A43 2012. URL http://www.sif.it/attivita/
saggiatore/recensioni/e_amaldi.

Asquith:1980:PPB

[AG80] Peter D. Asquith and Ronald N. (Ronald Nelson) Giere, editors. PSA 1980: proceedings of the 1980 biennial meeting of the Philosophy of science association. Contributed papers, volume 1. Philosophy of Science Association, East Lansing, MI, USA, 1980. ISBN 0-917586-14-X (hardcover), 0-917586-13-1 (paperback). LCCN ????

Asquith:1981:PPB

[AG81] Peter D. Asquith and Ronald N. (Ronald Nelson) Giere, editors. PSA 1980: proceedings of the 1980 Biennial Meeting of the Philosophy of Science Association, volume 2. Philosophy of Science Association, East Lansing, MI, USA, 1981. ISBN 0-917586-16-6. LCCN ????

Abbey:2004:BTP

[AH04] Cherie D. Abbey and Kevin Hillstrom, editors. Biography today: profiles of people of interest to young readers, volume 9 of Scientists and inventors. Omnigraphics, Detroit, MI, USA, 2004. ISBN 0-7808-0711-1. 222 pp. LCCN T39 .B4964 2004. URL http://www.loc.

gov/catdir/toc/fy0804/2008297730.html.

Albright:2010:JET

[Alb10] Eric D. Albright. Judging Edward Teller: A closer look at one of the most influential scientists of the Twentieth Century. Library Journal, 135(17):99–100, ???? 2010. CODEN LIBJA7. ISSN 0363-0277.

Allan:1956:RBH

[All56] H. R. Allan. Review of The Hydrogen Bomb: The Men, The Menace, the Mechanism. International Affairs (Royal Institute of International Affairs 1944–), 32(1):132–133, ???? 1956. ISSN 0020-5850 (print), 1468-2346 (electronic). URL http://www.jstor.org/

stable/2607946.

Ashkin:1943:TPS

[ALT43] Julius Ashkin, Willis E. (Willis Eugene) Lamb, and Edward Teller. Two problems in the statistical mechanics of crystals. I. The propagation of order in crystal lattices, by Julius Ashkin and Willis E. Lamb, Jr.. II. The statistics of two-dimensional lattices with four components, by Julius Ashkin and Edward Teller. PhD thesis, Columbia University, New York, NY, USA, 1943. 159–184 pp. URL http:/

/books.google.com/books?id=wpMESQAACAAJ. Reprinted from the

Physical Review 64(5-6), September 1–15, 1943.

Arnheim:1988:SFR

[ALW^{+}88] Rudolf Arnheim, Thomas E. Lovejoy, David Gordon Wilson, Freeman Dyson, Jane Goodall, Ian Shelton, Kenneth H. Olsen, Irene C. Peden, Richard W. Hamming, Thomas Eisner, Preston Cloud, Matt Cartmill, Samuel C. Florman, Jeremy Bernstein, George A. Miller, Robert M. May, G. Evelyn Hutchinson, Jerome Bruner, Priscilla C. Grew, William Bevan, Elisabeth S. Vrba, Myrdene Anderson, Kevin Padian, Harry Shipman, Victor F. Weisskopf, Walter A. Hill, Patricia D. Moehlman, Melvin Kranzberg, Malak Kotb, Raymond Kurzweil, Marcia McNutt, Masakazu Konishi, Miriam Rothschild, Edward Teller, Alison Jolly, H. Jane Brockmann, Keith Stewart Thomson, Peter J. Denning, Benoît B. Mandelbrot, Abraham Pais, Paul MacCready, Kip S. Thorne, Ruth Sager, Gerald J. Wasserburg, Neal E. Miller, Rita Levi-Montalcini, Stephen Jay Gould, Edwin H. Land, Michel Boudart, Anne Kernan, Douglas R. Hofstadter, Rosalyn S. Yalow, Bruce H. Tiffney, Mimi Koehl, Walter E. Massey, David P. Billington, John A. W. Kirsch, Abner Shimony, J. Donald Fernie, Brian J. Skinner, Lynn Margulis, Sheldon Lee Glashow, Michael LaBarbera, J. Tuzo Wilson, E. R. Ward Neale, Rudolf Peierls, Roald Hoffmann, Mary L. Good, Donald R. Griffin, Vaclav Smil, Michael S. Turner, Sarah Ann Woodin, Luis Alvarez, George A. Bartholomew, and George B. Schaller. Seventy-five reasons to become a scientist: American Scientist celebrates its seventyfifth anniversary. American Scientist, 76(5):450–463, September 1988. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/27855384;   http:
//www.jstor.org/stable/pdfplus/27855384.pdf.

Alpher:2023:GGR

[AM23] Victor S. Alpher and Simon A. Mitton. George Gamow and Ralph Alpher: a review of their cosmological collaboration as mentor and protégé 1942–1955. European Physical Journal H, 48(1):??, December 2023. CODEN EPJHAD. ISSN 2102-6459 (print), 2102-6467

(electronic). URL https://link.springer.com/article/10.1140/
epjh/s13129-023-00057-0.

Anderson:1973:TQA

[And73] Herbert Anderson. Three questions about the sustained nuclear chain reaction. The University of Chicago Magazine, 65(5):3–7, March/

April 1973. ISSN 0041-9508. URL http://library.ucsd.edu/dc/
object/bb3960604v.

Anderson:2008:WWR

[And08] Philip Anderson. Who or what is RVB? Physics Today, 61(4):8– 9, April 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://www.physicstoday.org/resource/

1/phtoad/v61/i4/p8_s1. According to a communication from the author reported in [Har10, pages 115, 502], Anderson was unaware of Teller’s work in [NST40] when he wrote this article.

Anonymous:1938:PAS

[Ano38] Anonymous. Professor Arnold Sommerfeld. Journal of Applied Physics, 9(12):754–755, December 1938. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. Note for the 70th birthday of Arnold Sommerfeld.

Anonymous:1955:NAW

[Ano55] Anonymous. National affairs: The work of many men. Time Magazine, ??(??):??, March 7, 1955. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). The article begins: “Edward Teller, father of the hydrogen bomb, last week published his eagerly awaited account of the explorations that made the weapon possible.”. The account referred to is [Tel55b].

Anonymous:1957:DTF

[Ano57a] Anonymous. Dr. Teller finds block to science is ‘square’. New York Times, ??(??):??, February 16, 1957. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1957:OBA

[Ano57b] Anonymous. Opinion: Beware the atomic bootlegger. Time Magazine, ??(??):??, December 30, 1957. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://content.time.

com/time/magazine/article/0,9171,810184,00.html.

Anonymous:1957:SET

[Ano57c] Anonymous. Scientist Edward Teller. Time Magazine, ??(??): ??, November 18, 1957. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://content.time.com/

time/covers/0,16641,19571118,00.html.

Anonymous:1958:FDD

[Ano58a] Anonymous. Fallout and disarmament: A debate between Linus Pauling and Edward Teller. Dædalus, 87(2):147–163, Spring 1958. CO-DEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL

http://www.jstor.org/stable/20026443.

Anonymous:1958:FII

[Ano58b] Anonymous. Fusion of identity irks non-atomic Dr. Teller. New York Times, ??(??):??, September 5, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1960:DTC

[Ano60a] Anonymous. Dr. Teller cites teaching ‘crisis’: H-bomb physicist proposes better science courses — pupil talks open here. New York Times, ??(??):??, October 12, 1960. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1960:NIR

[Ano60b] Anonymous. New issue raised on Test Ban Pact. New York Times, ?? (??):??, March 18, 1960. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1961:KCE

[Ano61a] Anonymous. Kennedy confers with experts on nuclear research. New York Times, ??(??):??, December 1, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1961:SRW

[Ano61b] Anonymous. Space race warning: Teller says Russians have power to shape world. New York Times, ??(??):??, November 5, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://wvw.jfklibrary.org/Historica+Resources/
Archives/Reference+Desk/Speeches/JFK/003POF03NationalNeeds05251961.
htm.

Anonymous:1962:DTR

[Ano62] Anonymous. Dr. Teller receives $50,000 Enrico Fermi Award. New York Times, ??(??):??, December 4, 1962. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1963:ETT

[Ano63a] Anonymous. Excerpts from testimony by Teller. New York Times, ?? (??):??, August 15, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1963:TTB

[Ano63b] Anonymous. Teller test-ban warning. New York Times, ??(??):??, February 1, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1965:FHB

[Ano65] Anonymous. ‘Father of H-bomb agrees to rally scientific talent. New York Times, ??(??):??, December 31, 1965. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1971:NWD

[Ano71] Anonymous. New world: Dissent blooms at AAAS circus. Nature, 229(5280):81–82, January 8, 1971. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v229/n5280/pdf/229081a0.pdf.

Anonymous:1973:TCT

[Ano73] Anonymous. Teller consults Thais on nuclear-dug canal. New York Times, ??(??):3, January 14, 1973. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1979:DTA

[Ano79a] Anonymous. Dr. Teller attributes ill health to foes of nuclear power. New York Times, ??(??):B4, August 1, 1979. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/search.proquest.com/hnpnewyorktimes/docview/120765871/.

Anonymous:1979:TTH

[Ano79b] Anonymous. Teller tested in hospital. New York Times, ??(??):??, May 12, 1979. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Anonymous:1987:TLL

[Ano87] Anonymous. Teller and the laser lobby. Bulletin of the Atomic Scientists, 43(8):21, October 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:HPH

[Ano89a] Anonymous. How Point Hope, Alaska took on Edward Teller. Bulletin of the Atomic Scientists, 45(10):[cover], December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1989:JO

[Ano89b] Anonymous. Just an oversight. Bulletin of the Atomic Scientists, 45 (6):12, July/August 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBD

[Ano90a] Anonymous. Bulletin: In brief: Did he take a pill?; too cheap to meter; starry eyed [Edward Teller]; think no small thoughts; good timing?; no preconceptions. Bulletin of the Atomic Scientists, 46 (7):4, September 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1990:BBMb

[Ano90b] Anonymous. Bulletin: In brief: The more things change …; whew! past the hard part; still a solution to global warming? Pebbles not so brilliant. Bulletin of the Atomic Scientists, 46(8):6, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1991:ETV

[Ano91] Anonymous. Edward Teller visit to Hungary. Energia es Atomtechnika, 44(1):14, ???? 1991. ISSN 0013-7316.

Anonymous:1993:CGB

[Ano93] Anonymous. Corrections: [German bomb program and Team B]. Bulletin of the Atomic Scientists, 49(4):6, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Bet93, CP93].

Anonymous:1995:BBD

[Ano95] Anonymous. Bulletin: In brief: Don’t phone home; Hitler’s hideaway as hot potato; Teller wants more; peace dividend update; secret — or dumb? counting (radiation-free) sheep; widening the customer base; bait and switch?; taking his word for it. Bulletin of the Atomic Scientists, 51(3):7–8, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1996:AL

[Ano96] Anonymous. Alexander Langsdorf. Bulletin of the Atomic Scientists, 52(4):10, July/August 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1998:TEE

[Ano98] Anonymous. Teller Ede 90 éves. Fizikai Szemle (Budapest), ??(??): ??, 1998. CODEN FISZA6. ISBN 963-8051-77-9. ISSN 0015-3257 (print), 1588-0540 (electronic).

Anonymous:19xx:BMP

[Anoxxa] Anonymous. The birth of modern physics. Los Alamos National Laboratory Web document, 19xx. URL http://www.lanl.gov/

history/road/birthofmodernphysics.shtml. Section The Einstein

Letter discusses the Einstein–Roosevelt letter of 2 August 1939, and says “drafted mostly by Szilard.”.

Anonymous:19xx:DET

[Anoxxb] Anonymous. The debate: Edward Teller and J. Robert Oppenheimer. Report, Los Alamos National Laboratory, Los Alamos,

NM, USA, 19xx. URL http://www.lanl.gov/history/postwar/
debate.shtml.

Anonymous:2000:MT

[Ano00] Anonymous. The moment in time. [Los Alamos National] Laboratory produced documentary on the Manhattan Project.,

April 20, 2000. URL http://www.governmentattic.org/11docs/
NNSAfilmVideoList_2014.pdf;   https://www.youtube.com/watch?

v=xwpgmEvlRpM. The hour-long program features interviews with some of those who worked on the Project Hans Bethe, Norman Ramsey, Phil Morrison, Edward Teller and others. The program profiles the creation of the Manhattan Project the obstacles that were faced, life in the “secret city,’ and events that led up to and followed the Trinity Test.

Anonymous:2002:EET

[Ano02] Anonymous. Encounters with Edward Teller. Bulletin of the Atomic Scientists, 58(1):4, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:2003:ETA

[Ano03a] Anonymous. Edward Teller awarded Presidential Medal of Freedom during White House ceremony. Web press release, July 23,

2003. URL https://www.llnl.gov/news/newsreleases/2003/NR-
03-07-07.html.

Anonymous:2003:GET

[Ano03b] Anonymous. Goodbye: Edward Teller. Nature, 426(6968):757, December 18, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-

4687 (electronic). URL http://www.nature.com/nature/journal/
v426/n6968/full/426757b.html.

Anonymous:2003:RHU

[Ano03c] Anonymous. Recipients of highest US civilian honor include Teller. Physics Today, 56(10):77–78, October 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

physicstoday.org/resource/1/phtoad/v56/i10/p77_s3.

Anonymous:2003:TE

[Ano03d] Anonymous. Teller Ede 1908–2003. Fizikai Szemle (Budapest), ??(??): ??, September 2003. CODEN FISZA6. ISSN 0015-3257 (print), 1588-

0540 (electronic). URL http://www.kfki.hu/fszemle/archivum/
fsz0309/tede0309.html.

Anonymous:2005:BRB

[Ano05a] Anonymous. Book review: Edward Teller: the Real Dr. Strangelove, by Peter Goodchild. Bulletin of the Atomic Scientists, 61(2):69, March/April 2005. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/1G1-

129710745.html. See [Goo04a, Goo04b].

Anonymous:2005:SPG

[Ano05b] Anonymous. SA Perspectives: Great Teller’s Ghost. Scientific American, 293(2):8, August 2005. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http:/

/www.nature.com/scientificamerican/journal/v293/n2/full/
scientificamerican0805-8.html;   http://www.nature.com/scientificamerican/
journal/v293/n2/pdf/scientificamerican0805-8.pdf.

Anonymous:2005:TE

[Ano05c] Anonymous. Teller, Edward. In Mitcham [Mit05], page ?? ISBN 0-02-865831-0 (set ), 0-02-865832-9 (vol. 1), 0-02-865833-7 (vol. 2), 0-02-865834-5 (vol. 3), 0-02-865901-5 (vol. 4). LCCN Q175.35 .E53

2005. URL http://www.highbeam.com/doc/1G2-3434900676.html.

Four volumes.

Anonymous:2008:ETS

[Ano08a] Anonymous. Edward Teller: Scientist and teacher. Web site,

May 9, 2008. URL https://www.llnl.gov/news/edward-teller-
scientist-and-teacher.

Anonymous:2008:TCH

[Ano08b] Anonymous. Teller centennial highlights. Web document, 2008.

URL https://str.llnl.gov/str2009teller.html. In 2007–2008,

Science and Technology Review published the following series to mark the 100th anniversary of Edward Teller’s birth on January 15, 1908. . ..

Anonymous:2013:OSH

[Ano13] Anonymous. The Oppenheimer security hearing. Web document.,

Anonymous:2015:NMP

[Ano15] Anonymous. New memoir by participant in U.S. H-bomb program sheds light on the making of the first test device. First-hand perspectives on Edward Teller and other leading figures, and on dispute over who originated key idea of radiation implosion. Book by Kenneth W. Ford being published over objections of Department of Energy. First full account of Project Matterhorn, pioneering effort in use of computer technologies for nuclear weapons development. National Security Archive Electronic Briefing Book 507, National Security Archive, George Washington University, Washington, DC, USA, March 24,

2015. URL http://nsarchive.gwu.edu/nukevault/ebb507/.

Anonymous:20xx:ET

[Anoxxa] Anonymous. Edward Teller. English-language Web site in Hungary.,

20xx. URL http://jewish.hu/teller-ede/.

Anonymous:20xx:WCT

[Anoxxb] Anonymous. Washington conferences on theoretical physics. Web

document, 20xx. URL http://home.gwu.edu/~kargaltsev/HEA/

washington-conferences.html. Undated. The page includes a photograph of a plaque with the preface “The most famous event at this 5th Washington Conference on Theoretical Physics came from the announcement by Niels Bohr at the 1939 conference, in the Hall of Government, Room 209, that the nucleus of uranium had been split by bombardment with neutrons, with significant energy released. This was the dawn of the atomic age.” and the engraving: “In this room, January 26, 1939, Niels Bohr made the first pubic announcement of the successful disintegration of uranium into barium with the attendant release of approximately two hundred million electron volts of energy per disintegration. This announcement was heard by the physicists listed below who where attending the fifth of the conferences on theoretical physics which are sponsored jointly by the Carnegie Institution of Washington and The George Washington University.” The participant listed on the plaque are: L. H. Adams; Donald Hatch Andrews; Ferdinand G. Brickwedde; Gerhard Heinrich Dieke; George A. Gamow; Maria Goeppert-Mayer; M. H. Hebb; Karl Ferdinand Herzfeld; J. H. Hibben; J. H. Hoge; D. R. Inglis; F. G. Keyes; F. C. Kracek; R. Myers; H. M. O’Bryan; E. Posnjak; A. E. Ruark; R. B. Scott; Francis B. Silsbee; C. Starr; Otto Stern; Edward Teller; Harold C. Urey; and B. D. van Evera.

Alfven:1949:OCR

[ART49] H. Alfven, R. D. Richtmyer, and E. Teller. On the origin of cosmic rays. Physical Review, 75(5):892–893, March 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v75/i5/p892_2.

Asimov:1962:LES

[Asi62] Isaac Asimov. Letter to the Editor: Scientist, writer. Bulletin of the Atomic Scientists, 18(7), September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mor62].

Ashkin:1943:STD

[AT43a] J. Ashkin and E. Teller. Statistics of two-dimensional lattices with four components. Physical Review, 64(5–6):178–184, September 1, 1943. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v64/i5-6/
p178_1.

Axilrod:1943:IVW

[AT43b] B. M. Axilrod and E. Teller. Interaction of the van der Waals type between three atoms. Journal of Chemical Physics, 11(6):299– 300, June 1943. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/doi/10.1063/
1.1723844.

Argo:1944:NEP

[AT44] M. Argo and E. Teller. Neutron emission by polonium oxide layers. Report LAMS-121, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 8, 1944. Declassified 1956.

Auer:2003:AAP

[Aue03] Catherine Auer. As American as atomic pie. Bulletin of the Atomic Scientists, 59(2):6–8, March 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-99147968.html.

Bethe:1975:SSN

[BAA^{+}75] Hans A. Bethe, Luis Alvarez, Peter Auer, William O. Baker, John Bardeen, Robert F. Bacher, Felix Bloch, Norris E. Bradbury, Harold Brown, Richard H. Chamberlain, Cyril L. Comar, Arthur Kantrowitz, Ralph E. Lapp, Joshua Lederberg, Willard F. Libby, Franklin A. Long, Edwin M. McMillan, Kenneth S. Pitzer, Edward M. Purcell, I. I. Rabi, Norman Rasmussen, Roger Revelle, Glenn T. Seaborg, Frederick Seitz, Edward Teller, James A. Van Allen, Warren Weaver, Alvin Weinberg, Victor F. Weisskopf, Edward Wenk, Jr., Eugene Wigner, and Richard Wilson. 32 scientists speak out: “No alternative to nuclear power”. Bulletin of the Atomic Scientists, 31(3):4–5, March 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See correction [Wei75] about the incorrect inclusion of Victor Weisskopf in the authors. See comments [Gir75, Smi75, Sau75, Gul75].

Badash:1977:BRBb

[Bad77a] Lawrence Badash. Book review: Energy and Conflict. The Life and Times of Edward Teller by Stanley A. Blumberg; Gwinn Owens. Isis, 68(3):497–498, September 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/231367.

Badash:1977:BRBa

[Bad77b] Lawrence Badash. Book review: The Advisors. Oppenheimer, Teller, and the Superbomb by Herbert York. Isis, 68(1):165–167, March 1977. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic).

URL http://www.jstor.org/stable/230424.

Ball:1986:JDA

[Bal86] Howard Ball. Justice downwind: America’s atomic testing program in the 1950’s. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1986. ISBN 0-19-503672-7. xviii + 280 pp. LCCN U264 .B35

1986. US$21.95. URL http://www.loc.gov/catdir/enhancements/
fy0636/85008777-d.html.

Ball:2003:HBI

[Bal03] Philip Ball. H-bomb inventor Edward Teller dies: Controversial life in nuclear-weapons research and policy ends at 95. Nature News, ?? (??):??, September 11, 2003. URL http://www.nature.com/news/

2003/030911/full/news030908-6.html.

Bankston:2002:ETD

[Ban02] John Bankston. Edward Teller and the development of the hydrogen bomb. Unlocking the secrets of science. Mitchell Lane Publishers, Bear, DE, 2002. ISBN 1-58415-108-0. 56 pp. LCCN QC16.T37 B36 2002.

Barker:1965:MCC

[Bar65] A. A. Barker. Monte Carlo calculations of the radial distribution functions for a proton-electron plasma. Australian Journal of Physics, 18 (??):119–133, ???? 1965. CODEN AUJPAS. ISSN 0004-9506 (print), 1446-5582 (electronic).

Barinaga:1988:GRV

[Bar88] Marcia Barinaga. GAO report vindicates Teller but critics disagree. Nature, 334(6180):282, July 28, 1988. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v334/n6180/pdf/334282a0.pdf.

Barkoczy:1990:HKT

[Bar90] Barkóczy Péter. Három kivándorlás: Teller Ede az Amerika Hangja stúdiójában. (Hungarian) [Three emigrations: Edward Teller, the Voice of America studio], 1990. One 64-minute VHS cassette.

Bazarko:1990:ETG

[Baz90] A. O. Bazarko. Edward Teller — giant of the golden age of physics — S. A. Blumberg, L. G. Panos. Journal of International Affairs, 44 (1):269–270, ???? 1990. ISSN 0022-197X.

Balazs:2000:ONC

[BBLS00] Nandor L. Balazs, John C. Browne, James D. Louck, and Daniel S. Strottman. Obituary: Nicholas Constantine Metropolis. Physics Today, 53(10):100, October 2000. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.aip.org/pt/vol-

53/iss-10/p100.html.

Bloch:1946:LSN

[BCC^{+}46] Felix Bloch, Robert F. Christy, C. Louis Critchfield, Edwin M. McMillan, Emilio Segrè, Edward Teller, Victor Frederick Weisskopf, John H. Williams, Hans A. Bethe, Robert E. Marshak, Lyman G. Parratt, Bruno Rossi, John W. Trischa, Rudolf Peierls, Klaus Fuchs, John Manley, Bernard T. Feld, Julian Schwinger, Joseph Keller, Julius Askin, Chaim Richman, George B. Kistiakowsky, Leonard Schiff, Eugene M. Baroody, and Emil J. Konopinski. Lecture series on nuclear physics, 1943–1944. Report LA-24, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1946. 329 pp. URL https://www.

pbagalleries.com/view-auctions/catalog/id/352/lot/106762/
LA-Report-24-Lecture-Series-on-Nuclear-Physics.

Baltimore:1997:WCB

[BCF^{+}97] David Baltimore, Robert Coles, John B. Fagan, Immanuel Jakobovits, Bill Frist, Leon R. Kass, John O’Connor, Robert Silverberg, and Edward Teller. Will cloning beget disaster? Wall Street Journal (Eastern edition), ??(??):A14, May 2, 1997. ISSN 0099-9660.

Baliberra:1982:FRE

[BD82] Mario S. Baliberra and V. A. Dulock, Jr. Film review: Edward Teller, An Early Time (film). American Journal of Physics, 50(3):286, March 1982. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://scitation.aip.org/content/aapt/journal/
ajp/50/3/10.1119/1.13042.

Brunauer:1940:TVW

[BDDT40] Stephen Brunauer, Lola S. Deming, W. Edwards Deming, and Edward Teller. Theory of the van der Waals adsorption of gases. Journal of the American Chemical Society, 62(7):1723–1732, July 1940. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Beichman:2004:TPE

[Bei04] Arnold Beichman. Tribute to a patriot [Edward Teller]. Hoover Digest, 2004(1):??, January 30, 2004. ISSN 1088-5161. URL http://

www.hoover.org/publications/hoover-digest/article/8096.

Bernstein:1978:BBC

[Ber78] Barton J. Bernstein. Books: Energy and Conflict: The Life and Times of Edward Teller, by Stanley Blumberg and Gwinn Owens. Bulletin of the Atomic Scientists, 34(5):51–53, May 1978. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BO76].

Bernstein:1980:HBP

[Ber80] Jeremy Bernstein. Hans Bethe, prophet of energy. Basic Books, New York, NY, USA, 1980. ISBN 0-465-02903-5. xii + 212 pp. LCCN QC16.B46 B47. US$11.95.

Bersuker:1984:JTEb

[Ber84a] I. B. (Isaak Borisovich) Bersuker. The Jahn–Teller effect: a bibliographic review. IFI/Plenum, New York, NY, USA, 1984. ISBN 0-306-65206-4. ix + 589 pp. LCCN Z5524.P6 B47 1984.

Bersuker:1984:JTEa

[Ber84b] I. B. (Isaak Borisovich) Bersuker. The Jahn–Teller effect and vibronic interactions in modern chemistry. Modern inorganic chemistry. Plenum Press, New York, NY, USA; London, UK, 1984. ISBN 0-306-41319-1. xiv + 319 pp. LCCN QD461 .B46 1984.

Bernstein:1987:LIB

[Ber87] Jeremy Bernstein. The life it brings: one physicist’s beginnings. Ticknor and Fields, New York, NY, USA, 1987. ISBN 0-89919-470-2. xv + 171 + 8 pp. LCCN QC16.B458 A3 1987. US$16.95.

Berman:1988:NGR

[Ber88] B. L. Berman. Nuclear giant resonances — a historical review. In Mark and Wood [MW88], pages 49–85. ISBN 90-277-2775-9. LCCN QC71 .E62 1988. US$.

Bernstein:1990:BRB

[Ber90] Barton J. Bernstein. Book review: Better a Shield Than a Sword: Perspectives on Defense and Technology by Edward Teller. Technology and Culture, 31(4):846–861, October 1990. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL http://www.

jstor.org/stable/pdfplus/3105912.pdf. See [Tel87e].

Bernstein:2003:RAG

[Ber03a] Barton J. Bernstein. Reconsidering the “Atomic General”: Leslie R. Groves. The Journal of Military History, 67(3):883–920, July 2003. ISSN 0899-3718 (print), 1543-7795 (electronic). URL http://www.

jstor.org/stable/3397331.

Bernstein:2003:CBR

[Ber03b] J. Bernstein. Comment on book review of “Brotherhood of the bomb: The tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller,” by Gregg Herken [Am. J. Phys. 71(4), 411-415 (2003)] Bernstein’s response ’. American Journal of Physics, 71(7):648, ???? 2003. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Bernstein:2003:BRB

[Ber03c] Jeremy Bernstein. Book review: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken. American Journal of Physics, 71(4):411–415, April 2003. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/
1/ajpias/v71/i4/p411_s1;   http://link.aip.org/link/?AJP/

71/411/1. See comment [Her03].

Bernstein:2003:DWH

[Ber03d] Jeremy Bernstein. The drawing, or why history is not mathematics. Physics in Perspective (PIP), 5(3):243–261, 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Bernstein:2004:ETRb

[Ber04a] B. J. Bernstein. Edward Teller: the real Dr Strangelove. Nature, 430 (6997):293–294, July 15, 2004. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Bernstein:2004:ETRa

[Ber04b] J. Bernstein. Edward Teller: The real Dr. Strangelove. The Nation, 279(14):31–34, ???? 2004. ISSN 0027-8378.

Bersuker:2006:JTE

[Ber06] I. B. (Isaak Borisovich) Bersuker. The Jahn–Teller effect. Cambridge University Press, Cambridge, UK, 2006. ISBN 0-521-82212-2 (hardcover). xvi + 616 pp. LCCN QD461 .B458 2006. URL

http://www.loc.gov/catdir/enhancements/fy0668/2006297795-
d.html;   http://www.loc.gov/catdir/enhancements/fy0668/
2006297795-t.html;   http://www.loc.gov/catdir/enhancements/
fy0732/2006297795-b.html.

Bernstein:2010:JNK

[Ber10] Jeremy Bernstein. John von Neumann and Klaus Fuchs: an unlikely collaboration. Physics in Perspective (PIP), 12(1):36–50, March 2010. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://link.springer.com/article/10.1007/
s00016-009-0001-1;   http://www.springerlink.com/content/
y54804274h88v546/.

Brunauer:1938:AGM

[BET38] Stephen Brunauer, Paul Hugh Emmett, and Edward Teller. Adsorption of gases in multimolecular layers. Journal of the American Chemical Society, 60(2):309–319, February 1938. CODEN JACSAT. ISSN 0002-7863 (print), 1520-5126 (electronic), 1943-2984.

Bethe:1946:CAW

[Bet46] Hans A. Bethe. Can air or water be exploded? Bulletin of the Atomic Scientists, 1(7):2, 14, March 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Bethe:1982:CHH

[Bet82] Hans A. Bethe. Comments on the history of the H-bomb. Los Alamos Science, 6:42–53, Fall 1982. CODEN LASCDI. ISSN 0273-7116. URL

http://library.lanl.gov/cgi-bin/getfile?06-03.pdf. This ar-

ticle corrects a few errors in one book [Goo80, Goo81], and points out numerous errors, fallacies, misrepresentations, and misunderstandings, in an earlier book [SB54]. It also contains a substantial discussion of the role of Edward Teller in the work as Los Alamos, and later, at Lawrence Livermore National Laboratory.

Bethe:1991:RAP

[Bet91] Hans Albrecht Bethe. The road from Los Alamos: Profound perspective and personal viewpoints on atomic weapons, nuclear power, and science. Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1991. ISBN 0-88318-707-8, 0-671-74012-1. xvii + 286 pp. LCCN U264 .B455 1991. US$24.95.

Bethe:1993:LBG

[Bet93] Hans A. Bethe. Letter: Bethe on the German bomb program. Bulletin of the Atomic Scientists, 49(1):53–54, January/February 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See corrections [Ano93].

Bethe:1997:SWH

[Bet97] Hans A. (Hans Albrecht) Bethe. Selected works of Hans A. Bethe: with commentary, volume 18 of World Scientific series in 20th century physics. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1997. ISBN 981-02-2876-7. viii + 605

pp. LCCN QC780 .B452 1996. URL http://www.worldscibooks.
com/physics/3295.html.

Bethe:2001:ETL

[Bet01] Hans A. Bethe. Edward Teller: A long look back: Book review: Edward Teller, Memoirs: A Twentieth-Century Journey in Science and Politics. Physics Today, 54(11):55–56, November 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v54/i11/p55/s1.

Biro:2005:MET

[BG05] Tamás Sándor Biró and Walter Greiner. In memoriam Edward Teller. Acta Physica Hungarica New Series: Heavy Ion Physics, 22(1–2):1– 2, ???? 2005. CODEN APHAC8. ISSN 1219-7580 (print), 1588-

2675 (electronic). URL http://www.springerlink.com/content/
t32g7u2121251325/.

Brink:1989:IET

[BH89] Jean R. Brink and Roland Haden. Interviews with Edward Teller and Eugene P. Wigner. Annals of the History of Computing, 11(3):177– 178, July/September 1989. CODEN AHCOE5. ISSN 0164-1239. URL

http://dlib.computer.org/an/books/an1989/pdf/a3177.pdf;
http://www.computer.org/annals/an1989/a3177abs.htm.

Bowes:1991:LMF

[BHM91] Stanley A. Bowes, Daniel Hirsch, and William G. Mathews. Letters: More fallout from H-bomb controversy. Bulletin of the Atomic Scientists, 47(9):45–46, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90] and comment [Han92b].

Badash:1986:NFR

[BHT86] Lawrence Badash, Elizabeth Hodes, and Adolph Tiddens. Nuclear fission: Reaction to the discovery in 1939. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 130(2):196–231, June 1986. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.

org/stable/987181.

Bird:1967:TBN

[Bir67] David Bird. Teller backs new smog curbs as a new source of broad benefits. New York Times, ??(??):??, September 14, 1967. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Bird:2002:BRB

[Bir02] Kai Bird. Book review: Bomb thrower: Memoirs: A Twentieth-Century Journey in Science and Politics by Edward Teller and Judith L. Shoolery. The Wilson Quarterly (1976–2012), 26(1):109–111, Winter 2002. CODEN ???? ISSN 0363-3276. URL http://www.

jstor.org/stable/pdfplus/40260575.pdf.

Bothe:1941:ATN

[BJ41] W. Bothe and P. Jensen. Die Absorption thermischer Neutronen in Elektrographit. (German) [The absorption of thermal neutrons in electrographite]. Report G-71, ????, ????, January 20, 1941. Captured German report cited in [Wea76, page 29] that contained incorrect results on the neutron capture cross-section of carbon. Those results caused the German Uranium Project scientists to switch from graphite to the hard-to-get deuterium (in heavy water obtained from the Rjukan Plant in Norway) as a neutron moderator, and likely, substantially delayed their progress in achieving a working nuclear reactor as a precursor to an atomic bomb.

Bachar:1988:LBW

[BJCP88] John Bachar, Jack R. Jennings, Antonie K. Churg, and Sheldon C. Plotkin. Letter: Behind the Woodruff disclosure. Bulletin of the Atomic Scientists, 44(7):60, September 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu88a].

Bloom:1969:AKM

[BJT69] Stewart D. Bloom, Montgomery H. Johnson, and Edward Teller. Absorption of K^{−}-mesons in nuclear surfaces. Physical Review Letters, 23 (1):28–30, July 7, 1969. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145.

Bloom:1972:KMA

[BJT72] Stewart D. Bloom, Montgomery H. Johnson, and Edward Teller. Kmesic atoms and capture in the nuclear surface. In ????, editor, ????, pages 89–111. W. H. Freeman and Company, New York, NY, USA, 1972. ISBN ???? LCCN ????

Brunings:1941:MLH

[BKT41] J. H. M. Brunings, J. K. Knipp, and E. Teller. On the momentum loss of heavy ions. Physical Review (2), 60(9):657–660, November 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v60/i9/p657_1.

Blount:1971:JTT

[Blo71] E. I. Blount. The Jahn–Teller theorem. Journal of Mathematical Physics, 12(9):1890–1896, September 1, 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Bloom:1984:GTS

[Blo84] Stewart D. Bloom. Gamow–Teller strength functions with the Lanczos algorithm. Progress in Particle and Nuclear Physics, 11(??):505– 528, ???? 1984. CODEN PPNPDB. ISSN 0146-6410 (print), 1873-2224 (electronic).

Bloom:1988:PGT

[Blo88] Stewart D. Bloom. On the prevalence of the Gamow–Teller transition. In Mark and Wood [MW88], pages 15–37. ISBN 90-277-2775-9. LCCN QC71 .E62 1988.

Blumberg:1972:PTS

[Blu72a] Stanley A. Blumberg. Physicist Teller sounds alarm. News American (Baltimore, Maryland), ??(??):??, June 18, 1972.

Blumberg:1972:SSI

[Blu72b] Stanley A. Blumberg. Secrecy in science: An interview with Dr. Edward Teller. Johns Hopkins Magazine, 24(??):8–16, Winter 1972. ISSN 0021-7255.

Blum:1988:WSL

[Blu88a] Deborah Blum. Weird science: Livermore’s X-ray laser flap. Bulletin of the Atomic Scientists, 44(6):7–13, July/August 1988. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [BJCP88, DeW88, Cun88].

Blumenthal:1988:WGM

[Blu88b] Sidney Blumenthal. When giants meet: H-bomb fathers Sakharov & Teller’s SDI dialogue. Washington Post, ??(??):??, November 17, 1988. ISSN 0190-8286.

Brown:2004:ETP

[BM04] Harold Brown and Michael May. Edward Teller in the public arena. Physics Today, 57(8):51–53, August 2004. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//physicstoday.org/resource/1/phtoad/v57/i8/p51_s1.

Boorse:1989:ASB

[BMW89] Henry A. (Henry Abraham) Boorse, Lloyd Motz, and Jefferson Hane Weaver. The atomic scientists: a biographical history. Wiley science editions. John Wiley, New York, NY, USA, 1989. ISBN 0-471-50455-6. vii + 472 pp. LCCN QC773 .B66 1989.

Blumberg:1972:OSW

[BO72] Stanley A. Blumberg and Gwinn Owens. Oil shale: Will it yield its treasure to a nuclear blast? The Evening Sun (Baltimore, MD), ?? (??):??, December 12, 1972.

Blumberg:1976:ECL

[BO76] Stanley A. Blumberg and Gwinn Owens. Energy and conflict: the life and times of Edward Teller. Putnam, New York, NY, USA, 1976. ISBN 0-399-11551-X. xvii + 492 + 4 pp. LCCN QC16.T37 B58

1976. US$12.95. URL http://alsos.wlu.edu/information.aspx?
id=367.

Boffey:1969:EJS

[Bof69] Philip M. Boffey. Ernest J. Sternglass: Controversial prophet of doom. Science, 166(3902):195–200, October 10, 1969. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://

www.sciencemag.org/content/166/3902/195.extract.

Boffey:1971:ACR

[Bof71] Philip M. Boffey. AAAS Convention: Radicals harass the establishment. Science, 171(3966):47–49, January 1971. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

sciencemag.org/content/171/3966/47.extract.

Boffey:1985:DSS

[Bof85] Philip M. Boffey. Dark side of ‘Star Wars’: System could also attack. New York Times, ??(??):??, March 7, 1985. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Blumberg:1989:TUH

[BOG89] Stanley A. Blumberg, Gwinn Owens, and Egri György. A Trefort utcától a hidrogénbombáig: [Edward Teller élete és kora]. (Hungarian) [Energy and conflict: The life and times of Edward Teller]. M. Világ, Budapest, Hungary, 1989. ISBN 963-7815-12-0. 190 pp. LCCN ????

Borgese:1948:WMC

[Bor48] Elizabeth Mann Borgese. Why a maximalist constitution? Bulletin of the Atomic Scientists, 4(7):199–204, July 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com48].

Born:1957:MA

[Bor57] Max Born. Man and the atom. Bulletin of the Atomic Scientists, 13 (6):186–194, June 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Bor63].

Born:1963:MA

[Bor63] Max Born. Man and the atom. In Grodzins and Rabinowitch [GR63], pages 590–601. LCCN D842 .B78. Reprint of [Bor57].

Borcherds:2003:BRB

[Bor03] P. Borcherds. Book review: Conversations on the Dark Secrets of Physics, by Edward Teller. European Journal of Physics, 24(4):495, ???? 2003. CODEN EJPHD4. ISSN 0143-0807 (print), 1361-6404 (electronic).

Borhi:2012:IHJ

[Bor12] László Borhi. István Hargittai, Judging Edward Teller: a Closer Look at One of the Most Influential Scientists of the Twentieth Century. New York: Prometheus Books, 2010. 575 pp. Journal of Cold War Studies, 14(4):263–266, Fall 2012. CODEN ???? ISSN 1520-3972 (print), 1531-3298 (electronic). URL http://muse.jhu.

edu/article/500353;   http://muse.jhu.edu/article/500353/
pdf;   http://muse.jhu.edu/journals/journal_of_cold_war_
studies/v014/14.4.borhi.pdf;   https://direct.mit.edu/jcws/
article-pdf/14/4/263/698191/jcws_r_00300.pdf.

Boyer:1995:BRR

[Boy95] Paul D. Boyer. Book review: The rise and fall of Project Chariot: The Firecracker Boys, by Daniel T. O’Neill. Bulletin of the Atomic Scientists, 51(3):62–65, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

highbeam.com/doc/1G1-16888670.html.

Boyer:2005:NTB

[Boy05] Paul Boyer. The nostalgia trap: Book review: Atomic Culture: How We Learned to Stop Worrying and Love the Bomb, edited by Scott C. Zeman and Michael A. Amundson. Bulletin of the Atomic Scientists, 61(1):67–69, January/February 2005. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/61/1/67.full.pdf+html;   http://www.highbeam.
com/doc/1G1-126748359.html. See [ZA04].

Bersuker:1983:VVV

[BP83] I. B. (Isaak Borisovich) Bersuker and V. Z. (Victor Zigfridovich) Polinger. Vibronnie vzaimodeistviya v molekulakh i kristallahk. (Russian) [Vibronic interactions in molecules and crystals]. Nauka, Moscow, USSR, 1983. ISBN ???? 336 pp. LCCN QD461 .B463 1983.

Bersuker:1989:VIM

[BP89] I. B. (Isaak Borisovich) Bersuker and V. Z. (Viktor Zigfridovich) Polinger. Vibronic interactions in molecules and crystals, volume 49 of Springer series in chemical physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1989. ISBN 0-387-19259-X. xii + 422 pp. LCCN QD461 .B46313 1989.

Blumberg:1990:ETG

[BP90] Stanley A. Blumberg and Louis G. Panos. Edward Teller: giant of the golden age of physics: a biography. Scribner’s, New York, NY, USA, 1990. ISBN 0-684-19042-7. xiv + 306 + 8 pp. LCCN QC16.T37 B57 1990. US$24.95 (CAN$34.95). URL http://alsos.wlu.edu/

information.aspx?id=47.

Bersuker:2012:VIM

[BP12] Isaak B. Bersuker and Victor Z. Polinger. Vibronic interactions in molecules and crystals. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-83481-7. xii + 422 pp. LCCN ????

Bradford:1979:CTM

[Bra79] Peter A. Bradford. The ‘casualty’ of Three Mile Island. Washington Post, ??(??):??, October 25, 1979. ISSN 0190-8286.

Breit:1953:NTA

[Bre53] Gregory Breit. Notes for talk on atmospheric ignition. Report ????, Los Alamos Scientific Laboratory of the University of California, Los Alamos, NM, USA, 1953. This is a talk presented before the first hydrogen-bomb test on 1 November 1952. The work showed that it was impossible for the bomb explosion to ignite the atmosphere or oceans, as had earlier work by Bethe and by Konopinski, Marvin, and Teller, for the atomic bomb. According to [Har10, page 224], this report remained classified until about 1975, when it was mentioned in an newspaper story [Sul75].

Brink:1965:NF

[Bri65] David Maurice Brink. Nuclear Forces, volume 354 of Commonwealth and international library: Selected readings in physics. Pergamon, New York, NY, USA, 1965. ISBN 0-08-011034-7. viii + 232 pp. LCCN QC173 .B8513 1965.

Brinkley:2000:IWT

[Bri00] Alan Brinkley. An idea whose time will not go: Ronald Reagan’s missile shield is still alive, though it shows no sign of consciousness. New York Times, ??(??):??, April 16, 2000. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Broad:1982:RHH

[Bro82] William J. Broad. Rewriting the history of the H-bomb: Nobel laureate Hans Bethe says technical errors by Edward Teller, not political opposition by Robert Oppenheimer, hindered work on the superbomb. Science, 218(4574):769–772, November 19, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1689737;   http:
//www.sciencemag.org/content/218/4574/769.extract. See re-

sponse [TB82].

Broad:1992:TWT

[Bro92] William J. Broad. Teller’s war: the top-secret story behind the Star Wars deception. Simon and Schuster, New York, NY, USA, 1992. ISBN 0-671-70106-1. 350 pp. LCCN UG486 .B76 1992. US$25.00.

URL http://alsos.wlu.edu/information.aspx?id=2018.

Broad:1996:THU

[Bro96] William J. Broad. Teller háborúja: az USA csillagháborús tévútjának szigorúan titkos története. (Hungarian) [Teller’s war: the top-secret story behind the Star Wars deception]. Osiris könyvtár. Modern újságírás. Osiris, Budapest, Hungary, 1996. ISBN 963-379-128-6. ISSN 1219-7017. 341 + 6 pp. LCCN ????

Boag:1990:KCM

[BRS90] J. W. Boag, P. E. Rubinin, and D. (David) Shoenberg, editors. Kapitza in Cambridge and Moscow: life and letters of a Russian physicist. North-Holland Publishing Co., Amsterdam, The Netherlands, 1990. ISBN 0-444-98753-3 (hardcover), 0-444-98749-5 (paperback). xvi + 429 pp. LCCN QC16.K25 A4 1990.

Bruckner:1962:DB

[Bru62] Karl Bruckner. The day of the bomb. Van Nostrand, Princeton, NJ, USA, 1962. 189 pp. LCCN PZ7.B8283 Day 1963.

Beichl:2000:MA

[BS00] Isabel Beichl and Francis Sullivan. The Metropolis algorithm. Computing in Science and Engineering, 2(1):65–69, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (elec-

tronic). URL http://dlib.computer.org/cs/books/cs2000/pdf/
c1065.pdf;   http://www.computer.org/cse/cs1999/c1065abs.
htm.

Broad:2003:ETF

[BS03a] William J. Broad and Walter Sullivan. Edward Teller, a fierce architect of the hydrogen bomb, is dead at 95. New York Times, ??(??): ??, September 11, 2003. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL http://www.nytimes.com/2003/09/
11/us/edward-teller-a-fierce-architect-of-the-hydrogen-bomb-
is-dead-at-95.html.

Bush:2003:ETP

[BS03b] G. W. Bush and W. Schulz. Edward Teller — physicist and key developer of the thermonuclear reaction is dead at 95. Chemical and Engineering News, 81(37):9, ???? 2003. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605x (electronic).

Brush:1966:MCS

[BST66] S. G. Brush, H. L. Sahlin, and E. Teller. Monte Carlo study of a one-component plasma. I. Journal of Chemical Physics, 45(6):2102– 2118, ???? 1966. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/doi/10.1063/
1.1727895.

Bartholome:1932:MBE

[BT32] E. Bartholomé and E. Teller. Modellmässige Berechnung von Eigenschwingungen organischer Kettenmoleküle. (German) [Calculation of the characteristic vibrations of organic chain-molecule models]. Zeitschrift f ür Physikalische Chemie, Abteilung B: Chemie der Elementarprozesse, Aufbau der Materie, 19(B5):366–388, ???? 1932. CO-DEN ZPCBAL. ISSN 0372-9664.

Breit:1940:MHH

[BT40] G. Breit and E. Teller. Metastability of hydrogen and helium levels. Astrophysical Journal, 91(2):215–238, March 1940. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Bethe:1941:DTE

[BT41] H. A. Bethe and E. Teller. Deviations from thermal equilibrium in shock waves. Memorandum ATI-18278 (NP-4898, BRL-X-117), Ballistic Research Laboratories, Aberdeen, MD, USA, ???? 1941. 154

pp. URL http://books.google.de/books?id=2qPvAAAAMAAJ. Un-

dated, perhaps 1941 or 1942. Classified until after World War II, and reprinted in [Bet97, pages 296ff]. Discussed briefly in [Ber80, page 64], which says work in 1940. Bethe, in Goodstein’s interview [Goo99, pages 256–258], also says 1940, and describes it as arising from a suggestion given to the authors by aerodynamicist Theodore von Kármán.

Bethe:1945:DTE

[BT45] H. A. Bethe and E. Teller. Deviations from thermal equilibrium in shock waves. Report X-117, Ballistic Research Laboratories, Aberdeen, MD, USA, 1945.

Bethe:1951:DTEa

[BT51a] Hans A. Bethe and Edward Teller. Deviations from thermal equilibrium in shock waves. Report AM-562, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1951. iv + 77 pp.

Bethe:1951:DTEb

[BT51b] Hans A. (Hans Albrecht) Bethe and Edward Teller. Deviations from thermal equilibrium in shock waves. Technical report, Engineering Research Institute, University of Michigan, Ann Arbor, MI,

USA, 1951. iv + 77 pp. URL http://pi.lib.uchicago.edu/1001/

cat/bib/ocm23330822. First reproduced for limited distribution by the Ballistic Research Laboratories, Aberdeen Proving Ground, Aberdeen, Maryland.

Bethe:1953:DTE

[BT53] Hans A. Bethe and Edward Teller. Deviations from thermal equilibrium in shock waves. Report BRL-X-117 (NP-4898), United States Department of Energy, Oak Ridge, TN, USA, 1953. 82 pp. URL

http://www.osti.gov/servlets/purl/4420349/.

Bloch:1955:RS

[BT55] Robert Bloch and Edward Teller. On responsibility of scientists. Science, 122(3170):600–601, September 30, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/pdfplus/1750582.pdf.

Bethe:1961:FNT

[BT61] Hans A. (Hans Albrecht) Bethe and Edward Teller. The future of nuclear tests: Nuclear tests, volume 145 of Headline series. Foreign Policy Association — World Affairs Center, New York, NY, USA, 1961. 61 pp. LCCN E744 .H43 no. 145.

Bernhardt:1977:ISR

[BT77] Anthony F. Bernhardt and Edward Teller. Isotope separation by resonance scattering. Applied Physics Letters, 30(10):550–552, ???? 1977.

CODEN APPLAB. ISSN 0003-6951 (print), 1077-3118 (electronic), 1520-8842.

Bernhardt:1978:AEC

[BT78] A. F. Bernhardt and E. Teller. Adiabatic excitation of colliding molecules and its application to isotope separation. Journal of the Optical Society of America, 68(5):685, ???? 1978. CODEN JOSAAH. ISSN 0030-3941.

Bethe:1982:NFW

[BT82] Hans A. Bethe and E. Teller. The nuclear freeze: Will it really make us safer? Los Angeles Times, October 17, 1982.

Broyles:1973:GH

[BTW73] A. A. Broyles, E. Teller, and E. Wigner. A glass house. Survive, 6 (5):1–??, ???? 1973. CODEN ???? ISSN 0039-6354.

Broyles:1990:TSH

[BTW90] A. A. Broyles, E. Teller, and B. G. Wilson. Theory of superconductivity at high temperatures. Journal of Superconductivity, 3(2):161–169, ???? 1990. CODEN JOUSEH. ISSN 0896-1107 (print), 1572-9605 (electronic).

Budiansky:1983:USP

[Bud83] Stephen Budiansky. US strategy: President Reagan opts for antimissile defence. Nature, 302(5907):365, March 31, 1983. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v302/n5907/pdf/
302365a0.pdf.

Budiansky:1984:SWS

[Bud84] Stephen Budiansky. Star Wars: Strategic weaknesses made plain. Nature, 310(5978):530, August 16, 1984. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v310/n5978/pdf/310530a0.pdf.

Bulkeley:1983:MET

[Bul83] Rip Bulkeley. Myths of Edward Teller. Bulletin of the Atomic Scientists, 39(8):64–65, October 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH83a].

Bunzel:2004:MET

[Bun04] John H. Bunzel. In memoriam: Edward Teller: A personal remembrance. Hoover Digest, 2004(1):??, January 30, 2004. ISSN

1088-5161. URL http://www.hoover.org/publications/hoover-
digest/article/6661.

Burns:1984:APC

[Bur84] Grant Burns. The atomic papers: a citizen’s guide to selected books and articles on the bomb, the arms race, nuclear power, the peace movement, and related issues. Scarecrow Press, Metuchen, NJ, USA, 1984. ISBN 0-8108-1692-X. xiv + 309 pp. LCCN Z6464.D6 B85 1984.

Bussey:2011:JET

[Bus11] Peter J. Bussey. Judging Edward Teller, by Istvan Hargittai, Scope: general interest. Level: general readership. Contemporary Physics, 52(4):382–383, 2011. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bohr:1939:MNF

[BW39] Niels Bohr and John Archibald Wheeler. The mechanism of nuclear fission. Physical Review (2), 56(5):426–450, September 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v56/i5/p426_1.

Blecher:2005:ETL

[BW05] Jens Blecher and Gerald Wiemers. Edward Teller in Leipzig 1928 bis 1930. (German) [Edward Teller in Leipzig 1928–1930]. In Kleint et al. [KRW05], pages 114–121. ISBN 3-7776-1402-5. ISSN 0365-6470.

LCCN QC16.H35 W474 2005. URL http://www.loc.gov/catdir/
toc/fy0613/2006402966.html.

Byrd:2001:DET

[Byr01] Sheri Byrd. Dr. Edward Teller honored with revived Hungarian Corvin Medal. Web site, August 17, 2001.

Carlisle:1991:BRB

[Car91] Rodney Carlisle. Book review: Edward Teller — Giant of the Golden-Age of Physics by S. A. Blumberg and L. G. Panos, (New York: Scribner’s, 1990. xiv + 306 pp. $24.95). The Journal of American History, 78(1):390–391, December 1991. ISSN 0021-8723 (print), 1945-2314

(electronic). URL http://jah.oxfordjournals.org/content/
78/1/390.full.pdf+html;   http://jah.oxfordjournals.org/
content/78/1/390.short.

Carlson:2003:HUS

[Car03] Bengt Carlson. How Ulam set the stage: history has not given enough credit to the main man behind the H-bomb. Bulletin of the Atomic Scientists, 59(4):46–51, July/August 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//bos.sagepub.com/content/59/4/46.full.pdf+html;   http:/
/www.highbeam.com/doc/1G1-105163616.html.

Cass:2001:SWE

[Cas01] Stephen Cass. A scientific warrior [Edward Teller, Memoirs, Perseus, Cambridge, Mass., $35, 592 pp., ISBN: 0-7382-0532-X] [books]. IEEE Spectrum, 38(10):65–66, October 2001. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Conquest:1996:PIC

[CD96] Robert Conquest and Dušan J. Djordjevich, editors. Political and ideological confrontations in Twentieth-Century Europe: essays in honor of Milorad M. Drachkovitch. St. Martin’s Press, New York, NY, USA, 1996. ISBN 0-312-12373-6. xiv + 290 pp. LCCN D415 .P6 1996. URL

http://www.loc.gov/catdir/toc/fy0714/95051502.html.

Chib:1995:UMH

[CG95] Siddhartha Chib and Edward Greenberg. Understanding the Metropolis–Hastings algorithm. The American Statistician, 49 (4):327–335, November 1995. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic). URL http://www.jstor.org/

stable/2684568.

Campbell:1996:HED

[CHL^{+}96] E. M. Campbell, N. C. Holmes, S. B. Libby, B. A. Remington, and E. Teller. High energy-density physics: From nuclear testing to the superlasers. AIP Conference Proceedings, 370(??):21–26, ???? 1996. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Campbell:1997:EHE

[CHL^{+}97] E. M. Campbell, N. C. Holmes, S. B. Libby, B. A. Remington, and E. Teller. The evolution of high energy-density physics: From nuclear testing to the superlasers. Laser and Particle Beams, 15(4):607–626, ???? 1997. CODEN LPBEDA. ISSN 0263-0346 (print), 1469-803X (electronic).

Cantelon:1991:AAD

[CHW91] Philip L. (Philip Louis) Cantelon, Richard G. Hewlett, and Robert Chadwell Williams, editors. The American atom: a documentary history of nuclear policies from the discovery of fission to the present. University of Pennsylvania Press, Philadelphia, second edition, 1991. ISBN 0-8122-3096-5 (hardcover), 0-8122-1354-8 (paper). xviii + 369 pp. LCCN UA23 .A597 1991.

Castell:2004:TQI

[CI04] Lutz Castell and Otfried Ischebeck. Time, quantum and information. Letters to Carl Friedrich von Weizsäcker on the occasion of his 90th birthday. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. xiii + 454 pp.

Chapline:1973:HEN

[CJTW73] G. F. Chapline, Montgomery H. Johnson, Edward Teller, and Morton S. Weiss. Highly excited nuclear matter. Physical Review D (Particles and Fields), 8(12):4302–4308, December 15, 1973. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500.

Caldirola:1971:PHE

[CK71] Piero Caldirola and Heinz Knoepfel, editors. Physics of high energy density, volume 48 of Its Proceedings, course. Academic Press, New York, USA, 1971. ISBN 0-12-368848-5. xiv + 418 pp. LCCN QC281 .V28 1971.

Cole:1994:CAM

[Col94] Leonard A. Cole. Clearing the air in Minneapolis. Bulletin of the Atomic Scientists, 50(9):5–6, November 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

highbeam.com/doc/1G1-15879963.html.

CFWC:1948:PDW

[Com48] Committee to Frame a World Constitution. Preliminary draft of a World Constitution. Bulletin of the Atomic Scientists, 4(5):145–150, May 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See explanation [Bor48] and comment [Tel48a].

Compton:1956:AQP

[Com56] Arthur Holly Compton. Atomic quest, a personal narrative. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1956. xix + 370 pp. LCCN QC773.A1 C65.

Cook:1980:BRB

[Coo80] Earl Cook. Book review: Energy from Heaven and Earth by Edward Teller. Quarterly review of biology, 55(3):274, September 1980. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

URL http://www.jstor.org/stable/pdfplus/2824751.pdf. See

[TW79b].

Coughlan:1954:DET

[Cou54] Robert Coughlan. Dr. Edward Teller’s magnificent obsession. Life, ??(??):61–??, July 13, 1954.

Coughlan:1963:OTT

[Cou63] Robert Coughlan. Oppenheimer and Teller: The tangled drama and private hells of two famous scientists. Life, ??(??):87A–110, Decem-

ber 13, 1963. URL http://www.oldlife.net/search.php?pndate=
1963-12-13.

Cousins:1979:QRM

[Cou79] N. Cousins. Q and A that raises more Qs than As + Edward Teller; newspaper advertisement on Three-Mile-Island Accident and benefits of nuclear power. Saturday Review, 6(19):10, ???? 1979. ISSN 0361-1655.

Cowen:2003:LAM

[Cow03] Robert C. Cowen. A long-ago moment with Edward Teller. The Christian Science Monitor, ??(??):11–??, September 12, 2003. ISSN 0882-7729 (print), 1540-4617 (electronic). URL http://search.

proquest.com/docview/405679692.

Cahn:1993:TBT

[CP93] Anne Hessing Cahn and John Prados. Team-B: the trillion dollar experiment. Bulletin of the Atomic Scientists, 49(3):22–31, April 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-13610752.html.

See corrections [Ano93].

Crease:2002:ETF

[Cre02] R. P. Crease. Edward Teller: friend and foe. Physics World, 15(6):19, June 2002. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Crease:2010:BEP

[Cre10] Robert P. Crease. Biography: Envy and power: Book review: Judging Edward Teller: A Closer Look at One of the Most Influential Scientists of the Twentieth Century. Nature, 468(7324):629–630, December 1, 2010. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/v468/
n7324/full/468629a.html.

Critchfield:1938:SNF

[CT38] C. Critchfield and E. Teller. On the saturation of nuclear forces. Physical Review, 53(10):812–818, May 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v53/i10/p812_1.

Critchfield:1941:ADA

[CT41] Charles L. Critchfield and Edward Teller. On the angular distribution of alpha-particles produced in the Li^{7}–proton reaction. Physical Review (2), 60(1):10–17, July 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.

org/abstract/PR/v60/i1/p10_1.

Canavan:1990:SD

[CT90a] Gregory Canavan and Edward Teller. Strategic defence for the 1990s. Nature, 344(6268):699–704, April 19, 1990. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.

com/nature/journal/v344/n6268/pdf/344699a0.pdf.

Canavan:1990:SEN

[CT90b] Gregory H. Canavan and Edward Teller. Survivability and effectiveness of near-term strategic defense. Report LA-11345-MS, UC-700, Los Alamos National Laboratory, Los Alamos, NM, USA, January

1990. 23 pp. URL http://handle.dtic.mil/100.2/ADA344756.

Canavan:1991:LLS

[CT91] Gregory Canavan and Edward Teller. Low-level satellites expand distributed remote sensing. SIGNAL Magazine, 45(12):99–103, August 1991. CODEN ???? ISSN ????

Cunningham:1988:LTH

[Cun88] Clarence M. Cunningham. Letter: Teller’s heavy hand. Bulletin of the Atomic Scientists, 44(8):52, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu88a].

Dean:1987:FAS

[DA87] Gordon E. Dean and Roger M. Anders. Forging the atomic shield: excerpts from the office diary of Gordon E. Dean. University of North Carolina Press, Chapel Hill, NC, USA, 1987. ISBN 0-8078-1714-7. xxxii + 309 pp. LCCN QC773.3.U5 D43 1987.

Davis:1991:BSB

[Dav91] Burtron H. Davis. B, E, & T: The scientists behind surface science. CHEMTECH, 21(1):18–25, January 1991. CODEN CHTEDD. ISSN 0009-2703. URL ????.

Davis:1994:BET

[Dav94] Burtron H. Davis. Brunauer, Emmett and Teller — the personalities behind the BET method. Energeia, 5(6):1, 4–5, ???? 1994. CODEN

???? ISSN ???? URL http://www.caer.uky.edu/energeia/PDF/

vol5_6.pdf. Part 1: see [Dav95] for part 2.

Davis:1995:BET

[Dav95] Burtron H. Davis. Brunauer, Emmett and Teller — the personalities behind the BET method. Energeia, 6(1):1, 3–4, ???? 1995. CODEN

???? ISSN ???? URL http://www.caer.uky.edu/energeia/PDF/

vol6-1.pdf. Part 2: see [Dav94] for part 1.

Day:1977:NWL

[Day77] Samuel H. Day, Jr. The nuclear weapons labs. Bulletin of the Atomic Scientists, 33(4):21–26, 28–32, April 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Wel77].

Day:1997:BRB

[Day97] Samuel H. Day, Jr. Book review: An Enemy of the State: The Life of Erwin Knoll, by Bill Lueders. Bulletin of the Atomic Scientists, 53(3): 57–59, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/1G1-
19368191.html.

Durr:1971:QFP

[DB71] Hans-Peter Dürr and Erich Bagge, editors. Quanten und Felder: Physikalische und philosophische Betrachtungen zum 70. Geburtstag von Werner Heisenberg. (German) [Quanta and fields: physical and philosophical reflections on the 70th birthday of Werner Heisenberg]. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. ISBN 3-528-08317-4. 366 pp. LCCN ???? URL

http://www.springer.com/physics/book/978-3-528-08317-5.

Atkinson:1929:FAE

[dEAH29] R. d E. Atkinson and F. G. Houtermans. Zur Frage der Aufbaumöglichkeit der Elemente in Sternen. (German) [On the question of the possibility of building elements in stars]. Zeitschrift für Physik, 54(9–10):656–665, September 1929. CODEN ZEPYAA. ISSN

0044-3328. URL http://link.springer.com/article/10.1007/
BF01341595.

Demortier:2015:QEQ

[Dem15] Guy Demortier. Qui est qui dans ^{≪}the letter to Roosevelt ^{≫}?. (French) [Who is who in “The letter to Roosevelt”?]. Revue des Questions Scientifiques, 186(1):125–152, ???? 2015. CODEN RQS-

CAN. ISSN 0035-2160. URL https://www.unamur.be/sciences/
philosoc/revueqs/textes-en-ligne/rqs_186_1_demortier.

Devlin:1962:TMO

[Dev62] John C. Devlin. Teller is mute on Oppenheimer; has TV show delete question. New York Times, ??(??):??, April 16, 1962. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

DeWitt:1979:UGD

[DeW79] Hugh E. DeWitt. Has US government disclosed the secret of the Hbomb? Bulletin of the Atomic Scientists, 35(6):60–62, June 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

DeWitt:1988:LXR

[DeW88] Hugh E. DeWitt. Letter: X-ray laser hype confirmed. Bulletin of the Atomic Scientists, 44(8):52, October 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Blu88a].

Drell:1984:RSD

[DFH84] Sidney D. (Sidney David) Drell, Philip J. Farley, and David Holloway. The Reagan Strategic Defense Initiative: a technical, political, and arms control assessment. International Strategic Institute at Stanford, Stanford, CA, USA, July 1984. ISBN ???? xi + 147 pp. LCCN ???? A special report of the Center for International Security and Arms Control, Stanford University.

Drell:1985:RSD

[DFH85] Sidney D. (Sidney David) Drell, Philip J. Farley, and David Holloway. The Reagan Strategic Defense Initiative: a technical, political, and arms control assessment. Ballinger, Cambridge, MA, USA, 1985.

ISBN 0-88730-064-2 (paperback). xiii + 152 pp. LCCN UG743 .D74 1985.

Daintith:1999:DS

[DG99] John Daintith and Derek Gjertsen, editors. A Dictionary of Scientists. Oxford paperback reference. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-585-11047-6 (e-book), 0-19-280086-8 (paperback). 586 pp. LCCN Q141 .D52 1994. URL http:

//www.loc.gov/catdir/enhancements/fy0637/99488304-d.html;
 http://www.loc.gov/catdir/enhancements/fy0637/99488304-
t.html;   http://www.netLibrary.com/urlapi.asp?action=summary&
v=1&bookid=12306.

Davis:1983:HCS

[DH83] Burtron H. Davis and William P. Hettinger, editors. Heterogeneous catalysis: selected American histories, volume 222 of ACS symposium series, 0097-6156. American Chemical Society, Washington, DC, USA, 1983. ISBN 0-8412-0778-X. LCCN QD505 .H47 1983.

URL http://pubs.acs.org/isbn/9780841207783.

deHoffmann:1950:MHS

[dHT50] Frederic de Hoffmann and E. Teller. Magneto-hydrodynamic shocks. Physical Review (2), 80(4):692–703, November 15, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v80/i4/p692_1.

Dickson:1979:ENC

[Dic79] David Dickson. End nuclear complacency — 3 Mile Island report. Nature, 282(5735):120–121, November 8, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v282/n5735/pdf/282120a0.pdf.

Dow:1974:BRB

[Dow74] John D. Dow. Book review: The Jahn–Teller Effect in Molecules and Crystals, by R. Englman. Physics Today, 27(3):57, March 1974. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.org/resource/1/phtoad/v27/

i3/p57_s1. See [Eng72].

Doyle:2002:BBT

[Doy02] J. Doyle. Brotherhood of the bomb: The tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. Library Journal, 127(12):96, ???? 2002. CODEN LIBJA7. ISSN 0363-0277.

Drewe:2002:ETH

[Dre02] J. Drewe. Edward Teller and the H-bomb (comment). TLS — The Times Literary Supplement, ??(5182):17, July 26, 2002. ISSN 0307-661X.

Drury:1958:TSP

[Dru58] Allen Drury. Teller sees peril if atom tests end. New York Times, ?? (??):??, April 17, 1958. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Duerr:1956:IAN

[DT56] Hans-Peter Duerr and Edward Teller. Interaction of antiprotons with nuclear fields. Physical Review, 101(1):494–495, January 1, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v101/i1/p494_1.

Donnan:1937:FCS

[DTT37] F. G. Donnan, E. Teller, and B. Topley. Fluctuations of composition in a system of molecules in chemical equilibrium. Philosophical Magazine, 24(??):981–1001, ???? 1937. CODEN PHMAA4. ISSN 0031-8086.

Dudley:1975:UCT

[Dud75] H. C. Dudley. The ultimate catastrophe. Bulletin of the Atomic Scientists, 31(9):21–24, November 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Durr:2003:MET

[Dür03] Hans-Peter Dürr. In memoriam Edward Teller. Physik Journal, 2 (12):60–61, ???? 2003. CODEN PJHOB2. ISSN 1617-9439 (print),

1619-6597 (electronic). URL http://www.pro-physik.de/details/
articlePdf/1107593/issue.html.

Durr:2005:ETW

[Dür05] Hans-Peter Dürr. Von Edward Teller zu Werner Heisenberg — Erinnerungen an meine Zusammenarbeit mit Heisenberg. (German) [From Edward Teller to Werner Heisenberg — memories of my collaboration with Heisenberg]. In Kleint et al. [KRW05], pages 321–331. ISBN 3-7776-1402-5. ISSN 0365-6470. LCCN QC16.H35 W474 2005. URL

http://www.loc.gov/catdir/toc/fy0613/2006402966.html.

Duxbury:2010:JTE

[Dux10] Professor Geoffrey Duxbury. The Jahn–Teller Effect, by I. B. Bersuker, Scope: monograph. Level: advanced/research. Contemporary Physics, 51(5):455, 2010. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

deVore:2008:ETC

[dV08a] Lauren de Vore. Edward Teller Centennial Symposium coming to Bankhead Theater. Web site, May 9, 2008. URL https://www.

llnl.gov/news/edward-teller-centennial-symposium-coming-
bankhead-theater.

deVore:2008:SRE

[dV08b] Lauren de Vore. Symposium reviews Edward Teller’s varied contributions to science. Web site, May 30, 2008. URL https:/

/www.llnl.gov/news/symposium-reviews-edward-teller%E2%80%
99s-varied-contributions-science.

Dyson:1958:BRBb

[Dys58] Freeman J. Dyson. Book review: Our Nuclear Future: Facts, Dangers and Opportunities. By Edward Teller and Albert L. Latter. 184 pp. Criterion Books, New York, 1958. $3.50. Physics Today, 11(8):38–40, August 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/11/38/3;
http://physicstoday.org/resource/1/phtoad/v11/i8/p38_s3.

Dyson:1976:HBD

[Dys76] F. J. Dyson. The hydrogen-bomb decision: a reappraisal: Review of The Advisors: Oppenheimer, Teller and The Superbomb, by Herbert F. York. Science, 193(4254):668–669, August 20, 1976. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/193/4254/668.full.pdf.

Dyson:1979:DU

[Dys79a] Freeman J. Dyson. Disturbing the Universe. Harper & Row, New York, NY, USA, 1979. ISBN 0-06-011108-9. x + 283 pp. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1979:PFM

[Dys79b] Freeman J. Dyson. Prelude in E-flat minor. In Disturbing the Universe [Dys79a], pages 84–93. ISBN 0-06-011108-9. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:2002:BRB

[Dys02a] Freeman J. Dyson. Book review: Memoirs: A Twentieth-Century Journey in Science and Politics. Edward Teller with Judith Shoolery. 602 pp. Perseus, Cambridge, MA, 2001. Price: $35.00 ISBN 0-7382-0532-X. American Journal of Physics, 70(4):462–463, April 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v70/i4/p462_s1.

Dyson:2002:POT

[Dys02b] George Dyson. Project Orion: the true story of the atomic spaceship. Henry Holt and Company, Incorporated, New York, NY, USA, 2002. ISBN 0-8050-5985-7. xv + 345 pp. LCCN TL783.5 .D95 2002. URL

http://www.loc.gov/catdir/bios/hol051/2001046500.html;
http://www.loc.gov/catdir/description/hol021/2001046500.
html.

Dyson:2006:ETB

[Dys06a] Freeman J. Dyson. Edward Teller’s Memoirs. In The scientist as rebel [Dys06b], pages 167–172. ISBN 1-59017-216-7.

LCCN Q158.5 .D977 2006. URL http://www.loc.gov/catdir/
enhancements/fy0665/2006022081-b.html;   http://www.loc.
gov/catdir/enhancements/fy0665/2006022081-d.html;   http:
//www.loc.gov/catdir/toc/ecip0616/2006022081.html. Collec-

tion of book reviews.

Dyson:2006:SRa

[Dys06b] Freeman J. Dyson. The scientist as rebel. New York Review Books, New York, NY, USA, 2006. ISBN 1-59017-216-7. xvi + 360 pp.

LCCN Q158.5 .D977 2006. URL http://www.loc.gov/catdir/
enhancements/fy0665/2006022081-b.html;   http://www.loc.
gov/catdir/enhancements/fy0665/2006022081-d.html;   http:
//www.loc.gov/catdir/toc/ecip0616/2006022081.html. Collec-

tion of book reviews.

Dyson:2007:ET

[Dys07] Freeman J. Dyson. Edward Teller (1908–2003). National Academy of Sciences, ??(??):1–21, ???? 2007. URL http://www.nap.edu/html/

biomems/eteller.pdf.

Dyson:2009:ETJ

[Dys09] Freeman J. Dyson. Edward Teller: January 15, 1908–September 9, 2003. Biographical memoirs — National Academy of Sciences of the United States of America, pages 413–430, 2009. CODEN BMNSAC. ISBN 0-309-12148-5. ISSN 0077-2933.

Dyson:2012:TCO

[Dys12a] George Dyson. Turing’s cathedral: the origins of the digital universe. Pantheon Books, New York, NY, USA, 2012. ISBN 0-375-42277-3 (hardcover). xxii + 401 pp. LCCN QA76.17 .D97 2012.

Dyson:2012:UD

[Dys12b] George Dyson. Ulam’s demons. In Turing’s cathedral: the origins of the digital universe [Dys12a], chapter 11, pages 200–224. ISBN 0-375-42277-3 (hardcover). LCCN QA76.17 .D97 2012. Pages 191–192 describe the origin of the Monte Carlo method.

Edgerton:1991:BRA

[Edg91] David Edgerton. Book reviews: Atomic Energy for Military Purposes, with a new foreword by Philip Morrison and an essay by Henry De-Wolf Smyth. Stanford: Stanford University Press, 1990. Pp. xvi + 324. ISBN 0-8047-1721-4, $39.50 (hardcover); 0-8047-1722-2, $12.95 (paperback). Herbert F. York. The Advisors: Oppenheimer, Teller and the Superbomb, with a historical essay by Hans A. Bethe. Stanford: Stanford University Press, 1989. Pp. xiv + 201. ISBN 0-8047-1713-3, $32.50 (hardcover); 0-8047-1714-1, $8.95 (paperback). British Journal for the History of Science, 24(4):476–477, December 1991. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic).

URL http://www.jstor.org/stable/4027137.

Eliezer:1986:IES

[EGH86] Shalom Eliezer, A. K. (Ajoy K.) Ghatak, and Heinrich Hora. An introduction to equations of state: theory and applications. Cambridge University Press, Cambridge, UK, 1986. ISBN 0-521-30389-3. xii + 366 pp. LCCN QC173.4.E65 E45 1986. URL http://www.loc.gov/

catdir/description/cam031/85029094.html.

Eyring:1970:PCA

[EHJ70] Henry Eyring, Douglas Henderson, and Wilhelm Jost, editors. Physical Chemistry, An Advanced Treatise: Valency, volume V. Academic Press, New York, USA, 1970. ISBN 0-12-245605-X. xix + 732 pp. LCCN ????

Einstein:1939:AEL

[Ein39] Albert Einstein. Albert Einstein’s letter to President Franklin Delano Roosevelt. World-Wide Web document, August 2, 1939. URL

http://hypertextbook.com/eworld/einstein.shtml;   http:/
/www.anl.gov/Science_and_Technology/History/Anniversary_
Frontiers/aetofdr.html;   http://www.lanl.gov/history/road/
pdf/Einstein.pdf.

Evans:1986:SWW

[EN86] Rowland Evans and Robert Novak. SDI: What will Reagan decide now? Washington Post, ??(??):??, October 13, 1986. ISSN 0190-8286.

Englman:1972:JTE

[Eng72] R. Englman. The Jahn–Teller effect in molecules and crystals. Wiley monographs in chemical physics. Wiley-Interscience, New York, NY, USA, 1972. ISBN 0-471-24168-7. xx + 350 pp. LCCN QD461 .E53.

Epstein:1992:BRB

[Eps92] Lewis Carrol Epstein. Book review: Conversations on the Dark Secrets of Physics, by Edward Teller. American Journal of Physics, 60(6):572, June 1992. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v60/i6/p572_s1.

Ermenc:1989:ABS

[Erm89] Joseph J. Ermenc, editor. Atomic bomb scientists: memoirs, 1939– 1945: interviews with Werner Karl Heisenberg, Paul Harteck, Lew Kowarski, Leslie R. Groves, Aristid von Grosse, C. E. Larson. Meckler Corp., 11 Ferry Lane West, Westport, CT 06880, USA, 1989. ISBN 0-88736-267-2. 385 pp. LCCN QC774.A2 A86 1989. US$120.00.

Ehrlich:1984:CDW

[ESKR84] Paul R. Ehrlich, Carl Sagan, Donald Kennedy, and Walter Orr Roberts, editors. The cold and the dark: the world after nuclear war: the Conference on the Long-Term Worldwide Biological Consequences of Nuclear War. W. W. Norton & Co., New York, NY, USA, 1984. ISBN 0-393-01870-9. LCCN QH545.N83 C66 1983. Foreword by Lewis Thomas.

Emmett:1940:RIS

[ET40] P. H. Emmett and E. Teller. The role of ions in surface catalysis. In National Research Council (US). Committee on Catalysis [Nat40], pages 68–81. LCCN QD505.N38x 1940.

Einstein:2020:WSS

[ET20] Albert Einstein and Edward Teller. 1950: What the scientists are saying. Bulletin of the Atomic Scientists, 76(6):350–352, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.1080/
00963402.2020.1847477.

Feld:1979:EPD

[Fel79a] Bernard T. Feld. Editorial: A progressive decision. Bulletin of the Atomic Scientists, 35(9):4–5, November 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79b].

Feld:1979:EBS

[Fel79b] Bernard T. Feld. Editorial: The Progressive secret. Bulletin of the Atomic Scientists, 35(5):5–6, May 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [San79, Han79, MSDP80] and outcome [Fel79a].

Fermi:1957:LEN

[Fer57] Laura Fermi. Letter to the Editor: The nature of nuclear warfare. Bulletin of the Atomic Scientists, 13(6):232, June 1957. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel57d].

Fermi:1968:III

[Fer68] Laura Fermi. Illustrious immigrants; the intellectual migration from Europe, 1930/41. University of Chicago Press, Chicago, IL, USA, 1968. xi + 440 pp. LCCN E184.A1 F47.

Findlay:1990:NDP

[Fin90] Trevor Findlay. Nuclear dynamite: the peaceful nuclear explosions fiasco. Brassey’s Australia/Pergamon Press, Sydney, NSW, Australia, 1990. ISBN 0-08-034436-4. xv + 339 pp. LCCN TA748.

Firor:2005:FSR

[Fir05] John Firor. Former student remembers Teller and Fermi with gratitude. Physics Today, 58(2):10, February 2005. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//physicstoday.org/resource/1/phtoad/v58/i2/p10_s1.

Fisch:2010:PPC

[Fis10] Nathaniel J. Fisch. Plasma physics and controlled nuclear fusion. In Libby and Van Bibber [LV10], pages 112–127. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Fitzpatrick:1998:TTN

[Fit98] Anne Fitzpatrick. Teller’s technical nemeses: The American hydrogen bomb and its development within a technological infrastructure. Society for Philosophy and Technology, 3(3):??, Spring 1998. URL http:

//scholar.lib.vt.edu/ejournals/SPT/v3n3/FITZPATR.html.

FitzGerald:2000:WTB

[Fit00] Frances FitzGerald. Way out there in the blue: Reagan, Star Wars, and the end of the Cold War. Simon and Schuster, New York, NY, USA, 2000. ISBN 0-684-84416-8 (hardcover), 0-7432-0023-3 (paperback). 592 pp. LCCN E876 .F58 2001.

French:1985:NBC

[FK85] A. P. (Anthony Philip) French and P. J. Kennedy, editors. Niels Bohr: a centenary volume. Harvard University Press, Cambridge, MA, USA, 1985. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). xiv + 403 pp. LCCN QC16.B63 N49 1985. US$27.50.

Flournoy:1988:WRI

[Flo88] Michele Flournoy. Washington report: INF: the first 60 days; air force may have twins; South Africa still in IAEA; Teller’s telltale letters; what the GAO didn’t let you read. Bulletin of the Atomic Scientists, 44(9):3–5, November 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Finkelnburg:1934:MKG

[FMRT34] Wolfgang Finkelnburg, Reinhard Mecke, Otto Reinkober, and Edward Teller, editors. Molekül- und Kristallgitterspektren. (German) [Molecular and crystal lattice spectra], volume 9 of Handund Jahrbuch der chemischen Physik. Die Spektren Entstehung und Zusammenhang mit der Struktur der Materie, Abschnitt 2. Akademische Verlagsgesellschaft, Leipzig, Germany, 1934. x + 408 pp. LCCN QD453 .E72 bd. 9, abschnitt 2.

Feynman:1947:ESE

[FMT47] Richard P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Report OSTI 4417654, Los Alamos Scientific Laboratory, Atomic Energy Commission, Los Alamos, NM, USA, 1947.

Feynman:1949:ESEa

[FMT49a] R. P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Technical Report AECD-2448, Technical Information Branch, Oak Ridge Operations, AEC, Oak Ridge, TN, USA, January 20, 1949. 41

pp. URL http://www.osti.gov/accomplishments/documents/
fullText/ACC0107.pdf;   http://www.osti.gov/servlets/purl/
4417654-BCgOtj/native/.

Feynman:1949:ESEb

[FMT49b] R. P. Feynman, N. Metropolis, and E. Teller. Equations of state of elements based on the generalized Fermi–Thomas theory. Physical Review (2), 75(10):1561–1573, May 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

Ford:2015:BHB

[For15] Kenneth William Ford. Building the H bomb: a personal history. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2015. ISBN 981-4632-07-4 (hardcover), 981-4618-79-9 (paperback). 240 (est.) pp. LCCN UG1282.A8 F67

2015. URL http://www.worldscientific.com/worldscibooks/
10.1142/9269.

Franklin:1981:BRB

[Fra81] Allan Franklin. Book review: The Pursuit of Simplicity, by Edward Teller. American Journal of Physics, 49(9):893, September 1981. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v49/i9/p893_s1.

Frank:2005:ERG

[Fra05] Tibor Frank. Ever ready to go: The multiple exiles of Leo Szilard. Physics in Perspective (PIP), 7(2):204–252, June 2005. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL http://adsabs.harvard.edu/abs/2005PhP.....7..204F;
http://www.springerlink.com/content/1422-6960/.

Frank:2008:MHV

[Fra08] Tibor Frank. Memory and history: Visiting Edward Teller. Hungarian Quarterly, 49(189):123–128, ???? 2008. ISSN 0028-5390.

Frisch:1962:BRB

[Fri62] David Frisch. Book review: The Legacy of Hiroshima, by Edward Teller and Allen Brown. Physics Today, 15(7):50–51, July 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://physicstoday.org/resource/1/phtoad/v15/

i7/p50_s1. See [TB62, TB75].

Fermi:1995:PBP

[FS95] Rachel Fermi and Esther Samra. Picturing the Bomb: Photographs from the Secret World of the Manhattan Project. H. N. Abrams, New York, NY, USA, 1995. ISBN 0-8109-3735-2. 232 pp. LCCN QC773.A1 F47 1995. Introduction by Richard Rhodes.

Franck:1932:BPD

[FST32] J. Franck, H. Sponer, and E. Teller. Bemerkungen über Prädissoziationsspektren dreiatomiger Moleküle. (German) [Remarks on the predissociation spectra of triatomic molecules]. Zeitschrift für physikalische Chemie (Leipzig), 18(B1):88–101, ???? 1932. CODEN ZPCLAH. ISSN 0323-4479.

Franck:1938:MPA

[FT38] James Franck and Edward Teller. Migration and photochemical action of excitation energy in crystals. Journal of Chemical Physics, 6 (12):861–872, ???? 1938. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://link.aip.org/link/doi/10.
1063/1.1750182.

Fermi:1947:CNM

[FT47] E. Fermi and E. Teller. The capture of negative mesotrons in matter. Physical Review (2), 72(5):399–408, September 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v72/i5/p399_1.

Fermi:1947:DNM

[FTW47] Enrico Fermi, Edward Teller, and Victor F. Weisskopf. The decay of negative mesotrons in matter. Physical Review, 71(5):314–315, March 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://inspirehep.net/record/45186;
 http://prola.aps.org/abstract/PR/v71/i5/p314_1. Reprinted

in [FTW65].

Fermi:1965:DNM

[FTW65] E. Fermi, E. Teller, and V. Weisskopf. The decay of negative mesotrons in matter. In Nuclear Forces [Bri65], pages 227–229. ISBN 0-08-011034-7. LCCN QC173 .B8513 1965. Reprint of [FTW47].

Goldman:1986:EWWa

[GA^{+}86a] Peter Louis Goldman, Pamela Abramson, et al. The end of the world that was: six lives in the atomic age. A Newsweek book. Dutton, New York, NY, USA, 1986. ISBN 0-525-24428-X, 0-452-25806-5 (paperback). xiv + 129 + 8 pp. LCCN CT220 .G6 1986.

Goldman:1986:EWWb

[GA^{+}86b] Peter Louis Goldman, Pamela Abramson, et al. The end of the world that was: six lives in the atomic age. A Newsweek book. New American Library, New York, NY, USA, 1986. ISBN 0-452-25806-5 (paperback). xiv + 129 + 8 pp. LCCN QC773 .G58 1986.

Garwin:1983:RR

[Gar83] Richard L. Garwin. Reagan’s riskiness. New York Times, ??(??): ??, March 30, 1983. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Galison:1989:LSD

[GB89] Peter Galison and Barton Bernstein. In any light: Scientists and the decision to build the superbomb, 1952–1954. Historical Studies in the Physical and Biological Sciences, 19(2):267–347, ???? 1989. CODEN HSPSEW. ISSN 0890-9997 (print), 1533-8355 (electronic).

URL http://www.jstor.org/stable/27757627.

Gamato:1980:HMD

[GBS^{+}80] George Gamato, Alan Berman, Edward Salkovitz, Edward Teller, and George Wald. How much does the Defense Department advance science? In Proceedings of an AAAS Symposium on January 8, 1980, page 37. ????, ????, September 24, 1980. ISBN ???? LCCN ???? URL

http://en.scientificcommons.org/18799903.

Green:1977:ASD

[GCB^{+}77] L. Green, R. L. Carle, A. Badruzzaman, ?. Dejou, ?. Saphier, C. Newstead, D. B. Trauger, K. H. Beckurts, ?. Oosterkamp, ?. Rastoin, ?. Block, J. S. Foster, E. Critoph, and E. Teller. Advanced systems — discussion. Transactions of the American Nuclear Society, 25(??): 201–203, ???? 1977. CODEN TANSAO. ISSN 0003-018X.

Gould:1965:PDM

[GEM^{+}65] L. M. Gould, L. Eiseley, P. B. Medawar, L. Norstad, E. Teller, A. Larson, C. A. Doxiadis, and O. M. Wilson. Panel discussion — mirror to man: Man’s adaptation to his expanding environment. Mayo Clinic Proceedings, 40(??):91–111, ???? 1965. CODEN MACPAJ. ISSN 0025-6196 (print), 1942-5546 (electronic).

Gerrish:1975:LTM

[Ger75] Laina Gerrish. Letters: Too much. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Galbiata:1992:SEE

[GER^{+}92] Domenico Galbiata, Padre Eligio, R. A. Ricci, et al., editors. Scienza ed etica alle soglie del terzo millennio: Varenna, Villa Monastero, 28–30 settembre 1992, volume 36. Società italiana di fisica, Bologna, Italy, 1992. ISBN 88-7794-051-4. LCCN ????

Gamow:1976:GGB

[GG76] George Gamow and Barbara Gamow. George Gamow and Barbara Gamow papers, 1915–1975, 1976. URL http://hdl.loc.gov/loc.

mss/eadmss.ms010191;   http://hdl.loc.gov/loc.mss/eadmss.
ms010191.3.

Girdner:1975:LNA

[Gir75] William Girdner. Letters: No alternative? Bulletin of the Atomic Scientists, 31(6):4–5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Galpern:1987:LDR

[GM87] Daniel M. Galpern and Gerald E. Marsh. Letters: Don’t rely on “Navy’s MX”. Bulletin of the Atomic Scientists, 43(9):53–54, November 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar87, Mar97].

Gehrels:1994:HDC

[GMS94] Tom Gehrels, Mildred Shapley Matthews, and A. M. Schumann, editors. Hazards due to comets and asteroids. Space science series. University of Arizona Press, Tucson, AZ, USA and London, UK, 1994. ISBN 0-8165-1505-0. xiii + 1300 pp. LCCN QB721 .H29 1994. With the editorial assistance of M.S. Matthews and A.M. Schumann, and 120 collaborating authors.

Goble:1949:BRB

[Gob49] A. T. Goble. Book review: The Structure of Matter, by Francis Owen Rice and Edward Teller. American Journal of Physics, 17(8): 523, November 1949. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/1/
ajpias/v17/i8/p523_s2. See [RT49a].

GoeppertMayer:1988:MGM

[Goe88] Maria Goeppert Mayer. Maria Goeppert Mayer papers, 1925–1973, MSS 20. Collection at the University of California, San Diego.,

1988. URL http://libraries.ucsd.edu/speccoll/findingaids/
mss0020.html.

Goldman:1984:CTS

[Gol84] Louis Goldman. Commentary: Is there a Soviet bomb in Wichita? Bulletin of the Atomic Scientists, 40(1):54–56, January 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Goldberger:1999:AEF

[Gol99] Marvin L. Goldberger. In appreciation: Enrico Fermi (1901– 1954): The complete physicist. Physics in Perspective (PIP), 1 (3):328–336, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.springer.com/

article/10.1007/s000160050024;   http://www.springerlink.
com/content/y1h02gr3rv7wurrl/.

Goldberg:2004:HS

[Gol04] Jonah Goldberg. Hitler’s scientists. National Review, ??(??): ??, March 11, 2004. ISSN 0028-0038. URL http://www.

nationalreview.com/corner/76704/hitlers-scientists/jonah-
goldberg.

Goncharov:1996:TM

[Gon96a] German A. Goncharov. Thermonuclear milestones. Physics Today, 49 (11):44, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMA

[Gon96b] German A. Goncharov. Thermonuclear milestones: (1) the American effort. Physics Today, 49(11):45–48, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMB

[Gon96c] German A. Goncharov. Thermonuclear milestones: (2) beginnings of the Soviet H-bomb program. Physics Today, 49(11):50–54, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goncharov:1996:TMR

[Gon96d] German A. Goncharov. Thermonuclear milestones: (3) the race accelerates. Physics Today, 49(11):56–61, November 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Goodchild:1980:JRO

[Goo80] Peter Goodchild. J. Robert Oppenheimer: shatterer of worlds. British Broadcasting Corporation, London, UK, 1980. ISBN 0-563-17781-0. 301 pp. LCCN ???? Published in conjunction with the BBC/WGBH television series Oppenheimer, produced by Peter Goodchild and written by Peter Prince. See important corrections [Bet82, page 53].

Goodchild:1981:JRO

[Goo81] Peter Goodchild. J. Robert Oppenheimer: shatterer of worlds. Houghton-Mifflin, Boston, MA, USA, 1981. ISBN 0-395-30530-6. 301 pp. LCCN ???? Published in conjunction with the BBC/WGBH television series Oppenheimer, produced by Peter Goodchild and written by Peter Prince. See important corrections [Bet82, page 53].

Goodstein:1999:CHB

[Goo99] Judith Goodstein. A conversation with Hans Bethe. Physics in Perspective (PIP), 1(3):253–281, October 1999. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://link.

springer.com/article/10.1007/s000160050022;   http://www.
springerlink.com/content/k4f0pllveakla47q/.

Goodchild:2004:ETRa

[Goo04a] Peter Goodchild. Edward Teller: the real Dr Strangelove. Weidenfeld and Nicolson, London, UK, 2004. ISBN 0-297-60734-0. 467 pp. LCCN QC16.T37 G66 2004b.

Goodchild:2004:ETRb

[Goo04b] Peter Goodchild. Edward Teller, the real Dr. Strangelove. Harvard University Press, Cambridge, MA, USA, 2004. ISBN 0-674-01669-6. xxv + 469 + 16 pp. LCCN QC16.T37 G66 2004. URL http://

alsos.wlu.edu/information.aspx?id=2237.

Gordon:1992:SWF

[Gor92] Michael R. Gordon. Star Wars’ final X-ray laser weapon dies as its final test is canceled. New York Times, ??(??):??, July 21, 1992. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Gordin:2009:RCD

[Gor09] Michael D. Gordin. Red cloud at dawn: Truman, Stalin, and the end of the atomic monopoly. Farrar, Straus and Giroux, New York, NY, USA, 2009. ISBN 0-374-25682-9. xii + 402 pp. LCCN U264 .G67 2009.

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Greuling:1945:TWT

[Gre45] E. Greuling. Theory of water-tamped water boiler. Report LA-399, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, September 27, 1945. Work done by E. Teller and others. Declassified 1958.

Greuling:1946:CDW

[Gre46] E. Greuling. Critical dimensions of water-tamped slabs and spheres of active material. Report LA-609, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 6, 1946. Work done by E. Teller and others. Declassified 1956.

Green:1977:OCS

[Gre77] Harold Green. The Oppenheimer case: a study in the abuse of law. Bulletin of the Atomic Scientists, 33(7):12–16, 56–61, September 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Groves:1962:NIC

[Gro62] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Harper & Row, New York, NY, USA, 1962. xiv + 464 pp. LCCN QC773.A1 G7.

Groves:1983:NIC

[Gro83] Leslie R. Groves. Now it can be told: the story of the Manhattan Project. Da Capo Press, New York, NY, USA and Cambridge, MA, USA, 1983. ISBN 0-306-80189-2, 0-7867-4822-2 (e-book). xiv + 464 + 8 pp. LCCN QC773.A1 G7.

Grove:2004:MWW

[Gro04] J. W. Grove. The man who was afraid of ignorance (Edward Teller). Queens Quarterly, 111(4):563–572, ???? 2004. ISSN 0033-6041.

Greiner:1989:NES

[GS89] Walter Greiner and Horst Stöcker, editors. The nuclear equation of state, volume 216A, 216B of NATO ASI series. Series B, Physics. Plenum Press, New York, NY, USA; London, UK, 1989. ISBN 0-306-43486-5 (part A), 0-306-43487-3 (part B). LCCN QC793.3.N8 N38 1989.

Gamow:1936:SRD

[GT36] G. Gamow and E. Teller. Selection rules for the β-disintegration. Physical Review (2), 49(12):895–899, June 15, 1936. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v49/i12/p895_1.

Gamow:1937:LES

[GT37] G. Gamow and E. Teller. Letter to editor: Some generalizations of the β-transformation theory. Physical Review, 51(4):289, February 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://prola.aps.org/abstract/PR/v51/i4/p289_
1.

Gamow:1938:LER

[GT38a] George Gamow and Edward Teller. Letter to editor: The rate of selective thermonuclear reactions. Physical Review, 53(7):608–609, April 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/v53/
i7/p608_1.

Gamow:1938:NCS

[GT38b] George Gamow and Edward Teller. On the neutron core of stars. Physical Review, 53(11):929–930, June 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/vtoc/PR/v53/i11/p913_1.

Gamow:1939:EUOa

[GT39a] G. Gamow and E. Teller. The expanding universe and the origin of the Great Nebulæ. Nature, 143(3612):116–117, January 21, 1939. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v143/n3612/pdf/
143116a0.pdf.

Gamow:1939:EUOb

[GT39b] G. Gamow and E. Teller. The expanding universe and the origin of the Great Nebulæ. Nature, 143(3618):375, March 4, 1939. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v143/n3618/pdf/
143375a0.pdf.

Gamow:1939:LEE

[GT39c] G. Gamow and E. Teller. Letter to editor: Energy production in Red Giants. Physical Review, 55(8):791, April 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v55/i8/p791_1.

Gamow:1939:OGN

[GT39d] G. Gamow and E. Teller. On the origin of Great Nebulae. Physical Review, 55(7):654–657, April 1, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v55/i7/p654_1.

Goldhaber:1948:NDV

[GT48] M. Goldhaber and E. Teller. On nuclear dipole vibrations. Physical Review, 74(9):1046–1049, November 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v74/i9/p1046_1.

Griggs:1956:DUT

[GT56] D. Griggs and E. Teller. Deep underground test shots. Report UCRL-4659, University of California, Lawrence Radiation Laboratory, Livermore, CA, USA, ???? 1956.

Gotsman:1985:GPH

[GT85] Errol Gotsman and Gerald E. Tauber, editors. From SU(3) to gravity: papers in honor of Yuval Ne’eman. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30784-8. xiv + 457 pp. LCCN QC793.28 .F76 1985.

Gwynne:1987:ETS

[GT87] P. Gwynne and E. Teller. Edward Teller on SDI, competitiveness. The Scientist (Philadelphia, PA), 1(21):14–15, September 1987. ISSN 0890-3670 (print), 1945-5127 (electronic). URL http://www.the-

scientist.com/?articles.view/articleNo/8950/title/Teller-
on-SDI--Competitiveness/.

Goodchild:2006:RDS

[GTS^{+}06] Peter Goodchild, Simon Templeman, Joe Spano, John De Lancie, and Matt August. The real Dr. Strangelove. Audio theatre collection. L. A. Theatre Works, Venice, CA, USA, 2006. ISBN 1-58081-357-7. LCCN RZC 1211. 2 sound discs (122 minutes).

Guinnessy:2004:UCO

[Gui04] Paul Guinnessy. US celebrates Oppenheimer’s centenary. Physics Today, 57(6):35–36, June 2004. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/link/?

PTO/57/35/1;   http://physicstoday.org/resource/1/phtoad/
v57/i6/p35_s1.

Gulbransen:1975:LSE

[Gul75] Earl A. Gulbransen. Letters: Not safe enough. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Haberer:1969:PCS

[Hab69] Joseph Haberer. Politics and the Community of Science. Van Nostrand Reinhold, New York, NY, USA, 1969. vi + 337 pp. LCCN Q125 .H23 1969.

Hacker:2004:BRB

[Hac04] Barton C. Hacker. Book review: Memoirs: A Twentieth-Century Journey in Science and Politics by Edward Teller and Judith L. Shoolery. Technology and Culture, 45(2):460–461, April 2004. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL http://muse.jhu.edu/journals/technology_
and_culture/v045/45.2hacker.pdf;   http://www.jstor.org/
stable/pdfplus/40060780.pdf.

Halloran:1970:TSU

[Hal70] Richard Halloran. Teller says unrest on campus perils defense research. New York Times, ??(??):??, July 25, 1970. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Ham:1971:JTE

[Ham71] Frank S. Ham. The Jahn–Teller effect. International Journal of Quantum Chemistry, 5(S5):191–199, January 18–23, 1971. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic). Supplement: Proceedings of the International Symposium on Atomic, Molecular and Solid-State Theory and Quantum Biology.

Hansen:1979:CBC

[Han79] Chuck Hansen. Commentary: That Progressive case again. Bulletin of the Atomic Scientists, 35(8):59–60, October 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79b].

Hansen:1992:BRD

[Han92a] Chuck Hansen. Book review: The dark side: Teller’s War: The Top-Secret Story Behind the Star Wars Deception, by William J. Broad. Bulletin of the Atomic Scientists, 48(4):45–46, May 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hansen:1992:LUS

[Han92b] Chuck Hansen. Letter: Ulam’s shock waves. Bulletin of the Atomic Scientists, 48(1):45, January/February 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BHM91].

Hansen:1995:BRB

[Han95] Chuck Hansen. Book review: The bomb, part 2: Dark Sun: The Making of the Hydrogen Bomb, by Richard Rhodes. Bulletin of the Atomic Scientists, 51(5):52–53, September/October 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.highbeam.com/doc/1G1-17290233.html.

Hansen:1997:LMF

[Han97] Chuck Hansen. Letter: More fallout on fallout. Bulletin of the Atomic Scientists, 53(5):3, 58, September/October 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM97].

Harper:2001:AGG

[Har01] Eamon Harper. In appreciation: George Gamow: Scientific amateur and polymath. Physics in Perspective (PIP), 3(3):335–372, September 2001. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960

(electronic). URL http://adsabs.harvard.edu/abs/2001PhP.
....3..335H;   http://link.springer.com/article/10.1007/
PL00000536;   http://www.springerlink.com/content/m0htmt9ww9b9jjv1/
;   http://www.springerlink.com/openurl.asp?genre=journal&
issn=1422-6944.

Hargittai:2004:TET

[Har04] Hargittai István. Teller Ede tragédiája. (Hungarian) [The tragedy of Edward Teller], volume 12 of Studia physica Savariensia. BDF, Szombathely, Hungary, 2004. ISBN 963-9531-25-1. ISSN 1219-2678. LCCN ????

Hargittai:2005:BRB

[Har05] István Hargittai. Book review: Stalin’s Great Science: The Times and Adventures of Soviet Physicists. By Alexei B. Kojevnikov (University of Georgia, Athens, USA, and Institute of History of Science and Technology, Moscow, Russia). Imperial College Press, London. 2004. xxiii + 360 pp. GPB 43.00. ISBN 1-86094-419-1. Structural Chemistry, 16(2):173–174, April 2005. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic). URL http://link.springer.com/

article/10.1007/s11224-005-2849-2.

Hargittai:2006:OVM

[Har06a] István Hargittai. Az öt világformáló marslakó. (Hungarian) [The five world-shaping Martians]. Vince, Budapest, Hungary, 2006. ISBN 963-9552-77-1. 397 + 32 pp. LCCN QC15 .H27155 2006.

Hargittai:2006:MSF

[Har06b] István Hargittai. The Martians of Science: five physicists who changed the twentieth century. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-517845-9 (hardcover). xxiv + 313 + 32 pp. LCCN QC15 .H27 2006. URL

http://www.loc.gov/catdir/enhancements/fy0636/2005029427-
d.html;   http://www.loc.gov/catdir/enhancements/fy0724/
2005029427-b.html;   http://www.loc.gov/catdir/toc/ecip061/
2005029427.html.

Hargittai:2008:ETG

[Har08a] István Hargittai. Edward Teller — guardian of freedom or Dr. Strangelove? Hungarian Quarterly, 49(189):108–122, ???? 2008. ISSN 0028-5390.

Hargittai:2008:MCO

[Har08b] István Hargittai. The Martians as chemists (on the occasion of the Teller centenary). Structural Chemistry, 19(1):3–4, February 2008. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic).

URL http://link.springer.com/article/10.1007/s11224-008-

9300-4. Based on a “Heritage Hour” presentation titled The Martians as Chemists on April 26, 2007, at the Chemical Heritage Foundation in Philadelphia, Pennsylvania. Edward Teller was born in Budapest on January 15, 1908.

Hargittai:2009:VIM

[Har09] Magdolna Hargittai. Vibronic interactions in metal halide molecules. Structural Chemistry, 20(1):21–30, February 2009. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic). URL http:

//link.springer.com/article/10.1007/s11224-008-9403-y.

Hargittai:2010:JET

[Har10] István Hargittai. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Prometheus Books, Amherst, NY, USA, 2010. ISBN 1-61614-221-9 (hardcover). 575 pp. LCCN QC16.T37 H37 2010. Afterword by Richard Garwin.

Hargittai:2011:BRS

[Har11a] Balazs Hargittai. Book review: Stephen B. Libby and Karl A. van Bibber (eds.): Edward Teller centennial symposium: modern physics and the scientific legacy of Edward Teller. Structural Chemistry, 22(4):957–960, August 2011. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic). URL http://link.springer.com/

article/10.1007/s11224-011-9821-0.

Hargittai:2011:ETC

[Har11b] István Hargittai. Edward Teller: Chemist, physicist, “Father of the hydrogen bomb”. Abstracts of Papers of the American Chemical Society, 241(??):??, March 27, 2011. CODEN ACSRAL. ISSN 0065-7727.

Hargittai:2011:T

[Har11c] Hargittai István. Teller. Akad. K., Budapest, Hungary, 2011. ISBN 963-05-8924-9. 563 + 16 pp. LCCN ???? Translated by Gács János.

Hastings:1970:MCS

[Has70] W. K. Hastings. Monte Carlo sampling methods using Markov chains and their applications. Biometrika, 57(1):97–109, April 1970. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (elec-

tronic). URL http://www.jstor.org/stable/2334940;   http:/
/www.probability.ca/hastings/. This paper introduces what is

now known as the Metropolis–Hastings algorithm, a generalization of the work in [MRR^{+}53]. See [Hit03, page 255].

Haxton:2010:ETN

[Hax10] Wick Haxton. Edward Teller and nuclei: along the trail to the neutrino. In Libby and Van Bibber [LV10], pages 80–95. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Hoddeson:2002:TGL

[HD02] Lillian Hoddeson and Vicki Daitch. True genius: the life and science of John Bardeen: the only winner of two Nobel Prizes in physics. Joseph Henry Press, Washington, DC, USA, 2002. ISBN 0-309-08408-3, 0-309-09511-5 (paperback). xi + 467 + 8 pp. LCCN QC16.B27 H63 2002.

Hall:1946:NSC

[HdH46] T. Hall and F. de Hoffman. Neutron spectrum from a cold parahydrogen radiator. Report LADC-111, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, February 11, 1946. Edited to conform to security regulations.

Hecker:2010:ETR

[Hec10] Siegfried S. Hecker. Edward Teller returns to Los Alamos. In Libby and Van Bibber [LV10], pages 4–12. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Hedberg:1949:BRB

[Hed49] Kenneth W. Hedberg. Book review: The Structure of Matter, by Francis Owen Rice and Edward Teller. Journal of Physical Chemistry, 53(8):1308, August 1949. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

Heimoff:1990:CET

[Hei90] Steven H. Heimoff. A conversation with Edward Teller. East Bay Express, 12(47):1, 12–14, 16, 18–19, ???? 1990.

Heller:1998:SSS

[Hel98] Arnie Heller. Still shaking up science today. Web document, July

1998. URL https://www.llnl.gov/str/Shaker.html.

Henriksen:2005:BRP

[Hen05] Paul Henriksen. Book review: Peter Goodchild: Edward Teller: The Real Dr. Strangelove. Isis, 96(4):672–673, December 2005. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http:

//www.jstor.org/stable/10.1086/501407.

Herbig:1976:KDA

[Her76] Jost Herbig. Kettenreaktion: das Drama der Atomphysiker. (German) [Nuclear reaction: the drama of the nuclear physicist]. Hanser-Verlag, München, Germany, 1976. ISBN ???? 514 + 8 pp. LCCN ???? Neumann János (1903–1957), Teller Ede (1908–2003), Szilárd Leó (1898-1964), Wigner Jenő (1902–1995).

Herzberg:1985:MSP

[Her85] Gerhard Herzberg. Molecular spectroscopy: A personal history. Annual Review of Physical Chemistry, 36:1–30, October 1985. CODEN ARPLAP. ISSN 0066-426X (print), 1545-1593 (elec-

tronic). URL http://www.annualreviews.org/doi/abs/10.1146/
annurev.pc.36.100185.000245.

Herken:1990:BRT

[Her90] Gregg Herken. Book review: Teller retells: Edward Teller: Giant of the Golden Age of Physics, by Stanley A. Blumberg and Louis G. Panos. Bulletin of the Atomic Scientists, 46(8):40–41, October 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Herken:2001:BRM

[Her01] Gregg Herken. Book review: a Martian’s chronicles. Memoirs: A Twentieth-Century Journey in Science and Politics by Edward Teller and Judy Shoolery. Science, 294(5547):1657–1658, November 23, 2001. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/3085278.
pdf.

Herken:2002:BBT

[Her02] Gregg Herken. Brotherhood of the bomb: the tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. Henry Holt and Company, Incorporated, New York, NY, USA, 2002. ISBN 0-8050-6588-1 (hardcover). xiv + 448 pp. LCCN QC16.O62 H47 2002.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;
http://alsos.wlu.edu/information.aspx?id=1264;   http://
www.loc.gov/catdir/bios/hol051/2002017219.html;   http://
www.loc.gov/catdir/description/hol021/2002017219.html.

Herken:2003:CBR

[Her03] Gregg Herken. Comment on book review of Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken [Am. J. Phys. 71(4), 411–415 (2003)]. American Journal of Physics, 71(7):647–648, July 2003. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://ajp.aapt.org/resource/1/ajpias/v71/
i7/p647_s1;   http://link.aip.org/link/?AJP/71/647/1. See

[Ber03c]. Includes a response by reviewer Jeremy Bernstein.

Hawkins:1987:TLW

[HGS87] Helen S. Hawkins, G. Allen Greb, and Gertrud Weiss Szilard. Toward a livable world: Leo Szilard and the crusade for nuclear arms control, volume 3 of Collected works of Leo Szilard. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-19260-8, 0-262-08162-8. lxxiv + 499 pp. LCCN QC3 .S97 vol. 3.

Hargittai:1997:ET

[HH97] István Hargittai and Magdolna Hargittai. Edward Teller. The Chemical Intelligencer, 3(1):14–23, January 1997. CODEN CHEIFK. ISSN 0947-0662.

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711

pp. LCCN QC15 .H295 2004. URL http://www.worldscibooks.
com/physics/p304.html.

Hargittai:2004:ET

[HH04b] Magdolna Hargittai and István Hargittai. Edward Teller. In Candid science IV: conversations with famous physicists [HH04a], pages 404– 423. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). LCCN QC15

.H295 2004. URL http://www.worldscibooks.com/physics/p304.
html.

Hargittai:2004:EPW

[HH04c] Magdolna Hargittai and István Hargittai. Eugene P. Wigner. In Candid science IV: conversations with famous physicists [HH04a], pages 1–19. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). LCCN

QC15 .H295 2004. URL http://www.worldscibooks.com/physics/
p304.html.

Hargittai:2008:LDL

[HH08] István Hargittai and Magdolna Hargittai. Lev D. Landau (1908– 1968): in memoriam. Structural Chemistry, 19(2):181–184, April 2008. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001

(electronic). URL http://link.springer.com/article/10.1007/
s11224-008-9323-x.

Hargittai:2009:HJR

[HH09] Magdolna Hargittai and István Hargittai. Hermann Jahn and Rudolf Renner of the Jahn–Teller and Renner–Teller effects. Structural Chemistry, 20(4):537–540, August 2009. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic). URL http://link.

springer.com/article/10.1007/s11224-009-9479-z.

Herzberg:1937:HSG

[HHT37] G. Herzberg, K. F. Herzfeld, and E. Teller. Heat of sublimation of graphite. Journal of Physical Chemistry, 41(2):325–331, February 1937. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

Howes:1999:TDS

[HHW99] Ruth (Ruth Hege) Howes, Caroline L. Herzenberg, and Ellen C. Weaver. Their day in the sun: women of the Manhattan Project. Labor and social change. Temple University Press, Philadelphia, PA, USA, 1999. ISBN 1-56639-719-7 (hardcover), 1-59213-192-1 (paperback), 0-585-38881-4 (e-book). viii + 264 pp. LCCN QC773.3.U5 H68 1999.

Higinbotham:1976:BRH

[Hig76] William A. Higinbotham. Book review: H. York, The Advisors: Oppenheimer, Teller and the Superbomb. Physics Today, 29(12):49–50, 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Hiltzik:2015:BSE

[Hil15] Michael A. Hiltzik. Big Science: Ernest Lawrence, the Cyclotron, and the Birth of the Military–Industrial Complex. Simon and Schuster, New York, NY, USA, 2015. ISBN 1-4516-7575-5 (hardcover), 1-4516-7576-3 (paperback), 1-4516-7603-4 (e-book). x + 512 pp. LCCN QC787.C8 H55 2015.

Hirsch:1988:BRB

[Hir88] Daniel Hirsch. Book review: Better a Shield than a Sword: Perspectives on Defense and Technology, by Edward Teller. Bulletin of the Atomic Scientists, 44(4):48–49, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Hitchcock:2003:HMH

[Hit03] David B. Hitchcock. A history of the Metropolis–Hastings algorithm. The American Statistician, 57(4):254–257, November 2003. CODEN ASTAAJ. ISSN 0003-1305 (print), 1537-2731 (electronic).

URL http://oberon.ingentaselect.com/cgi-bin/linker?ini=
asa&reqidx=/cw/asa/00031305/v57n4/s7/p254;   http://www.
jstor.org/stable/30037292.

Horowitz:1984:NW

[HLL^{+}84] Dan Horowitz, Robert J. Lieber, Edward N. Luttwak, Patrick Clawson, Russell Seitz, and Carl Sagan. Nuclear winter. Foreign Affairs, 62(4):995–1002, Spring 1984. CODEN FRNAA3. ISSN 0003-0554.

URL http://www.jstor.org/stable/20041916.

Hirsch:1990:HBW

[HM90] Daniel Hirsch and William G. Mathews. The H-bomb: Who really gave away the secret? Bulletin of the Atomic Scientists, 46 (1):22–30, January/February 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Kno90, MHM90, BHM91, HM97].

Hirsch:1997:LHB

[HM97] Daniel Hirsch and William G. Mathews. Letter: H-bomb secrets. Bulletin of the Atomic Scientists, 53(4):3, 58, July/August 1997. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90, KAS96] and comment [Han97].

Hora:2005:ETL

[HM05] Heinrich Hora and George H. (George Hunter) Miley, editors. Edward Teller lectures: lasers and inertial fusion energy. Imperial College Press, London, UK, 2005. ISBN 1-86094-468-X, 1-86094-727-1. xi + 365 pp. LCCN QC791.775 .L37 E39 2005. URL http://ebooks.

worldscinet.com/ISBN/9781860947278/9781860947278.shtml.

Hnatowich:2003:MFE

[Hna03] D. J. Hnatowich. Many faces of Edward Teller. Chemical and Engineering News, 81(42):4, ???? 2003. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605x (electronic).

Hollinger:2002:BBT

[Hol02] David A. Hollinger. Brotherhood of the bomb: the tangled lives and loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller. New York Times Book Review, ??(??):10, September 15, 2002. ISSN 0028-7806.

Holliday:2004:LEB

[Hol04] Bobbie Holliday. Letter to the Editor: Blast from the past. Bulletin of the Atomic Scientists, 60(6):5–??, November/December 2004. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://www.highbeam.com/doc/1G1-124135855.html.

Holzrichter:2010:ETH

[Hol10] John F. Holzrichter. Edward Teller and higher education. In Libby and Van Bibber [LV10], pages 144–156. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Horvath:1997:TKE

[Hor97] T. Horvath. Theodore Kármán, [Eugene] Paul Wigner, John [von] Neumann, Leo Szilard, Edward Teller and their ideas of ultimate reality and meaning. Ultimate Reality and Meaning, 20(2–3):123–146, June/September 1997. ISSN 0709-549X.

Houghton:2003:DEE

[Hou03] Susan Houghton. Director Emeritus Edward Teller, distinguished physicist and education advocate, dies at 95. Web

site., September 9, 2003. URL https://www.llnl.gov/news/
director-emeritus-edward-teller-distinguished-physicist-
and-education-advocate-dies-95.

Harper:1997:GGS

[HPA97] Eamon Harper, William Carleton Parke, and G. D. (George David) Anderson, editors. The George Gamow Symposium: sponsored by the George Washington University and the Carnegie Institution of Washington, 12 April 1997 [i.e., 1996], volume 129 of Astronomical Society of the Pacific conference series. Astronomical Society of the Pacific, San Francisco, CA, USA, 1997. ISBN 1-886733-49-X.

LCCN QC16.G37 G46 1996. URL http://adsabs.harvard.edu/
abs/1997ASPC..129.....H.

Hafemeister:1983:PTN

[HS83] David W. Hafemeister and Dietrich Schroeer, editors. Physics, technology, and the nuclear arms race, volume 104 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1983. ISBN 0-88318-203-3. LCCN U264 .P5 1983.

Hayen:2019:BSG

[HS19] L. Hayen and N. Severijns. Beta spectrum generator: High precision allowed β spectrum shapes. Computer Physics Communications, 240(??):152–164, July 2019. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.sciencedirect.

com/science/article/pii/S0010465519300645.

Herzberg:1933:SEM

[HT33] G. Herzberg and E. Teller. Schwingungsstruktur der Elektronenübergänge bei mehratomigen Molekülen. (German) [Oscillation structure of electron-transfer in polyatomic molecules]. Zeitschrift für physikalische Chemie (Leipzig), 21(B5–6):410–446, ???? 1933. CODEN ZPCLAH. ISSN 0323-4479.

Heitler:1936:TEM

[HT36] W. Heitler and E. Teller. Time effects in the magnetic cooling method. I. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 155A(??):629–639, July 1, 1936. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Hafstad:1938:APM

[HT38a] L. R. Hafstad and E. Teller. The alpha-particle model of the nucleus. Physical Review, 54(9):681–692, November 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v54/i9/p681_1.

Herzfeld:1938:VPI

[HT38b] Karl F. Herzfeld and Edward Teller. The vapor pressure of isotopes. Physical Review (2), 54(11):912–915, December 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v54/i11/p912_1.

Haupt:1939:SHD

[HT39a] R. F. Haupt and E. Teller. Specific heat and double minimum problem of NH_{3}. Journal of Chemical Physics, 7(10):925–927, ???? 1939. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/doi/10.1063/1.1750345.

Hibben:1939:REC

[HT39b] James Herbert Hibben and Edward Teller. The Raman effect and its chemical applications, volume 80 of American Chemical Society. Monograph series. Reinhold Publishing Corporation, New York, NY, USA, 1939. 544 pp. LCCN QC454 .H624.

Hendricks:1942:XRI

[HT42] Sterling Hendricks and Edward Teller. X-ray interference in partially ordered layer lattices. Journal of Chemical Physics, 10(3): 147–167, March 1942. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://link.aip.org/link/doi/10.
1063/1.1723678.

Ionson:1986:LSD

[IPB^{+}86] J. Ionson, K. Patel, C. Brau, C. Rhodes, R. Roberts, R. Sepucha, L. Marquet, G. Yonas, and E. Teller. Lasers in strategic defense — a panel discussion. Laser Focus-Electro-Optics, 22(2):124–??, ???? 1986. ISSN 0275-1399.

Isaacs:1993:BTO

[Isa93] John Isaacs. Bulletin: Three outrages. Bulletin of the Atomic Scientists, 49(5):6–7, June 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.com/

doc/1G1-13918630.html.

Inglis:1939:IDS

[IT39] D. R. Inglis and E. Teller. Ionic depression of series limits in oneelectron spectra. Astrophysical Journal, 90(3):439–448, October 1939. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Inglis:1940:PTO

[IT40] D. R. Inglis and E. Teller. On a proposed thermoelectric origin of the Earth’s magnetism. Physical Review (2), 57(12):1154–1155, June 15, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065

(electronic). URL http://prola.aps.org/abstract/PR/v57/i12/
p1154_1.

Jacobs:1965:BRB

[Jac65] Walter Darnell Jacobs. Book review: The Reluctant Revolutionary by Edward Teller. World Affairs, 128(3):206, December 1965. CODEN

???? ISSN ???? URL http://www.jstor.org/stable/pdfplus/

20670697.pdf. See [Tel64b].

Jacobs:2010:FHN

[Jac10] Robert A. Jacobs, editor. Filling the hole in the nuclear future: art and popular culture respond to the bomb. AsiaWorld. Lexington Books, Lanham, MD, USA, 2010. ISBN 0-7391-3556-2 (hardcover), 0-7391-3557-0 (paperback), 0-7391-3558-9 (e-book). xiv + 276 pp. LCCN NX180.S6 F55 2010.

Jain:1994:THM

[Jai94] Kewal K. Jain. Textbook of hyperbaric medicine. Hogrefe and Huber Publishers, Toronto, ON, Canada, 1994. ISBN 3-8017-0735-0. ???? pp. LCCN ????

Johnston:2003:RET

[Joh03] Don Johnston. Reflecting on Edward Teller’s life. Lawrence Livermore National Laboratory Newsline, 28(45):1, 4, November 7, 2003. URL

https://www.llnl.gov/file/26221/download?token=MSvJ2IZV.

Johnson:2013:CMG

[Joh13] Jeffrey Allan Johnson. The case of the missing German quantum chemists: On molecular models, mobilization, and the paradoxes of modernizing chemistry in Nazi Germany. Historical Studies in the Natural Sciences, 43(4):391–452, September 2013. CODEN ???? ISSN 1939-1811 (print), 1939-182X (electronic). URL http://www.

jstor.org/stable/10.1525/hsns.2013.43.4.391.

Jones:1962:BRB

[Jon62] Vincent C. Jones. Book review: Now It Can be Told: The Story of the Manhattan Project by Leslie R. Groves. Men and Decisions by Lewis L. Strauss. The Legacy of Hiroshima by Edward Teller and Allen Brown. Military Affairs, 26(4):178–179, Winter 1962. CODEN

???? ISSN ???? URL http://www.jstor.org/stable/pdfplus/

1985613.pdf. See [Gro62, Str63, TB62, TB75].

Jones:1985:MAA

[Jon85] Vincent C. Jones. Manhattan, the Army and the atomic bomb. United States Army in World War II. Special studies. Center of Military History, U.S. Army, Washington, DC, USA, 1985. xx + 660 + 4 pp. LCCN QC773.3.U5 J65 1985.

Jahn:1937:SPM

[JT37] H. A. Jahn and E. Teller. Stability of polyatomic molecules in degenerate electric states. I. Orbital degeneracy. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, A161(??):220–235, July 15, 1937. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Johnson:1954:LEP

[JT54] Montgomery H. Johnson and Edward Teller. Letter to the editor: Proton distribution in heavy nuclei. Physical Review, 93(2):357–358, January 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/v93/
i2/p357_1.

Johnson:1955:CFT

[JT55] Montgomery H. Johnson and Edward Teller. Classical field theory of nuclear forces. Physical Review (2), 98(3):783–787, May 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v98/i3/p783_1.

Johnson:1982:ICD

[JT82] Montgomery H. Johnson and Edward Teller. Intensity changes in the Doppler effect. Proceedings of the National Academy of Sciences of the United States of America, 79(4):1340, February 15, 1982. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http:

//www.jstor.org/stable/pdfplus/11852.pdf.

Johnson:1984:EOG

[JT84] Montgomery H. Johnson and Edward Teller. On the extragalactic origin of gamma-ray bursts. Report UCRL-91744 (Rev. 1), University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 1984. Presented at the Quantum Theory and the Structures of Time and Space, 6th Symposium, Tutzing, Germany, July 2–5, 1984.

Khariton:1996:WIW

[KAS96] Yuli Khariton, Viktor Adamskii, and Yuri Smirnov. The way it was. Bulletin of the Atomic Scientists, 52(9):54–59, November/December 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-18855733.html.

See comment [HM97].

Kauffman:2012:BRI

[Kau12] George B. Kauffman. Book review: István Hargittai: Judging Edward Teller: A closer look at one of the most influential scientists of the twentieth century. Foundations of Chemistry, 14(1):99–101, April 2012. CODEN FOCHFL. ISSN 1386-4238 (print), 1572-8463

(electronic). URL http://link.springer.com/article/10.1007/
s10698-011-9133-x.

Koppel:2009:JTE

[KBY09] Horst Köppel, Heinz Barentzen, and David R. Yarkony. The Jahn– Teller Effect: Fundamentals and Implications for Physics and Chemistry, volume 97 of Springer Series in Chemical Physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. ISBN 3-642-03432-2. xxi + 915 pp. LCCN QD461 .J346 2009.

Kearny:1980:NWS

[Kea80] Cresson H. Kearny. Nuclear war survival skills. NWS Research Bureau, Coos Bay, OR, USA, revised and updated edition, 1980. ISBN 0-939002-02-7 (paperback). 232 pp. LCCN UF767 .K4 1980. US$9.95. Foreword by Edward Teller.

Kenworthy:1954:DHB

[Ken54] E. W. Kenworthy. The drama of the hydrogen bomb — and Dr. Oppenheimer’s key role: Security case focuses attention on disputes that preceded first successful test of H-bomb at Pacific Proving Ground. New York Times, ??(??):E5, April 18, 1954. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL https://search.

proquest.com/hnpnewyorktimes/docview/113104238/.

Kenworthy:1963:PVU

[Ken63] E. W. Kenworthy. President vows U.S. will step up atomic readiness: Disputes Teller on danger to security in Test Ban — Schriever against pact. New York Times, ??(??):??, August 21, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Kevles:2003:SCR

[Kev03] Daniel J. Kevles. The strange case of Robert Oppenheimer: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller by Gregg Herken, Henry Holt, 448 pp., $30.00. Pandora’s Keepers: Nine Men and the Atomic Bomb by Brian VanDeMark Little, Brown, 339 pp. (withdrawn). New York Review of Books, 50(19):37– 40, December 4, 2003. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/2003/
dec/04/the-strange-case-of-robert-oppenheimer/.

Kuckuck:2003:ETR

[KGA^{+}03] Bob Kuckuck, Bruce Goodwin, Bob Andrews, Greg Herken, Hans Mark, Sig Hecker, and Judith L. Shoolery. Edward Teller remembered by friends and colleagues. Newsline (Lawrence Livermore National Laboratory), 28(36):3, September 12, 2003. URL https://newsline.

llnl.gov/employee/articles/2003/09-12-03-newsline.pdf.

Kihss:1967:BUT

[Kih67] Peter Kihss. A-blasts urged to tap gas wells: Teller says plan would put new sources into use. New York Times, ??(??):??, January 11, 1967. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Kirsch:1998:EPE

[Kir98] S. Kirsch. Experiments in progress: Edward Teller’s controversial, geographies. Ecumene, 5(3):267–285, ???? 1998. ISSN 0967-4608.

Kirsch:2005:PGP

[Kir05] Scott Kirsch. Proving grounds: Project Plowshare and the unrealized dream of nuclear earthmoving. Rutgers University Press, New Brunswick, NJ, USA, 2005. ISBN 0-8135-3666-9 (hardcover). xi + 257 pp. LCCN TA748 .K57 2005. URL http://www.loc.gov/

catdir/toc/ecip057/2005002579.html.

Kirz:2010:TE

[Kir10] Janos Kirz. Teller on energy. In Libby and Van Bibber [LV10], pages 140–143. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Klein:1962:BRB

[Kle62] Stanley Klein. Book review: The Legacy of Hiroshima by Edward Teller and Allen Brown. Technology and Culture, 3(3):359–362, Summer 1962. CODEN TECUA3. ISSN 0040-165X (print), 1097-

3729 (electronic). URL http://www.jstor.org/stable/pdfplus/
3100847.pdf;   https://muse.jhu.edu/pub/1/article/895155/

pdf. See [TB62, TB75].

Kuran:1997:TBA

[KMN^{+}97] Peter Kuran, Alan Munro, Scott Narrie, Don Pugsley, William Shatner, Edward Teller, Frank H. Shelton, Barbu Marian, William T. Stromberg, Jacqueline Zietlow, and William S. Conner. Trinity and beyond: the atomic bomb movie. Goldhil Video, Thousand Oaks, CA, USA, 1997. ISBN ???? LCCN ???? 1 videocassette (95 minutes).

Konopinski:1946:IAN

[KMT46] Emil Konopinski, Cloyd Marvin, Jr., and Edward Teller. Ignition of the atmosphere with nuclear bombs. Report LA-602, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, August 14, 1946. Declassified 1974.

Konopinski:1943:CAU

[KMTW43] Emil Konopinski, Nick Metropolis, Edward Teller, and Leona Woods. Critical amounts of uranium compounds. Technical Report CF-548, United States Atomic Energy Commission, Washington, DC, USA, March 19, 1943. 13 pp. URL http://www.

osti.gov/accomplishments/pdf/DE04384835/DE04384835.pdf;
http://www.osti.gov/servlets/purl/4384835-NQvnzm/native/.

Knight:2003:CLM

[Kni03] Jonathan Knight. California laboratory mourns loss of h-bomb pioneer. Nature, 425(6955):226, September 18, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.

nature.com/nature/journal/v425/n6955/full/425226b.html.

Knoll:1990:LHB

[Kno90] Erwin Knoll. Letter: H-bomb secret never was. Bulletin of the Atomic Scientists, 46(3):53–54, April 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90].

Koestler:1940:DN

[Koe40] Arthur Koestler. Darkness at noon. J. Cape, London, UK, 1940. 254 pp. LCCN PZ3.K8194 PR6021.O4. URL https://en.wikipedia. org/wiki/Arthur_Koestler. Translated by Daphne Hardy in Paris from the original German manuscript, and smuggled out of France to Britain just before the Nazi invasion. Its first publication was thus in English. The book has since been republished numerous times in English and several other languages.

Kursunoglu:1973:IBR

[KPB73] Behram Kurşunoğlu, Arnold Perlmutter, and John Bardeen. Impact of basic research on technology, volume 1 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1973. ISBN 0-306-36901-X. xv + 301 pp. LCCN QC30 .K82.

Kleint:2005:WHB

[KRW05] Christian Kleint, Helmut Rechenberg, and Gerald Wiemers, editors. Werner Heisenberg, 1901–1976: Beiträge, Berichte, Briefe: Festschrift zu seinem 100. Geburtstag. (German) [Werner Heisenberg, 1901–1976: Contributions, Reports, and Letters: Celebration of his 100th Birthday], volume 62 of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse; Aufsatzsammlung. Verlag der Sächsischen Akademie der Wissenschaften zu Leipzig, Leipzig, Germany, 2005. ISBN 3-7776-1402-5. ISSN 0365-6470. 424 pp. LCCN

QC16.H35 W474 2005. URL http://www.loc.gov/catdir/toc/
fy0613/2006402966.html.

Kalckar:1934:RMM

[KT34] F. Kalckar and E. Teller. Ratio of the magnetic moments of proton and diplon. Nature, 134(3379):180, August 4, 1934. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

URL http://www.nature.com/nature/journal/v134/n3379/pdf/
134180a0.pdf.

Kalckar:1935:TCO

[KT35] F. Kalckar and E. Teller. On the theory of the catalysis of the ortho-para transformation by paramagnetic gases. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 150A(??):520–533, July 1, 1935. CODEN PRLAAZ. ISSN 0080-4630 (print), 2053-9169 (electronic).

Kuper:1940:SFE

[KT40] J. B. Horner Kuper and Edward Teller. Scattering of fast electrons in helium. Physical Review, 58(7):602–603, October 1, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v58/i7/p602_1.

Knipp:1941:ELH

[KT41] Julian Knipp and Edward Teller. On the energy loss of heavy ions. Physical Review, 59(8):659–669, April 15, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v59/i8/p659_1.

Konopinski:1948:TCC

[KT48] Emil J. Konopinski and Edward Teller. Theoretical considerations concerning the D + D reactions. Physical Review, 73(8):822–830, April 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/v73/
i8/p822_1.

Kursunoglu:1973:FIP

[Kur73] Behram Kurşunoğlu, editor. Fundamental Interactions in Physics: proceedings of the Coral Gables Conference on Fundamental Interactions, January 22–26, 1973, volume 2 of Studies in the natural sciences. Plenum Press, New York, NY, USA; London, UK, 1973. ISBN 0-306-36902-8. LCCN ????

Kurzman:1986:DBC

[Kur86] Dan Kurzman. Day of the bomb: countdown to Hiroshima. McGraw-Hill, New York, NY, USA, 1986. ISBN 0-07-035683-1. xiv + 546 + 24 pp. LCCN D767.25.H6 K865 1986.

Kleint:1993:WHL

[KW93] Christian Kleint and Gerald Wiemers, editors. Werner Heisenberg in Leipzig, 1927–1942, volume 58(2) of Abhandlungen der Sächsischen Akademie der Wissenschaften zu Leipzig, Mathematisch-Naturwissenschaftliche Klasse, 0365-6470. Akademie-Verlag, Berlin, Germany, 1993. ISBN 3-05-501585-1. 263 pp. LCCN QC16.H35 W48 1933. EUR$607 (!).

Kaplan:2001:VIJ

[KZ01] Michael D. Kaplan and George O. Zimmerman, editors. Vibronic interactions: Jahn–Teller effect in crystals and molecules, volume 39 of NATO science series. Sub-series II, Mathematics, physics and chemistry. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 1-4020-0044-8. LCCN QD461 .V48 2001.

Lakoff:1978:ECL

[Lak78] S. A. Lakoff. Energy and conflict — life and times of Edward Teller — S. A. Blumberg and G. Owens. International Journal, 33(3):624–629, ???? 1978. ISSN 0020-7020.

Lanouette:1983:DM

[Lan83] William Lanouette. Dream machine. Atlantic Monthly, ??(??):35–52, 85–86, April 1983. ISSN 1072-7825 (print), 2151-9463 (electronic). A 20-page cover story about the history of the nuclear breeder reactor, a power plant designed to make more fuel than it consumes, including Szilard’s role both devising and naming the nuclear “chain reaction” in 1933 and the “breeder” in 1943. Letters and author’s replies in June, pages 6-7; July, pages 6 & 8; and August.

Langone:1984:ETB

[Lan84] J. Langone. Edward Teller — of bombs and brickbats. Discover, 5 (7):63–??, ???? 1984. ISSN 0274-7529.

Lanouette:1993:HGT

[Lan93a] William Lanouette. Hungarian geniuses and their bomb. Seminar on Science and Social Responsibility, about Szilard, Teller, von Neumann, and Wigner. Georgetown University, Washington, DC, USA., May 10, 1993.

Lanouette:1993:WFM

[Lan93b] William Lanouette. Where the four Martians of the Manhattan Project landed: The national laboratory connections of Teller, Wigner, von Neumann, and Szilard. Talk at the American Physical Society, Washington, DC, USA., April 13, 1993.

Lanouette:1999:WFM

[Lan99] William Lanouette. Where the four Martians of the Manhattan Project landed: The National Laboratory connections of Teller, Wigner, von Neumann, and Szilard. A Talk at a Conference on Laboratory History and Sociology, State University of New York at Stony Brook & Brookhaven National Laboratory, June 10, 1999.

Lanouette:2011:FHM

[Lan11] William Lanouette. The Four Hungarians of the Manhattan Project. Washington Magyar Club. Described the roles of Leo Szilard, Eugene Wigner, John von Neumann, and Edward Teller in the Manhattan Project., January 15, 2011.

Lapp:1962:KOS

[Lap62] Ralph Eugene Lapp. Kill and overkill: the strategy of annihilation. Basic Books, New York, NY, USA, 1962. 197 pp. LCCN UA11 .L3.

Laurence:1965:WYM

[Lau65] William L. Laurence. Would you make the bomb again? The New York Times Magazine, ??(??):8–??, August 1, 1965. ISSN 0362-1308.

Lawren:1988:GBB

[Law88] William Lawren. The general and the bomb: a biography of General Leslie R. Groves, director of the Manhattan Project. Dodd, Mead, New York, NY, USA, 1988. ISBN 0-396-08761-2. xii + 324 pp. LCCN UG128.G76 L39 1988.

Lemonick:2003:MSE

[LBB^{+}03] Michael D. Lemonick, Harriet Barovick, Elizabeth L. Bland, Daren Fonda, Kate Novack, and Richard Corliss. Milestones Sep. 22, 2003: Edward Teller. Time Magazine, ??(??):??, September 22, 2003. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (elec-

tronic). URL http://content.time.com/time/magazine/article/
0,9171,1005751,00.html.

Linhart:1975:EEC

[LBF^{+}75] J. G. Linhart, O. Buneman, H. T. Furth, J. Luce, E. Teller, L. Motz, ?. Winterberg, R. Sudan, Ryutov, Schumaker, J. G. Lominadze, ?. Consoli, H. Sahlin, ?. Bostick, ?. Fleischmann, and ?. Rosenbluth. Electrostatic and electromagnetic confinement of plasmas and phenomenology of relativistic electron-beams — panel discussion. Annals of the New York Academy of Sciences, 251(MAY8):690–711, ???? 1975. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Leith:1988:CPG

[Lei88] C. E. Leith. The computational physics of the global atmosphere. In Mark and Wood [MW88], pages 161–173. ISBN 90-277-2775-9. LCCN QC71 .E62 1988.

Lennick:2005:ET

[Len05] Michael Lennick. Edward Teller. American Heritage, ??(??):54–63, June/July 2005.

Leskov:1993:DGF

[Les93] Sergei Leskov. Dividing the glory of the fathers. Bulletin of the Atomic Scientists, 49(4):37–39, May 1993. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

highbeam.com/doc/1G1-13869959.html.

Lewis:1985:SWC

[Lew85] Flora Lewis. A ‘Star Wars’ cover-up. New York Times, ??(??):??, December 3, 1985. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Libby:2011:BRB

[Lib11] Stephen B. Libby. Book review: Judging Edward Teller: A Closer Look at One of the Most Influential Scientists of the Twentieth Century, by István Hargittai. Physics Today, 64(2):49, February 2011. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/64/49/1;   http:
//link.aip.org/link/phtoad/v64/i2/p49/s1. See [Har10].

Lindner:1998:AEE

[Lin98] Konrad Lindner. Ein Atomphysiker erzählt: Edward Teller zwischen Leipzig und Livermore. (German) [A nuclear physicist says: Edward Teller between Leipzig and Livermore]. Interview, Universität Leipzig, Leipzig, Germany, 1998. 41 pp.

Larson:1984:IET

[LL84] Clarence Larson and Jane Larson. Interview with Edward Teller. Video program, 1984.

Latter:1959:SSL

[LMT59] A. L. Latter, E. A. Martinelli, and E. Teller. Seismic scaling law for underground explosions. Physics of Fluids, 2(3):280–282, ???? 1959. CODEN PHFLE6. ISSN 1070-6631.

Loeber:2005:BRB

[Loe05] Charles R. Loeber. Book review: Edward Teller: The Real Dr. Strangelove, by Peter Goodchild. American Journal of Physics, 73(9): 893–894, September 2005. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://ajp.aapt.org/resource/
1/ajpias/v73/i9/p893_s1;   http://link.aip.org/link/?AJP/
73/893/1.

Longmire:1958:PPT

[Lon58] Conrad L. Longmire, editor. Plasma physics and thermonuclear research, volume 11 of Progress in nuclear energy. Pergamon Press, New York, NY, USA, 1958. ISSN 0555-4136. ???? pp. LCCN QC791 .P5.

Lourie:2002:SB

[Lou02] Richard Lourie. Sakharov: a biography. Brandeis University Press, Waltham, MA, USA, 2002. ISBN 1-58465-207-1 (hardcover). 465 pp. LCCN DK275.S25 L68 2002.

Libby:2009:ETB

[LS09] Stephen B. Libby and Andrew M. Sessler. Edward Teller biographical memoir. In Libby and Van Bibber [LV10], pages 13–61. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008. URL https://e-reports-ext.llnl.gov/pdf/376159.pdf. Lawrence Livermore National Laboratory report LLNL-CONF-415308.

Lyddane:1941:PVA

[LST41] R. H. Lyddane, R. G. Sachs, and E. Teller. On the polar vibrations of alkali halides. Physical Review, 59(8):673–676, April 15, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v59/i8/p673_1.

Landau:1936:TSG

[LT36] L. Landau and E. Teller. Zur Theorie der Schalldispersion. (German) [On the theory of sound dispersion]. Physikalische Zeitschrift der Sowjetunion, 10(1):34–43, ???? 1936. CODEN PHZSAL. ISSN 0369-9811.

Lord:1937:SBP

[LT37] R. C. Lord and E. Teller. Structure of benzene. Pt. X. intensities of Raman lines in benzene and hexadeuterobenzene. Chem. Soc. J., ?? (??):1728–1737, ???? 1937. CODEN ???? ISSN ????

Lennick:2005:WKI

[LT05] M. Lennick and E. Teller. “We knew that if we succeeded, we could at one blow destroy the city” (final interview with Edward Teller). American Heritage, 56(3):54–63, ???? 2005. ISSN 0002-8738.

Libby:2010:ETC

[LV10] Stephen B. Libby and Karl A. Van Bibber, editors. Edward Teller Centennial Symposium: modern physics and the scientific legacy of Edward Teller: Livermore, CA, USA, 28 May 2008. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2010. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Libby:2004:CTS

[LW04a] Stephen B. Libby and Morton S. Weiss. Clarifying Teller’s science story. Physics Today, 57(12):17, December 2004. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//physicstoday.org/resource/1/phtoad/v57/i12/p17_s1.

Libby:2004:ETS

[LW04b] Stephen B. Libby and Morton S. Weiss. Edward Teller’s scientific life. Physics Today, 57(8):45–50, August 2004. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//link.aip.org/link/?PTO/57/45/1;   http://physicstoday.
org/resource/1/phtoad/v57/i8/p45_s1.

Liu:1995:CNY

[LY95] C. S. (Chao Shiuan) Liu and Shing-Tung Yau, editors. Chen Ning Yang: a great physicist of the twentieth century. International Press, Cambridge, MA, USA, 1995. ISBN 1-57146-001-2. LCCN QC16.Y364 A3 1995.

MacPherson:1986:TBF

[Mac86] Malcolm MacPherson. Time bomb: Fermi, Heisenberg, and the race for the atomic bomb. Dutton, New York, NY, USA, 1986. ISBN 0-525-24409-3. xiii + 316 + 8 pp. LCCN QC773 .M24 1986. URL http:

//catalog.hathitrust.org/api/volumes/oclc/12751401.html.

MacPherson:1987:TBF

[Mac87] Malcolm MacPherson. Time bomb: Fermi, Heisenberg, and the race for the atomic bomb. Berkley Books, New York, NY, USA, 1987. ISBN 0-425-10423-0. xi + 284 + 8 pp. LCCN QC773 .M24 1987.

Macrae:1992:JNSa

[Mac92] Norman Macrae. John von Neumann: The Scientific Genius Who Pioneered the Modern Computer, Game Theory, Nuclear Deterrence, and Much More. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-41308-1. x + 405 pp. LCCN QA29.V66 M33 1992. US$25.00 (US$31.50 Can.).

Maddox:2003:OET

[Mad03] John Maddox. Obituary: Edward Teller (1908–2003). Nature, 425 (6956):362–363, September 25, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v425/n6956/full/425362a.html.

Magraw:1988:TCB

[Mag88] Katherine Magraw. Teller and the “clean bomb” episode. Bulletin of the Atomic Scientists, 44(4):32, 34–37, May 1988. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Noemi:2004:NET

[Mál04] Málovics Noémi. A nukleáris energia tudósa, Teller Ede. (Hungarian) [Edward Teller, a nuclear scientist]. ????, ????, 2004. ISBN ???? 209–217 pp. LCCN ????

Margenau:1981:BRB

[Mar81] Henry Margenau. Book review: The Pursuit of Simplicity by Edward Teller. American Scientist, 69(3):349, May 1981. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://www.

jstor.org/stable/pdfplus/27850514.pdf. See [Tel80c].

Marsh:1987:PDL

[Mar87] Gerald E. Marsh. Perspective: Dangers of limited SDI. Bulletin of the Atomic Scientists, 43(2):13–14, March 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [RC87, SLM87, GM87].

March:1992:BRB

[Mar92] Robert March. Book review: Conversations on the Dark Secrets of Physics, by Edward Teller and Wendy Teller and Wilson Talley. Physics Today, 45(1):74, January 1992. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//physicstoday.org/resource/1/phtoad/v45/i1/p74_s1.  See

[TTT02].

Marx:1996:MMG

[Mar96] George Marx. The myth of the Martians and the golden age of Hungarian science. Science & Education (Springer), 5(3):225–234, July 1996. CODEN SCEDE9. ISSN 0926-7220 (print), 1573-1901 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1996Sc%26Ed...5.
.225M.

Marsh:1997:BRD

[Mar97] Gerald Marsh. Book review: Details, details: The Swords of Armageddon, by Chuck Hansen. Bulletin of the Atomic Scientists, 53(4): 54, 56, July/August 1997. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/1G1-

19580893.html. See comment [GM87].

Marx:1998:ETN

[Mar98] George Marx. Edward Teller ninety. Acta Physica Hungarica New Series: Heavy Ion Physics, 7(2):139–153, December 1, 1998. CODEN APHPFM. ISBN 963-05-7269-9. ISSN 1219-7580 (print), 1588-2675 (electronic).

Marton:2006:GEN

[Mar06] Kati Marton. The great escape: nine Jews who fled Hitler and changed the world. Simon and Schuster, New York, NY, USA, 2006. ISBN 0-7432-6115-1 (hardcover), 0-7432-6116-X (paperback). 271 + 16 pp. LCCN DS135.H93 A153 2006. URL http://catdir.loc.

gov/catdir/enhancements/fy0665/2006049162-d.html;   http:
//catdir.loc.gov/catdir/enhancements/fy0668/2006049162-s.
html;   http://catdir.loc.gov/catdir/enhancements/fy0668/
2006049162-t.html;   http://catdir.loc.gov/catdir/enhancements/
fy0912/2006049162-b.html.

Marx:2010:VMH

[Mar10] Marx György. The voice of the Martians: Hungarian scientists who shaped the 20th century in the West. Akad. K., Budapest, Hungary, third edition, 2010. ISBN 963-05-7830-1. 427 pp. LCCN ???? US$. Epilogue by Edward Teller.

Matthews:1998:GGR

[Mat98] Robert Matthews. Giants go on the record. New Scientist, ??(??): ??, May 9, 1998. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-

8500 (electronic). URL https://www.newscientist.com/article/
mg15821336-300-giants-go-on-the-record/.

Mathews:1992:SWE

[MB92] G. J. (Grant J.) Mathews and Stewart Dave Bloom, editors. Strong, weak, and electromagnetic interactions in nuclei, atoms, and astrophysics: Livermore, CA 1991, volume 242 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1992. ISBN 0-88318-943-7. LCCN QC793.9 .S77 1992. URL http://scitation.

aip.org/content/aip/proceeding/aipcp/242.

Maniscalco:1983:FBE

[MBM^{+}83] J. A. Maniscalco, D. H. Berwald, R. W. Moir, J. D. Lee, and Edward Teller. The fusion breeder — an early application of nuclear fusion. Report UCRL-87801, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, June 1983. For submission to Science.

Maniscalco:1984:FBE

[MBM^{+}84] J. A. Maniscalco, D. H. Berwald, R. W. Moir, J. D. Lee, and E. Teller. The fusion breeder — an early application of nuclear-fusion. Fusion Technology, 6(3):584–596, ???? 1984. CODEN FUSTE8. ISSN 0748-1896.

McGrory:1985:SNS

[McG85] Mary McGrory. Senate nonsense on Star Wars. Washington Post, ?? (??):??, June 9, 1985. ISSN 0190-8286.

McGrory:1986:SWG

[McG86] Mary McGrory. A ‘Star Wars’ gathering. Washington Post, ??(??): ??, June 24, 1986. ISSN 0190-8286.

McKay:1984:MAA

[McK84] H. A. C. (Herbert Alwyn Cochrane) McKay. The making of the atomic age. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1984. ISBN 0-19-289174-X (paperback), 0-19-219193-4. xii + 153 + 8 pp. LCCN QC773 .M35 1984.

McMillan:1990:BRE

[McM90] Priscilla Johnson McMillan. Books: Review of Edward Teller: Giant of the Golden Age of Physics, by Stanley A. Blumberg and Louis G. Panos. Scientific American, 262(5):130–135, May 1990. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087

(electronic). URL http://www.nature.com/scientificamerican/
journal/v262/n5/pdf/scientificamerican0590-130.pdf.

McMillan:1994:SFF

[McM94] Priscilla Johnson McMillan. The Sudoplatov file: Flimsy memories. Bulletin of the Atomic Scientists, 50(4):30–33, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://www.highbeam.com/doc/1G1-15587910.html.

McMillan:2003:ETO

[McM03] Priscilla Johnson McMillan. Edward Teller: Obituary. Bulletin of the Atomic Scientists, 59(6):10–11, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.highbeam.com/doc/1G1-110621486.html.

McMillan:2005:RJR

[McM05] Priscilla Johnson McMillan. The ruin of J. Robert Oppenheimer, and the birth of the modern arms race. Viking, New York, NY, USA, 2005. ISBN 0-670-03422-3. viii + 373 + 16 pp. LCCN QC16.O62 M36 2005.

Medvedev:1979:NDUa

[Med79a] Zhores A. Medvedev. Nuclear disaster in the Urals. Norton, New York, NY, USA, 1979. ISBN 0-393-01219-0. vii + 214 pp. LCCN TD196.R3 M42 1979. Translated by George Saunders from the Russian original.

Medvedev:1979:NDUb

[Med79b] Zhores A. Medvedev. Nuclear disaster in the Urals. Vintage Books, New York, NY, USA, 1979. ISBN 0-394-74445-4. vii + 214 pp. LCCN TD196.R3 M42 1980. Translated by George Saunders from the Russian original.

Medvedev:1979:NDUc

[Med79c] Žores Aleksandrovič Medvedev. Nuclear disaster in the Urals. Angus and Robertson, London, UK, 1979. ISBN 0-207-95896-3. vii + 214 pp. LCCN ???? Translated by George Saunders from the Russian original.

Meyers:1987:EPS

[Mey87] Robert A. (Robert Allen) Meyers, editor. Encyclopedia of physical science and technology. Academic Press, New York, USA, 1987. ISBN 0-12-226901-2 (vol. 1), 0-12-226902-0 (vol. 2), 0-12-226903-9 (vol. 3), 0-12-226904-7 (vol. 4), 0-12-226905-5 (vol. 5), 0-12-226906-3 (vol. 6), 0-12-226907-1 (vol. 7), 0-12-226908-X (vol. 8), 0-12-226909-8 (vol. 9), 0-12-226910-1 (vol. 10), 0-12-226911-X (vol. 11), 0-12-226912-8 (vol. 12), 0-12-226913-6 (vol. 13), 0-12-226914-4 (vol. 14). ???? pp. LCCN Q123 .E497 1987.

Mark:1969:PMU

[MF69] Hans Mark and Sidney Fernbach, editors. Properties of matter under unusual conditions: in honor of Edward Teller’s 60th birthday. Interscience Publishers, New York, NY, USA, 1969. ISBN 0-470-56990-5. ix + 389 pp. LCCN Q171 .P98.

Miley:1992:LIR

[MH92] George H. Miley and Heinrich Hora, editors. Laser interaction and related plasma phenomena. Plenum Press, New York, NY, USA; London, UK, 1992. ISBN 0-306-44353-8. LCCN QC717.6 L343 1991 V10.

Moss:1990:LWF

[MHM90] Norman Moss, Daniel Hirsch, and William G. Mathews. Letters: What Fuchs didn’t know, and when he didn’t know it. Bulletin of the Atomic Scientists, 46(4):51–52, May 1990. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [HM90].

Miller:1986:UCD

[Mil86] Richard L. (Richard Lee) Miller. Under the cloud: the decades of nuclear testing. Free Press, New York, NY, USA, 1986. ISBN 0-02-921620-6. xii + 547 + 8 pp. LCCN U264 .M55 1986.

Miller:2002:TT

[Mil02] John J. Miller. Truth Teller. National Review, ??(??):38–40, September 30, 2002. ISSN 0028-0038.

Miller:2003:ETR

[Mil03a] John J. Miller. Edward Teller, RIP. National Review, ??(??): ??, September 10, 2003. ISSN 0028-0038. URL http://www.

nationalreview.com/corner/68774/edward-teller-rip/john-j-
miller.

Miller:2003:MT

[Mil03b] John J. Miller. More Teller. National Review, ??(??):??, September

10, 2003. ISSN 0028-0038. URL http://www.nationalreview.com/
corner/68775/more-teller/john-j-miller.

Miller:2003:TTNa

[Mil03c] John J. Miller. Truth Teller: The nuclear scientist the Left loves to hate. National Review, ??(??):??, July 23, 2003. ISSN 0028-

0038. URL http://www.nationalreview.com/articles/207587/

truth-teller/john-j-miller. Comments on Edward Teller’s receipt of the US Presidential Medal of Freedom.

Miller:2003:TTNb

[Mil03d] John J. Miller. Truth Teller: The nuclear scientist the Left loves to hate. National Review, ??(??):??, September 10, 2003. ISSN 0028-

0038. URL http://www.nationalreview.com/articles/207978/

truth-teller/john-j-miller. Comments on the death of Edward Teller on 9 September 2003.

Mitchell:1997:ASD

[Mit97] Gordon Mitchell. Another strategic deception initiative: after 14 years of secrecy and misinformation on missile defense research, Congress and the public deserve openness. Bulletin of the Atomic Scientists, 53(3):22–23, May/June 1997. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-19201489.html.

Mitcham:2005:EST

[Mit05] Carl Mitcham, editor. Encyclopedia of Science, Technology, and Ethics. Macmillan Reference, Detroit, MI, USA, 2005. ISBN 0-02-865831-0 (set ), 0-02-865832-9 (vol. 1), 0-02-865833-7 (vol. 2), 0-02-865834-5 (vol. 3), 0-02-865901-5 (vol. 4). cxiv + 2378 pp.

LCCN Q175.35 .E53 2005. URL http://www.loc.gov/catdir/toc/
ecip059/2005006968.html. Four volumes.

Maroncelli:2002:TGN

[MK02] James M. Maroncelli and Timothy L. Karpin. The traveler’s guide to nuclear weapons: a journey through America’s cold war battlefields. Historical Odysseys Pub., Silverdale, WA, USA, 2002. ISBN 0-9725051-0-5. ???? pp. LCCN U264.3. One CD-ROM.

Molella:2023:AAP

[MK23] Arthur Molella and Robert Kargon. Atomville: Architects, planners, and how to survive the Bomb. Technology and Culture, 64(3):823–844, July 2023. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729

(electronic). URL https://muse.jhu.edu/pub/1/article/903974.

Metropolis:1987:NDP

[MKR87] N. (Nicholas) Metropolis, Donald M. Kerr, and Gian-Carlo Rota, editors. New Directions in Physics: The Los Alamos 40th Anniversary Volume. Academic Press, New York, USA, 1987. ISBN 0-12-492155-8. xii + 292 pp. LCCN QC44 .N49 1987. URL http://www.osti.gov/

energycitations/product.biblio.jsp?osti_id=6120718&query_
id=0.

Middleton:2007:EIQ

[ML07] Carolin Middleton and Stephen B. Libby. Exchanging insights on quantum behavior: Teller’s contributions to condensed-matter physics. Science & Technology Review (Lawrence Livermore National Laboratory), ??(??):26–27, May 2007.

Mladjenovic:1998:DYN

[Mla98] Milorad Mladjenović. The Defining Years in Nuclear Physics, 1932– 1960s. IOP Publishing, Bristol, UK, 1998. ISBN 0-7503-0472-3 (hardcover). xx + 441 pp. LCCN QC773 .M54 1998.

Mohr:1983:SDM

[Moh83] Charles Mohr. Scientists dubious over missile plan: Some consider it technically unworkable — others view it as strategic danger. New York Times, ??(??):??, March 25, 1983. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Moncrieffe:1964:BRB

[Mon64] Anthony Moncrieffe. Book review: The Legacy of Hiroshima by Edward Teller and Allen Brown; Kill and Overkill: The Strategy of Annihilation by Ralph E. Lapp. International Affairs (Royal Institute of International Affairs 1944–), 40(1):99–100, January 1964. CODEN

???? ISSN ???? URL http://www.jstor.org/stable/pdfplus/

2609716.pdf. See [TB62, TB75, Lap62].

Moore:1995:BRG

[Moo95] Mike Moore. Book review: The gadgets made real: Picturing the Bomb: Photographs from the Secret World of the Manhattan Project, by Rachel Fermi and Esther Samra, introduction by Richard Rhodes. Bulletin of the Atomic Scientists, 51(5):65–66, November/December 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (elec-

tronic). URL http://www.highbeam.com/doc/1G1-17459604.html.

See [FS95].

Moore:2000:BRB

[Moo00] Mike Moore. Book review: Atomic Fragments: A Daughter’s Questions, by Mary Palevsky. Bulletin of the Atomic Scientists, 56(9): 74–76, November/December 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-67151470.html.

Moore:2002:TTB

[Moo02] Mike Moore. Twinkle, twinkle: Book review: Sputnik: The Shock of the Century, by Paul Dickson Walker. Bulletin of the Atomic Scientists, 58(3):65–67, May/June 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.

com/content/58/3/65.full.pdf+html;   http://www.highbeam.
com/doc/1G1-86128161.html.

Morgenthau:1962:BBL

[Mor62] Hans J. Morgenthau. Books: The Legacy of Hiroshima, by Edward Teller with Allen Brown. Bulletin of the Atomic Scientists, 18(6):34– 36, June 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See response [Asi62].

Moss:1968:MWPb

[Mos68a] Norman Moss. Men who play God: the story of the H-bomb and how the world came to live with it. Harper & Row, New York, NY, USA, 1968. 352 pp. LCCN UF767 .M68 1968b.

Moss:1968:MWPa

[Mos68b] Norman Moss. Men who play God: the story of the hydrogen bomb. Gollancz, London, UK, 1968. ISBN 0-575-00145-3. 352 pp. LCCN UF767 .M68 1968.

Moss:1972:MWP

[Mos72] Norman Moss. Men who play God: the story of the hydrogen bomb. Penguin Books, Harmondsworth, UK, 1972. ISBN 0-14-003112-X. 382 pp. LCCN UF767 .M68 1972.

Medawar:2001:HGT

[MP01] J. S. Medawar and David Pyke. Hitler’s Gift: the True Story of the Scientists Expelled by the Nazi Regime. Arcade Publishing, New York, NY, USA, 2001. ISBN 1-55970-564-7. xx + 268 pp. LCCN Q141 .M385 2001. Foreword by Max Perutz.

Metropolis:1952:ESC

[MRR^{+}52] Nicholas Constantine Metropolis, Arianna W. Rosenbluth, Marshall N. Rosenbluth, Augusta H. Teller, and Edward Teller. Equation of state calculations by fast computing machines. Report AECU-2435; LADC-1359, Los Alamos Scientific Laboratory, Los Alamos,

NM, USA, 1952. URL http://www.osti.gov/energycitations/
product.biblio.jsp?osti_id=4390578&query_id=0.

Metropolis:1953:ESC

[MRR^{+}53] Nicholas Metropolis, Arianna W. Rosenbluth, Marshall N. Rosenbluth, Augusta H. Teller, and Edward Teller. Equation of state calculations by fast computing machines. Journal of Chemical Physics, 21(6):1087–1092, June 1953. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http:/

/link.aip.org/link/doi/10.1063/1.1699114;   http://pubs.
acs.org/cgi-bin/chemport/version=1.0&coi=1:CAS:528:DyaG3sXltlKhsw%
253D%253D&pissn=0095-2338&pyear=2005&&md5=2eba6ee7d50b361b924c3ff8efeda4b1.

This article introduces the Metropolis algorithm, which the journal Computing in Science and Engineering cited in the top 10 algorithms having the “greatest influence on the development and practice of science and engineering in the 20th Century.” See [BBLS00, BS00], and the Hasting–Metropolis generalization in [Has70]. See also [Bar65, Pes73, Hit03]. According to [Har10, page 263], this paper has been cited more than 10,000 times.

Mark:1995:CLC

[MRTM95] Carson Mark, Louis Rosen, Edward Teller, and Roger Meade. Charles Louis Critchfield [obituary]. Physics Today, 48(2):70–71, February 1995. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://physicstoday.org/resource/1/
phtoad/v48/i2/p70_s2.

Morton-Smith:2015:O

[MS15] Tom Morton-Smith. Oppenheimer. Oberon Books Ltd., London, UK, 2015. ISBN 1-78319-198-8. ???? pp. LCCN

???? URL http://stratford-upon-avon-theatre.blogspot.com/
2015/01/rsc-oppenheimer-review-stratford-upon.html.

Marsh:1980:CFB

[MSDP80] G. E. Marsh, G. S. Stanford, A. De Volpi, and T. A. Postol. Commentary: Fallout from the Progressive H-bomb. Bulletin of the Atomic Scientists, 36(4):64–65, April 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79b].

Mulliken:1942:IMI

[MT42] Robert S. Mulliken and Edward Teller. Interpretation of the methyl iodide absorption bands near λ2000. Physical Review, 61(5–6):283– 296, March 1, 1942. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://prola.aps.org/abstract/PR/
v61/i5-6/p283_1.

McMillan:1947:PMN

[MT47] W. G. McMillan and E. Teller. On the production of mesotrons by nuclear bombardment. Physical Review (2), 72(1):1–6, July 1, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v72/i1/p1_1.

Mayer:1949:OE

[MT49] Maria G. (Maria Goeppert) Mayer and Edward Teller. On the origin of elements. Physical Review (2), 76(8):1226–1231, October 15, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v76/i8/p1226_1.

Mayer:1950:AOE

[MT50] M. G. Mayer and E. Teller. On the abundance and origin of elements. In ????, editor, Les particules élémentaires: Institut International de Chimie Solvay: Huitiéme Conseil de Physique tenu á l’université de Bruxelles du 27 septembre au 2 octobre 1948: Rapports et discussions, pages 59–88. R. Stoop, Bruxelles, Belgium, 1950. LCCN ????

McMillan:1951:ABT

[MT51a] W. G. McMillan and E. Teller. The assumptions of the B.E.T. theory. Journal of Physical Chemistry, 55(1):17–20, January 1951. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

McMillan:1951:RST

[MT51b] W. G. McMillan and E. Teller. The role of surface tension in multilayer gas adsorption. Journal of Chemical Physics, 19(25): 25–32, 1951. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/doi/10.1063/
1.1747984.

More:1987:CS

[MT87] R. More and E. Teller. On the chemistry of superconductivity. Report UCRL-98684, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 1987. Presented at the International Conference on Lasers ’87, December 1987.

Morrison:1994:IHI

[MT94] David Morrison and Edward Teller. The impact hazard: Issues for the future. In Gehrels et al. [GMS94], pages 1135–1143. ISBN 0-8165-1505-0. LCCN QB721 .H29 1994. With the editorial assistance of M.S. Matthews and A.M. Schumann, and 120 collaborating authors.

Moir:2005:TFU

[MT05] Ralph W. Moir and Edward Teller. Thorium-fueled underground power plant based on molten salt technology. Nuclear Technology, 151 (3):334–340, September 2005. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic). URL http://www.ans.org/pubs/

journals/nt/a_3655.

Marshak:1946:DCI

[MTK46] J. Marshak, E. Teller, and L. R. Klein. Dispersal of cities and industries. Bulletin of the Atomic Scientists, 1(9):13–15, 20, April 15, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Mullins:1950:BRB

[Mul50] L. J. Mullins. Book review: The Structure of Matter by Francis Owen Rice and Edward Teller. Quarterly review of biology, 25(1): 104, March 1950. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-

7718 (electronic). URL http://www.jstor.org/stable/pdfplus/

2810232.pdf. See [RT49a].

Mark:1988:EPW

[MW88] Hans Mark and Lowell Wood, editors. Energy in physics, war and peace: a festschrift celebrating Edward Teller’s 80th birthday. Fundamental theories of physics. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1988. ISBN 90-277-2775-9. vi + 403 pp. LCCN QC71 .E62 1988.

Norris:1991:NNP

[NA91] Robert S. Norris and William M. Arkin. Nuclear notebook: Papa Teller says; if they exist, there are 100 of them; weapons watch; Silk Purse abandoned. Bulletin of the Atomic Scientists, 47(7):48, September 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Norris:1993:NNRb

[NA93] Robert S. Norris and William M. Arkin. Nuclear notebook: Russian/ Soviet weapons secrets revealed. Bulletin of the Atomic Scientists, 49(3):48, April 1993. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://www.highbeam.com/doc/1G1-
13610870.html.

Nagy:1989:JNC

[NAH^{+}89] D. Nagy, W. Aspray, P. Horvath, E. Teller, N. Vonneuman, and E. P. Wigner. John von Neumann — a case-study of scientific creativity — discussion. Annals of the History of Computing, 11(3):165–169, ???? 1989. CODEN AHCOE5. ISSN 0164-1239.

NRCCC:1940:TRC

[Nat40] National Research Council (US). Committee on Catalysis, editor. Twelfth report of the Committee on Catalysis, National Research Council. John Wiley, New York, NY, USA, 1940. x + 388 pp. LCCN QD505.N38x 1940.

Nier:1940:NFS

[NBDG40] Alfred O. Nier, E. T. Booth, J. R. Dunning, and A. V. Grosse. Nuclear fission of separated uranium isotopes. Physical Review, 57 (6):546, March 1940. CODEN PHRVAO. ISSN 0031-899X (print),

1536-6065 (electronic). URL http://link.aps.org/doi/10.1103/

PhysRev.57.546. This paper reports experiments that confirm the Bohr–Wheeler prediction [BW39] that U-235 is much more fissile than U-238. That information was critical for the production of both a nuclear reactor, and a nuclear bomb.

Neeman:1998:ETH

[Ne’98] Y. Ne’eman. Edward Teller, the Hungarian, American and Jew, too. Acta Physica Hungarica New Series: Heavy Ion Physics, 7(2):155– 165, ???? 1998. CODEN APHPFM. ISSN 1219-7580 (print), 1588-2675 (electronic).

Staff:2003:HRG

[New03a] Newsline Staff. Hungarian roots grow into strong ties. Newsline (Lawrence Livermore National Laboratory), 28(36):2, September

12, 2003. URL https://newsline.llnl.gov/employee/articles/
2003/09-12-03-newsline.pdf.

Staff:2003:LAE

[New03b] Newsline Staff. LabTV to air Edward Teller films. Newsline (Lawrence Livermore National Laboratory), 28(36):2, September 12, 2003.

URL https://newsline.llnl.gov/employee/articles/2003/09-
12-03-newsline.pdf.

Staff:2003:LTE

[New03c] Newsline Staff. The life and times of Edward Teller. Newsline (Lawrence Livermore National Laboratory), 28(36):4–5, September

12, 2003. URL https://newsline.llnl.gov/employee/articles/
2003/09-12-03-newsline.pdf.

Staff:2003:TLSa

[New03d] Newsline Staff. Teller lived and shaped events of the 20th century. Newsline (Lawrence Livermore National Laboratory), 28(36):1, 8,

September 12, 2003. URL https://newsline.llnl.gov/employee/
articles/2003/09-12-03-newsline.pdf.

Staff:2003:TMS

[New03e] Newsline Staff. Teller made scientific contributions in diverse fields. Newsline (Lawrence Livermore National Laboratory), 28(36):6,

September 12, 2003. URL https://newsline.llnl.gov/employee/
articles/2003/09-12-03-newsline.pdf.

Staff:2003:THD

[New03f] Newsline Staff. Teller’s honorary degrees and awards. Newsline (Lawrence Livermore National Laboratory), 28(36):6, September

12, 2003. URL https://newsline.llnl.gov/employee/articles/
2003/09-12-03-newsline.pdf.

Staff:2003:TLSb

[New03g] Newsline Staff. Teller’s last scientific paper. Newsline (Lawrence Livermore National Laboratory), 28(36):6, September 12, 2003.

URL https://newsline.llnl.gov/employee/articles/2003/09-
12-03-newsline.pdf.

Nichols:1987:RTP

[Nic87] Major General Kenneth D. (Kenneth David) Nichols. The road to Trinity: a personal account of how America’s nuclear policies were made. Morrow, New York, NY, USA, 1987. ISBN 0-688-06910-X. 401 pp. LCCN QC774.N45 A3 1987.

Nemes:1992:ASH

[NK92] Laszlo Nemes and Istvan Kovács, editors. Advances in spectroscopy in Hungary, volume 48A(1A) of Spectrochimica acta. Part A, Molecular spectroscopy. Pergamon Press, Oxford [etc.], 1992. ISSN 0584-8539 (print), 1873-3824 (electronic). 132 pp. LCCN ????

Norris:2003:NP

[NK03] Robert S. Norris and Hans M. Kristensen. Nuclear pursuits. Bulletin of the Atomic Scientists, 59(5):71–72, September/October 2003. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL

http://bos.sagepub.com/content/59/5/71.full.pdf+html.

Munson:1999:TR

[nLM99] Edward Teller nd Lee Munson. Teller reflects. Hoover Digest, 1999 (2):??, April 30, 1999. ISSN 1088-5161. URL http://www.hoover.

org/publications/hoover-digest/article/8009.

Norman:1983:RET

[Nor83] Colin Norman. The reascendancy of Edward Teller (contd.). Science, 220(4602):1132, June 10, 1983. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). Edward Teller was awarded the US National Medal of Science on 24 May 1983.

Norris:2002:RBG

[Nor02] Robert S. (Robert Stan) Norris. Racing for the bomb: General Leslie R. Groves, the Manhattan Project’s indispensable man. Steerforth Press, South Royalton, VT, USA, 2002. ISBN 1-58642-039-9. xxi + 722 pp. LCCN UG128.G76 N67 2002. URL http://www.loc.gov/

catdir/toc/fy036/2001057629.html.

Nordheim:1940:NUA

[NST40] G. Nordheim, H. Sponer, and E. Teller. Note on the ultraviolet absorption systems of benzene vapor. Journal of Chemical Physics, 8 (6):455–458, June 1940. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://link.aip.org/link/doi/10.
1063/1.1750688.

Northrop:1960:SAM

[NT60] Theodore G. Northrop and Edward Teller. Stability of the adiabatic motion of charged particles in the Earth’s field. Physical Review (2), 117(1):215–225, January 1, 1960. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v117/i1/p215_1.

Oliphant:1969:BRB

[Oli69] Thomas A. Oliphant. Book review: Properties of Matter under Unusual Conditions. In Honor of Edward Teller’s 60th Birthday. Hans Mark and Sidney Fernbach, Eds. Interscience (Wiley), New York, 1969. x + 390 pp., illus. $19.50. Science, 166(3910):1261–1262, December 5, 1969. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See [MF69].

Oltmans:1974:IET

[Olt74a] Willem L. Oltmans. Interview with Edward Teller. In On growth [Olt74b], chapter 45, pages 305–312. ISBN 0-399-11233-2.

LCCN HC59 .O563. URL http://catalog.hathitrust.org/api/
volumes/oclc/979019.html.

Oltmans:1974:G

[Olt74b] Willem L. Oltmans, editor. On growth. Capricorn Books, New York, 1974. ISBN 0-399-11233-2. xii + 493 pp. LCCN HC59

.O563.  URL http://catalog.hathitrust.org/api/volumes/
oclc/979019.html.

ONeill:1989:PCH

[O’N89a] Dan O’Neill. Project Chariot: how Alaska escaped nuclear excavation. Bulletin of the Atomic Scientists, 45(10):28–37, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ONeill:1989:SM

[O’N89b] Dan O’Neill. Shooting the moon. Bulletin of the Atomic Scientists, 45 (10):31, December 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

ONeill:1994:FB

[O’N94] Dan (Daniel T.) O’Neill. The Firecracker Boys. St. Martin’s Press, New York, NY, USA, 1994. ISBN 0-312-11183-5, 0-312-11086-3. 388 + 16 pp. LCCN UG1282.A8 O39 1994.

ONeill:2007:FBH

[O’N07] Dan (Daniel T.) O’Neill. The Firecracker Boys: H-bombs, Inupiat Eskimos and the roots of the environmental movement. Basic Books, New York, NY, USA, 2007. ISBN 0-465-00348-6 (paperback). x + 418 + 16 pp. LCCN UG1282.A8 O39 2007. URL

http://www.loc.gov/catdir/enhancements/fy0830/2007030702-
b.html;   http://www.loc.gov/catdir/enhancements/fy0830/
2007030702-d.html;   http://www.loc.gov/catdir/toc/ecip0723/
2007030702.html.

Orear:1958:BBN

[Ore58] Jay Orear. Books: Our Nuclear Future. Facts, Dangers, and Opportunities, by Edward Teller. Bulletin of the Atomic Scientists, 14 (6):235–236, June 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ottaway:1985:SOS

[Ott85] David B. Ottaway. SDI Office seeks funds to speed atomic tests: Program’s ‘non-nuclear’ nature questioned. Washington Post, ??(??): ??, June 9, 1985. ISSN 0190-8286.

Palevsky:2000:AFD

[Pal00] Mary Palevsky. Atomic Fragments: a Daughter’s Questions. University of California Press, Berkeley, CA, USA, 2000. ISBN 0-520-22055-2 (hardcover), 0-250-22055-2 (hardcover). xiv + 289 pp. LCCN

QC773.3.U5 P35 2000. URL http://catdir.loc.gov/catdir/
bios/ucal052/99087422.html;   http://catdir.loc.gov/catdir/
description/ucal042/99087422.html.

Palevsky:2002:TTB

[Pal02] Mary Palevsky. Teller on Teller: Book review: Memoirs: A Twentieth-Century Journey in Science and Politics, by Edward Teller and Judith L. Shoolery. Bulletin of the Atomic Scientists, 58(1):64– 67, January/February 2002. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://bos.sagepub.com/

content/58/1/64.full.pdf+html;   http://www.highbeam.com/
doc/1G1-81891405.html.

Park:1992:SWS

[Par92] R. L. Park. Star warrior on sky patrol: Edward Teller wants to nuke asteroids. New York Times, ??(??):??, March 25, 1992. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Park:1999:ASW

[Par99] Robert L. Park. Another ‘Star Wars’ sequel. New York Times, ??(??): ??, February 15, 1999. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Pawlikowski:2002:ALA

[Paw02] Ellen Pawlikowski. The airborne laser (ABL) program. Defense Horizons, ??(??):4–??, April 2002.

Polanyi:1962:BRB

[PB62] John C. Polanyi and Allen Brown. Book review: The Legacy of Hiroshima by Edward Teller. International Journal, 18(1):98–100, Winter 1962. CODEN ???? ISSN ???? URL http://www.jstor.org/

stable/pdfplus/40198594.pdf. See [TB62, TB75].

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp.

LCCN QC16.O62 P35 2006. URL http://www.loc.gov/catdir/
enhancements/fy0636/2005002173-d.html;   http://www.loc.
gov/catdir/enhancements/fy0723/2005002173-b.html;   http:
//www.loc.gov/catdir/toc/ecip056/2005002173.html.

Peierls:1980:BRE

[Pei80] Sir Rudolf Peierls. Book review: Energy from Heaven and Earth, by Edward Teller. Bulletin of the Atomic Scientists, 36(5):53–54, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Peierls:1987:BRC

[Pei87] Rudolf Peierls. Book review: The case for the defence: Better a Shield Than a Sword: Perspectives on Defense and Technology, by Edward Teller. Free Press, New York/Collier Macmillan, London: 1987, pp.

257. US$19.95, £17.25. Nature, 328(6131):583, August 13, 1987. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Penner:1979:BRB

[Pen79] S. S. Penner. Book review: Energy from Heaven and Earth by Edward Teller. American Scientist, 67(6):710, November 1979. CODEN AM-SCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http:/

/www.jstor.org/stable/pdfplus/27849541.pdf. See [TW79b].

Peskun:1973:OMC

[Pes73] P. H. Peskun. Optimum Monte–Carlo sampling using Markov chains. Biometrika, 60(3):607–612, December 1973. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.

org/stable/2335011.

Piccard:1962:BRB

[Pic62] Paul J. Piccard. Book review: The Legacy of Hiroshima, by Edward Teller and Allen Brown. The Journal of Politics, 24(4):754–756, November 1962. CODEN ???? ISSN 1468-2508. URL http://www.

jstor.org/stable/pdfplus/2128047.pdf. See [TB62, TB75].

Ponnuru:2003:EGE

[Pon03] Ramesh Ponnuru. Evil geniuses, evil fools. National Review, ?? (??):??, September 15, 2003. ISSN 0028-0038. URL http://www.

nationalreview.com/corner/69000/evil-geniuses-evil-fools/
ramesh-ponnuru.

Pool:1989:TCB

[Poo89] Robert Pool. Teller, Chu “Boost” cold fusion. Science, 246(4929): 449, October 27, 1989. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Powell:2004:PAW

[Pow04] Dennis E. Powell. The politics of Armageddon: Why presidents gray. National Review, ??(??):??, March 29, 2004. ISSN 0028-

0038. URL http://www.nationalreview.com/articles/210065/
politics-armageddon/dennis-e-powell.

Preston:1962:BBP

[Pre62] Richard S. Preston. Books: Project Plowshare: The Development of the Peaceful Uses of Nuclear Explosions, by Ralph Sanders; Foreword by Willard F. Libby. Bulletin of the Atomic Scientists, 18(10):25–26, December 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Placzek:1932:RRMa

[PT32a] G. Placzek and E. Teller. Die Rotationsstruktur der Ramanbanden mehratomiger Moleküle. (German) [Rotation structure of the Raman bands of polyatomic molecules]. Zeitschrift für Physik, 81(3–4):209– 258, March 1932. CODEN ZEPYAA. ISSN 0044-3328. URL http:/

/link.springer.com/article/10.1007/BF01338366. See [PT32b].

Placzek:1932:RRMb

[PT32b] G. Placzek and E. Teller. Die Rotationsstruktur der Ramanbanden mehratomiger Moleküle. (German) [Rotation structure of the Raman bands of polyatomic molecules]. Zeitschrift für Physik, 81(11–12):839, November 1932. CODEN ZEPYAA. ISSN 0044-3328. URL http://

link.springer.com/article/10.1007/BF01342084. See [PT32a].

Poschl:1933:BQA

[PT33] G. Pöschl and E. Teller. Bemerkungen zur Quantenmechanik des anharmonischen Oszillators. (German) [Remark on the quantum mechanics of anharmonic oscillators]. Zeitschrift für Physik, 83(3–4): 143–151, March 1933. CODEN ZEPYAA. ISSN 0044-3328. URL

http://link.springer.com/article/10.1007/BF01331132.

Panofsky:1979:DSI

[PT79] W. K. H. Panofsky and Edward Teller. Debate on SALT II. Physics Today, 32(6):32–38, June 1979. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

org/resource/1/phtoad/v32/i6/p32_s1.

Patterson:1981:NES

[PT81] Clair C. Patterson and Edward Teller. A new era of science? Physics Today, 34(12):15, 69, December 1981. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://physicstoday.

org/resource/1/phtoad/v34/i12/p15_s1.

Pupyshev:2005:JTT

[Pup05] Vladimir I. Pupyshev. Jahn–Teller theorem and nodal points of wave functions. International Journal of Quantum Chemistry, 104(2):157– 166, ???? 2005. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Pupyshev:2007:SPJ

[Pup07] Vladimir I. Pupyshev. Simplest proof of the Jahn–Teller theorem for molecular systems. International Journal of Quantum Chemistry, 107 (6):1446–1453, ???? 2007. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Perlin:1984:DJT

[PW84] Yu. E. (Yuri Evgen’evich) Perlin and M. (Max) Wagner, editors. The Dynamical Jahn–Teller effect in localized systems, volume 7 of Modern problems in condensed matter sciences. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-86779-1. x + 908 pp. LCCN QD461 .D94 1984.

Rabinowitch:1956:LAE

[Rab56] Eugene I. Rabinowitch. Letter to art editor. Bulletin of the Atomic Scientists, 12(2):57, February 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). From the introduction: “In my article Ten Years That Changed the World, I mentioned that in 1947, Dr. Teller ‘sketched the now familiar design of the Bulletin cover.’ ”.

Radnai:2004:TEE

[Rad04] Radnai Gyula. Teller Ede az Eötvös-versenyen. (Hungarian) [Edward Teller at the Eötvös competition]. Debreceni Szemle, 12(??):548–560, ???? 2004.

Rao:2003:ET

[Rao03] K. R. Rao. Edward Teller (1908–2003). Current Science, 85(9):1372– 1374, ???? 2003. CODEN CUSCAM. ISSN 0011-3891.

Ronay:2002:TEH

[RB02] Maria Ronay and Hans A. Bethe. Teller ‘exiled’ himself for science. Physics Today, 55(3):11–12, March 2002. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//physicstoday.org/resource/1/phtoad/v55/i3/p11_s2.

Roy-Chaudhury:1987:LAE

[RC87] Rahul Roy-Chaudhury. Letter: Aiming at empty silos. Bulletin of the Atomic Scientists, 43(6):60–61, July/August 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar87].

Redei:2005:JNS

[Réd05] Miklós Rédei. John von Neumann selected letters, volume 27 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2005. ISBN 0-8218-3776-1 (hardcover). ISSN 0899-2428. xxv + 301 pp. LCCN QA29.V66 A4 2005.

Reines:1972:CFOa

[Rei72a] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Colorado Associated University Press, Boulder, CO, USA, 1972. ISBN 0-87081-025-1. xiv + 320 pp. LCCN QC780 .C65.

Reines:1972:CFOb

[Rei72b] Frederick Reines, editor. Cosmology, fusion and other matters: George Gamow memorial volume. Adam Hilger Ltd., Bristol, UK, 1972. ISBN 0-85274-223-1. xiv + 320 pp. LCCN QC780.

URL http://adsabs.harvard.edu/abs/1972cht..conf.....R;
http://adsabs.harvard.edu/abs/1972cofu.book.....R.

Renner:1934:TWZ

[Ren34] Rudolph Renner. Zur Theorie der Wechselwirkung zwischen Elektronen- und Kernbewegung bei dreiatomigen, stabförmigen Molekülen. (German) [On the theory of the interaction between electron and nuclear motion in triatomic, rod-shaped molecules]. Zeitschrift für Physik, 92:172–193, March 1934. CODEN ZEPYAA.

ISSN 0044-3328. URL http://link.springer.com/article/. This

paper, by Edward Teller’s Ph.D. student, is the original description of the Renner–Teller effect (or analysis, or interaction, or split). See [Tel82d] for historical background.

Rheaume:1995:ISB

[Rhe95] C. Rheaume. Internationalism of scientists and the bomb — the case of Robert Oppenheimer, Edward Teller, Igor Kourtchatov, and Andrei Sakharov. Social Science Information = Sur les sciences sociales, 34(2):269–285, ???? 1995. ISSN 0539-0184 (print), 1461-7412 (electronic).

Rheaume:2011:EAS

[Rhé11] Charles Rhéaume. Entretiens avec des scientifiques à propos de Andreï Sakharov. (French) [Interviews with scientists about Andrei Sakharov]. 64 compact disks from Les Presses de l’université Laval, Québec., 2011. Interviews 27 and 28 with Edward Teller, 23 March 1997.

Rhodes:1986:MAB

[Rho86] Richard Rhodes. The Making of the Atomic Bomb. Simon and Schuster, New York, NY, USA, 1986. ISBN 0-671-44133-7 (paperback), 0-671-65719-4 (hardcover). 886 + 42 pp. LCCN QC773 .R46 1986.

Rhodes:1990:ETG

[Rho90] R. Rhodes. Edward Teller — giant of the golden-age of physics — S. A. Blumberg, L. G. Panos. New York Times Book Review, ??(??): 3–??, February 11, 1990. ISSN 0028-7806.

Rhodes:2005:ETR

[Rho05] R. Rhodes. Edward Teller — the real Dr. Strangelove. New York Times Book Review, ??(??):7–8, May 15, 2005. ISSN 0028-7806.

Richelson:2000:SM

[Ric00] Jeffrey T. Richelson. Shootin’ for the moon. Bulletin of the Atomic Scientists, 56(5):22–27, September/October 2000. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

highbeam.com/doc/1G1-65230029.html.

Rival:1996:ASH

[Riv96] Michel Rival. Les apprentis sorciers: Haber, von Braun, Teller. (French) [The sorcerer’s apprentices: Haber, von Braun, Teller]. Science ouverte. Seuil, Paris, France, 1996. ISBN 2-02-021515-2. 234 pp. LCCN Q141 .R48 1996.

Robinson:2004:ET

[Rob04] Peter Robinson. Edward Teller. National Review, ??(??):??, March

11, 2004. ISSN 0028-0038. URL http://www.nationalreview.com/
corner/76740/edward-teller/peter-robinson.

Ropp:1962:BRB

[Rop62] Theodore Ropp. Book review: The Legacy of Hiroshima by Edward Teller and Allen Brown. Annals of the American Academy of Political and Social Science, 343(??):143, September 1962. CODEN ???? ISSN 0002-7162 (print), 1552-3349 (electronic). URL http://www.jstor.

org/stable/pdfplus/1033620.pdf. See [TB62, TB75].

Rose:1971:PJS

[Ros71] Hilary Rose. Pangloss and Jeremiah in science. Nature, 229(5285): 459–462, February 12, 1971. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v229/n5285/pdf/229459a0.pdf.

Ross:1980:BRB

[Ros80] Thomas E. Ross. Book review: Energy from Heaven and Earth by Edward Teller. Economic Geography, 56(1):84–85, January 1980.

CODEN ???? ISSN ???? URL http://www.jstor.org/stable/
pdfplus/143034.pdf. See [TW79b].

Rose:2010:SHE

[Ros10] Steven J. Rose. Surprises in high energy density physics. In Libby and Van Bibber [LV10], pages 96–111. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Ross:2022:IFP

[Ros22] Andrew J. Ross. An icy feud in planetary science: Carl Sagan, Edward Teller, and the ideological roots of the nuclear winter debates, 1980–1984. Historical Studies in the Natural Sciences, 52(2):190–??, April 2022. CODEN ???? ISSN 1939-1811 (print), 1939-182X (elec-

tronic). URL https://online.ucpress.edu/hsns/article/52/
2/190/144553/An-Icy-Feud-in-Planetary-ScienceCarl-Sagan-
Edward.

Rothstein:1994:BSK

[Rot94] Linda Rothstein. Bulletin: The ‘soft kill’ solution. Bulletin of the Atomic Scientists, 50(2):4–6, March/April 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rice:1938:RFR

[RT38] F. O. Rice and E. Teller. The role of free radicals in elementary organic reactions. Journal of Chemical Physics, 6(8):489–496, ???? 1938. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/doi/10.1063/

1.1750298. See corrections [RT39].

Rice:1939:CPR

[RT39] F. O. Rice and E. Teller. Corrections to paper “The Role of Free Radicals in Elementary Organic Reactions”. Journal of Chemical Physics, 7(3):199, 1939. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://link.aip.org/link/doi/10.
1063/1.1750412. See [RT38].

Rice:1949:SM

[RT49a] Francis Owen (Francis Owen) Rice and Edward Teller. The structure of matter. Structure of Matter Series. John Wiley, New York, NY, USA, 1949. xiii + 361 pp. LCCN QC173 .R495.

Richtmyer:1949:OCR

[RT49b] R. D. Richtmyer and Edward Teller. On the origin of cosmic rays. Physical Review, 75(11):1729–1731, June 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v75/i11/p1729_1.

Reid:20xx:ESM

[RTS^{+}xx] R. Reid, Richard Thomas, John Stockbridge, Robert Rietty, James Langahm, and Murray Kash. Einstein: The story of the man by his friends. Motion picture, 20xx. J. Robert Oppenheimer, Thomas Bucky, Helen Dukas, Patrick Gordon Walker, Banesh Hoffman, Otto Nathan, Leo Szilard, Edward Teller and A. R. Ubbelohde.

Rumer:1932:VGB

[RTW32] G. Rumer, E. Teller, and H. Weyl. Eine für die Valenztheorie geeignete Basis der binären Vektorinvarianten. (German) [A basis for binary vector invariants suitable for the valence theory]. Nachr. Ges. Wiss. Göttingen, Math-Physik Klasse, 5(??):499–504, 1932. CODEN

???? ISSN ???? URL http://gdz.sub.uni-goettingen.de/dms/
load/img/?PPN=GDZPPN002508982.

Sagan:1983:NWC

[Sag83] Carl Sagan. Nuclear war and climatic catastrophe: Some policy implications. Foreign Affairs, 62(2):257–292, Winter 1983. CODEN

FRNAA3. ISSN 0003-0554. URL http://www.jstor.org/stable/
20041818.

Sakharov:1995:M

[Sak95] Andrei D. Sakharov, editor. Memoirs. Random House, New York, NY, USA, 1995. ISBN 0-517-13765-8. LCCN ????

Sanders:1977:AOT

[San77] Ralph Sanders. The Advisors: Oppenheimer, Teller and the Superbomb by Herbert York (review). Technology and Culture, 18 (1):129–131, January 1977. CODEN TECUA3. ISSN 0040-165X (print), 1097-3729 (electronic). URL https://muse.jhu.edu/pub/

1/article/891895/pdf.

Sandor:1979:BC

[San79] David Sandor. That Progressive case. Bulletin of the Atomic Scientists, 35(7):64, September 1979. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Fel79b].

Sandoval:2003:PET

[San03] Steve Sandoval. Physicist Edward Teller dies Tuesday in California. Los Alamos News Bulletin, ??(??):??, September 11, 2003.

URL http://www.lanl.gov/orgs/pa/newsbulletin/2003/09/11/
text03.shtml.

Sauberman:1975:LCL

[Sau75] Nat. H. Sauberman. Letters: Credibility lost. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Shepley:1954:HBM

[SB54] James R. Shepley and Clay Blair, Jr. The hydrogen bomb: the men, the menace, the mechanism. Jarrolds, London, UK, 1954. 216 pp. LCCN UF767 .S4 1955. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56], and a longer paper [Tel55b].

Shepley:1955:WKB

[SB55a] James R. Shepley and Clay Blair, Jr. Die Wasserstoffbombe: der Konflikt, die Bedrohung, die Konstruktion. (German) [The hydrogen bomb: the conflict, the threat construction]. Steingrüben Verl., Stuttgart, Germany, 1955. 288 pp. LCCN ???? Translation of [SB54] to German by Hans Dieter Müller. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56].

Shepley:1955:GSA

[SB55b] James R. Shepley and Clay Blair, Jr. La guerre secrète autour de la superbombe: la conception, la compétition, le risque mondial. (French) [The hydrogen bomb: conception, competition, and world risk]. Plon, Paris, France, 1955. x + 272 pp. LCCN ???? Translation of [SB54] to French by Anne Frauger. Preface by Yves Rocard. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56].

Shepley:1971:HBM

[SB71] James R. Shepley and Clay Blair, Jr. The hydrogen bomb: the men, the menace, the mechanism. Jarrolds, London, UK, 1971. ISBN 0-8371-5235-6. vii + 244 pp. LCCN UF767 .S4 1955. This book is seriously flawed; see the lengthy now-declassified rebuttal in [Bet82], as well as the comments in a short review [All56].

Squire:1939:FWC

[SBTT39] C. F. Squire, F. G. Brickwedde, E. Teller, and M. A. Tuve. The Fifth Washington Conference on Theoretical Physics. Science, 89 (2304):180–182, February 24, 1939. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Schroeer:1980:FRB

[Sch80] Dietrich Schroeer. Film reviews: A is for Atom, B is for Bomb: A Portrait of Dr. Edward Teller, produced by Brian Kaufman. Bulletin of the Atomic Scientists, 36(5):55–56, May 1980. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Schaaf:2001:HHB

[Sch01] Michael Schaaf. Heisenberg, Hitler und die Bombe: Gespräche mit Zeitzeugen. (German) [Heisenberg, Hitler and the Bomb: Conversations with witnesses]. Verlag für Geschichte der Naturwissenschaften und der Technik, Berlin, Germany, 2001. ISBN 3-928186-60-4 (paperback). 148 pp. LCCN QC16.H35 S32 2001.

Schweber:2003:BRB

[Sch03a] Silvan Schweber. Book review: Brotherhood of the Bomb: The Tangled Lives and Loyalties of Robert Oppenheimer, Ernest Lawrence, and Edward Teller, by Gregg Herken. Physics Today, 56(5):59–60, May 2003. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://physicstoday.org/resource/1/phtoad/
v56/i5/p59_s1.

Schwoegler:2003:YSL

[Sch03b] David Schwoegler. Young scientist looked to Teller for enlightenment. Newsline (Lawrence Livermore National Laboratory), 28(36):3,

September 12, 2003. URL https://newsline.llnl.gov/employee/
articles/2003/09-12-03-newsline.pdf.

Seaver:2003:CHD

[Sea03a] Lynda Seaver. Commemoration to honor Dr. Edward Teller.

Web site, October 22, 2003. URL https://www.llnl.gov/news/
commemoration-honor-dr-edward-teller.

Seaver:2003:ETL

[Sea03b] Lynda Seaver. Edward Teller (1908–2003): a life dedicated to science. Web document, October 2003. URL https://www.llnl.gov/str/ October03/Teller.html. Dr. Edward Teller, world-renowned physicist, co-founder of Lawrence Livermore National Laboratory, and a lifelong advocate for education, died September 9, 2003. He was 95.

Seaver:2003:ETA

[Sea03c] Lynda Seaver. Edward Teller awarded Presidential Medal of Freedom during White House ceremony. Web site, July 23, 2003. URL https:/

/www.llnl.gov/news/edward-teller-awarded-presidential-medal-
freedom-during-white-house-ceremony.

Seaver:2003:NME

[Sea03d] Lynda Seaver. Nation mourns Edward Teller. Newsline (Lawrence Livermore National Laboratory), 28(36):1, 7, September 12, 2003.

URL https://newsline.llnl.gov/employee/articles/2003/09-

12-03-newsline.pdf. Teller is remembered for his devotion to the advancement of science and education.

Seidel:1990:BRB

[Sei90] Robert Seidel. Book reviews: Books on the bomb. Atomic Bomb Scientists: Memoirs, 1939-1945 by Joseph J. Ermenc. The End of the World That Was: Six Lives in the Atomic Age by Peter Goldman. Manhattan: The Army and the Atomic Bomb by Vincent C. Jones. Day of the Bomb: Countdown to Hiroshima by Dan Kurzman. The General and the Bomb: A Biography of General Leslie R. Groves, Director of the Manhattan Project by William Lawren. Time Bomb:

Fermi, Heisenberg, and the Race for the Atomic Bomb by Malcolm C. MacPherson. The Making of the Atomic Age by Alwyn McKay. The Road to Trinity: A Personal Account of How America’s Nuclear Policies Were Made by K. D. Nichols. The Making of the Atomic Bomb by Richard Rhodes. Stallion Gate by Martin Cruz Smith. The Atomic Scientists: A Biographical History by Henry A. Boorse. Lloyd Motz, and Jefferson Hane Weaver. Forging the Atomic Shield: Excerpts from the Office Diary of Gordon E. Dean by Gordon E. Dean and Roger M. Anders. The Nuclear Oracles: A Political History of the General Advisory Committee of the Atomic Energy Commission, 1947-1977 by Richard T. Sylves. Better a Shield Than a Sword by Edward Teller. Klaus Fuchs, Atom Spy by Robert Chadwell Williams. Justice Downwind: America’s Atomic Testing Program in the 1950s by Howard Ball. The Atomic Papers: A Citizen’s Guide to Selected Books and Articles on the Bomb, the Arms Race, Nuclear Power, the Peace Movement, and Related Issues by Grant Burns. Physics, Technology and the Nuclear Arms Race by D. W. Hafemeister and D. Schroeer. Under the Cloud: The Decades of Nuclear Testing by Richard L. Miller. Bombs in the Backyard: Atomic Testing and American Politics by A. Costandina Titus. Nuclear Fear: A History of Images by Spencer R. Weart. Isis, 81(3):519–537, September 1990. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/233429.

Seifritz:1998:ETC

[Sei98] W. Seifritz. Edward Teller celebrates his ninetieth birthday. Kerntechnik, 63(1–2):24, ???? 1998. CODEN KERNEU. ISSN 0932-3902.

Seidel:1999:GJL

[Sei99] Robert W. Seidel. The Golden Jubilees of Lawrence Berkeley and Los Alamos National Laboratories. Osiris (Series 2), 14(??):187– 202, ???? 1999. CODEN OSIRE3. ISSN 0369-7827 (print), 1933-8287

(electronic). URL http://www.jstor.org/stable/301968.

Selove:1958:BRB

[Sel58] W. Selove. Book review: Our Nuclear Future. Facts, Dangers and Opportunities by Edward Tellerand Albert L. Latter. Science, 127(3305): 1042, May 2, 1958. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/pdfplus/

1755479.pdf. See [TL58b].

Sella:2003:MFE

[Sel03] A. Sella. Many faces of Edward Teller. Chemical and Engineering News, 81(42):4, ???? 2003. CODEN CENEAR. ISSN 0009-2347 (print), 1520-605x (electronic).

Semenov:1992:ETM

[Sem92] A. S. Semenov. Edward Teller Medal — international prize for outstanding work on laser controlled thermonuclear fusion. Soviet Journal of Quantum Electronics, 22(4):377, 1992. CODEN SJQEAF.

ISSN 0049-1748. URL http://stacks.iop.org/0049-1748/22/i=
4/a=N17.

Settle:2002:BRB

[Set02] Frank Settle. Book review: Memoirs: A Twentieth Century Journey in Science and Politics by Edward Teller and Judith Shoolery. The Journal of Military History, 66(2):635–636, April 2002. CODEN ???? ISSN 0899-3718 (print), 1543-7795 (electronic). URL http://www.

jstor.org/stable/pdfplus/3093150.pdf.

Settle:2007:ETR

[Set07] Frank A. Settle. Edward Teller: The real Dr. Strangelove. The Journal of Military History, 71(1):279–281, January 2007. ISSN 0899-3718 (print), 1543-7795 (electronic).

Sewell:1988:BLL

[Sew88] Duane C. Sewell. The branch laboratory at Livermore during the 1950’s. In Mark and Wood [MW88], pages 321–?? ISBN 90-277-2775-9. LCCN QC71 .E62 1988.

Schwarz:1974:LIR

[SH74] Helmut J. Schwarz and Heinrich Hora, editors. Laser interaction and related plasma phenomena. Plenum Press, New York, NY, USA; London, UK, 1974. ISBN 0-306-37143-X, 0-306-37150-2. LCCN ????

Stober:2001:CSW

[SH01] Dan Stober and Ian Hoffman. A convenient spy: Wen Ho Lee and the politics of nuclear espionage. Simon and Schuster, New York, NY, USA, 2001. ISBN 0-7432-2378-0. 384 + 8 pp. LCCN UB271.C62

L47 2001. URL http://www.loc.gov/catdir/bios/simon052/
2001054945.html;   http://www.loc.gov/catdir/description/
simon032/2001054945.html;   http://www.loc.gov/catdir/toc/
fy02/2001054945.html.

Sherwin:1976:ECL

[She76] M. Sherwin. Energy and conflict — life and times of Edward Teller — S. A. Blumberg and G. Owens. New York Times Book Review, 81 (32):1–2, ???? 1976. ISSN 0028-7806.

Shils:1991:RUC

[Shi91] Edward Shils, editor. Remembering the University of Chicago: teachers, scientists, and scholars. University of Chicago Press, Chicago, IL, USA, 1991. ISBN 0-226-75335-2. xxi + 593 pp. LCCN LD920 .R46

1991. URL http://www.loc.gov/catdir/description/uchi051/
91016741.html;   http://www.loc.gov/catdir/enhancements/
fy0608/91016741-t.html.

Shore:1992:LON

[Sho92] Lloyd G. Shore. Letter: Outlawing nuclear weapons. Bulletin of the Atomic Scientists, 48(2):47–48, March 1992. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel91k].

Shoolery:2001:ETJ

[Sho01] J. L. Shoolery. Edward Teller’s journey. New York Times Book Review, ??(??):4, December 23, 2001. ISSN 0028-7806.

Segre:1969:GMP

[SKST69] Emilio G. Segrè, Joseph Kaplan, Leonard I. Schiff, and Edward Teller. Great men of physics: the humanistic element in scientific work. The University of California letters and science extension series. Tinnon-Brown, Los Angeles, CA, USA, 1969. vi + 110 pp. LCCN ???? With an introduction by Marvin L. Chachere.

Slayton:2012:JET

[Sla12] Rebecca Slayton. Judging Edward Teller: a closer look at one of the most influential scientists of the Twentieth Century. Technology and Culture, 53(2):509–510, April 2012. CODEN TECUA3. ISSN 0040-165x (print), 1097-3729 (electronic). URL http://muse.jhu.edu/

journals/technology_and_culture/v053/53.2.slayton.pdf.

Sobel:1987:LPM

[SLM87] Michael I. Sobel, Steven R. Lampman, and Gerald E. Marsh. Letters: Protecting missiles with SDI. Bulletin of the Atomic Scientists, 43 (4):54–, May 1, 1987. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Mar87].

Solomon:1986:MIN

[SM86] Fredric Solomon and Robert Q. Marston, editors. The Medical implications of Nuclear War. National Academy Press, Washington, DC, USA, 1986. ISBN 0-309-03636-4 (paperback), 0-309-03692-5 (hardcover). LCCN RA648.3 .M445 1986.

Smith:1965:PHS

[Smi65] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. University of Chicago Press, Chicago, IL, USA, 1965. xiv + 591 pp. LCCN Q127.U6 S6.

Smith:1970:PHS

[Smi70] Alice Kimball Smith. A peril and a hope; the scientists’ movement in America, 1945–47. MIT Press, Cambridge, MA, USA, 1970. ISBN 0-262-69026-8. xvi + 398 pp. LCCN Q127.U6 S6 1971.

Smith:1975:LWN

[Smi75] Anthony Wayne Smith. Letters: We need restraint. Bulletin of the Atomic Scientists, 31(6):5, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BAA^{+}75].

Smith:1985:ECD

[Smi85a] R. J. Smith. Experts cast doubts on X-ray laser: The jewel of the ‘Star Wars’ missile defense program fails to glitter. Science, 230 (4726):646, November 8, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.org/

content/230/4726/646.extract.

Smith:1985:LOS

[Smi85b] R. Jeffrey Smith. Lab officials squabble over X-ray laser: Edward Teller lobbies for $100 million in new research funds, but others say it isn’t warranted. Science, 230(4728):923, November 22, 1985. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/230/4728/923.citation.

Smith:1986:SG

[Smi86] Martin Cruz Smith. Stallion Gate. Random House, New York, NY, USA, 1986. ISBN 0-394-53006-3. 321 pp. LCCN PS3569.M5377 S7

1986. US$17.95. URL http://www.loc.gov/catdir/description/
random0412/85024444.html.

Smiley:2010:MWI

[Smi10] Jane Smiley. The man who invented the computer: the biography of John Atanasoff, digital pioneer. Doubleday, Garden City, NY, USA, 2010. ISBN 0-385-52713-6, 0-385-53372-1 (e-book), 1-299-11995-6 (ebook). 246 + 8 pp. LCCN QA76.2.A75 S64 2010. US$25.95.

Szilard:1960:SWS

[SMM^{+}60] Leo Szilard, Thomas Edward Murray, John B. (John Bruce) Medaris, Edward Teller, and Jerome B. (Jerome Bert) Wiesner, editors. Small World: Szilard, Murray, Medaris, Teller, Wiesner. Columbia Broadcasting System, ????, 1960. LCCN ???? 28-minutes 16mm black-and-white film.

Sponer:1939:ANU

[SNST39] H. Sponer, G. Nordheim, A. L. Sklar, and E. Teller. Analysis of the near ultraviolet electronic transition of benzene. Journal of Chemical Physics, 7(4):207–220, ???? 1939. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.aip.org/link/

doi/10.1063/1.1750419.

Sahlin:1971:EPC

[SST71] H. Sahlin, R. Speed, and E. Teller. An effective potential calculation for the ground state of excited states of H_{2}. Report UCRL-73338, University of California, Lawrence Radiation Laboratory, Livermore, CA, USA, April 1971. Submitted to the Journal of Chemical Physics.

Schwinger:1937:LES

[ST37a] Julian Schwinger and E. Teller. Letter to editor: The scattering of neutrons by ortho- and para-hydrogen. Physical Review, 51(9):775, May 1, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://prola.aps.org/abstract/PR/v51/
i9/p775_1.

Schwinger:1937:SNO

[ST37b] Julian Schwinger and E. Teller. The scattering of neutrons by orthoand parahydrogen. Physical Review, 52(4):286–295, August 15, 1937. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v52/i4/p286_1.

Seeger:1938:EBA

[ST38] R. J. Seeger and E. Teller. On the electrical breakdown of the alkali halides. Physical Review, 54(7):515–519, October 1, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v54/i7/p515_1.

Seeger:1939:RDB

[ST39a] R. J. Seeger and E. Teller. Remarks on the dielectric breakdown. Physical Review (2), 56(4):352–354, August 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v56/i4/p352_1.

Sponer:1939:LEA

[ST39b] H. Sponer and E. Teller. Letter to the editor: On the application of the Franck–Condon principle to the absorption spectrum of HgCl_{2}. Journal of Chemical Physics, 7(5):382, 1939. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.aip.

org/link/doi/10.1063/1.1750454.

Seeger:1940:LEE

[ST40] R. J. Seeger and E. Teller. Letter to editor: Electric breakdown of alkali halides. Physical Review, 58(3):279–280, August 1, 1940. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v58/i3/p279_1.

Sachs:1941:SSN

[ST41a] R. G. Sachs and E. Teller. The scattering of slow neutrons by molecular gases. Physical Review (2), 60(1):18–27, July 1, 1941. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://prola.aps.org/abstract/PR/v60/i1/p18_1.

Sponer:1941:ESP

[ST41b] H. Sponer and E. Teller. Electronic spectra of polyatomic molecules. Reviews of Modern Physics, 13(2):75–170, April 1, 1941. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Seeger:1942:ISN

[ST42] R. J. Seeger and E. Teller. On the inelastic scattering of neutrons by crystal lattices. Physical Review, 62(1–2):37–40, July 1, 1942. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://prola.aps.org/abstract/PR/v62/i1-2/p37_1.

Szilard:1960:DDP

[ST60] Leo Szilard and Edward Teller. Debate: Is disarmament possible and desirable? NBC television broadcast., 1960.

Shaich:1981:FE

[ST81] Harry Shaich and Edward Teller. February editorial. Physics Today, 34(7):76, July 1981. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.org/resource/
1/phtoad/v34/i7/p76_s1. See [Tel81h].

Stix:1999:PIH

[ST99] Gary Stix and Edward Teller. Profile: Infamy and honor at the Atomic Café: Edward Teller has no regrets about his contentious career. Scientific American, 281(4):42–44, October 1999. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL

http://www.nature.com/scientificamerican/journal/v281/n4/
pdf/scientificamerican1099-42.pdf.

Sterba:1973:HBB

[Ste73] James P. Sterba. H-bombs blast to free gas. New York Times, ?? (??):??, May 20, 1973. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Sterba:1974:OSL

[Ste74] James P. Sterba. Oil shale: Leasing to begin on vast reserves in West. New York Times, ??(??):??, January 6, 1974. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Stix:2003:ET

[Sti03] Gary Stix. Edward Teller, 1908–2003. Scientific American, 289 (5):32, November 2003. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v289/n5/pdf/scientificamerican1103-
32b.pdf.

Stoicheff:2002:GHI

[Sto02] B. P. (Boris Peter) Stoicheff. Gerhard Herzberg: an illustrious life in science. NRC press biography series. NRC Press, Ottawa, ON, Canada, 2002. ISBN 0-660-18757-4, 0-660-19258-6 (e-book). ISSN 1701-1833. xiii + 468 pp. LCCN QC16.H47 S86 2002.

Strauss:1963:MDa

[Str63] Lewis L. Strauss. Men and decisions. Doubleday, Garden City, NY, USA, 1963. viii + 468 + 16 pp. LCCN E741 .S78.

Sullivan:1975:EDV

[Sul75] Walter Sullivan. Experts doubt view that atom blast could end all life. New York Times, ??(??):??, November 23, 1975. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:

//search.proquest.com/hnpnewyorktimes/docview/120557618/.

This story comments on the recently-declassified report [Bre53] on whether a hydrogen bomb explosion could ignite the atmosphere or oceans, or blow the atmosphere into space. See also [Dud75].

Sullivan:2003:ETD

[Sul03] Walter Sullivan. Edward Teller is dead at 95; fierce architect of hbomb. New York Times, ??(??):??, September 10, 2003. CODEN NY-TIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http:/

/www.nytimes.com/2003/09/10/obituaries/10TELL.html.

Sylves:1987:NOP

[Syl87] Richard Terry Sylves. The nuclear oracles: a political history of the General Advisory Committee of the Atomic Energy Commission, 1947–1977. Iowa State University Press, Ames, 1987. ISBN 0-8138-0062-5. xviii + 319 + 16 pp. LCCN QC773.3.U5 S95 1987.

Szilard:1987:DSA

[Szi87] Leo Szilard. Draft of a statement about Edward Teller (August 23, 1963). In Toward a livable world: Leo Szilard and the crusade for nuclear arms control [HGS87], pages 405–406. ISBN 0-262-19260-8, 0-262-08162-8. LCCN QC3 .S97 vol. 3.

Tarter:1998:HBE

[Tar98] Bruce Tarter. Happy birthday, edward!, July 1998. URL https://

www.llnl.gov/str/Com798.html.

Tarter:2010:BRE

[Tar10] C. Bruce Tarter. Brief reflections on Edward Teller’s scientific life at Livermore. In Libby and Van Bibber [LV10], pages 1–3. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Tarter:2018:ALI

[Tar18] C. Bruce Tarter. The American Lab: An Insider’s History of the Lawrence Livermore National Laboratory. Johns Hopkins Nuclear History and Contemporary Affairs. The Johns Hopkins University Press, Baltimore, MD, USA, 2018. ISBN 1-4214-2531-9 (hardcover), 1-4214-2532-7 (e-book). ix + 453 pp. LCCN U394.L58 T37 2018.

Tauscher:2003:TTT

[Tau03] Ellen Tauscher. Tauscher tribute to Teller in U.S. House of Representatives. Newsline (Lawrence Livermore National Laboratory), 28(36):7, September 12, 2003. URL https://newsline.llnl.gov/

employee/articles/2003/09-12-03-newsline.pdf.

Taylor:1996:GOC

[Tay96] Theodore B. Taylor. Guest opinion: Circles of destruction. Bulletin of the Atomic Scientists, 52(1):4, January/February 1996. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.highbeam.com/doc/1G1-17773681.html.

Teller:1962:LH

[TB62] Edward Teller and Allen Brown. The legacy of Hiroshima. Doubleday, Garden City, NY, 1962. 325 pp. LCCN UF767 .T4 (LC); UF767 .T4.

Teller:1975:LH

[TB75] Edward Teller and Allen Brown. The legacy of Hiroshima. Greenwood Press, 88 Post Road West, Westport, CT 06881, USA, 1975. ISBN 0-8371-8344-8. ix + 325 pp. LCCN UF767 .T274 1975. URL http:/

/alsos.wlu.edu/information.aspx?id=142.

Teller:1982:HBH

[TB82] Edward Teller and Hans A. Bethe. Hydrogen bomb history. Science, 218(4579):1270, December 24, 1982. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/ stable/1690175. Response to [Bro82].

Teller:1984:FND

[TB84] Edward Teller and Hans Bethe. Face-off on nuclear defense. Technology Review (M.I.T.), 84(3):??, ???? 1984. CODEN TEREAU. ISSN 0040-1692.

Teller:2002:HSU

[TB02] Teller Ede and Bekk Mária. Huszadik századi utazás: Tudományban és politikában. (Hungarian) [Twentieth Century Travel: Science and politics]. Huszadik Század Intézet, Budapest, Hungary, 2002. ISBN 963-9406-55-4. 595 pp. LCCN QC16.T37 M5516 2001.

Teller:1951:EDT

[TdHK51] Edward Teller, Frederic de Hoffmann, and Arnold Kramish. An estimate of [deleted] temperatures. Report ????, Los Alamos Scientific Laboratory of the University of California, Los Alamos, NM, USA, February 4, 1951.

Teller:1930:BTF

[Tel30a] E. Teller. Bemerkung zur Theorie des Ferromagnetismus. (German) [Remark on the theory of ferromagnetism]. Zeitschrift für Physik, 62 (1–2):102–105, January 1930. CODEN ZEPYAA. ISSN 0044-3328.

URL http://link.springer.com/article/10.1007/BF01340408.

Teller:1930:WGH

[Tel30b] Eduard Teller. Über das Wasserstoffmolekülion. (German) [On the hydrogen molecule ion]. Zeitschrift für Physik, 61(7–8):458–480, July 1930. CODEN ZEPYAA. ISSN 0939-7922 (print), 1431-5831

(electronic). URL http://link.springer.com/article/10.1007/

BF01330302. Submitted 6 February 1930 as Ph.D. dissertation at Universität Leipzig, under the direction of Professor Werner Heisenberg. Library registration 7 March 1930. Oral examination on 28 February 1930. Doctoral certificate awarded 19 May 1930.

Teller:1931:DFE

[Tel31] E. Teller. Der Diamagnetismus von freien Elektronen. (German) [Diamagnetism of free electrons]. Zeitschrift für Physik, 67(5–6):311–319, January 31, 1931. CODEN ZEPYAA. ISSN 0044-3328. URL http:

//link.springer.com/article/10.1007/BF01392167.

Teller:1934:TKG

[Tel34a] E. Teller. Theorie der Kirstallgitterspektren. (German) [Theory of crystal lattice spectra]. In Finkelnburg et al. [FMRT34], chapter 3, pages 161–188. LCCN QD453 .E72 bd. 9, abschnitt 2.

Teller:1934:TLM

[Tel34b] E. Teller. Theorie der langwelligen Molekülspektren. (German) [Theory of long-wave spectra of molecules]. In Finkelnburg et al. [FMRT34], chapter 2, pages 43–160. LCCN QD453 .E72 bd. 9, abschnitt 2.

Teller:1937:CPS

[Tel37] E. Teller. The crossing of potential surfaces. Journal of Physical Chemistry, 41(1):109–116, January 1937. CODEN JPCHAX. ISSN 0022-3654 (print), 1541-5740 (electronic).

Teller:1941:AVE

[Tel41] E. Teller. Asymmetric vibrations excited by an electronic transition. Annals of the New York Academy of Sciences, XLI(3):173–186, June 1941. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Teller:1946:PFP

[Tel46a] Edward Teller. Prima facie proof of the super. Classified report LA-551, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, April 15, 1946.

Teller:1946:SWP

[Tel46b] Edward Teller. Scientists in war and peace. Bulletin of the Atomic Scientists of Chicago, 1(6):10–11, March 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1946:SDR

[Tel46c] Edward Teller. The State Dep’t report — ‘a ray of hope’. Bulletin of the Atomic Scientists, 1(8):10, 13, April 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1946:SAA

[Tel46d] Edward Teller. A suggested amendment to the Acheson Report. Bulletin of the Atomic Scientists, 1(12):5, June 1, 1946. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1947:AST

[Tel47a] Edward Teller. The atomic scientists have two responsibilities. Bulletin of the Atomic Scientists, 3(12):355–356, December 1947. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Tel63c].

Teller:1947:HEM

[Tel47b] Edward Teller. The high-energy machines: their design and application. Bulletin of the Atomic Scientists, 3(7):171–172, July 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1947:HDA

[Tel47c] Edward Teller. How dangerous are atomic weapons? Bulletin of the Atomic Scientists, 3(2):35–36, February 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1947:PDF

[Tel47d] Edward Teller. Politics, domestic and foreign. Bulletin of the Atomic Scientists, 3(3):100, March 1947. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1948:CDW

[Tel48a] Edward Teller. Comments on the “Draft of a World Constitution”. Bulletin of the Atomic Scientists, 4(7):204, July 1948. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Com48].

Teller:1948:FYA

[Tel48b] Edward Teller. The first year of the Atomic Energy Commission. Bulletin of the Atomic Scientists, 4(1):5–6, January 1948. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1948:CPC

[Tel48c] Edward Teller. On the change of physical constants. Physical Review, 73(7):801–802, April 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v73/i7/p801_1.

Teller:1949:BRL

[Tel49a] Edward Teller. Book review: Lang’s Early Tales of the Atomic Age, by Daniel Lang, New York: Doubleday, 1948. Bulletin of the Atomic Scientists, 5(2):64, February 1949. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1949:OCR

[Tel49b] Edward Teller. The origin of cosmic radiation. Physics Today, 2 (8):6–13, August 1949. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.org/resource/
1/phtoad/v2/i8/p6_s1.

Teller:1949:SM

[Tel49c] Edward Teller. The Structure of Matter. ????, ????, 1949. ?? pp. LCCN ????

Teller:1950:BL

[Tel50a] Edward Teller. Back to the laboratories. Bulletin of the Atomic Scientists, 6(3):71–72, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1950:DTB

[Tel50b] Edward Teller. On the development of thermonuclear bombs. Classified report LA-643, Los Alamos Scientific Laboratory, Los Alamos, NM, USA, February 1950. 16 pp.

Teller:1951:NTD

[Tel51a] Edward Teller. A new thermonuclear device. Report LAMS-1230, Los Alamos Scientific Laboratory of the University of California, Los Alamos, NM, USA, April 4, 1951. This report is based on work by Frederic de Hoffmann, and was likely largely written by him, but at his request, was issued under Teller’s name alone. See [Tel55b, page 273] and [Har10, page 221].

Teller:1951:S

[Tel51b] Edward Teller. To Sagittarius. Bulletin of the Atomic Scientists, 7 (1):22, January 1951. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1954:CPS

[Tel54a] E. Teller. Comments on plasma stability and on a constant-pressure thermonuclear reactor. In ????, editor, Conference on Thermonuclear Reactions, Princeton University, October 26–27, 1954, page ?? ????, ????, 1954. LCCN ???? Declassified January 23, 1959.

Teller:1954:LAA

[Tel54b] Edward Teller. Letter: The accomplishments of Los Alamos. Bulletin of the Atomic Scientists, 10(9):359, November 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1954:TOC

[Tel54c] Edward Teller. Theory of origin of cosmic rays. Reports on Progress in Physics, 17(1):154–172, ???? 1954. CODEN RPPHAG. ISSN 0034-4885 (print), 1361-6633 (electronic).

Teller:1955:ORS

[Tel55a] E. Teller. Opening remarks — symposium on plasma instabilities. In ????, editor, Conference on Thermonuclear Reactions, University of California Radiation Laboratory, February 1955, page ?? ????, ????, 1955. LCCN ???? Declassified January 23, 1959.

Teller:1955:WMP

[Tel55b] Edward Teller. The work of many people. Science, 121(3139):267–275, February 25, 1955. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/1681529.

Teller:1956:GPCa

[Tel56a] E. Teller. General problems of the controlled thermonuclear process. Nuclear Science and Engineering, 1(??):313–324, 1956. CODEN NSE-NAO. ISSN 0029-5639 (print), 1943-748x (electronic).

Teller:1956:GPCb

[Tel56b] E. Teller. General problems of the controlled thermonuclear process. In ????, editor, Conference on Controlled Thermonuclear Reactions, June 1956, page ?? ????, ????, 1956. LCCN ???? Declassified January 23, 1959.

Teller:1956:AW

[Tel56c] Edward Teller. The atom at work. New York Times, ??(??):??, December 9, 1956. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1956:CTR

[Tel56d] Edward Teller. Controlled thermonuclear reactions. Bulletin of the Atomic Scientists, 12(7):271–272, September 1956. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1957:DKP

[Tel57a] Edward Teller. Defense: Knowledge is power. Time Magazine, ?? (??):??, November 18, 1957. CODEN TYMEA9. ISSN 0040-781x (print), 2169-1665 (electronic). URL http://content.time.com/

time/magazine/article/0,9171,868002,00.html.

Teller:1957:JND

[Tel57b] Edward Teller. John von Neumann [28 December 1903–8 February 1957]. Bulletin of the Atomic Scientists, 13(4):150–151, April 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1957:NNWa

[Tel57c] Edward Teller. The nature of nuclear warfare. Air Force magazine, 40(1):??, January 1957. CODEN AFORCO. ISSN 0730-6784 (print), 1943-4782 (electronic). URL http://www.

airforce-magazine.com/MagazineArchive/Pages/1957/January%
201957/0157nuclear.aspx.

Teller:1957:NNWb

[Tel57d] Edward Teller. The nature of nuclear warfare. Bulletin of the Atomic Scientists, 13(5):162–165, May 1957. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted from Air Force Magazine, January 1957. See comment [Fer57].

Teller:1957:NLW

[Tel57e] Edward Teller. A new look at war-making. New York Times, ?? (??):??, July 7, 1957. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1957:SNU

[Tel57f] Edward Teller. The size and nature of the universe. The theory of relativity. General Dynamics Corporation, New York, NY, USA, 1957. LCCN ???? Phonodisc.

Teller:1958:PUF

[Tel58a] E. Teller. Peaceful uses of fusion. In Longmire [Lon58], pages 56–65. ISSN 0555-4136. LCCN QC791 .P5.

Teller:1958:AS

[Tel58b] Edward Teller. Alternatives for security. Foreign Affairs, 36 (2):201–208, January 1958. CODEN FRNAA3. ISSN 0015-

7120. URL http://www.jstor.org/stable/20029276;   http://
www.jstor.org/stable/pdfplus/20029276.pdf.

Teller:1958:CNN

[Tel58c] Edward Teller. The compelling need for nuclear tests. Life Magazine, ??(??):??, ???? 1958.

Teller:1958:RC

[Tel58d] Edward Teller. The Russian challenge. Bulletin of the Atomic Scientists, 14(2):87–89, February 1958. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1959:PUA

[Tel59] E. Teller. Peaceful uses of atomic explosives. In ????, editor, Fifth World Petroleum Congress Proceedings, Symposium on the Applications of Atomic Energy to the Petroleum Industry, New York, pages 11–20. ????, ????, 1959. LCCN ????

Teller:1960:BCP

[Tel60a] Edward Teller. Basic concepts of physics. California Book Co., Berkeley, 1960. 199 pp. LCCN QC23 .T4.

Teller:1960:FAC

[Tel60b] Edward Teller. The feasibility of arms control and the principle of openness. Dædalus, 89(4):781–799, Fall 1960. CODEN DAEDAU. ISSN 0011-5266 (print), 1548-6192 (electronic). URL http://www.

jstor.org/stable/20026615.

Teller:1960:GP

[Tel60c] Edward Teller. The goal of peace. Air Force magazine, 43(6):90, 100– 101, 103, June 1960. CODEN AFORCO. ISSN 0730-6784 (print),

1943-4782 (electronic). URL http://www.airforce-magazine.com/
MagazineArchive/Pages/1960/June%201960/0660peace.aspx.

Teller:1960:IP

[Tel60d] Edward Teller. The issue of peace. Bulletin of the Atomic Scientists, 16(6):201–204, June 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1960:PIS

[Tel60e] Edward Teller. Perilous illusion: Secrecy means security. New York Times, ??(??):??, November 13, 1960. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1961:QME

[Tel61a] E. Teller. Der quantenmechanische Meßprozess und die Entropie. (German) [The quantum-mechanical measurement process and entropy]. In ????, editor, Werner Heisenberg und die Physik unserer Zeit, pages 90–92. ????, ????, 1961. LCCN ????

Teller:1961:NPP

[Tel61b] E. Teller. Nuclear power potential for space vehicles. In ????, editor, Proceedings of the Bureau of Naval Weapons Missiles and Rockets Symposium, Concord, CA, pages 15–17. ????, ????, April 1961. LCCN ????

Teller:1961:LET

[Tel61c] Edward Teller. Letter to the Editor: Teller advocates testing: Physicist says Soviet explosions make our effort essential. New York Times, ??(??):??, November 12, 1961. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1962:AOP

[Tel62a] Edward Teller. An address on Operation Plowshare. The Journal of the Arkansas Medical Society, 58(??):538–541, May 1962.

Teller:1962:HRa

[Tel62b] Edward Teller. The hazards of radiation. Journal of Nuclear Medicine, 3(5):1–9, January 1962. CODEN JNMEAQ. ISSN 0161-5505 (print), 1535-5667 (electronic).

Teller:1962:HRb

[Tel62c] Edward Teller. The hazards of radiation. California Medicine, 97(5): 257–262, ???? 1962.

Teller:1962:HRc

[Tel62d] Edward Teller. The hazards of radiation. New Physician, 11(??): 82–87, ???? 1962.

Teller:1962:HBP

[Tel62e] Edward Teller. The hydrogen bomb project. In Fremont E. Kast and James E. Rosenzweig, editors, Science, Technology, and Management, pages 41–48. McGraw-Hill, New York, NY, USA, 1962. LCCN ????

URL http://alsos.wlu.edu/information.aspx?id=1249.

Teller:1962:I

[Tel62f] Edward Teller. Introduction. In Now it can be told: the story of the Manhattan Project [Gro62], pages iii–ix. LCCN QC773.A1 G7.

Teller:1962:SRH

[Tel62g] Edward Teller. On the speed of reactions at high pressures. Journal of Chemical Physics, 36(4):901–903, 1962. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.aip.

org/link/doi/10.1063/1.1732687.

Teller:1962:SMT

[Tel62h] Edward Teller. On the stability of molecules in the Thomas–Fermi theory. Reviews of Modern Physics, 34(4):627–631, October 1, 1962.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

Teller:1963:Pa

[Tel63a] E. Teller. Plowshare. Report UCRL-7222, University of California Lawrence Radiation Laboratory, Berkeley, CA, USA, February 4, 1963.

Teller:1963:Pb

[Tel63b] E. Teller. Plowshare. Nuclear News, 6(??):1–13, March 1963. CODEN NUNWA8. ISSN 0029-5574.

Teller:1963:TRS

[Tel63c] Edward Teller. The two responsibilities of scientists. In Grodzins and Rabinowitch [GR63], pages 121–124. LCCN D842 .B78. Reprint of [Tel47a].

Teller:1964:PEN

[Tel64a] Edward Teller. Physical effects of nuclear attack. The North American Review, 249(3):11–14, Autumn 1964. CODEN ???? ISSN 0029-2397.

URL http://www.jstor.org/stable/pdfplus/25115991.pdf.

Teller:1964:RR

[Tel64b] Edward Teller. The Reluctant Revolutionary. The Paul Anthony Brick lectures, 3d ser. University of Missouri Press, Columbia, MO, USA, 1964. xi + 71 pp. LCCN Q147 .T4. Foreword by A. Cornelius Benjamin.

Teller:1965:EON

[Tel65a] E. Teller. Energy from oil and from the nucleus. Journal of Petroleum Technology, ??(??):505–508, May 1965. CODEN JPTCAZ. ISSN 0022-3522.

Teller:1965:PNA

[Tel65b] E. Teller. Progress in the nuclear age. Mayo Clinic Proceedings, 40(??):47–55, January 1965. CODEN MACPAJ. ISSN 0025-6196 (print), 1942-5546 (electronic).

Teller:1965:SMP

[Tel65c] Edward Teller. Die Situation der modernen Physik. (German) [The situation in modern physics], volume 147 of Natur-, Ingenieur- und Gesellschaftswissenschaften. Westdeutscher Verlag, Köln, 1965. 37 pp. LCCN Q49.C95 A8 Heft 147.

Teller:1965:HHH

[Tel65d] Teller Ede. Hiroshima hagyatéa. (Hungarian) [Legacy of Hiroshima]. Kossuth Kiadó, Budapest, Hungary, 1965. 217 pp. LCCN ????

Teller:1966:TQ

[Tel66a] E. Teller. On a theory of quasars. In R. E. Marshak, editor, Perspectives in Modern Physics, Essays in Honor of Hans A. Bethe, pages 449–462. Interscience Publishers, New York, NY, USA, 1966. LCCN ????

Teller:1966:RAS

[Tel66b] Edward Teller. The role of applied science. Bulletin of the Atomic Scientists, 22(3):15–19, March 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Teller:1967:ACL

[Tel67a] E. Teller. On the adiabatic compression of a light packet. Science of Light (Tokyo), 16(1):22–25, May 1967. CODEN SCLIAL. ISSN 0080-7583.

Teller:1967:EPA

[Tel67b] Edward Teller. The evolution and prospects for applied physical science in the United States. In Applied Science and Technological Progress, A Report to the Committee on Science and Astronautics, U. S. House of Representatives, pages 365–397. National Academy of Sciences, Washington, DC, USA, June 1967. LCCN ???? URL

http://books.google.com/books?id=BTcrAAAAYAAJ.

Teller:1968:PSS

[Tel68a] E. Teller. Prehistoric safety studies. In ????, editor, Proceedings of the Livermore Array Symposium, Livermore, CA, pages 1–5. ????, ????, 1968. LCCN ????

Teller:1968:SRT

[Tel68b] E. Teller. Some remarks on time reversal. In ????, editor, Proceedings of Symmetry Principles at High Energy, Fifth Coral Gables Conference, University of Miami, pages 1–6. ????, ????, 1968. LCCN ????

Teller:1968:AES

[Tel68c] Edward Teller. Atomic explosives: Solved and unsolved problems. The Physics Teacher, 6(5):207, May 1968. CODEN PHTEAH. ISSN 0031-921X (print), 1943-4928 (electronic).

Teller:1968:SAD

[Tel68d] Edward Teller. Should APS discuss public issues: Ignorance and pressure groups. Physics Today, 21(1):17, 19, January 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://physicstoday.org/resource/1/phtoad/v21/i1/p17_s2.

Teller:1969:ICP

[Tel69a] E. Teller. Internal conversion in polyatomic molecules. Israel Journal of Chemistry, 7(2):227–235, ???? 1969. CODEN ISJCAT. ISSN 0021-2148.

Teller:1969:TE

[Tel69b] E. Teller. Thermonuclear energy. Report CTS-HS-69-2, Center for Theoretical Studies, University of Miami, Miami, FL, USA, ???? 1969.

Teller:1969:EW

[Tel69c] Edward Teller. Exotic weapons. Survive, 2(3):3–4, 1969. ISSN 0039-

6354. URL http://nektar.oszk.hu/en/manifestation/1466163.

Teller:1969:NBI

[Tel69d] Edward Teller. Niels Bohr and the idea of complementarity. In Great men of physics: the humanistic element in scientific work [SKST69], pages 76–97. LCCN ???? With an introduction by Marvin L. Chachere.

Teller:1970:CPC

[Tel70a] Edward Teller. Can a progressive be a conservationist? New Scientist, 45(??):346–349, February 19, 1970. CODEN NWSCAL. ISSN 0262-4079, 0028-6664.

Teller:1970:URN

[Tel70b] Edward Teller. University role in nuclear explosives engineering research. In ????, editor, Education for Peaceful Uses of Nuclear Explosives, pages 293–298. University of Arizona Press, Tucson, AZ, USA, 1970. ISBN ???? LCCN ????

Teller:1971:VWG

[Tel71a] E. Teller. Die Verantwortlichkeit des Wissenschaftlers in der Gesellschaft. (German) [The responsibility of the scientist in society]. In Dürr and Bagge [DB71], pages 93–100. ISBN 3-528-08317-4.

LCCN ???? URL http://www.springer.com/physics/book/978-
3-528-08317-5.

Teller:1971:RHS

[Tel71b] E. Teller. Relativistic hydrodynamics in supernovae. In Caldirola and Knoepfel [CK71], pages 402–418. ISBN 0-12-368848-5. LCCN QC281 .V28 1971.

Teller:1971:STA

[Tel71c] E. Teller. Some thoughts about high energy densities. In Caldirola and Knoepfel [CK71], pages 1–6. ISBN 0-12-368848-5. LCCN QC281 .V28 1971.

Teller:1971:EBS

[Tel71d] Edward Teller. The era of big science. Bulletin of the Atomic Scientists, 27(4):34–36, April 1971. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1971:HMS

[Tel71e] Edward Teller. How many secrecies? New York Times, ??(??):??, December 1, 1971. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1971:RS

[Tel71f] Edward Teller. Reopen the skies. New York Times, ??(??):??, April 25, 1971. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1972:CC

[Tel72a] E. Teller. Are the constants constant? In Reines [Rei72a], pages 60–66. ISBN 0-87081-025-1. LCCN QC780 .C65.

Teller:1972:CSC

[Tel72b] E. Teller. Concerning strong convergent compression of thermonuclear fuel in laser-fusion CTR schemes. Report UCRL-74117, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, ???? 1972. Annual Meeting of the Division of Plasma Physics of the American Physical Society (1972).

Teller:1972:FIC

[Tel72c] E. Teller. A future internal combustion engine. Report UCRL-74487, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, ???? 1972. For presentation at the 7th International Quantum Electronics Conference, Montreal, Canada (May 1972).

Teller:1973:GAG

[Tel73a] E. Teller. Die grundsätzliche Antwort. (German) [The basic answer]. In Erhard Scheibe and George Süssmann, editors, Einheit und Vielheit: Festschrift für Carl Friedrich v. Weizsäcker zum 60. Geburtstag. (German) [Unity and multiplicity: Festschrift for Carl Friedrich v. Weizsäcker’s 60th Birthday], page ?? Vandenhoeck & Ruprecht in Göttingen, Göttingen, Germany, 1973. ISBN 3-525-30124-3. LCCN ????

Teller:1973:DH

[Tel73b] E. Teller. The Dirac hypothesis. In Kurşunoğlu [Kur73], pages 351– 352. ISBN 0-306-36902-8. LCCN ????

Teller:1973:LC

[Tel73c] E. Teller. Lasers in chemistry. In ????, editor, Proceedings of The Robert A. Welch Foundation Conferences on Chemical Research, XVI. Theoretical Chemistry, Houston, TX (November 20–22, 1972), pages 205–228. The Robert A. Welch Foundation, Houston, TX, USA, 1973. ISBN ???? LCCN ????

Teller:1973:TE

[Tel73d] E. Teller. Thermonuclear energy. In Impact of basic research on technology [KPB73], pages 181–192. ISBN 0-306-36901-X. LCCN QC30 .K82.

Teller:1973:FIT

[Tel73e] Edward Teller. A future ICE (thermonuclear, that is!). IEEE Spectrum, 10(1):60–64, January 1973. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teller:1973:KSH

[Tel73f] Edward Teller. Kicking the secrecy habit. New York Times, ??(??): ??, May 27, 1973. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1974:FHI

[Tel74] E. Teller. Futurology of high intensity lasers. In Schwarz and Hora [SH74], pages 3–10. ISBN 0-306-37143-X, 0-306-37150-2. LCCN ????

Teller:1975:ED

[Tel75a] E. Teller. Energy disease. Harpers, 250(1497):16–??, ???? 1975. ISSN 0017-789X.

Teller:1975:PNE

[Tel75b] E. Teller. Potential of nuclear-energy. Forensic Quarterly, 49(4):491– 496, ???? 1975. ISSN 0196-304X.

Teller:1975:RLE

[Tel75c] E. Teller. Relevance of lasers in energy shortage. IEEE Journal of Quantum Electronics, 11(9):D83, ???? 1975. CODEN IEJQA7. ISSN 0018-9197 (print), 1558-1713 (electronic).

Teller:1975:EPA

[Tel75d] Edward Teller. Energy: a Plan for Action. Commission on Critical Choices for Americans, New York, NY, USA, April 2, 1975. ISBN ???? 80 pp. LCCN ????

Teller:1976:ASI

[Tel76a] E. Teller. Advanced systems — introduction. Transactions of the American Nuclear Society, 24(??):13, November 19, 1976. CODEN TANSAO. ISSN 0003-018X.

Teller:1976:SRE

[Tel76b] E. Teller. Social responsibility of expert. CHEMTECH, 6(9):539–??,

???? 1976. CODEN CHTEDD. ISSN 0009-2703.

Teller:1976:LLL

[Tel76c] Edward Teller. Lawrence Livermore Laboratory. In Energy and conflict: the life and times of Edward Teller [BO76], pages 453–458. ISBN 0-399-11551-X. LCCN QC16.T37 B58 1976. US$12.95. URL

http://alsos.wlu.edu/information.aspx?id=367.

Teller:1976:OWH

[Tel76d] Edward Teller. Obituary: Werner Heisenberg. Nature, 260(5552): 657–658, April 15, 1976. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v260/n5552/pdf/260657a0.pdf.

Teller:1977:ASI

[Tel77a] E. Teller. Advanced systems — introduction. Transactions of the American Nuclear Society, 25(??):159–160, ???? 1977. CODEN TANSAO. ISSN 0003-018X.

Teller:1977:ECA

[Tel77b] E. Teller. Energy-crisis — alternate solutions by 1982. Journal of Environmental Sciences, 20(3):16–17, ???? 1977. ISSN 0022-0906.

Teller:1977:NRT

[Tel77c] E. Teller. New reactor technologies. Naturwissenschaften, 64(9):445– 447, ???? 1977. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Teller:1977:NRW

[Tel77d] E. Teller. Nuclear reactors — why — where — when? Nuclear Engineering International, 22(256):29–30, ???? 1977. ISSN 0029-5507.

Teller:1977:EG

[Tel77e] Edward Teller. Electromagnetism and gravitation. Proceedings of the National Academy of Sciences of the United States of America, 74(7):2664–2666, July 1977. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://www.jstor.org/

stable/pdfplus/67217.pdf;   http://www.ncbi.nlm.nih.gov/
pmc/articles/PMC431235/.

Teller:1977:SSD

[Tel77f] Edward Teller. In search of solutions for defense and for energy: remarks. Hoover Institution, Stanford University, Stanford, CA, USA, 1977. 23 pp. LCCN UA646.3 .T44.

Teller:1977:NED

[Tel77g] Edward Teller. Nuclear Energy in the Developing World. Mitre Corporation, Metrek Division, McLean, VA, USA, 1977. ISBN ???? ???? pp. LCCN ????

Teller:1977:PPN

[Tel77h] Edward Teller. Pollution by poverty: the need for nuclear energy. Science and Public Policy, 4(3):234–238, June 1977. CODEN ???? ISSN 0302-3427 (print), 1471-5430 (electronic).

Teller:1978:AES

[Tel78a] E. Teller. Albert Einstein — 3 successes, 3 failures — unconventional thoughts about an unconventional man. Interdisciplinary Science Reviews, 3(4):265–266, ???? 1978. ISSN 0308-0188.

Teller:1978:HE

[Tel78b] E. Teller. Human element. CHEMTECH, 8(12):739, ???? 1978. CODEN CHTEDD. ISSN 0009-2703.

Teller:1978:RTC

[Tel78c] E. Teller. Remarks on the thorium cycle. Annals of Nuclear Energy, 5(8–10):287–296, ???? 1978. CODEN ANENDJ. ISSN 0306-4549 (print), 1873-2100 (electronic). Thorium and Gas Cooled Reactors. A seminar for the 65th birthday of Dr. Peter Fortescue.

Teller:1978:STA

[Tel78d] E. Teller. Science and technology in America — chronicle and some predictions. Combustion, 49(7):8–11, ???? 1978. ISSN 0010-2172.

Teller:1978:TIE

[Tel78e] E. Teller. Technical initiatives on early sufficiency in energy. International Journal of Energy Research, 2(1):3–7, ???? 1978. CODEN IJERDN. ISSN 0363-907X (print), 1099-114X (electronic).

Teller:1979:FF

[Tel79a] E. Teller. Fission or fusion? Transactions of the American Nuclear Society, 33(??):214–215, November 1979. CODEN TANSAO. ISSN 0003-018X.

Teller:1979:FPN

[Tel79b] E. Teller. Future prospects of nuclear energy. Siemens Review, 46(1): 2–5, ???? 1979. ISSN 0302-2528.

Teller:1979:NEI

[Tel79c] E. Teller. Nuclear energy and the interdependence of nations. Technology in Society, 1(1):37–41, ???? 1979. ISSN 0160-791x (print), 1879-3274 (electronic).

Teller:1979:WOV

[Tel79d] Edward Teller. I was the only victim of Three-Mile Island. Wall Street Journal, ??(??):??, July 31, 1979. CODEN WSJOAF. ISSN 0099-9660. Advertisement for Dresser Industries, makers of the valve that stuck and contributed to the reactor meltdown at Three-Mile Island, PA, on 28 March 1979.

Teller:1980:F

[Tel80a] Edward Teller. Foreword. In Nuclear war survival skills [Kea80], page i. ISBN 0-939002-02-7 (paperback). LCCN UF767 .K4 1980. US$9.95. Foreword by Edward Teller.

Teller:1981:PPB

[Tel80b] Edward Teller. The projection postulate and Bohr’s interpretation of quantum mechanics. In Asquith and Giere [AG81], pages 201–223. ISBN 0-917586-16-6. LCCN ????

Teller:1980:PS

[Tel80c] Edward Teller. The pursuit of simplicity. Pepperdine University Press, Malibu, CA, USA, 1980. ISBN 0-932612-02-4, 0-932612-03-2 (paperback). 173 pp. LCCN Q158.5 .T45 1980; Q172.T44. US$10.95.

Teller:1981:I

[Tel81a] E. Teller. Introduction. In Teller [Tel81f], pages 1–29. ISBN 0-12-685201-4. LCCN QC791 .F87. URL http://www.sciencedirect.

com/science/book/9780126852011.

Teller:1981:NDU

[Tel81b] E. Teller. Nuclear disaster in the Urals. Russian Review, 40(1):69–70, ???? 1981. ISSN 0036-0341.

Teller:1981:SRN

[Tel81c] E. Teller. Secrecy — the road to nowhere. Technology Review (M.I.T.), 84(1):12–??, ???? 1981. CODEN TEREAU. ISSN 0040-1692.

Teller:1981:WWL

[Tel81d] E. Teller. What have we learned? In Erice 1981, Proceedings, The Unity Of The Fundamental Interactions, pages 775–789. ????, ????, 1981.

Teller:1981:BRB

[Tel81e] Edward Teller. Book review: Nuclear Disaster in the Urals by Zhores A. Medvedev and George Saunders. Russian Review, 40(1):69–70, January 1981. CODEN ???? ISSN ???? URL http://www.jstor.

org/stable/pdfplus/128746.pdf. See [Med79a, Med79b].

Teller:1981:FMCa

[Tel81f] Edward Teller, editor. Fusion: Magnetic confinement: Part A, volume 1A. Academic Press, New York, USA, 1981. ISBN 0-12-685201-4. xiii + 491 pp. LCCN QC791 .F87. URL http://www.

sciencedirect.com/science/book/9780126852011.

Teller:1981:FMCb

[Tel81g] Edward Teller, editor. Fusion: Magnetic confinement: Part B, volume 1B. Academic Press, New York, USA, 1981. ISBN 0-12-685241-3. xi + 529 pp. LCCN QC791 .F87.

Teller:1981:RP

[Tel81h] Edward Teller. Role of physicists in the 1980’s. Physics Today, 34 (2):136, February 1981. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://physicstoday.org/resource/

1/phtoad/v34/i2/p136_s1. See rebuttal and reply [ST81].

Teller:1982:DMA

[Tel82a] Edward Teller. Dangerous myths about nuclear arms. Reader’s Digest, ??(??):139–143, November 1982. CODEN RDIGAU. ISSN 0034-0375. See comment [vH83a], in which this article is reprinted.

Teller:1982:EPN

[Tel82b] Edward Teller. Electromagnetic pulses from nuclear explosions. IEEE Spectrum, 19(10):65, October 1982. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Teller:1982:EH

[Tel82c] Edward Teller. In the eye of a hurricane. Progress in Scientific Culture, 7(1):1–2, ???? 1982. CODEN ???? ISSN ???? URL ????.

Teller:1982:JTE

[Tel82d] Edward Teller. The Jahn–Teller effect — its history and applicability. Physica A, 114(1–3):14–18, 1982. CODEN PHYADX. ISSN 0378-4371 (print), 1873-2119 (electronic). Proceedings of the Tenth International Colloquium on Group-Theoretical Methods in Physics (Canterbury, 1981).

Teller:1982:WWS

[Tel82e] Edward Teller. Why we should resist a nuclear freeze. Vital history cassettes. Encyclopedia Americana/CBS News Audio Resource Library, 10821, ????, October 1982. LCCN ???? 1 cassette.

Teller:1983:LD

[Tel83a] E. Teller. Lasers and defense. Laser Focus with Fiberoptic Technology, 19(6):8, ???? 1983. ISSN 0190-1451.

Teller:1983:MUP

[Tel83b] E. Teller. A mistake under pressure. Power, 127(3):146, ???? 1983.

ISSN 0032-5929.

Teller:1983:HBP

[Tel83c] Edward Teller. The history of the BET paper. In Davis and Hettinger [DH83], chapter 18, pages 227–231. ISBN 0-8412-0778-

X. LCCN QD505 .H47 1983. URL http://pubs.acs.org/isbn/
9780841207783.

Teller:1983:AYA

[Tel83d] Edward Teller. Los Alamos 40 years after. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque, NM 87185-5400, 1983. URL http:

//www.governmentattic.org/11docs/NNSAfilmVideoList_2014.

pdf. Film Y352 (??). Index version of 10 February 2014.

Teller:1983:RC

[Tel83e] Edward Teller. Reagan’s courage. New York Times, ??(??):??, March 30, 1983. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1983:SHR

[Tel83f] Edward Teller. Seven hours of reminiscences. Los Alamos Science, 7:190–195, Winter/Spring 1983. CODEN LASCDI. ISSN 0273-7116.

URL http://library.lanl.gov/cgi-bin/getfile?07-22.pdf.

Teller:1984:TWP

[Tel84a] E. Teller. Technology, war and the prevention of war. Naturwissenschaften, 71(3):109–114, ???? 1984. CODEN NATWAY. ISSN 0028-1042 (print), 1432-1904 (electronic).

Teller:1984:CDC

[Tel84b] Edward Teller. Civil defense is crucial. New York Times, ??(??): ??, January 3, 1984. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1984:WAE

[Tel84c] Edward Teller. Widespread after-effects of nuclear war. Nature, 310 (5979):621–624, August 23, 1984. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v310/n5979/pdf/310621a0.pdf.

Teller:1985:CS

[Tel85a] E. Teller. The case for space. Policy Review, 32(??):4, ???? 1985. ISSN 0146-5945.

Teller:1985:HA

[Tel85b] E. Teller. Hiroshima alternatives. Policy Review, 34(??):7, Fall 1985. ISSN 0146-5945.

Teller:1985:HRM

[Tel85c] E. Teller. The human-rights movement — providing for peace. Center Magazine, 18(1):22–26, ???? 1985. ISSN 0008-9125.

Teller:1985:RSC

[Tel85d] E. Teller. Remarks on stellar clusters. In Gotsman and Tauber [GT85], pages 321–324. ISBN 0-521-30784-8. LCCN QC793.28 .F76 1985.

Teller:1985:SWS

[Tel85e] E. Teller. Star Wars and the scientists — Robert Jastrow and critics. Commentary, 79(3):4, ???? 1985. ISSN 0010-2601.

Teller:1985:CCN

[Tel85f] Edward Teller. Climatic change with nuclear war. Nature, 318 (6042):99, November 14, 1985. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v318/n6042/pdf/318099a0.pdf. See [TTA^{+}90].

Teller:1985:FM

[Tel85g] Edward Teller. A few memories. In French and Kennedy [FK85], pages 181–182. ISBN 0-674-62415-7, 0-674-62416-5 (paperback). LCCN QC16.B63 N49 1985. US$27.50.

Teller:1985:PMP

[Tel85h] Edward Teller. Progress made in protective-defense research. New York Times, ??(??):??, December 29, 1985. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1985:WYT

[Tel85i] Edward Teller. Whatever you think, think again. Quest for Peace, ??

(??):??, ???? 1985. URL http://www.lib.uci.edu/quest/index.
php?page=teller.

Teller:1986:F

[Tel86a] E. Teller. Foreword. In An introduction to equations of state: theory and applications [EGH86], page ?? ISBN 0-521-30389-3.

LCCN QC173.4.E65 E45 1986. URL http://www.loc.gov/catdir/
description/cam031/85029094.html.

Teller:1986:SGVa

[Tel86b] E. Teller. Some generalizations of the virial theorem. In Weingartner and Dorn [WD86], page ?? ISBN 3-209-00630-X. LCCN QC5.56 .F68 1986.

Teller:1986:URR

[Tel86c] E. Teller. The ultimate renewable resource. CHEMTECH, 16(2): 94–95, ???? 1986. CODEN CHTEDD. ISSN 0009-2703.

Teller:1986:WHS

[Tel86d] E. Teller. What helps superpower Star-Wars stability — following on from the Charlton Series. Encounter, 67(3):68–69, ???? 1986. ISSN 0013-7073.

Teller:1986:DNA

[Tel86e] Edward Teller. Defense and the NATO Alliance. The National Interest, ??(??):??, Spring 1986. CODEN ???? ISSN 0884-9382 (print), 1938-1573 (electronic).

Teller:1986:WGA

[Tel86f] Edward Teller. Die Waffenabwehr. (German) [The anti-weapons]. Physikalische Blätter (Physikalische und technische Aspekte von Teilchenstrahl- und Laserwaffen), 42(8):289–292, ???? 1986. CO-DEN ???? ISSN ????

Teller:1986:WEB

[Tel86g] Edward Teller. In the worst event: Book review: The Medical Implications of Nuclear War, and A World Beyond Healing: The Prologue and Aftermath of Nuclear War. In Solomon and Marston [SM86], page ?? ISBN 0-309-03636-4 (paperback), 0-309-03692-5 (hardcover). LCCN RA648.3 .M445 1986.

Teller:1986:NSD

[Tel86h] Edward Teller. The necessity of strategic defense. Annals of the New York Academy of Sciences, 489(1):146–151, December 1986. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic). Presented to The New York Academy of Sciences Conference on the High Technologies and Reducing the Risk of War.

Teller:1986:SGVb

[Tel86i] Edward Teller. Some generalizations of the virial theorem. Foundations of Physics, 16(5):423–435, ???? 1986. CODEN FNDPA4. ISSN 0015-9018 (print), 1572-9516 (electronic). URL http://www.osti.

gov/energycitations/product.biblio.jsp?osti_id=5408332;
http://www.springerlink.com/content/m543kn8473665585/. Part

IV. Invited Papers Dedicated to John Archibald Wheeler.

Teller:1986:WA

[Tel86j] Edward Teller. Why I am in America. San Jose Mercury News, July 4, 1986. CODEN ???? ISSN 0747-2099.

Teller:1987:LL

[Tel87a] E. Teller. The lunar laboratory. In Metropolis et al. [MKR87], pages 77–85. ISBN 0-12-492155-8. LCCN QC44 .N49 1987. URL http:/

/www.osti.gov/energycitations/product.biblio.jsp?osti_id=
6120718&query_id=0.

Teller:1987:RAF

[Tel87b] E. Teller. Remarks about fields of high intensity. Nuclear Instruments and Methods in Physics Research (Netherlands), 24–25(1):1–2, April 3, 1987. CODEN NIMRD9. ISSN 0167-5087 (print), 1872-9606 (electronic). Submitted April 1987.

Teller:1987:SSK

[Tel87c] E. Teller. Should the scientist be king? The Scientist (Philadelphia, PA), 1(24):24, November 1987. ISSN 0890-3670 (print), 1945-5127

(electronic). URL http://www.the-scientist.com/?articles.
view/articleNo/9072/title/Should-the-Scientist-Be-King-/.

Teller:1987:WBH

[Tel87d] E. Teller. A world beyond healing — the prologue and aftermath of nuclear-war. Nature, 328(6125):23–24, July 2, 1987. CODEN NAT-UAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Teller:1987:BST

[Tel87e] Edward Teller. Better a shield than a sword: perspectives on defense and technology. Free Press, New York, NY, USA, 1987. ISBN 0-02-932461-0. xiv + 257 pp. LCCN UA23 .T39 1987. URL http://

alsos.wlu.edu/information.aspx?id=2449.

Teller:1987:DBD

[Tel87f] Edward Teller. Defense is the best defense. The New York Times Magazine, pages 47–48, April 5, 1987. CODEN ???? ISSN 0362-1308.

Teller:1987:EFD

[Tel87g] Edward Teller. Explosive fusion devices. In Meyers [Mey87], pages 723–726. ISBN 0-12-226901-2 (vol. 1), 0-12-226902-0 (vol. 2), 0-12-226903-9 (vol. 3), 0-12-226904-7 (vol. 4), 0-12-226905-5 (vol. 5), 0-12-226906-3 (vol. 6), 0-12-226907-1 (vol. 7), 0-12-226908-X (vol. 8), 0-12-226909-8 (vol. 9), 0-12-226910-1 (vol. 10), 0-12-226911-X (vol. 11), 0-12-226912-8 (vol. 12), 0-12-226913-6 (vol. 13), 0-12-226914-4 (vol. 14). LCCN Q123 .E497 1987.

Teller:1987:WEB

[Tel87h] Edward Teller. In the worst event: Book review: The Medical Implications of Nuclear War, and A World Beyond Healing: The Prologue and Aftermath of Nuclear War. Nature, 328(6125):23–24, July 2, 1987. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/v328/
n6125/pdf/328023a0.pdf. See [SM86, Wad87].

Teller:1987:QSD

[Tel87i] Edward Teller. The question of strategic defense. Information Week, 136:72, September 28, 1987. CODEN INFWE4. ISSN 8750-6874.

Teller:1987:SND

[Tel87j] Edward Teller. Scientists and national defense, an open letter to Hans Bethe. Policy Review, 39(??):18–23, Winter 1987. CODEN ???? ISSN ????

Teller:1988:DDD

[Tel88a] Edward Teller. Deterrence? defense? disarmament? the many roads toward stability. In Anderson and Bark [AB88], pages 21–32. ISBN 0-8179-8751-7, 0-8179-8752-5 (paperback). LCCN E876 .T47 1988.

Teller:1988:HB

[Tel88b] Edward Teller. The H-bomb. In Encyclopedia Americana, volume 14, pages 654–?? Grolier, Inc., Danbury, CT, USA, 1988.

Teller:1989:ATS

[Tel89a] E. Teller. Adaptation of the theory of superconductivity to the behavior of oxides. In Greiner and Stöcker [GS89], pages 485–495. ISBN 0-306-43486-5 (part A), 0-306-43487-3 (part B). LCCN QC793.3.N8 N38 1989.

Teller:1989:STP

[Tel89b] Edward Teller. Science, technology and policy in the 1980’s. Harvard International Review, ??(??):114–118, ???? 1989. CODEN ???? ISSN 0739-1854. 10th Anniversary Issue.

Teller:198x:AYA

[Tel8xa] Edward Teller. Breakthrough in HDR technology. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque, NM 87185-5400, 198x. URL http:

//www.governmentattic.org/11docs/NNSAfilmVideoList_2014.

pdf. Box MP108 (film). Index version of 10 February 2014.

Teller:198x:ET

[Tel8xb] Edward Teller. An early time. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque,

NM 87185-5400, 198x. URL http://www.governmentattic.org/
11docs/NNSAfilmVideoList_2014.pdf. Boxes MP233, MP237,

MP264, film Y-378. Index version of 10 February 2014.

Teller:1990:PCF

[Tel90a] Edward Teller. Press conference following Edward Teller’s colloquium. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque, NM 87185-5400, 1990. URL

http://www.governmentattic.org/11docs/NNSAfilmVideoList_

2014.pdf. Document 90-062. Index version of 10 February 2014.

Teller:1990:QNY

[Tel90b] Edward Teller. Questions for the New York State Institute on Superconductivity Newsletter: Answers provided by Dr. Edward Teller. New York State Institute on Superconductivity Newsletter, ??(??):??, October 2, 1990. URL ????.

Teller:1990:RLE

[Tel90c] Edward Teller. The role of LLNL in the education process. The Quarterly at Lawrence Livermore National Laboratory, 21(2):15–17, November 1990. CODEN ???? ISSN ???? Looking Toward the 21st Century: Laboratory Leaders Share Their Vision of the Future.

Teller:1990:UN

[Tel90d] Edward Teller. An unsinkable navy. Naval Engineers Journal, 102 (1):20–23, January 1990. CODEN ???? ISSN ????

Teller:1991:MAT

[Tel91a] E. Teller. Military applications of technology — a new turn — guest comment. American Journal of Physics, 59(10):873, ???? 1991. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Teller:1991:NTS

[Tel91b] E. Teller. Nuclear techniques in structural chemistry — foreword. Structural Chemistry, 2(3–4):R7, ???? 1991. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001 (electronic).

Teller:1991:ETA

[Tel91c] Edward Teller. Edward Teller addresses students. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque, NM 87185-5400, December 19, 1991. URL http:

//www.governmentattic.org/11docs/NNSAfilmVideoList_2014.

pdf. Document 91-126. Index version of 10 February 2014.

Teller:1991:ETB

[Tel91d] Edward Teller. Electron transfer between layers in high-temperature superconductivity. AIP Conference Proceedings, 219(??):65–74, 1991. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Teller:1991:F

[Tel91e] Edward Teller. Foreword. Structural Chemistry, 2(3–4):vii, ???? 1991. CODEN STCHES. ISSN 1040-0400 (print), 1572-9001

(electronic). URL http://link.springer.com/article/10.1007/
BF00672215.

Teller:1991:BPB

[Tel91f] Edward Teller. From Brilliant Pebbles to brilliant eyes. New York Times, ??(??):??, June 23, 1991. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1991:FDT

[Tel91g] Edward Teller. The future of defense and technology. Report UCRL-JC-105925, CTS-11-90, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, January 10, 1991. Formal paper presented at the Lawrence Livermore National Laboratory workshop: The Role of Nuclear Weapons in the Year 2000 sponsored by the Center for Technical Studies on Security, Energy and Arms Control.

Teller:1991:GCM

[Tel91h] Edward Teller. Guest comment: Military applications of technology: a new turn. American Journal of Physics, 59(10):873, October 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

Teller:1991:HTS

[Tel91i] Edward Teller. High temperature superconductivity: the early 1991 edition. AIP Conference Proceedings, 242(1):217–222, December 10, 1991. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465.

Teller:1991:JF

[Tel91j] Edward Teller. James Franck. In Shils [Shi91], chapter 10, pages 130–137. ISBN 0-226-75335-2. LCCN LD920 .R46 1991. URL http:

//www.loc.gov/catdir/description/uchi051/91016741.html;
http://www.loc.gov/catdir/enhancements/fy0608/91016741-t.
html.

Teller:1991:SNG

[Tel91k] Edward Teller. START: Nuclear glasnost. Bulletin of the Atomic Scientists, 47(9):34–35, November 1991. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comment [Sho92].

Teller:1991:HHC

[Tel91l] Teller Ede. Hazajöttem. (Hungarian) [I came home]. Fizikai Szemle (Budapest), 41(1):1–??, ???? 1991. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic).

Teller:1992:ASH

[Tel92a] E. Teller. Advances in spectroscopy in Hungary — introduction. Spectrochimica Acta Part A — Molecular and Biomolecular Spectroscopy, 48(1):1, ???? 1992. ISSN 0584-8539 (print), 1873-3824 (electronic).

Teller:1992:I

[Tel92b] E. Teller. Introduction. In Nemes and Kovács [NK92], pages 1–?? ISSN 0584-8539 (print), 1873-3824 (electronic). LCCN ????

Teller:1992:KFS

[Tel92c] E. Teller. Keeping fewer secrets. Issues in Science and Technology, 9 (1):6, ???? 1992. ISSN 0748-5492 (print), 1938-1557 (electronic).

Teller:1992:LCE

[Tel92d] E. Teller. Lecture in connection with The Edward Teller Medal Award. In Miley and Hora [MH92], pages 1–4. ISBN 0-306-44353-8. LCCN QC717.6 L343 1991 V10.

Teller:1992:MS

[Tel92e] E. Teller. The miracle of superconductivity. Materials Science Forum, 105–110:161–170, 1992. CODEN MSFOEP. ISSN 0255-5476 (print), 1662-9752 (electronic), 1662-9760.

Teller:1992:SDSa

[Tel92f] E. Teller. Sull’importanza della scienza e dell’Etica. I. (Italian) [On the importance of science and ethics. I]. In Galbiata et al. [GER^{+}92], pages 39–46. ISBN 88-7794-051-4. LCCN ????

Teller:1992:SDSb

[Tel92g] E. Teller. Sull’importanza della scienza e dell’Etica. II. (Italian) [On the importance of science and ethics. II]. In Galbiata et al. [GER^{+}92], pages 47–56. ISBN 88-7794-051-4. LCCN ????

Teller:1992:WC

[Tel92h] E. Teller. Weapons culture. The Sciences (New York), 32(5):3, ???? 1992. CODEN SCNCAD. ISSN 0036-861X.

Teller:1992:FTR

[Tel92i] Edward Teller. Foreign trip report. Lawrence Livermore National Laboratory, Hoover Archives, September 15, 1992. In Hoover archives.

Teller:1992:FDE

[Tel92j] Edward Teller. Fusion devices, explosive. In Meyers [Mey87], pages 1– 4. ISBN 0-12-226901-2 (vol. 1), 0-12-226902-0 (vol. 2), 0-12-226903-9 (vol. 3), 0-12-226904-7 (vol. 4), 0-12-226905-5 (vol. 5), 0-12-226906-3 (vol. 6), 0-12-226907-1 (vol. 7), 0-12-226908-X (vol. 8), 0-12-226909-8 (vol. 9), 0-12-226910-1 (vol. 10), 0-12-226911-X (vol. 11), 0-12-226912-8 (vol. 12), 0-12-226913-6 (vol. 13), 0-12-226914-4 (vol. 14). LCCN Q123 .E497 1987.

Teller:1993:AC

[Tel93a] Edward Teller. 50th anniversary concert. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque, NM 87185-5400, 1993. URL http:

//www.governmentattic.org/11docs/NNSAfilmVideoList_2014.

pdf. Document 93-061. Index version of 10 February 2014.

Teller:1993:EBH

[Tel93b] Edward Teller. Erinnerungen und Bilder: Heisenberg in Leipzig. (German) [Memories and pictures: Heisenberg in Leipzig]. In Kleint and Wiemers [KW93], pages 108–109. ISBN 3-05-501585-1. LCCN QC16.H35 W48 1933. EUR$607 (!).

Teller:1993:LAA

[Tel93c] Edward Teller. The laboratory of the Atomic Age. Los Alamos Science, 21:32–37, 1993. CODEN LASCDI. ISSN 0273-7116.

Teller:1993:NWI

[Tel93d] Edward Teller. Nuclear weapons & international relations at the 100th anniversary at LANL. Video recording at Director’s Colloquium, 10 June 1993, Los Alamos National Laboratory., June 10, 1993.

Teller:1993:FNM

[Tel93e] Teller Ede. A fizika nagyszerű, mert egyszerű. (Hungarian) [Physics is great, because it is simple]. Akad. K., Budapest, Hungary, 1993. ISBN 963-05-6469-6. x + 281 pp. LCCN ???? Conversation with Wendy Teller and Wilson Talleyl.

Teller:1994:F

[Tel94a] E. Teller. Foreword. In Textbook of hyperbaric medicine [Jai94], pages xi–xii. ISBN 3-8017-0735-0. LCCN ????

Teller:1994:FTR

[Tel94b] Edward Teller. Foreign trip report. Lawrence Livermore National Laboratory, Hoover Archives, October 31, 1994.

Teller:1994:KAE

[Tel94c] Edward Teller. Keynote address: The Edward Teller Lecture. AIP Conference Proceedings, 318(1):3–7, October 5, 1994. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-

0465. URL http://www.deepdyve.com/lp/american-institute-
of-physics/keynote-address-the-edward-teller-lecture-NJzHDbn6LI.

Teller:1994:UNF

[Tel94d] Edward Teller. U.S. needs fresh approach to nuclear energy. The Christian Science Monitor, ??(??):66–67, ???? 1994. CODEN ???? ISSN 0882-7729 (print), 1540-4617 (electronic).

Teller:1994:WAS

[Tel94e] Edward Teller. Who’s afraid of science? and why? National Review, ??(??):72, October 10, 1994. CODEN ???? ISSN 0028-0038.

Teller:1995:EPW

[Tel95a] E. Teller. Eugene Paul Wigner, 1902–1995. Physics World, 8(2): 62, February 1995. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

Teller:1995:HPC

[Tel95b] E. Teller. In honor of Professor C. N. Yang’s 70th birthday. In Liu and Yau [LY95], pages 251–254. ISBN 1-57146-001-2. LCCN QC16.Y364 A3 1995.

Teller:1995:HPD

[Tel95c] Edward Teller. Hiroshima — the psychology of a decision. Fizikai Szemle (Budapest), 5(??):176–181, May 1995. CODEN FISZA6. ISSN 0015-3257 (print), 1588-0540 (electronic). URL http://www.kfki.

hu/fszemle/archivum/fsz9905/teller.html. Lecture of Edward

Teller, Honorary Member of the Eötvös Society (delivered at the Eötvös University in Budapest on 25th November 1993). Printed in Hungarian in the Fizikai Szemle 1994/1 p. 1. Includes copy of letter from US President Bill Clinton.

Teller:1996:LWT

[Tel96a] E. Teller. Not as late as we had thought. Fortune, 134(4):20, ???? 1996. CODEN FORTAP. ISSN 0015-8259.

Teller:1996:PPU

[Tel96b] Edward Teller. To promote peace: U.S. foreign policy in the mid-1990s. In Conquest and Djordjevich [CD96], pages 259–264. ISBN 0-312-12373-6. LCCN D415 .P6 1996. URL http://www.loc.gov/

catdir/toc/fy0714/95051502.html.

Teller:1997:NET

[Tel97a] E. Teller. Nuclear energy for the third millennium. Report UCRL-JC-129547, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 1997. Prepared for submittal to the International Conference on Environment and Nuclear Energy, Washington, D.C., 27–29 October 1997.

Teller:1997:SPM

[Tel97b] Edward Teller. Some personal memories of George Gamow. In Harper et al. [HPA97], pages 124–126. ISBN 1-886733-49-X. LCCN

QC16.G37 G46 1996. URL http://adsabs.harvard.edu/abs/
1997ASPC..129..123T.

Teller:1998:ETS

[Tel98a] Edward Teller. Edward Teller SOT. In Lists of motion picture films and videos at The National Nuclear Security Administration (NNSA), 2014. National Nuclear Security Administration, Albuquerque, NM 87185-5400, 1998. URL http://www.

governmentattic.org/11docs/NNSAfilmVideoList_2014.pdf. In-

terview with Dr. Teller, key figure in the Manhattan Project, who reminisces about the project and Trinity test, listed in index version of 10 February 2014.

Teller:1998:FD

[Tel98b] Edward Teller. Fateful decision. Hoover Digest, 1998(4):??, October 30, 1998. ISSN 1088-5161.

Teller:1998:IET

[Tel98c] Edward Teller. Interview with Edward Teller. Web transcript., November 15, 1998. URL http://alsos.wlu.edu/

information.aspx?id=3425;   http://www.gwu.edu/~nsarchiv/
coldwar/interviews/episode-8/teller1.html.

Teller:1998:SM

[Tel98d] Edward Teller. Science and morality. Science, 280(5367):1200–1201, May 22, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/pdfplus/
2896053.pdf.

Teller:1998:SPE

[Tel98e] Edward Teller. Sunscreen for Planet Earth. Hoover Digest, 1998(1): ??, January 30, 1998. ISSN 1088-5161. URL http://www.hoover.

org/publications/hoover-digest/article/6791.

Teller:1998:AHN

[Tel98f] Teller Ede. Atomfizika. (Hungarian) [Nuclear Physics]. Házitanítói Szolgálat Alapítvány, ????, 1998. ISBN 963-04-9849-9. LCCN ????

Teller:1999:OLC

[Tel99a] Edward Teller. Open letter to the citizens of the year 2100. Web

document, July 20, 1999. URL http://web.archive.org/web/
20090107004547/http://www.lhg.org/time/tellerletter.html.

Teller:1999:SSTa

[Tel99b] Edward Teller. Secret-stealing, then and now. New York Times, ?? (??):??, May 14, 1999. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1999:SSTb

[Tel99c] Edward Teller. Stealing secrets, then and now. Hoover Digest, 1999 (4):??, October 30, 1999. ISSN 1088-5161. URL http://www.hoover.

org/publications/hoover-digest/article/6869.

Teller:1999:TPN

[Tel99d] Edward Teller. Three problems: Nuclear energy, national defense and international cooperation. Report UCRL-JC-135608, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 1999. Prepared for submittal to the Second International Symposium on the “History of Atomic Projects–The 1950s: Sociopolitical, Environmental and Engineering Lessons Learned” (HISAP ’99), Laxenburg,, Austria, October 4–8, 1999.

Teller:19xx:HBA

[Telxx] Edward Teller. Heisenberg, Bohr and the atomic bomb. Video in-

terview (6m3s)., 19xx. URL https://www.youtube.com/watch?v=
ClhLnZtgcsE.

Teller:2001:RAJ

[Tel01a] E. Teller. Remarks about the Jahn–Teller effect. Journal of Superconductivity, 14(2):299, ???? 2001. CODEN JOUSEH. ISSN 0896-1107 (print), 1572-9605 (electronic).

Teller:2001:HAS

[Tel01b] Edward Teller. Heisenberg hat die Atombombe sabotiert. (German) [Heisenberg sabotaged the atomic bomb]. In Heisenberg, Hitler und die Bombe: Gespräche mit Zeitzeugen. (German) [Heisenberg, Hitler and the Bomb: Conversations with witnesses] [Sch01], page 114. ISBN 3-928186-60-4 (paperback). LCCN QC16.H35 S32 2001.

Teller:2002:LSS

[Tel02a] Edward Teller. Letter to Sara Stewart, editor of The Week (New York). Teller lists his favorite books: (1) Goethe’s Faust; (2) Arthur Koestler’s Darkness at Noon; (3) William Shakespeare’s Anthony and Cleopatra; (4) John Toland’s The Rising Sun, and (5) the pair Jules Verne’s Around the World in 80 Days and H. G. Well’s The Man Who Could Work Miracles., February 1, 2002.

Teller:2002:UD

[Tel02b] Edward Teller. The ultimate defense. Hoover Digest, 2002(1):??, January 30, 2002. ISSN 1088-5161. URL http://www.hoover.org/

publications/hoover-digest/article/6737.

Teller:2003:TBD

[Tel03a] Teller Ede. Teller a barátjának diktált. (Hungarian) [Teller dictated to his friend]. Népszabadság, ??(??):??, September 16, 2003. ISSN

0133-1752. URL http://nol.hu/archivum/archiv-126521.

Teller:2003:TEU

[Tel03b] Teller Ede. Teller Ede üzenete. (Hungarian) [Edward Teller’s message]. Népszabadság, ??(??):??, September 15, 2003. ISSN 0133-1752.

URL http://nol.hu/archivum/archiv-126353.

Teller:2008:UMV

[Tel08] Teller Ede. Üzenetek egy marslakótól: válogatott tanulmányok: [előadások, beszélgetések, gondolatok]. (Hungarian) [Message in a ????: selected studies: [lectures, conversations, thoughts]]. Lilli K., Budapest, Hungary, 2008. ISBN 963-9837-02-4. 294 pp. LCCN ????

Teller:2010:ETP

[Tel10a] Edward Teller. Edward Teller’s poem on theory and experiment on White House stationery. In Libby and Van Bibber [LV10], pages v–vi. ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Teller:2010:MNG

[Tel10b] Edward Teller. Message to the next generation. In Libby and Van Bibber [LV10], pages 157–?? ISBN 981-283-799-X. LCCN QC16.T37 E39 2008.

Taylor:1958:BRB

[TG58] Hugh Taylor and Guest Reviewers. Book review: The Immense Journey by Loren Eiseley. The Poet’s Way of Knowledge by C. Day Lewis. Great Essays in Science by Martin Gardner. The Inhabited Universe by Kenneth W. Gatland and Derek D. Dempster. The Black Cloud by Fred Hoyle. Paths Across the Earth by J. Lorus and Margery Milne. Our Nuclear Future: Facts, Dangers and Opportunities by Edward Teller and Albert L. Latter. American Scientist, 46(2):140A, June 1958. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/pdfplus/27827113.

pdf. See [TL58b].

Teller:1941:SAW

[TGF41] E. Teller, G. Gamow, and J. A. Fleming. The Seventh Annual Washington Conference of Theoretical Physics, May 22–24, 1941. Science, 94(2430):92–94, July 25, 1941. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Teller:1992:SPF

[TGF92] Edward Teller, Alexander J. Glass, and T. Kenneth Fowler. Space propulsion by fusion in a magnetic dipole. Fusion Technology, 22 (1):82–97, ???? 1992. CODEN FUSTE8. ISSN 0748-1896. Also presented at the 1st International A. D. Sakharov Conference on Physics, Moscow, USSR, May 27–31, 1991.

Thirring:1946:GAE

[Thi46] Hans Thirring. Die Geschichte der Atombombe; mit einer elementaren Einführung in die Atomphysik auf Grund der Originalliteratur gemeinverständlich dargestellt. (German) [The history of the atomic bomb, with an elementary introduction to atomic physics according to the original literature, understandably presented]. Wissenschaft für jedermann; Phönix-Bücherel. “Neues Österreich” Zeitungs- und Verlagsgesellschaft, Wien, Austria, 1946. 149 pp. LCCN QC778 .T47.

Thirring:1950:SB

[Thi50] Hans Thirring. The super bomb. Bulletin of the Atomic Scientists, 6(3):69–70, March 1950. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). English translation of a chapter of the author’s book [Thi46] (reviewed in the Bulletin in April 1948). According to [Har10, page 203], its English translation was delayed until after President Truman’s order of 1 February 1950 directing the Atomic Energy Commission (AEC) to work on the super bomb.

Thorndike:1980:BRB

[Tho80] Edward H. Thorndike. Book review: Energy from Heaven and Earth, by Edward Teller. American Journal of Physics, 48(6):504, June 1980. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://ajp.aapt.org/resource/1/ajpias/v48/i6/p504_s1.

Thorpe:2003:BRE

[Tho03] Charles Thorpe. Book review: Edward Teller with Judith Shoolery, Memoirs: A Twentieth-Century Journey in Science and Politics. Oxford: Perseus Press, 2001. Pp. xii + 628. ISBN 1-903985-12-9. £24.99 (hardback). British Journal for the History of Science, 36(2):250–251, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X

(electronic). URL http://www.jstor.org/stable/4028247.

Tierney:1994:MCE

[Tie94] Luke Tierney. Markov chains for exploring posterior distributions. Annals of Statistics, 22(4):1701–1728, December 1994. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176325750;   http://
www.jstor.org/stable/2242477.

Tisza:1998:ETN

[Tis98] L. Tisza. Edward Teller ninety. Acta Physica Hungarica New Series: Heavy Ion Physics, 8(3):179–184, ???? 1998. CODEN APHPFM. ISSN 1219-7580 (print), 1588-2675 (electronic).

Titus:1986:BBA

[Tit86] A. Costandina Titus. Bombs in the backyard: atomic testing and American politics, volume 25 of Nevada studies in history and political science. University of Nevada Press, Reno, NV, USA, 1986. ISBN 0-87417-107-5. xiv + 214 + 12 pp. LCCN U264 .T58 1986; U264 .T58 1986 (LC). US$16.95.

Titus:2001:BBA

[Tit01] A. Costandina Titus. Bombs in the backyard: atomic testing and American politics, volume 25 of Nevada studies in history and political science. University of Nevada Press, Reno, NV, USA, second edition, 2001. ISBN 0-87417-370-1. xvi + 242 pp. LCCN U264 .T58 2001.

Teller:1996:CAN

[TIW^{+}96] Edward Teller, Muriel Ishikawa, Lowell Wood, Roderick Hyde, and John Nuckolls. Completely automated nuclear reactors for longterm operation II: Toward a concept-level point-design of a hightemperature, gas-cooled central power station system. Report UCRL-JC-122708 Pt. 2, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, June 20, 1996. Prepared for submittal to 1996 International Conference on Emerging Nuclear Energy Systems (ICENES 96), Obninsk, Russian Federation; 24–28 June 1996.

Teller:1972:GSC

[TJ72] Edward Teller and Montgomery H. Johnson. General shock conditions. Geophysical Monograph, 16(??):313–318, ???? 1972. CODEN GPMGAD. ISSN 0065-8448.

Teller:1979:U

[TK79] Edward Teller and George A. (Jay) Keyworth. [unknown]. Memorandum, ????, September 20, 1979. Transcribed from a taped conversation between the authors about the Three-Mile Island nuclear reactor failure. Cited in [Har10, page 144], but attempts to locate it, or determine its title or other publication data, have so far failed.

Teller:2000:FI

[TK00] Edward Teller and Behram N. Kurşunoğlu. Five issues. Report UCRL-JC-142054, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 2000. Submitted to Global Warming and Energy Policy Conference, Fort Lauderdale, Florida, November 26–28, 2000.

Teller:1958:AKG

[TL58a] Edward Teller and Albert Latter. Ausblick in das Kernzeitalter. (German) [Outlook in the nuclear age]. Fischer Bücherie, Frankfurt am Main and Hamburg, Germany, 1958. 159 (est.) pp. LCCN ???? US$. German translation by Otto Maser.

Teller:1958:ONF

[TL58b] Edward Teller and Albert L. Latter. Our nuclear future: facts, dangers, and opportunities. Criterion Books, New York, NY, USA, 1958. 184 pp. LCCN QC778 .T4; QC778.T45x 1958. URL http://alsos.

wlu.edu/information.aspx?id=423.

Teller:2001:LRW

[TLCW01] Edward Teller, Cecil Leith, Gregory Canavan, and Lowell Wood. Long range weather prediction III: Miniaturized distributed sensors for global atmospheric measurements. Report UCRL-JC-146204, Part 3, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 2001. Prepared for presentation at the 26th International Symposium on Planetary Emergencies, Majorana Centre for Scientific Culture, Erice, Italy, 20–23 August 2001.

Teller:1976:PS

[TMF76] Edward Teller, Hans Mark, and John S. Foster. Power and security, volume 4 of Critical choices for Americans. Lexington Books, Lexington, MA, USA, 1976. ISBN 0-669-00416-2. xxiii + 204 pp. LCCN HD9502.U52 T44.

Teller:1997:BF

[TNR97] Edward Teller, Edward Neilan, and Peter M. Robinson. Back to the future. Hoover Digest, 1997(1):??, January 30, 1997. ISSN 1088-5161.

Toth:1963:TOT

[Tot63a] Robert C. Toth. Teller opposes Test Ban Treaty but 35 Nobel Laureates ask Senate to ratify accord. New York Times, ??(??):??, August 15, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Toth:1963:TSCa

[Tot63b] Robert C. Toth. Teller shows consistency in opposing test ban. New York Times, ??(??):??, August 23, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Toth:1963:TSCb

[Tot63c] Robert C. Toth. Teller suggests Congress guide growth of scientific projects. New York Times, ??(??):??, September 19, 1963. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1958:FD

[TP58] Edward Teller and Linus Pauling. Fallout and Disarmament. Fearon Publishers, San Francisco, CA, USA, 1958. 13 pp. URL http://

osulibrary.oregonstate.edu/specialcollections/coll/pauling/

peace/papers/1958p2.1.html. Transcript of debate broadcast by KQED-TV, San Francisco. Introduced and moderated by James Day.

Teller:1960:SWT

[TR60] Edward Teller and Bertrand Russell, editors. Small world: Teller, Russell [Motion picture]. Columbia Broadcasting System, ????, 1960. LCCN ???? Two black-and-white 16mm motion pictures (28 minutes each).

Treverton:1987:BRB

[Tre87] Gregory F. Treverton. Book review: Better a Shield than a Sword: Perspectives on Defense and Technology by Edward Teller. Foreign Affairs, 66(1):192, Fall 1987. CODEN FRNAA3. ISSN 0015-7120.

URL http://www.jstor.org/stable/pdfplus/20043305.pdf. See

[Tel87e].

Teller:1970:GRE

[TS70a] E. Teller and H. L. Sahlin. General remarks on electronic structure. In Eyring et al. [EHJ70], pages 1–34. ISBN 0-12-245605-X. LCCN ????

Teller:1970:HMI

[TS70b] E. Teller and H. L. Sahlin. The hydrogen molecular ion and the general theory of electron structure. In Eyring et al. [EHJ70], pages 35–124. ISBN 0-12-245605-X. LCCN ????

Teller:1982:MAR

[TS82] E. Teller and J. Shoolery. The myth about radiation. Progress in Clinical Cancer, 8(??):205–213, ???? 1982. URL http://www.ncbi.

nlm.nih.gov/pubmed/6760248.

Teller:1984:DWD

[TS84] Edward Teller and R. Jeffrey Smith. Defensive weapons development. Science, 223(4633):236–238, January 20, 1984. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.

jstor.org/stable/pdfplus/1693030.pdf.

Teller:2001:MTC

[TS01] Edward Teller and Judith L. Shoolery. Memoirs: a Twentieth-Century Journey in Science and Politics. Perseus Publishers, Cambridge, MA, USA, 2001. ISBN 0-7382-0532-X, 1-903985-12-9. xii + 628 pp. LCCN QC16.T37 M55 2001. URL http:/

/alsos.wlu.edu/information.aspx?id=784;   http://www.loc.
gov/catdir/enhancements/fy0833/2001097880-d.html;   http:
//www.loc.gov/catdir/toc/fy02/2001097880.html.

Teller:1931:MMG

[TT31] E. Teller and L. Tisza. Über mehratomige Moleküle. (German) [On polyatomic molecules]. Physikalische Zeitschrift, 32(??):219–??, 1931. CODEN PHZTAO. ISSN 0369-982X. URL http://babel.

hathitrust.org.

Teller:1932:DUS

[TT32] E. Teller and L. Tisza. Zur Deutung des ultraroten Spektrums mehratomiger Moleküle. (German) [On the interpretation of the infrared spectrum of polyatomic molecules]. Zeitschrift für Physik, 73(11–12):791–812, November 1932. CODEN ZEPYAA. ISSN

0044-3328. URL http://link.springer.com/article/10.1007/
BF01344226.

Teller:1935:EHR

[TT35a] Eduard Teller and Bryan Topley. 199. on the equilibrium and the heat of the reaction C_{2}H_{4} + H_{2} = C_{2}H_{6}. Journal of the Chemical Society, ??(??):876–885, July 1935. CODEN JCSOA9. ISSN 0368-1769 (print), 2050-5574 (electronic). URL http://pubs.rsc.org/

en/content/articlelanding/1935/jr/jr9350000876.

Teller:1935:VFE

[TT35b] Eduard Teller and Bryan Topley. 200. on the vibration frequencies of ethylene and ethane. Journal of the Chemical Society, ??(??): 885–889, ???? 1935. CODEN JCSOA9. ISSN 0368-1769 (print),

2050-5574 (electronic). URL http://pubs.rsc.org/en/content/
articlelanding/1935/jr/jr9350000885.

Teller:1940:SWC

[TT40] Edward Teller and M. A. Tuve. The Sixth Washington Conference on Theoretical Physics March 21–23, 1940. Science, 91(2374):621–623, June 28, 1940. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-

9203 (electronic). URL http://www.jstor.org/stable/pdfplus/
1667439.pdf.

Talley:1964:PEE

[TT64] W. K. Talley and E. Teller. Proposal of experiments on electromagnetic radiation from a distant nuclear explosion. In ????, editor, Proceedings of the Third Plowshare Symposium, Engineering with Nuclear Explosives, University of California/Davis, pages 31–35. ????, ????, April 1964. LCCN ????

Teller:1977:STA

[TT^{+}77] Edward Teller, Robb M. (Robb Milton) Thomson, et al. Science and technology in America: an assessment, volume 465 of NBS special publication. United States Government Printing Office, Washington, DC, USA, 1977. CODEN XNBSAV. ISSN 0083-1883. vi + 162 + 1 pp. LCCN QC100 .U57 no. 465. US$2.50. Distinguished lecture series published on the occasion of the 75th anniversary of the National Bureau of Standards.

Tedeschi:1994:PWP

[TT94] W. Tedeschi and E. Teller. A plan for worldwide protection against asteroid impacts. Space Policy, 10(3):183–184, ???? 1994. ISSN 0265-9646.

Turco:1990:CSA

[TTA^{+}90] R. P. Turco, O. B. Toon, T. P. Ackerman, J. B. Pollack, and Carl Sagan. Climate and smoke: An appraisal of nuclear winter. Science, 247(??):166–176, ???? ??, 1990. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). See [Tel85f].

Teller:1968:CUN

[TTHJ68] Edward Teller, Wilson K. Talley, Gary H. Higgins, and Gerald W. Johnson, editors. The Constructive Uses of Nuclear Explosives. Mc-Graw-Hill, New York, NY, USA, 1968. xiv + 320 pp. LCCN TK9153 .C6.

Teller:1991:CDS

[TTT91] Edward Teller, Wendy Teller, and Wilson Talley. Conversations on the dark secrets of physics. Plenum Press, New York, NY, USA; London, UK, 1991. ISBN 0-306-43772-4. xii + 247 pp. LCCN QC21.2 .T45 1991. US$23.95.

Teller:2002:CDS

[TTT02] Edward Teller, Wendy Teller, and Wilson Talley. Conversations on the dark secrets of physics. Perseus Publishers, Cambridge, MA, USA, 2002. ISBN 0-7382-0765-9. xii + 247 pp. Reprint of the 1991 original.

Teller:1951:HDH

[TU51] Edward Teller and Stanisław Ulam. On heterocatalytic detonations. I. Hydrodynamic lenses and radiation mirrors. Report LAMS-1225, Los Alamos Scientific Laboratory of the University of California, Los Alamos, NM, USA, March 9, 1951.

20 pp. URL http://blog.nuclearsecrecy.com/wp-content/
uploads/2012/02/1951-Teller-Ulam-LAMS1225.pdf;   http://
www.nuclearnonproliferation.org/LAMS1225.pdf. This now-

classified report is believed to contain the Teller–Ulam hydrogenbomb trigger design. See [Har10, page 221], [Mac92, page 344], [Goo04b, pages 179–181]. For a discussion of this report, see [For15, Chapter 1]. Copies found on the Web are heavily censored, with many pages completely blanked out.

Turner:1940:NF

[Tur40] Louis A. Turner. Nuclear fission. Reviews of Modern Physics, 12(1):1– 29, January 1940. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.12.1;   http://rmp.aps.org/abstract/RMP/

v12/i1/p1_1. This paper surveys the early work from 1934 to 1939 on nuclear fission, including more than 100 from 1939 alone following the news of the Hahn and Strassmann work that Niels Bohr brought to the USA in January 1939.

Turner:1979:WHB

[Tur79] Wallace Turner. Writer of H-bomb letter expresses surprise at uproar. New York Times, ??(??):??, September 19, 1979. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1983:FH

[TvH83] Edward Teller and Frank von Hippel. On facts and hopes. Bulletin of the Atomic Scientists, 39(4):42–44, April 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [vH83a].

Teller:1933:SWG

[TW33] E. Teller and K. Weigert. Die spezifische Wärme des gehemmten eindimensionalen Rotators. (German) [The specific heat of the restricted one-dimensional rotator]. Göttingen Nachrichten, Math. Phys. Klasse, 2(??):218–231, 1933. CODEN ???? ISSN ????

Teller:1938:RAN

[TW38] E. Teller and J. A. Wheeler. On the rotation of the atomic nucleus. Physical Review, 53(10):778–789, May 15, 1938. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v53/i10/p778_1.

Teller:1973:IR

[TW73] Edward Teller and Morton S. Weiss. Introductory remarks. In ????, editor, International Conference on Photonuclear Reactions and Applications, Asilomar Conference Grounds, Pacific Grove, CA, March 26-30, 1973, page ?? ????, ????, 1973. ISBN ???? LCCN ???? Lawrence Livermore Laboratory Report CONF-730301 (1973).

Teller:1979:ANB

[TW79a] Edward Teller and Eugene P. Wigner. America needs a better civil defense program. New York Times, ??(??):??, July 22, 1979. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Teller:1979:EHE

[TW79b] Edward Teller and Eugene Paul Wigner. Energy from Heaven and Earth: in which a story is told about energy from its origins 15,000,000,000 years ago to its present adolescence — turbulent, hopeful, beset by problems, and in need of help. W. H. Freeman and Company, New York, NY, USA, 1979. ISBN 0-7167-1063-3, 0-7167-1064-1 (paperback). xiv + 322 pp. LCCN TJ163.2 .T4.

Teller:1980:MMM

[TW80] Edward Teller and Morton S. Weiss. Is the muon a multipole meter? Transactions of the New York Academy of Sciences, Series II, 40(??): 222–229, ???? 1980. CODEN TNYAAE. ISSN 0028-7113 (print), 1939-2966 (electronic). A Festschrift for Maurice Goldhaber.

Teller:1992:TWS

[TW92] Edward Teller and Lowell Wood. Teller, Wood: Silent on Goldin to Quayle. Physics Today, 45(8):11, August 1992. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:

//physicstoday.org/resource/1/phtoad/v45/i8/p11_s1.

Teller:1973:CDWa

[TWB73a] E. Teller, E. Wigner, and A. Broyles. Without civil defense we are in a glass house. Technical report, Washington Report of the American Security Council, Washington, DC, USA, May 1973. ???? pp.

Teller:1973:CDWb

[TWB73b] E. Teller, E. Wigner, and A. Broyles. Without civil defense we are in a glass house. Survive, 6(5):1–2, ???? 1973. CODEN ???? ISSN 0039-6354.

Teller:1997:GWI

[TWH97] Edward Teller, Lowell Wood, and Roderick Hyde. Global warming and ice ages: I. Prospects for physics-based modulation of global change. Report UCRL-JC-128715, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 1997. Prepared for submittal to the 22nd International Seminar on Planetary Emergencies, Erice (Sicily), Italy, 20–23 August 1997.

Teller:1990:LHA

[TZ90] Teller Ede and Zeley László. Légiposta. (Hungarian) [Airmail]. Háttér Lap-és Könyvkiadó, Budapest, Hungary, 1990. ISBN 963-7403-66-3. 135 pp. LCCN QC16.T37 A32 1990.

Teller:1991:BBA

[TZ91] Teller Ede and Zeley László. A biztonság bizonytalansága: az atomkor — fél évszázad múltán. (Hungarian) [The uncertainty of the security: the nuclear age — fifty years later]. Relaxa, Budapest, Hungary, 1991. ISBN 963-7991-10-7. 286 pp.

Udall:1994:MAP

[Uda94] Stewart L. Udall. The myths of August: a personal exploration of our tragic Cold War affair with the atom. Pantheon Books, New York, NY, USA, 1994. ISBN 0-679-43364-3. xii + 399 pp. LCCN E840 .U33 1994.

Udall:1998:MAP

[Uda98] Stewart L. Udall. The myths of August: a personal exploration of our tragic Cold War affair with the atom. Rutgers University Press, New Brunswick, NJ, USA, 1998. ISBN 0-8135-2546-2. xii + 399

pp. LCCN E840 .U33 1998. URL http://www.loc.gov/catdir/
enhancements/fy1502/97053053-d.html.

VanDeMark:2003:PKN

[Van03a] Brian VanDeMark. Pandora’s keepers: nine men and the atomic bomb. Little, Brown and Co., Boston, MA, USA, 2003. ISBN 0-316-73833-6. xii + 399 + 16 pp. LCCN QC774.A2 V36 2003. URL

http://www.loc.gov/catdir/toc/fy045/2002043646.html.

Vanderbilt:2003:YTR

[Van03b] Tom Vanderbilt. Not your typical road trip: Book review: The Traveler’s Guide to Nuclear Weapons: A Journey Through America’s Cold War Battlefields, by James M. Maroncelli and Timothy L. Karpin. Bulletin of the Atomic Scientists, 59(6):74–75, November/December 2003. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://bos.sagepub.com/content/
59/6/74.full.pdf+html;   http://www.highbeam.com/doc/1G1-
110621498.html. See [MK02].

vonHippel:1983:METa

[vH83a] Frank von Hippel. The myths of Edward Teller. Bulletin of the Atomic Scientists, 39(3):6–12, March 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Tel82a], which is reprinted in this article, response [TvH83], and comments [Bul83, vH83b].

vonHippel:1983:METb

[vH83b] Frank von Hippel. On ‘The Myths of Edward Teller’. Bulletin of the Atomic Scientists, 39(9):60–61, November 1983. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [vH83a].

vonHippel:2001:WFS

[vH01] Frank N. von Hippel. Where Fermi stood. Bulletin of the Atomic Scientists, 57(5):26–29, September/October 2001. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

URL http://bos.sagepub.com/content/57/5/26.full;   http:
//bos.sagepub.com/content/57/5/26.full.pdf+html;   http:/
/www.highbeam.com/doc/1G1-78334537.html.

Vincze:2003:TET

[Vin03] Vincze Attila Tamás. Teller Ede: a tudós és világa: [a vele készült utolsó interjúk alapján]. (Hungarian) [Edward Teller: The scientist and the world: [the last interviews with him]]. Pallas Antikvárium, Gyöngyös, Hungary, 2003. ISBN 963-9504-56-4. 111 pp. LCCN ???? Preface by Czeizel Endre.

Volkman:2002:SGW

[Vol02] Ernest Volkman. Science goes to war: the search for the ultimate weapon, from Greek fire to Star Wars. John Wiley, New York, NY, USA, 2002. ISBN 0-471-41007-1. ix + 278 pp. LCCN U27

.V65 2002. URL http://www.loc.gov/catdir/bios/wiley044/
2002284341.html;   http://www.loc.gov/catdir/description/
wiley037/2002284341.html;   http://www.loc.gov/catdir/toc/
wiley023/2002284341.html.

vonWeizsacker:2002:LMH

[vW02] C. F. von Weizsäcker. Letters to Martin Heidegger, Niels Bohr, Hoimar von Ditfurth, Edward Teller, Joachim Illies, Klaus Scholder, Jürgen Habermas, Jürgen Kuczyński, Hans-Georg Gadamer. Sinn und Form, 54(2):161–176, ???? 2002. ISSN 0037-5756.

Wade:1987:WBH

[Wad87] Nicholas Wade. A world beyond healing: the prologue and aftermath of nuclear war. W. W. Norton, New York, NY, USA, 1987. ISBN 0-393-02335-4. 190 pp. LCCN U263 .W33 1987.

Wagner:1981:EPW

[Wag81] Francis S. Wagner. Eugene P. Wigner: an architect of the Atomic Age: highlights of a career with a comprehensive bibliography, volume 1 of Rákóczi Foundation bio-bibliographies. Rákóczi Foundation, Toronto, ON, Canada, 1981. ISBN 0-919545-00-9 (paperback). 80 pp. LCCN QC16.W52 W33. With the research assistance of Christina Maria T. Wagner-Jones. Epilogue by Edward Teller.

Wagner:1998:WJP

[Wag98] Wagner Ferenc. Wigner Jenő Pál: az atomkor egyik megalapítója: egy karrier fénypontjai, teljes bibliográfiával. (Hungarian) [Eugene Paul Wigner: one of the founders of the nuclear age: highlights of a career full bibliography], volume 4 of Studia physica Savariensia. BDTF, Szombathely, Hungary, 1998. ISBN 963-9017-36-1. ISSN 1219-2678. 78 pp. LCCN ???? With the assistance of Christina Maria T. Wagner-Jones. Epilog by Edward Teller.

Walter:1998:LS

[Wal98] Katie Walter. For the love of science, July 1998. URL https://www.

llnl.gov/str/Science.html.

Wald:2008:NAW

[Wal08] Matthew L. Wald. Nuclear agency weighs attack threat at plants. New York Times, ??(??):??, July 2, 2008. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Wang:2003:BRE

[Wan03] Zuoyue Wang. Book review: Edward Teller and Judith L. Shoolery: Memoirs: A Twentieth-Century Journey in Science and Politics. Isis, 94(2):419–420, June 2003. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.

org/stable/10.1086/379484;   http://www.jstor.org/stable/
pdfplus/10.1086/379484.pdf.

Wattenberg:2003:BRE

[Wat03] Albert Wattenberg. Book review: Edward Teller with Judith Shoolery, Memoirs: A Twentieth-Century Journey in Science and Politics. Cambridge, Massachusetts: Perseus Publishing, 2001, xii + 628 pages. $35.00 (cloth). Physics in Perspective (PIP), 5(3):349–352, September 2003. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

Wigner:1939:EPF

[WCT39] Eugene P. Wigner, Charles L. Critchfield, and Edward Teller. The electron-positron field theory of nuclear forces. Physical Review (2), 56(6):530–539, September 15, 1939. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.aps.org/

abstract/PR/v56/i6/p530_1.

M:1975:FAT

[W.D75] W.D.M. Fusion advocates tell off Teller. Science, 187(4172):146– 147, January 17, 1975. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic).

Weingartner:1986:FPS

[WD86] Paul Weingartner and Georg Dorn, editors. Foundations of physics: a selection of papers contributed to the Physics Section of the 7th International Congress of Logic, Methodology and Philosophy of Science. Hölder-Pichler-Tempsky, Vienna, Austria, 1986. ISBN 3-209-00630-X. LCCN QC5.56 .F68 1986.

Weart:1976:SS

[Wea76] Spencer R. Weart. Scientists with a secret. Physics Today, 29(2):23– 30, February 1976. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://www.physicstoday.org/resource/
1/phtoad/v29/i2/p23_s1.

Weart:2012:RNF

[Wea12] Spencer R. Weart. The rise of nuclear fear. Harvard University Press, Cambridge, MA, USA, 2012. ISBN 0-674-05233-1. ???? pp. LCCN TK9146 .W384 2012.

Weisskopf:1975:LCE

[Wei75] Victor Weisskopf. Letters: A clerical error. Bulletin of the Atomic Scientists, 31(4):3, April 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reports incorrect inclusion of Weisskopf’s name in [BAA^{+}75].

Weisman:1994:ERW

[Wei94a] Jonathan Weisman. Early retirement for weaponeers? Bulletin of the Atomic Scientists, 50(4):16–22, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.

highbeam.com/doc/1G1-15587904.html.

Weisman:1994:NR

[Wei94b] Jonathan Weisman. Nuckolls rapped. Bulletin of the Atomic Scientists, 50(4):18–19, July/August 1994. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://www.highbeam.

com/doc/1G1-15587906.html.

Weigel:2004:ETR

[Wei04] J. W. Weigel. Edward Teller: The real Dr. Strangelove. Library Journal, 129(15):80, ???? 2004. CODEN LIBJA7. ISSN 0363-0277.

Wells:1977:LUW

[Wel77] Merle W. Wells. Letter: University weapons work. Bulletin of the Atomic Scientists, 33(8):3–4, October 1977. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Day77].

Wellerstein:2012:ITU

[Wel12] Alex Wellerstein. The infamous Teller–Ulam report (1951). Web site.,

February 22, 2012. URL http://blog.nuclearsecrecy.com/2012/
02/22/weekly-document-15-the-infamous-teller-ulam-report-
1951/.

Westwick:2008:SOI

[Wes08] Peter J. Westwick. The Strategic Offense Initiative? The Soviets and Star Wars. Physics Today, 61(6):43–49, June 2008. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v61/i6/p43/s1.

Wellerstein:2017:SSH

[WG17] Alex Wellerstein and Edward Geist. The secret of the Soviet hydrogen bomb. Physics Today, 70(4):40–47, April 2017. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Wilt:1998:GFM

[WH98] Gloria Wilt and Bart Hacker. Gifts of a fertile mind, July 1998. URL

https://www.llnl.gov/str/Mind.html.

Wood:1995:CBV

[WHIT95] Lowell Wood, Rod Hyde, Muriel Ishikawa, and Edward Teller. Cosmic bombardment V: Threat object-dispersing approaches to active planetary defense. Report UCRL-JC-Draft, University of California, Lawrence Livermore Laboratory, Livermore, CA, USA, 1995. Paper prepared for submittal to the Planetary Defense Workshop, May 22-26, 1995, Livermore, CA.

Wigner:1962:FAA

[Wig62] Eugene Wigner. Fermi Award: AEC honors Teller for contributions to nuclear science. Science, 138(3545):1087–1088, December 7, 1962. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/1709475;   http:/
/www.sciencemag.org/content/138/3545/1087.full.pdf.

Wigner:1969:ABE

[Wig69] Eugene P. Wigner. An appreciation on the 60th birthday of Edward Teller. In Mark and Fernbach [MF69], chapter 1, pages 1–6. ISBN 0-470-56990-5. LCCN Q171 .P98.

Wigner:1976:BRB

[Wig76] Eugene P. Wigner. Book review: The Advisors: Oppenheimer, Teller, and the Superbomb, by Herbert York. American Scientist, 64(5):561, September 1976. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-

2786 (electronic). URL http://www.jstor.org/stable/27847469.

Wigner:1996:ABE

[Wig96a] E. P. Wigner. An appreciation on the 60th birthday of Edward Teller. In Wigner [Wig96b], pages 133–138. ISBN 3-540-56972-3. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wigner:1996:CWE

[Wig96b] Eugene Paul Wigner, editor. The collected works of Eugene Paul Wigner. Part A. The scientific papers. Vol. II. Nuclear physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-56972-3. x + 574 pp. Annotated by Herman Feshbach, Edited and with a preface by Arthur S. Wightman and Jagdish Mehra.

Wilson:1986:MRL

[Wil86] George C. Wilson. Mirror-reflected laser suggested to shield Allies. Washington Post, ??(??):??, March 3, 1986. ISSN 0190-8286.

Williams:1987:KFA

[Wil87] Robert Chadwell Williams. Klaus Fuchs, atom spy. Harvard University Press, Cambridge, MA, USA, 1987. ISBN 0-674-50507-7. vi + 267 + 10 pp. LCCN UB271.R9 F838 1987.

Wilt:1998:GEM

[Wil98] Gloria Wilt. Glimpses of an exceptional man. Web document, July

1998. URL https://www.llnl.gov/str/Exceptional.html.

Ward:2015:RWF

[WM15] Tim Ward and Domenic Mastrippolito. Race for the world’s first atomic bomb: A thousand days of fear. [Los Alamos National] Laboratory produced documentary on the Manhat-

tan Project., 2015. URL http://www.governmentattic.org/
11docs/NNSAfilmVideoList_2014.pdf;   https://www.youtube.

com/watch?v=xwpgmEvlRpM. Written by Michael Wilson. The hourlong program features interviews with some of those who worked on the Project Hans Bethe, Norman Ramsey, Phil Morrison, Edward Teller and others. The program profiles the creation of the Manhattan Project the obstacles that were faced, life in the “secret city,’ and events that led up to and followed the Trinity Test.

Wood:1988:DNE

[WN88] Lowell Wood and John Nuckolls. The development of nuclear explosives. In Mark and Wood [MW88], pages 317–?? ISBN 90-277-2775-9. LCCN QC71 .E62 1988.

Wolff:1967:TFY

[Wol67] Anthony Wolff. Twenty-five years with the bomb: A conversation with Nobel Prize-winner Eugene P. Wigner. Look Magazine, ??(??): 58, 60–61, December 26, 1967. URL http://library.ucsd.edu/dc/

object/bb5564714p.

Wolk:2009:MHB

[Wol09] Herman S. Wolk. Making the H-bomb. Air Force magazine, 95(3): 66–69, March 2009. CODEN AFORCO. ISSN 0730-6784 (print),

1943-4782 (electronic). URL http://www.airforce-magazine.com/
MagazineArchive/Pages/2009/March%202009/0309H-Bomb.aspx.

Wood:2003:ETK

[Woo03] Lowell Wood. The Edward Teller I knew. Newsline (Lawrence Livermore National Laboratory), 28(36):2, September 12, 2003.

URL https://newsline.llnl.gov/employee/articles/2003/09-
12-03-newsline.pdf.

Wigner:1992:REP

[WS92] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Plenum Press, New York, NY, USA; London, UK, 1992. ISBN 0-306-44326-0. xxiv + 335 pp. LCCN QC16.W52 A3 1992.

Wigner:2003:REP

[WS03] Eugene Paul Wigner and Andrew Szanton. The recollections of Eugene P. Wigner as told to Andrew Szanton. Basic Books, New York, NY, USA, 2003. ISBN 0-7382-0886-8. xxiv + 335 pp. LCCN QC16.W52.

Wigner:1979:ANB

[WT79] Eugene P. Wigner and Edward Teller. America needs a better civil defense program: Congressional Record Senate, August 2, 1979. In USA Congressional Record. Proc. and Debates of the 96th Congress, First Session, page ?? US Government Printing Office, Washington, DC, USA, 1979.

Wyden:1984:DOB

[Wyd84] Peter Wyden. Day One: Before Hiroshima and After. Simon and Schuster, New York, NY, USA, 1984. ISBN 0-671-46142-7 (hardcover). xxxii + 412 pp. LCCN D767.25.H6 W93 1984.

York:1975:OLL

[Yor75] Herbert F. York. The origins of the Lawrence Livermore Laboratory. Bulletin of the Atomic Scientists, 31(7):8–14, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

York:1976:AOT

[Yor76] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. W. H. Freeman and Company, New York, NY, USA, 1976. ISBN 0-7167-0718-7. x + 175 pp. LCCN UG1282.A8 Y67.

York:1987:MWT

[Yor87] Herbert F. (Herbert Frank) York. Making weapons, talking peace: a physicist’s odyssey from Hiroshima to Geneva. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-04338-0. xiv + 359 + 8 pp. LCCN JX1974.7 .Y575 1987. US$22.95.

York:1989:AOT

[Yor89] Herbert F. (Herbert Frank) York. The advisors: Oppenheimer, Teller, and the superbomb. Stanford nuclear age series. Stanford University Press, Stanford, CA, USA, 1989. ISBN 0-8047-1713-3, 0-8047-1714-1 (paperback). xiv + 201 pp. LCCN UG1282.A8 Y67

1989. URL http://alsos.wlu.edu/information.aspx?id=442;
http://www.loc.gov/catdir/description/cam023/88062671.html;
 http://www.loc.gov/catdir/toc/cam027/88062671.html.

York:1995:AP

[Yor95] Herbert F. (Herbert Frank) York. Arms and the physicist, volume 12 of Masters of modern physics. American Institute of Physics, Woodbury, NY, USA, 1995. ISBN 1-56396-099-0. xiii + 294 pp. LCCN QC16.Y67 A3 1995.

Zeman:2004:ACH

[ZA04] Scott C. Zeman and Michael A. Amundson, editors. Atomic culture: how we learned to stop worrying and love the bomb. Atomic history and culture. University Press of Colorado, Boulder, CO, USA, 2004. ISBN 0-87081-763-9 (hardcover), 0-87081-764-7 (paperback). x + 187

pp. LCCN E169.12 Z46 2004. URL http://www.loc.gov/catdir/
toc/ecip0413/2004001010.html.

Zacharias:1976:BPS

[Zac76] Jerrold R. Zacharias. Books: Pragmatism, secrecy and moral values: The Advisors/Oppenheimer, Teller, and the superbomb, by Herbert F. York. Bulletin of the Atomic Scientists, 32(10):57–59, December 1976. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Yor76, Yor89].

ZamoraCollina:1995:LD

[Zam95] Tom Zamora Collina. Livermore on the defensive. Bulletin of the Atomic Scientists, 51(3):42–45, May/June 1995. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Ziegler:1991:BRS

[Zie91] Charles Ziegler. Book review: Stanley A. Blumberg and Louis G. Panos: Edward Teller: Giant of the Golden Age of Physics. Herbert F. York: The Advisors: Oppenheimer, Teller, and the Superbomb. J. Robert Oppenheimer: Atom and Void: Essays on Science and Community. Isis, 82(3):589–590, September 1991. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/233287.

Zimmermann:1992:LSL

[Zim92] Uwe Zimmermann. Die Logik des Schreckens — Leben und Werk des Physikers Edward Teller. (German) [The logic of terror — the life and work of physicist Edward Teller]. 60-minute film, 1992.

Zorpette:2024:QRG

[Zor24] Glenn Zorpette. 5 questions: Richard Garwin: The designer of the first hydrogen bomb on creating “the sausage”. IEEE Spectrum, 61(9):22–23, September 2024. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Zuckerman:1988:NWB

[Zuc88] Lord Zuckerman. Nuclear wizards: Rabi: Scientist and Citizen by John S. Rigden. Toward a Livable World: Leo Szilard and the Crusade for Nuclear Arms Control edited by Helen S. Hawkins, edited by G. Allen Greb, edited by Gertrud Weiss Szilard. Better a Shield than a Sword: Perspectives on Defense and Technology by Edward Teller. New York Review of Books, 35(5):26–31, March 31, 1988. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/archives/1988/mar/31/nuclear-wizards/.